A Complete Bibliography of Publications in Studies in Applied Mathematics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

19 August 2023 Version 1.00

Title word cross-reference

(2 + 1) [DA05, ACR23, JTLW13, RA10]. (m, n) [Bri90]. (P ± i0)^{λ} [Tri99]. 1 [Dya20, LWC23]. 1 + 1 [KR22]. 1/2 [Vár13]. 2 [BW02, BD17, CI23, CLW23, LLS04, MWZ22, McG23, RRY22, SCFFLPA23]. 2 + 1 [VPE09]. 2π [Min86]. 3 [CCY23, LST06, MWZ22]. 3x + 1 [Ven92]. ^{†} [NLB10]. α [Gre74]. β [FT19, GY93, NC15b, Swa18, Zen91]. C(X)∗ [BD70]. PT [CCMH21, CZ15, DKK^{+}14, GP14, JA16, Kon14a, Kon14b, LKM14, MK14, Mos14, NY14, NY16]. d [LY13]. Δ [CZjZ21]. δ^{(k)}(m^{2} + P) [AT90]. e^{−np(x)} cos nx [Min86]. e^{−np(x)} sin nx [Min86]. ϵ [BHH^{+}11]. f [NC15a]. f(X_{1}) + ⋯ + f(X_{n}) = 0 [Ric84]. GP (m, n) [Yan14a]. ∞ [PP19]. K [Tri99]. L^{1} [Ark92]. L^{2} [Lu23, YRZ20]. L_{2} [Haj92]. L_{∞} [Haj92]. λ [WF12]. C^{4} [Bri90]. CP [HK04]. N [AFY15, CKL16, CFY17, EM71, GXW13, JTLW13, OWY11, RS08, SJV96, SRJ99, SY03, Tho69b, AT07a, BG83, Fow90b, JZ86, LTY22, MC18, Pec80c, Vas11]. N + 1 [Rog14a]. N = 1 [TW17]. N = 2 [FH10]. n ≥ 0 [Min86]. O(2) [Bri90, DB90]. p [SO20]. p(x) [Min86]. P_{4} [JO89a]. P_{5} [JO89c]. P_{II} [Nis07]. ψ_{xxx} + 6Qψ_{x} + 6R_{ψ} = λψ [Kau80b]. Q [GST14, Cao13, DIW19, IL20, Jos15, KSV01, KM18, KV06]. q_{t} = P (x, t, q, q_{x}, q_{xx}) [Sat99]. q_{xy} = F (q, q_{y}) [BN94]. qd [PCH23]. r [Fow84]. R^{2} [WW03]. R^{N} [Gra73]. S_{2} [KMM80b]. σ [WXZ18]. so(3, R) [MR02]. t [Nau92]. × [GWH11, Pat70, XF16]. U(1) [ACS18]. u′′′ + λ^{2}zu′ + (α − 1)λ^{2}u = 0 [Her76]. u_{tt} = F (u, u_{x}, u_{xx}, u_{t}) [JT87]. u_{xx} + 2xu_{x}/\sqrt{1 − αu} = 0 [IB15]. v [EM71]. X [Abr72, ABGML72, Gal71, Nau92]. Y [Abr72, ABGML72, Gal71]. Y′′ + G(z)Y = 0 [SV99]. Z [She82b].

-Algorithm [BHH^{+}11]. -Askey [KM18]. -Binomial [KSV01]. -component [LTY22]. -critical [YRZ20]. -Difference [Jos15]. -Dimensional [BG83, JTLW13, RA10, ACR23, KR22, Rog14a]. -dynamos [Gre74]. -Extension [Zen91]. -finite [SO20]. -Form [WXZ18]. -Free [JO89c]. -Function [GST14]. -Integrals [Cao13]. -Invariant [Bri90, LKM14, ACS18]. -Orthogonal [LY13, IL20]. -Plane [GY93, NC15a, NC15b, Swa18]. -Polylaplacian [PP19]. -polynomials [DIW19]. -Reducible [JO89a]. -Species [AT07a]. -Stirling [KV06]. -supercritical [YRZ20]. -Symmetric [CZ15, DKK^{+}14, GP14, MK14, Mos14, NY14, NY16, CCMH21]. -Symmetry [JA16, Kon14a, Kon14b]. -th [EM71]. -transform [Tri99]. -type [PCH23]. -Wave [Vas11, RS08]. -Waves [SRJ99].

1 [Cos06, NC15a, Rei74a, Sta71a, XK93]. 1-Dimensional [RP11, SAR13, NS98]. 1-K [YK94a]. 115 [Ano05b]. 122 [BMOT09a]. 19 [GKK21].

2 [BS22, BS23, Ben06, NC15b, Rei74b, Sta71b, WW21]. 2-D [LMS19, BS22, BS23, WW21].

4-d [KL00].

69 [UG84].

80th [CD21a, KP19a, KP19b].

Abel [OF23, Sch98]. Ablowitz [BF16a, BF16b, GDZ07, GHMT08, Kok20, OP19, PV16, ZC10a, ZC10b]. Ablowitz-Segur [Kok20]. ABS [ZZ13]. Absence [HY86, Mil79]. Absolute [FP05, TU05]. Absolutely [TG01]. Absorbance [BB87]. Absorption [AH95, BW79, CCD22, Gal10b]. Absorption/emission [CCD22]. Abyssal [Swa18]. Accelerated [YL08]. Access [WA83, Wan85]. accessory [XXZ21]. Accounting [RANC09]. Accuracy [MO21]. Ackerberg [Sri88, WY03]. Acoustic [AT07b]. Acoustical [SV87]. Acoustics [MRS88]. Across [BPS73, SS04, BH94, ZC21]. Action [CL91, Gub74, Mos14, LP23]. Activators [Kee78]. Acyclic [Coh91]. ad [MS08]. Addenda [ABT87].

Addendum [Ran96b, Ran96a]. Addition [Air81, FMRR77]. Adiabatic [Hu98, Kev82, KL84]. Adjacent [KKR71, Pec80b, Wes78]. Adjoint [SR98, XB22]. Admits [Ifi10]. Admitting [MR02, PS20b]. Advection [How86a, How86b, How90b, KO04, LO95, LO99, YLW21]. advective [CLW22]. Aedes [ELV21]. aegypti [ELV21]. Aerodynamic [Cro70]. Affine [Bry75]. Age [WLS17, GYZ21, LCTZ22]. Age-Structured [WLS17, GYZ21, LCTZ22]. Aggregation [LNVW05]. Aging [CW70, Rei70b]. Agreement [MS08]. Air [Mas72]. airway [FCP^{+}21]. Airy [CD18a, Cla16, FG12]. AKNS [CDLjZ18, DvdM23, Tro21, XL15, Yu15, ZC10b]. Algebra [Gre72c, Gre72a, MR02]. Algebraic [Ber92, FH09, FOX23, SLN16]. Algebraically [LF02]. Algebras [LRS04, Sat85, SV99, SS14, dB75]. Algorithm [BHH^{+}11, MS08, Sri92]. Algorithms [CKL16, CH18, RLW01, Ulr75]. Allied [Nem18]. Almost [New02, BC23]. almost-circular [BC23]. along [VA02]. Alpha [Kun78, STZZ14]. Alternate [GMM05]. Alternative [BC85, CC17, Lin02]. Always [New02]. American [Kne01]. Amplitude [BMM87, BG07, BC09, BK78b, Ben79b, BG82, BMOT09a, BMOT09b, BM12b, CS80a, GY93, GPT^{+}05, Hic83a, MM90, Mes96, MB84, Mor86, PR94, TCK82, Tur07, VS85, War83, ZS86b, MC21b, Phi98]. amplitudes [Wri20]. analog [Ryb23]. Analogues [Yu15]. Analyses [LW95, YBSM08b]. Analysis [AKNS74, AHNZ09, AMR91a, AMR91b, AKK^{+}09, AB10, Ben81, CV09, DN11, DS71a, FMRR77, FS69, Fow84, Gao95, GV03, HH94, JW06, Jos99, KMM82, KM12, KLW11, Lan87b, LS88, LM91, LS93, Le 04, Li05, LNW04, LK10, MBT00, MM90, MK09, MR80, NC15b, Nis07, Nou00, ÖM94, PS11, PAS88, PCZP00, STC09, Sch98, Tro17, TNM96, VA03, VM18, WWZ10, WHS17, War96, WM94, Yan14a, YM04, YM07, YKdM^{+}07, YBSM08a, YBSM08b, ZZ08, ZK13, ZHY09, CCF19, CLW22, DDAAB20, EWW22, FMX22, LSG22, LCTZ22, PP19, XHW22, ZYZ22]. Analytic [CR06, CM09, Har98, HM86, IR86, Ioa89, LT02, Lu89, SRJ99, SKB08, TG01, ZL17, DL23]. Analytical [CFY17, GGS95, GT90, Ran96c, SF85, Yan14a, Yan21, JFM21, Luo21]. Analyticity [CW21, Nic21]. Anger [Dun22]. angle [LBP21]. Angular [Ber92, Che69, HG82]. Anharmonic [TK69]. Animals [Hol80]. Anisotropic [QK94, RW71]. Annihilators [Li76]. Annular [GP87]. Annulus [LAHM17]. Antagonistic [MMPS13]. Antarctic [MQ22]. Anti [ACT87, BH16]. Anti-Self-Dual [ACT87, BH16]. antisymmetric [KHJK21]. any [Gro70]. Appearing [KD17]. Appl [BMOT09a, UG84]. Application [Cau18, CK09, Che11, Dun01, Dun04, Dun11, JQS14, KR03, LW91, Li06, Lin97, LP11, MM01a, RKS83, SRD04, Sta71a, Sta71b, FDL21, FLS23, MMRL23, PCH23, XYZY22]. Applications [Abl71, HLQ09, KL88, MS08, PS93, WK91, ASSV20, CD18b, ML20, Rat19, JlPG79]. APPLIED [Ano06]. Approach [ABGML71, ASSS12, Ash10, BC85, Cla78, Cro00, DK05, FN05b, JK92, KKM87, Kne00, KSW14, LT07, Lu89, Ran96c, Sha91, TG01, TR02, WW05, WTW13, Wil72, XZ13, ZZ13, CF22, DFG19, DFS22, IL20, LQYjZ22, zLWjZ22, Sin21, Tro21, WZ22]. Approaches [Lin02, CLV20]. Approximants [Bak76]. Approximate [Abl72, GGS95]. Approximated [LR69]. approximating [MO21]. Approximation [DLK16, EM71, GP14, JW99, KT99, KO13, LST06, Ors72, Sun93, FLS23, LLZ23, Wu21]. Approximations [ACZ12, FCGB95, GST18, KL88, Lip06, LT99, Mar18, Mil75, Rei72, Rei74a, Rei74b, RMA12, FA22]. Aqueous [ASSS12]. arbitrarily [OP19]. Arbitrary [DPHJ03, PT18, Ran98, Sch01, Sob04, Sob06, Yan02, ZL17, Dur22, Ryb22, YS21]. Arc [MPG17]. Argument [Nem18]. Arise [AO03]. Arises [KT99]. Arising [AM09, BS88, BS93, CV09, Kne01, MRS88, MPB00, PSS75, PMAD07, Van87, Wei84]. Array [SS81, PSAZ22]. arrays [ADH^{+}21, SCFFLPA23]. Arrival [Hel74, KM12]. Arrivals [Bar73]. Askey [KM18]. Aspects [Bri94, DB94, NAY17, SAR13]. assess [GLH22]. Associated [AV91, Ben79a, Dun04, FC08, FA83b, GWH11, GŠ06, JT87, NRGO01, RS14, AE22, FN22, HW20, VW23, XJN21, YY23, dlIJ23]. Asymmetric [JZ86, TWK00, WW02, WW03, GWCH23]. Asymptotic [Abd97, AS77, AHNZ09, AMR91a, AMR91b, BOGK04, BL87, BK95, BS16, BS89, BS99, BS00, CKLS96, Dun11, Dun20, DH96b, FR17, FLS23, GST14, GST18, Hwa97, JW06, JW69, JK92, Jos99, KD17, KM12, KLW11, Lan87b, Laq81, LW95, Lev69, LR69, LF02, LS04, LP11, Mac97, Mar18, Mil75, Mil79, MK09, NR77, NA88, Nau92, Nem18, NPT08, NC15b, ÖM94, PW12, PS10, RS08, Rei74a, Rei74b, RW96, SK23, Sri91, SH76, Sta76, Stu92, Tem93, TR02, VP10, Wan77, War96, WL92, WZ07, Woo95, Wor86, Xu90, YLP95, ZK13, ZZ07, Dun22, LS23, LCG21, LDLW20, MC21a, AK82]. Asymptotic-Numerical [YLP95]. Asymptotically [Ros76a, Fro19]. Asymptotics [BL70, BDGL02, CM08, CK04, DK09, DIW19, DZ22a, GLN14, Hu96, KSV01, KK91, KS91, KV06, LO99, LDLW20, MMPS13, NdJVA22, NAY13, NY14, NAY17, PW14, RS11, SER^{+}03, SRE05, SVS08, SAM17, SW17, WW05, WQW13, WB98, XZ11, XZ13, ZXZ15, AK20, BCD21, DL23, HW20, IL20, NMBD23, PRAV21, RS21, WX22, XF23b]. Atlanta [RKS83]. Atmosphere [Saf72]. Attached [SS77]. Attachment [AB03]. Attachment-Line [AB03]. Attracting [HN96]. Attractor [CJN04]. Attractors [You13]. August [Mal02]. Author [Ano73, Ano76a, Ano76b, Ano78, Ano80a, Ano80b, Ano81a, Ano82a, Ano83a, Ano84a, Ano85a, Ano86a, Ano87a, Ano88a, Ano89a, Ano90a, Ano91a, Ano92a, Ano94a, Ano95, Ano96, Ano98, Ano99, Ano00, Ano04, Ano05a, Ano06]. Author-Title [Ano73, Ano76a, Ano78]. Auto [RS98]. Automata [FPSA89, Kfo70, PAS88, PF89]. Automorphism [GY15]. Autonomous [LS75, LS77, BM22]. autoresonant [Sul20]. Auxiliary [JTLW13]. Avalanche [DARB11]. avascular [JFM21]. Averaged [DS71b, DS71a, Dre71, LBP21]. Avoiding [ELY03, LMY08]. Axially [Kro77]. axis [AK20]. Axisymmetric [Ack73, GMW98, LW88, Mar71, Wor86, EWW22]. Axisymmetrical [Xu90].

Background [BM14, CCMH21, FLT20, LFW20, OP19]. backgrounds [Ryb22]. Bäcklund [BCH95, BT88, BRT89, DRS02, Fan11, JQ13, JQS14, ML20, MCB97, QJW07, RS98, RS10a, Sch01, XL15]. Backward [BSGC99, BSCS00, GS23]. Balanced [Nou00]. Ball [Gra73]. Ballot [Oli88]. Banach [SV99]. Bands [CP12]. Baroclinic [BM04, KS96, MB84]. Barotropic [BM04, BMB02, LYZ20, SP88a]. Barring [She82a, She82b]. Based [DGW02, LK16, BK23, CHSV22, CH18, DS23]. Bases [VS95a]. Basic [KX18, BL22]. Basins [Tro17]. Basis [ASSS12]. batch [KHJK21]. Bathymetric [RANC09]. Bathymetry [LN13]. Baxter [APT06, CGP23, dB75]. Beam [Rei73b, BE22, FCCTR22]. Beds [DS71a, EM09, Hal06]. beetles [XYZY22]. Behavior [DNP17, Gaa82, KC82, Kee82, LM91, Lev69, Nau92, PS10, Rei70a, RW96, SS95, STC09, ST17, XK93, CKS23, – Dor22, KK22b, LCG21]. Behaviors [VP10, YBSM08a, CCY23, LS23]. Bell [JTLW13, WTW13]. Bell-Polynomial [WTW13]. Belonging [Gea85]. Belousov [SH76, Sta76]. Benard [SW81, Sar01, Wal82]. Bending [SW74, Wan70c, SK23]. bending-torsion [SK23]. benefits [Wan19]. Benjamin [ABM^{+}22, CEHS21, FA83a, GLS23, KM98, MW16, NS21, RRY22]. Benney [AAL02, AY17a, AY17b, CS15, Sta15]. Beppo [Tri92]. Berman [CK04]. Bernoulli [LT99, SY15]. Besov [Wu21]. Bessel [BP71, CFS92, Dun01, RV18]. Beta [Oxl82, RV18]. Between [Ben76, Ben77, RD04, Cau18, GB96, GG84, Gea85, GZ94, Hu98, KR78, LKK80, LY13, MB95, Mil95, Ste19, TNM96, WM94, YY18]. Beyond [Wal95, Hu96, KS91, LLZ23]. Bi [Bła11, HCC21]. bi-periodic [HCC21]. Bi-Presymplectic [Bła11]. Bialgebras [JR79]. Bianchi [BH16]. Bickley [BMB02, MD99]. Bidirectional [CJ18]. Bifurcating [DR77]. Bifurcation [CS85, CJM92, CJ18, DH02, EWW22, Kat81, Kee76, Lan80, LS77, LSG22, Lu93, Ma82, NY16, RC80, RC81, SW81, Wax80, GS23, GM21, XHW22]. Bifurcations [BZZ21, Fow90b, Fow90a, MR84b, ST17, Sul20, Yan12]. Biharmonic [DF15, KLW11]. Bijective [Sen89]. bilateral [dlIJ23]. Bilayer [Mos14]. Bilinear [Fan11, HLQ09, JTLW13, zLWjZ22]. Bilinearization [bXlZS23]. Binary [KS06, Kun86, SR98, Ryb22, Ryb23]. Binding [ACZ12]. Binomial [KSV01, Loe90, Rei77, Rei81, Sob05, Sob06, Sta87, Wag94]. Binomial-Type [Sob05]. Biochemical [Boa75]. biological [Wan19]. Biorthogonal [AMT18, JRS99, Tum06, VS95a, LT23]. Birefringent [Yan96, Yan97]. Birkhoff [CS80c]. Birth [ZGQ12, dlIJ23]. Birth-Death [ZGQ12]. birthday [CD21a, KP19a, KP19b]. bistable [MMO23]. BKP [KS02]. black [BHN21]. Bloch [ALZ21, CP12]. Blocking [TZHL09]. Blow [Gal08, Gal10a, Gal11, KS18, DWZ21, MEHS22, TYZZ21, YRZ20]. Blow-Up [Gal08, Gal10a, Gal11, KS18, DWZ21, MEHS22, TYZZ21, YRZ20]. Blowup [DFKZ14, DFK16]. Bodies [TW97, Wan93, Weg03]. Body [Cal16, TWK00, Vár09, YM04, YM07, YKdM^{+}07, MO21]. Bogoyavlenskii [VM18]. Bogoyavlensky [Wan12]. Boltzmann [Ark92]. Bona [ABM^{+}22, CEHS21, GLS23]. bonded [Gab21]. Boolean [FMRR77, Har75].

Boomeron [DRS02]. Borders [Mal01a]. Bore [YK94b]. Borel [BS99, BS00]. Bores [EGP01, EGK05, YGJ18, Pri23]. Bose [WTW13]. Boson [Fan11]. both [RF13]. Bottom [Ben92, O’N69, RP83, Swa18, BP20, CDF^{+}22]. Bottoms [MPS98]. Bound [Hic85, Sta75, Wan14]. Boundaries [Gre84, Kne00, Ors71a, Ric92, Tan86]. Boundary [AFLM18, AO70, AB03, AB07a, BMM87, BB12, BS01, CRW16, CKS23, CE73, DPvdMV13, DB96, DB03, DH96b, FR08, Gaj96, GLZ15, HHA86, HS84, Her82, HM15, How77, IR86, Isr72a, JS01, Kat85, KKM87, KC05, Kne01, Lak72, LM91, LW95, Lia06, LLC10, Lin97, Mal94, MQ15, OB94, OW03, OW04, PMAD07, PPV08, PP08, PS18, PBT11, PV16, RS80, Rus73, SBA05, SKB08, Saf71a, SS06, SY06, Tro17, TG01, TU05, Tur05a, Tur05b, Tur07, UG83, Vla08, Wad91, Wan73b, War92, WY02, WXZ18, XF16, YK99, YK02a, YK02b, Zau72, CCD22, CBKM20, Dya22, FMX22, FvdW21, Fro19, LMS19, MZ23, PV23, SA19, SK23, XF23b, Zha19, ZS22, FR01]. Boundary-Layer [HS84, IR86, Kat85, Lia06, Lin97, TU05, YK99]. Boundary-Value [How77, KKM87, OW03, OW04, RS80, YK02a, YK02b]. Bounded [Amu17, Her91, Rod03, WNH12, SA19]. boundedness [Jin23]. Bounding [Rod03]. Bounds [BSGC99, BSCS00, KSV01, NR77, Nem18, War78, Seg21]. Boussinesq [ABCD17, BM12b, EHS18, Gal08, Hic83b, HM15, JGM21, KMP14, KT19, Lou19, MWZ22, NS21, PW97, Rat18, RS14]. Boussinesq-KdV [Lou19]. Boussinesq-Klein [KT19]. Boussinesq-Type [KMP14]. Boutroux [JK01]. Box [BS77]. brackets [CLV20]. Bragg [DA05]. Braiding [Hal06]. Branch [DLK16, HG81, Tur05a, Tur05b]. Branching [New91]. Break [EM09]. Breakdown [Bri15, MR84b]. Breaking [EGK05, GHJ12, GZ13, JTLW13, JA16, Tov00, VM18, Cai22a, Cai22b, dSF20]. Breakup [KWW06]. Breather [CFiMO18, KR22, CF23, FLT20]. Breathers [CMKP16, GR23, RMMH22, Slu19]. Bricks [DDNP90, DDNP91]. Briggs [Lin97]. Bright [CFiMO18, CF23]. Brinkman [GK95, PS99a]. Brittle [JO89b]. Broer [RP11]. BSQ [ZZN12]. Bubble [Gol04, Gri12]. Bubbly [GT04]. Buckling [Gao95, RJM78]. Bulk [BK78a]. Bulk-Service [BK78a]. Bullets [DA05]. Buoyant [DB03]. Bureau [Cos00b, Cos06]. Burger [Vil04]. Burgers’ [Lea16, MR93, MB03, SLN16, Cro98, JS01, LO99, NdJVA22, NMBD23, RS08, RY10, RS11, RN15, SJV96, SRJ99, SER^{+}03, SRE05, SAM17, VA03, VK06, VP10, VJ11, Vag12]. Burnside [Pec79].

Cahn [Hol23, SW00]. Calapso [Sch01]. Calculating [TG01]. Calculation [Ran96c]. Calculations [Str86]. Calculus [JG79, JlPG79, JG80, Lap87, Mor80, PS93, SJB91]. Calls [Leb73]. Camassa [GXW13, GMNP22, GZ10, GZ13, Kau06, SYF22, SS14, dSF20]. Can [BK23, Yan14b, Gro70]. Cancer [WSK10, Jin23, JFM21]. Canonical [BLP88, Dec00, Dya20, DS23, MRS88, FLS23]. Cantilevered [LW90]. Cantor [BD70, HLS03]. Cap [GMW98]. Capacities [KM12]. Capacity [ZK13]. Capillarity [Cro00, RS14, MKP^{+}20]. Capillary [AM08, AM09, AN13, Bri90, CS80a, CS85, DB90, GMW21, WVBM14, CC19, Dya22, EWW22, JW20, MC21b]. Capillary-Gravity [AM09, Bri90, DB90, GMW21]. carcinoma [WCYW19]. Cardiovascular [YM07, YKdM^{+}07]. Carrier [MBT00, WZ08]. Cartesian [AFY15]. Case [BDGL02, Gri90, SV94, SV95, SR99, ZC10a, MMRL23, RRY22]. catalogue [FÖK22]. Catastrophe [KO13, FLS23]. Categorical [BRSV92]. Cat’s [Ber70]. Cauchy [KMP14, KT19, LQYjZ22, NR22, QK14, VA02, ZZ13, Zhe20]. Causal [Tri76]. Caused [Gre77, Vla08, Lou19]. Caustic [BC76, Kne00]. Caustics [HjS85]. CBC [GY15]. Cell [Wan14]. Cell-Surface [Wan14]. Cells [WLS17]. Cellular [FPSA89, PF89]. Center [DH02]. Central [KH81]. Central-Force [KH81]. Cerebellum [Gro69]. Certain [AS89, BW86, Cao13, ERK^{+}81, Hei71, Hu96, JTLW13, Nis79, PS93, SY15]. Chain [BR73, Bar77, DL23, dlIJ23]. Chains [PB03, LGZZ21]. Challenges [STWW03]. Chance [FY01]. Change [GPT98, Gri15]. changing [XHW22]. Channel [BH89, Ben92, DH96a, Hal12, HS92, KA88, KS96, RP83, SY06, VB76, YKH00, YY03]. channels [BCCM23]. Chaos [HTWW98, Li06, VGW15]. Chaotic [Kee82, SS95]. Chaplygin [Xu21]. characteristic [BR19]. Characteristics [BZ92, HH94, HN96, BR19]. Characterization [DL94, Mon74]. Characterizations [RJRTF17]. Charlier [LY89]. Chazy [ACH99, Bag08, CA10, Cos00a]. Chebyshev [GLN14, PW12, PW14]. Checkerboard [She82a, She82b]. Chemical [LW91, PSS75, WK91, ZGQ12]. Chemotactic [MMPS13, Lóp22]. Chemotaxis [PS09]. Chess [Iwa82]. Chimney [LR01]. Chlamydia [WE17]. Choquard [Lie77]. Chree [ZZ08]. Christoffel [AMT18]. Circle [PRNA01, AP23]. Circuits [Wes78]. Circular [BFM17, GJ84, MPG17, STC09, Wal01, Aik19, BC23, DDM21, PSAZ22]. Circular-Arc-Shaped [MPG17]. Circumpolar [MQ22]. Class [AL77, BB69c, CKL16, CS93, Cos00b, Cos06, DIT16, JO89c, JO89b, JO89a, Kau80b, LR69, PS10, PEE04, Ran94a, Sob04, SP88b, Sri92, SY15, CCK^{+}21, CCMX22, DER21, DWZ21, DZ22b, FOX23, XXZ21, YS21, ZSA22, ZS22]. Classes [Cos00a, Har75, Hei71, PS93]. Classical [BH14a, CS90, DKK^{+}14, HSS17, KO13, Stu92, CCMX22, JGM21]. Classification [Cos93a, Cos93b, CS93, Cos97, Fre72, HZ10, LPH03, LLL12, MNW07, MB09, Qu97, Sat99, Sob06, SS14, TW17, Yan12, Hol23, JFM21]. Classifications [Liu13]. Clear [Mas72]. cloaking [LLZ23]. Closed [Bea70, Ioa89, MR93]. Closed-Form [Ioa89, MR93]. Closure [SW07]. Closures [BN69, JMT08]. Cloverleaves [Wat79]. Cluster [DGW02, RKS83]. clustering [KHS22]. Clusters [LB69]. CMV [AMT18]. Cnoidal [APS98]. Coalescence [HJ96]. Coalgebras [JR79]. Coarsening [SW00, War94]. Coated [LG82]. Codazzi [CG16]. Codes [Gre76]. Coefficient [GP04, Gri15, HL11, Mos79, TZ14, WTL^{+}10, WCYW19]. Coefficients [BDS20, Kau06, KD17, KSV01, Lea16, NA88, SV92, Sri91, SLN16, Vag12, VS88a, Ven92, YK02b, DFS22, PV23]. cognitive [LWZ23].

Cohen [Sta75]. Cohen-Macaulay [Sta75]. Coherent [BBC^{+}05, MMRL23, Yan96, BO20]. Cold [KT89, KT90]. Cold-Fluid [KT89, KT90]. Cole [Vag12]. Collapse [BG83, Gol04, SB00]. Collapses [War94]. Collective [YBSM08a]. collision [RCMH21]. Colonies [MMPS13]. colored [CCY23, Gab21]. Column [FN05b]. Combinations [Iwa82]. Combinatorial [And71, BRSV92, CL91, CR70, Dou72, DRS74, ELY06, GR70, VS88b, VS91, CR70, JG79]. Combinatorics [JR79, LY89, Rei78, She99, Wat79]. Combined [OK10, CI23, Sul20]. Combustion [Gaa82, Tam88, Tam89a, Tam89b, TA94, YG01]. Comment [TNM96]. Communication [Leb73]. communities [PRAV21]. Commutativity [LWAO21]. Compactly [JRS99]. Comparison [Ors72]. Competition [CJN04, CLW22, GJLW22, KK22a, WPO22]. Complementary [GKM74]. Complete [Hol23, LLL12, Liu13, Ma78, RJM78, KHS22]. Completely [RCMH21]. Completeness [LM74, Min86]. Completion [RLW01]. Complex [BOGK04, BH14a, Che69, FMO17, Har98, KL81, Mar18, NY16, RS12, SFO16, SZWM16, SV99, CGP23, CZ20, GPFT22, KK22b, Lóp22, Yan21, ZSA22]. Complex-Lamellar [RS12]. Compliant [AB03]. Complicated [Gro70]. Component [GZ10, PBT11, XQZ15, AKZ16, LTY22, MZ23, RMMH22]. Components [Gro69, GZ13]. Composite [APT06, Rei72, SD16]. Compositionally [AM03, AM05]. Compositionally-Driven [AM03]. Compositions [Yeh95]. Compound [Lan83a]. Compressible [AFY15, CFY17, Gaj96, Gre84, How73, SP88a, SV95, SJH05, SVS08, TU05, Tur05a, Tur05b, Tur07, VS96, Wad91, AZ22, Wu21]. Computation [APT06, ASSS12, ET87, Fat74, KD17, PAS88, TA94, WTL^{+}10, YL08, DH19]. Computational [TEW91, CLV20]. Computer [Kne03, RD04]. Computer-Extended [RD04]. Computing [Yan15, CH18]. Concentration [RS10b, ASSV20]. Concerning [Bel77, Mon74, OAVS97]. condensate [Esc23]. Condensates [WTW13, BO20]. Condition [BM99, ERK^{+}81, Saf71a]. Conditional [JQ13, JQS14, Qu97, QJW07]. Conditioned [BW79, LW95]. Conditions [AFLM18, AB82, BB12, DPvdMV13, MM01a, PBT11, PV16, RC05, SY06, SS03a, BE22, BO20, CBKM20, FMX22, FvdW21, Gri19, LMS19, MZ23, XF23b]. conductive [FDL21]. Conductivity [CRW16, MWL23]. Cone [Bea70]. Cones [RV18]. configurational [WBZ21]. Configurations [Boo89, KKR71]. Confined [YW92]. Confinement [CCM12]. Confining [GP14]. Confluent [BHH^{+}11, DZ22a]. conformal [DS23]. Congruences [ELY06]. Conical [NA88, Wan70b]. Conjecture [Kal73, Pec80a, Plo02, Spe74, Sta75, AP23]. Conjectures [McL97]. Conjugate [CRW16]. Connected [Bel79, Mat89, Wan76a]. Connection [AK20, BH94, BS89, DDN91, GJ84, JK92, RP11, Rog14b, AZ22]. Connections [DRS02, SR98, Tov00, Wri06]. Consecutive [Tho69b, Pet22]. Consensus [CHSV22, KHJK21, KHS22]. Consensus-based [CHSV22]. Conservation [ABCD17, FCGB95, GHMT08, GM12, GM13, HM94a, JMT08, KB98, Miu74, MM01a, MM03, Nou00, Pen17, SS03a, SS04, SDS12, Wad78, YK02a, YK02b, ASSV20, BP20, Hol23, MP19, MP22]. Conservative [KDM22, ACS18, FT21]. Conserved [MM01b]. Consistency [BD70, JN21]. Consistent [Lou15]. Constancies [Gro73]. Constant [GG76, How90b, Wan80, DH19]. Constants [DDN91]. constraint [CZjZ21]. Constraints [AV91]. Constructing [Boy16]. Construction [Ber72, Fro19, Ran94b, Wan76a, BMD22, BDM22, BM22, Yan21]. Constructions [Iwa82]. Constructive [Ben81, Ran96c]. consumer [SSW22]. consumer-resource [SSW22]. Consumption [Wan80]. contact [LBP21]. contained [Pri23]. contaminated [WCTZ22]. Contamination [ASSS12]. Continuation [Har98, DL23]. continuity [Jin23]. Continuous [BG07, HP78, KPS80, Lan73, PS99a, Str11, YA03, BW19, Dur22, Ryb23, ZSA22]. Continuously [Swa12, TEW91]. Continuum [Lip06, MK14, Wan73a, bXlZS23]. Contour [Gro73]. Contributions [AAL02]. Control [GKK18, NPT08, WSK10, ZN17, CR21, GKK21, NR22, SSVS20, XYZY22]. Controllability [PPV08, Rus73]. Convection [BS77, Cha71, Fow85a, Fow85b, Fow93, GK95, Hal12, How90a, JQ13, Ken76, LR01, LS04, Lu93, Mor86, PR94, Qu97, SS06, SW81, Str11, Tam93, Wal82, YK99]. Convective [FP05, PW97, QK94, QC96, TW97, CBJCN20]. Convergence [Bak76, BP71, FT21, FCGB95, MB03, PS99a, PS99b]. Convergent [Dun01, Dun04, YL07]. Convex [BB69a, Che94, DF15]. conveying [FMX22]. Cooling [Fow85b]. Cooperative [AT07a]. Coordinates [Kau75, MRW07, KDM22, MQ22]. Copyright [Ano16a, Ano16i]. Core [Wan73b]. Cores [DG05]. corner [BS23]. Correction [Air81]. corrections [FT19]. Correlation [Abr72, YBSM08b]. Correlations [WF12]. Correspondence [DD11, MV13]. Corrigenda [ABT87]. Corrigendum [BMOT09a]. Corrosion [AKK^{+}09]. Cosmology [Mat89]. Cost [Ulr75]. Couette [CLZ91, Rei74a, Rei79]. Countable [PB03]. Counter [BSSS00]. counterpropagating [BLM21]. Counting [RT79]. Couple [Wax84]. Couple-Stress [Wax84]. Coupled [BM13, BLP01, CC17, CD18a, DNP17, DFKZ14, DFK16, FCL12, GLZ15, GA79, GA82, Gri86, Gri87, JA16, KM12, LKM14, LY11, Oht09, OWY11, PY05, Rog14b, Rog14a, Rub77, SFO16, STZZ14, WTW13, Aou20, BW19, CGP23, FN22, GPFT22, LC21, WZ22]. Coupling [AS89, AN96, PR94]. covered [DDM21]. COVID [GKK21]. COVID-19 [GKK21]. Coxeter [GY15]. CPA [Mos14]. CPA-Laser [Mos14]. Crapo [Oxl82]. Crenation [LG82, Lan83a]. Crest [BS01]. Criteria [Mar71, dSF20]. Criterion [CHR91, BHIS21]. Critical [Ben83, BN98, CG07, CM07, CG08, Cam04, FLM95, Hab72, KM70, Ken74, Le 04, MBM94, MC02, PS10, SB87, WW78, Cai22a, YRZ20]. Critical-Layer [Le 04]. Critical-Wall [FLM95]. Cross [DPHJ03, Gaj96, HF17, Wal01, Yan95a, ACM18, KT22, LWC23, SXZ22]. cross-diffusion [LWC23]. Cross-Flow [Gaj96]. Cross-Section [HF17, SXZ22]. Cross-Waves [Yan95a]. Crossing [FK77]. Crowning [YY04]. Crystal [BS93, KS91, SKP^{+}05, Xu90]. cubatures [CH18]. Cubic [CNW12, Kau80b, Ma79b, MA81, MV13, SZWM16]. Current [FM11, LM16, MQ22, SS85, CHS21, FMN19, GMW21]. current-topography [FMN19]. Currents [DMPS96, DMPS97, GL17, MM99b, MM01a, MPS98, Moo00, MP01, PMAD07, YC15]. Curvature [DM96, Gre72c, Gre72a]. curve [FN22]. Curved [BH89, DH96a, Gre84, RS21, SW74, Tan85]. Curves [CWT^{+}15, HK04, FOX23, GM21]. Cusped [HjS85]. Customer [KM12]. Cut [KA88, TNT21b, TNT21a]. Cut-off [KA88, TNT21b, TNT21a]. Cuts [DLK16]. Cycle [CJN04, Li06]. Cycle-Strange [CJN04]. Cycles [FK77, Pec80b]. Cyclic [CLY95, CGUGM20]. Cyclones [NC15a, NC15b]. Cylinder [Ben06, STC09, Seg21]. Cylinders [BB69a, Dun13, KR78]. Cylindrical [HF17, JW69, TW97, TWK00, Gri19].

D [LMS19, KL00, BS22, BS23, BW02, BD17, CCY23, CI23, CZjZ21, CLW23, DA05, Dya20, LLS04, LST06, LWC23, MWZ22, McG23, RRY22, SCFFLPA23, WW21]. DAE [RLW01]. DAHA [KM18]. D’Alembert [KT19]. D’Alembert-type [KT19]. Dam [EM09]. Dam-Break [EM09]. Damped [CFY17, Mil76, CBJCN20]. Damping [AAA08, CK13, GHO98, MG91, PY00, SVS08, VK06, VP10, DWZ21, DZ22b]. Darboux [ALZ21, GLL13, LY03, Ryb22, Ryb23, SR98, Vas11, XL15, Zho18]. Darboux-Manakov-Zakharov [Vas11]. Dark [CFiMO18, MZ23, OWY11, CF23, FMO17]. Data [Kau06, LLLZ17, LS04, MW16, NK05, RKS83, VA02, WK89, Lu23]. Davey [FKMS22, ISFH20, KS18, KM90, RCH17, Zho18]. David [AAL02, AY17a, AY17b]. Dbar [WZ22, Vil91]. Dbar-approach [WZ22]. Death [ZGQ12, dlIJ23]. Decay [LY11, PS09, PPV08, Saf71b]. Decaying [LF02, NMBD23]. Deck [Tum06]. decoding [PCH23]. Decomposition [GDZ07]. decompositions [FA22]. Decrease [Kat88, KK90]. Deep [AM08, AN13, AM05, Bry79a, CS80b, CS80a, CS85, DLK16, Kli03, KR03, Ma82, MB95, MP01, SS85, WVBM14, MC21b]. Deep-Water [KR03, Ma82, SS85]. Defect [FYC05]. Defocusing [BP14, DPvdMV13, FMO17, LN13, MQ15, DFG19, Fro19, OP19]. Deformation [Gre77, GMW98, GFGUMM22, NLB10]. Deformations [KL00, KA02, LW88]. deformed [DZ22a]. Degasperis [CI17]. Degeneracies [GS90]. Degeneracy [HM93a, Hoc72, Rat18, Tho69a, TK69]. Degenerate [Bri94, BGL97, KM18, XF23a, Xu21]. Degree [CS93, BCD21]. Degrees [Wan76a]. Delay [NPT08, GYZ21, LSG22]. Delays [AT07a]. Delbrück [ZGQ12]. Delbrück-Gillespie [ZGQ12]. Dendrites [PP86]. Denert [FZ90]. dengue [SSVS20]. Density [FL73, LL79, Nis79, Vár13, CBKM20]. Dependence [LZZ23, PS99a, Str11, Tam93]. Dependent [Ber92, CW96, FA83b, FP12, Fow93, HM94a, Isr70, Lea16, MM03, NC15a, NC15b, PS00, PP08, QC96, Ric92, SB87, SLN16]. Deposition [MPS98, MPB00]. Depth [AS80, DLW87, DLW89, GA11, HCB72, LHC04, Wal82, Yu10, ZS86b, GMW21, PTDS19, VMA21]. Derivation [BGT07, GM89, NR77, Hic83a]. Derivative [CFiMO18, GLL13, AE22, CZ20, EE19, Lu23, Wri20]. Derivatives [MB09, OAVS97, Sun93]. Descent [Lin97]. Descents [CFW87]. Describe [Rab89]. Described [GPT^{+}05]. Describing [FMM99, Gre72a, MR80, SDS12, Swa12, VS96, FdST22]. Description [AZ13, Fow93, CEHS19, ISFH20, Pet22]. Designable [HL11]. Designs [Gri78]. Detecting [AKK^{+}09]. Determinant [WF12, DZ22a]. Determinants [AGJ03, Kat88, KK90, ZXZ15, MC21a]. Determination [Ioa89]. Detonation [HS82b]. Developing [CHR91]. Development [DDAAB20, OB94, Sha91, Wad91, Wan19]. Developments [JG80]. Deviations [LK16]. Diagonal [Kat88, KK90, WF12]. Diagrams [Hsi73, Kle70, CGUGM20]. Dielectric [Dun13]. Difference [AL76, AL77, BL87, CR06, HM77, JO07, Jos15, LM91, Ors71b, Pen17, RH07, VS94, Wan73b, WL92, CCK^{+}21, IL20, JL19, KDM22, MC21a]. Differences [VS91, VS95b, Wag94]. Different [Gea85, KM12]. Differentiable [LY11]. Differential [And71, Bag08, BS88, BS89, BS93, Che84, Cos93a, Cos93b, CS93, Cos00a, DDMMSYD22, Dun01, Dun04, Dun11, DFS22, FS96, Fow90a, Her76, Hoc72, Kee82, KX18, LPH03, LM91, Lev69, LM74, LT07, LR69, MS16, Mat89, Mil01, MC21a, OK09, PSS75, Pen17, PG95, RLW01, Rus73, SV92, SBA05, Sob04, Sob05, Sob06, SV99, Tov00, VGW15, VS88a, VS94, Wan72, WZ07, Woo95, Zau72, ZL17, BE22, BJV22, BOPS21, Dem23, Dun20, Olv21, PS20b, SO20, YS21, YLW21, ZL21]. Differential-Difference [LM91, Pen17]. Differentially [BM82, O’N69]. Differentiation [WE17, IW23]. Diffraction [GRMOH01, GS15, MML17]. Diffusion [AH95, BDLP11, BL70, BGL97, CW96, Gal10a, GRMOH01, Gre72d, Her82, HK77, How86a, How86b, How90b, JQ13, JQS14, KQ10, KC82, Kee76, KO04, KWW05b, KH73, KH81, LO95, LO99, LL80, LS04, LK10, MR80, MTW16, Qu97, QJW07, STWW03, Vla17, War96, Wit96, WB98, CLW23, FT21, GSYG21, Jin23, LWZ23, LWC23, PRAV21, TNT21b, TNT21a, WCTZ22, YLW21]. Diffusion-Reaction [KO04]. Diffusive [GK95, LL80, Mas85, MM90, Mor89, WC95, GJLW22, SSW22]. Diffusivity [BS16, MM93]. Dike [GJ81]. Dilatational [Ber71]. Dimension [KWW05b, Sch01, SW00, Tho69a]. Dimensional [AZ13, Abr71, AN13, AFY15, BM72, BZ89, BG83, CS85, CBB87, CFY17, Cos93b, CLZ91, CS15, DHC00, FYC05, GMM05, HS82b, How90b, Ifi10, JTLW13, Kau80a, KWW05a, KWW06, Laq81, LAHM17, LCB99, Lun03, MK96, MW13, Mos79, NR77, Nec12, NAY13, OB94, Pec80c, Ran94a, Ran94b, Ran94c, Ran95a, Ran96b, Ran96a, Red80, Rei73b, RA10, RP11, Rog14a, Saf71b, ST84, STC09, SAR13, Sha91, Sta15, STWW03, TG06, TK69, TBKiM18, Vag12, VM18, War94, Wei85, YMBC04, ZGQ12, ZB92, Zho09, dB75, ACR23, BE22, BŠT23, BP20, ISFH20, JN21, KR22, KHF^{+}20, Lu23, MP19, NS98, PP19, RRY22, VSS98, XF23a, Yan21, YF21, YLW21]. Dimensionality [HS84]. Dimensions [EEP03, FR01, Fow90b, JZ86, Ran95b, Tho69b, Vár09, Vár13, NR22, TMPVB19]. Dimer [YY08, FH23]. Dimers [LKM14]. dipolar [Vaf20]. Dipole [PT18, YMBC04, KT22]. Dirac [BP14, LZZ23]. Direct [AE22, AB80, AB82, DvdM23, Gus86, Gus89, GS90, HK73, JK94, Ma84, Mas85, MW16, Mos78, NLB10, VA03, Wan73a, ZZN12]. Direct-Resonance [AB82]. Directed [Coh91]. Direction [Kat88, KK90, HCC21]. Dirichlet [BDLP11, DD11, FP12, Hwa14]. Dirichlet-to-Neumann [BDLP11, DD11, FP12]. disc [Dya22]. Discontinuities [CK14, LC21]. Discontinuity [PJ86]. Discontinuous [LS04, ASSV20]. Discrete [APT06, AKS^{+}08, Bar73, BP14, BW19, CCM12, CM13, GDZ07, HN96, HK04, LPH03, Lip06, NRGO01, PW12, PW14, Sch75, Sch01, STZZ14, TBKiM18, XZ13, XL15, Yu15, ZC10b, AE22, BV19, CCM19, CF22, FIM23, JN21, MFIGH22, PCH23, bXlZS23, XJN21]. Discrete-Time [Bar73]. Discretization [NS98, Oht09, Sch01]. Discretizations [GT80]. Discretized [LRS04]. Disease [GXLX16, GS23, ZYZ22]. disease-severity-structured [GS23]. Disintegration [APS98]. Disk [Hun69, Hun70, Mar71, Nis79, SS95, TG01, TU05, Tur05a, Tur05b, SO20]. Disks [RD04]. Disorder [Mal01a, Mal02]. Disordered [LN13]. dispersal [HMBYZ22, MMO23]. Dispersion [AKS^{+}08, Ben74, Bri15, DIT16, FG12, Gal11, HY86, Nis79, SZWM16, TR97, HP19, SYF22]. Dispersionless [SFO16]. Dispersive [AB70, Abl71, Abl72, Abl75, AB00, ALB09, AN96, AN98, BGT07, BG07, CEHS21, DDM13, GKT10, HjS85, How83, New91, NS21, SD15, TG06, VS85, Yan14a, CEHS19, HSS19]. Displacement [HY86, Rei73b]. Dissection [Hen72]. Dissipated [Hab88, SS04]. Dissipation [BH94, GA82, MTVE02, Nau92, YKH00, YY03, ZZ08]. Dissipative [HTWW98, VS85, Yan15, ZC21]. Distance [MN22]. Distinct [CM13, Pec80a]. Distinguished [Vla17]. Distributed [LNVW05, PPV08]. Distribution [Dou72, DKZ16, FIN94, HLS03, KS06, MC18, Wil72, YK94a, ACM18, GYZ21, SK23]. Distribution- [Wil72]. Distributional [AT90, Tél94]. Distributions [BFM17, Gez79, Gez80, Gzy00a, Hwa97, Kne00, RV18, Saf71b, Sch80, Tri76, KT22, MN22]. Disturbance [OH95, PCZP00]. Disturbances [AY95, Ben06, Ben84, Ber92, CLZ91, Gus89, Hic83a, Hic83b, MD99, Mal01b, SA04, SY06, Tur07, YK94b, YY03]. diversity [DDMMSYD22]. Divided [VS91, VS95b]. Dividers [AFN73]. Divisibility [Leb77]. division [GBSS21]. Dobinski [CY94]. Does [Ben06]. Domain [Car90, EHV14, HF17, KL84, PP08, PSG23, Xu21]. Domains [FP12, Fow90a, HHW94, Rod03, SD16, War96, WS99, PV23]. Double [GK95, Mas85, MC18, Rat18, CF22]. Double-Diffusive [GK95, Mas85]. Doubling [FM83b]. Doubly [Cha21, CS80c, FCL12]. Dove [WWZ10]. Down [Fri74]. Dpp} [LNWZ09]. Drag [QC96]. Dressing [CI17, Zha19, BW19]. Drift [LK10]. Drifts [Vla17]. Driven [AM03, AM05, Gub74, Moo00, MP01, MPD05, VM11, CCY23, GYZ21, LWZ23, WBZ21]. Droplet [Gre77, Lac82, Lan83a]. Droplets [LG82, NS91]. Drosophila [LNWZ09]. Ds [KL00]. DSII [Li05]. Dual [ACT87, BH16, CSS77, LQYjZ22]. Dualities [KM18]. Duality [ACH99, Gao95, GY95b, MW73]. Duals [JRS99]. duct [SXZ22]. ductal [WCYW19]. Due [GPT98, GHO98, LN13, Tur05a, Tur05b, KA19]. Dumbbell [WS99]. Dumbbell-Shaped [WS99]. Durfee [GUGM18]. during [YM04]. Dyck [ELY03, LMY08]. Dym [BM13]. Dynamic [EY91, GY95a, Gri10, JW69, Mar71, YF05, ZYZ22, LWAO21, LBP21, MKP^{+}20]. Dynamical [FN05b, LCTZ22, SR97a, SR97b, Wan73a]. Dynamics [AHNZ09, AZ13, ABGML71, ABGML72, BG23, Bar83, CZ21, CV09, CLJ97, CZ15, GOS11, Gro02, GWCH23, HMBYZ22, JC02, Kee85, Kon96, KL00, KA02, KO13, NC15a, NC15b, Peg88, RSCM05, SR99, SW00, STWW03, VPE09, WLS17, WS99, Wax84, Yan01, YLP95, ZN17, ACS18, BO20, CCY23, FCP^{+}21, JP22, KR23, LWZ23, Lóp22, PSAZ22, SSW22, WCTZ22, YRZ20]. Dynamo [Gub74, Ken76, Lu89, SS95]. dynamos [Gre74].

Earth [Fow86]. Eccentric [KR78]. ECM [Jin23]. eco [ZYZ22]. eco-epidemiological [ZYZ22]. Ectopic [LNWZ09]. Eddies [VB83]. Edge [Ack73, Gre70, NA88, Wan76a, XZ11, Yan95b, FT19]. Edge-Connected [Wan76a]. Edge-Loaded [NA88]. Effect [BM04, DR90, GA82, GK95, MM93, Mos79]. Effects [AM03, AM05, BFM17, CWT^{+}15, FSY78, Gre90, HS82a, HS84, HG01, LB95, MG91, MP01, Tur05a, Yu10, YM07, YKdM^{+}07, MKP^{+}20, PSG23, SA19, Ste19]. Efficient [Leb73, NR22]. Eigenfunction [HS92, Tum06, CZjZ21]. Eigenfunctions [FG12, FS86, Yan02]. Eigenproblem [BN98]. Eigenvalue [BW86, BSGC99, BSCS00, HLS03, Hoc72, KS14, Kau80b, KLW11, LNW04, MTW16, Tho69a, TK69, Wan75, WHS17, WK91, War78, LT23, TV22]. Eigenvalues [BOGK04, BP14, Gra73, HG81, HG82, Kne00, RJRTF17, Tho69b, LZZ23, SK23]. Eigenvector [WHS17]. Eight [Bar84]. Eight-Vertex [Bar84]. Einstein [Gre72a, WTW13]. Either [New02]. Ekman [DH96b, Gre84, Gre89]. Elastic [CLS86, GP87, GMW98, Hun92, NV17, Rei70a, RS10b, Ros10, Wan70b, Wri85, BE22, LZW23]. Elasticity [Rei73a, Rei75, SRD04]. Elastodynamic [Rog14a]. Elastodynamical [LLZ23]. Elastodynamics [RS10a, SW07]. electric [FDL21]. Electrical [AO03]. Electrolyte [BPS73]. Electromagnetic [Dun13, KT22, PZ03, PT18, HHZ20, Vaf20]. Electron [KL88, MWL23]. Element [FS69, SO20]. Elementary [Hwa97, MC18, Tri88b]. Elements [Kat88, KK90]. Elevation [CG08]. Eliminating [War92]. Elimination [RLW01]. Ellipse [LK10]. Ellipses [Boy16]. ellipsoid [Vaf20]. Ellipsoidal [DF09]. Elliptic [CM09, HHW94, HF17, LST06, SAR13, BZZ21, FLT20, FN22, LS23]. elliptic-localized [LS23]. elliptical [AZ22]. Embedded [BG07, MK14, Wax80, Yan01, YA03]. Embeddings [Bry79b]. Embryo [LNWZ09]. Embryonic [SW07]. Emergence [Bri15, Rat18]. emergent [KK22b]. Emission [Cro70, CCD22]. encompassing [ELV21]. Endofunction [Yeh95]. Endomorphisms [Mor80]. energies [BŠT23]. Energy [CP12, Gri86, LKK80, MDG10, PPV08, SS04, TYZZ21, YBSM08b, CC19, Esc20]. Enhanced [MM93, MPD05, ADH^{+}21]. Enhancement [Gro73]. Ensemble [MC18, WXZ18, BC23, FT19]. Ensembles [BFM17, CI98, CZ10, LC21]. Entrainment [FH78]. Entries [CS80c]. entropy [HS19]. Enumeration [Cla78, CKLS96, GJ84, Gre76]. Enumerative [JG79, JlPG79, JG80]. Envelope [AZ13, PG96, TB02]. Environment [ASSS12, GXLX16, CLW22, LP23, LCTZ22, Vaf20, WPO22, XHW22]. environments [WCTZ22]. Epidemic [GKK18, LWZ23, MSW07, ZN17, GS23, LP23, LCTZ22, SSVS20, WCTZ22]. epidemics [Pet22]. epidemiological [ZYZ22]. Epidermal [SW07]. Equal [Tan86]. Equation [Abd97, AS77, AK82, AYF83, ACT87, AV91, ACH99, AM09, ABCD17, Ark92, Atk78, BSGC99, BSCS00, BDLP11, BCH95, BG83, BGL15, BMOT09a, BMOT09b, BB12, BP14, BMS18, BS88, BS89, BS01, Cal84, CK09, CA10, Che84, CM08, Cla16, CI17, CNW12, CS15, Dec00, DN11, DNP17, Dha94, DM96, DFS97, Fai97, FH10, Fan11, FCL12, FR17, FMO17, FSY78, FA83a, FA83b, FS86, Fok09, Gal08, Gal10a, Gal11, GXW13, GLZ15, GGS95, Gra73, GM89, GT90, GM93, Gri95, GP04, GPT^{+}05, GH08, GHJ12, Gri15, GS16, GZ10, HJ96, Har11, HN96, Her82, HMM87, Her76, Hic83b, HM15, HG81, HG82, HJ15, Hwa14, IB15, JW99, JW06, JTLW13, JZ10, Jos99, JK01, Jos15, KF10, KC82, Kau75, KM98, Kau06, Kee85, KL84]. Equation [KMP14, KM90, KD92, LO99, Lan87a, Lea16, Len16, Li05, Li06, Lie77, Lin14, LLL12, LLLZ17, Ma79b, MA81, MV13, MR93, MM90, Mil79, Mil93, MB03, MQ15, MW16, MCB97, Moo78, Mos78, Mos79, NPT08, NRGO01, NS98, PS10, PSS75, PP08, PBT11, Qu97, RS08, RY10, RS11, RN15, Rat18, Rei72, Rei74a, Rei74b, Rei79, RJRTF17, Ric84, Rod03, RS98, SRJ99, SRE05, SKB08, SAM17, Sch01, SR12, Sch98, SD16, She99, SZWM16, SB00, SV99, Sri91, SLN16, SW00, TZHL09, TZ14, Tov00, VK06, VJ11, VBA97, Var97, VA02, Vil04, VPE09, VM18, Wad78, WTL^{+}10, War94, Wei84, WZ07, Woo84, WXZ18, XF16, YA03, Yu15, Zau72, ZZ08, Zho09, Zho18, ACR19, ACR23, ABM^{+}22, AZ22, BS23, BLM21, BBN22, CI23, CBJCN20, CCD22, CGP23, CKS23]. equation [CF22, CLW23, CEHS21, DvdM23, DFG19, FT21, FLT20, FdST22, Fro19, GLS23, GMNP22, GPFT22, Gri19, GM21, Hol23, IL20, JW20, JP22, JTW22, JL19, JN21, KT19, KR22, Kok20, LGZZ21, LFW20, LQYjZ22, LS23, LCG21, LW22, Lóp22, Luo21, MZ23, ML20, NdJVA22, NMBD23, NS21, PCH23, PS20a, RCMH21, Rat19, RRY22, RS21, Ryb23, SYF22, Sin21, TYZZ21, WZ22, Wri20, Xu21, XF23b, YRZ20, YZ20, Zha19, ZTYH20, dSF20]. Equations [AL77, ABT86, ABT87, AS89, AW97, ACH99, AM17, AFLM18, Air79, Air81, AM08, AHP02, AFY15, AB07a, ALB09, AB10, Bag08, BF85, BT88, BRT89, BN94, BH16, BL87, BH97, BW02, Bri94, BLP01, BS93, CD04, CKL16, CCM12, CRW16, CC17, Che11, CM13, CWT^{+}15, CDLjZ18, CT86, CK14, CR06, CFY17, CD18a, CG16, Cos93a, Cos93b, CS93, Cos00a, Cos00b, Cos06, DLW87, DPvdMV13, DFKZ14, DFK16, DF15, DS71b, DS71a, Dre71, Dun01, Dun04, Dun11, DDM13, DH96b, EY91, EGP01, EEP03, EHV18, FR08, FA89, FIN94, FR01, FM83b, Fow84, Fow90a, Gal06, Gal10b, Gao95, GWH11, GG76, Gri81, Gri85, Gri16, Gri10, GLL13, Hab75, Hag81, HH94, Hal06, HL11, Hic83a, Hir09, HM77, HK77, How77, How86a, How86b, How90b, HTWW98, HY00, HLQ09]. Equations [Hun70, JQ13, JZ10, Joh10, JT87, JS01, JK94, JPS00, JO07, KQ10, KT89, KT90, Kee76, Kee82, Kic99, KLPS18, KO04, KK91, KM90, KM04, KH73, KH81, KX18, Kup06, LO95, LPH03, Lan80, LL80, LS88, LM91, LS93, Lev69, LM74, Li02, LY03, LHC04, LT07, LR69, LY11, Liu13, LCB99, Ma78, Ma79a, MRS88, MRW07, MC06, MS16, Mat89, MNW07, Mil01, Miu74, MB09, MM01b, MR02, Nau92, NRGO01, NW07, OK09, Oht09, OWY11, ÖSV01, PP16, PR94, PS99a, PS99b, PW97, PY05, PS18, Pen17, PEE04, POvdMG18, PG95, QJW07, QK14, Rab89, Ran94a, Ran94b, Ran94c, Ran95a, Ran95b, Ran96b, Ran96a, RCH17, RH07, Red80, RW69a, Rei69, Rei75, Ros78, Rus73, SV92, SV94, SV95, SJV96, SER^{+}03, SBA05, SJH05, Saf74, SR12, SHC06]. Equations [SD15, SFO16, SRD04, Sob04, Sob05, Sob06, SS03b, Sta15, SS14, STZZ14, Swa07, TB02, TW17, TR97, VA03, VP10, VJ11, Vag12, VGW15, VS88a, VS96, Vas11, VS94, VM11, Wan69b, Wan70a, Wan70b, Wan72, WTW13, Wit96, WB98, WL92, Woo95, Wri06, Yan01, Yan02, YL07, Yan12, YY18, You13, ZC10b, ZZ13, ZL17, ALZ21, ACS18, Aou20, BS22, BS23, BP20, BJV22, BHIS21, BM20, BMD22, BDM22, BM22, BOPS21, CDF^{+}22, CCY23, CZ20, CCK^{+}21, CCMX22, CM21, CMP22, CEHS19, DWZ21, DZ22b, Dun20, Dya20, DFS22, Esc23, EE19, FIM23, FvdW21, GLS23, KDM22, LWAO21, LTY22, zLWjZ22, Lu23, MWZ22, MFIGH22, Olv21, OF23, PS20b, SO20, TV22, Wu21, XXZ21, bXlZS23, XJN21, Yan21, YF21, YLW21, ZL21, ZS22]. Equatorial [BM04, BN98]. Equilateral [KF10]. Equilibria [BLP01, KWW05a, LAHM17, WS99]. Equilibrium [ABGML71, Ark92, Hel74, Weg03]. Equivalence [Bag08, BM12a]. Equivalences [GUGM18]. Equivalent [ADM82, Ber72, Ros76a]. Erdös [GKK76, Pec80a]. Erdös-Ko [GKK76]. Erlang [YK08]. Ermakov [AKZ16, AZ22, RA10, RMA12, Rog14b, Rog14a]. Erodible [EM09]. Errata [Ano05b, KT90, Wan70a]. Erratum [DDNP91, Gez80, Sch80, UG84]. Error [NR77, Nem18, SAM17, KHJK21, NR22]. Essential [Cos97, YS21, BZZ21, SZ23]. estimate [FDL21]. Estimates [Abr77, FR08, Hwa97, SAM17, Tem93, KHJK21]. estimation [CBKM20]. Estimators [BCKW85]. Euclidean [GM13, MRE19]. Euler [AFY15, BS22, BS23, CKL16, CLW23, CFY17, Cro00, FZ90, LLS04, LY03, LT99, Mac94, SVS08, SR12, VA03, Wu21]. Euler-like [CLW23]. Euler-Mahonian [FZ90]. Eulerian [And71, GR70, Rei81]. Evaluation [AM05, CFW87, Isr72a, Sri92]. Evaluations [WF12]. Events [LK16]. eventual [Jin23]. Evidence [CS80b, MR84b]. Evolution [AB70, Abl01, AB82, ACMR01, BH06, BM75, Ben84, CDF^{+}22, Che11, CT86, CD90, DB03, EEP03, Fai97, FA89, FR01, FP05, GWH11, Gri81, Gri85, Hab75, Hal06, Hic83a, Hic83b, HM08, HTWW98, JQS14, JT87, Kau06, KM90, LN94, LHC04, LB81, Ma79a, MD99, MM99a, MB09, MB84, MSW07, NC15a, PF89, PS18, PEE04, Red80, Rod03, Ros78, SY06, Tam88, Var97, WW78, War83, WK89, XF16, Yan95b, YL08, YY03, YGJ18, CD21b, EE19, FvdW21, TNT21b, TNT21a, TNT21a]. Evolutionary [TW17, DDMMSYD22]. Exact [AKZ16, BB87, BAM98, CNW12, FMM99, Kau78, LHC04, LLL12, LCB99, Luo21, MQ22, Nis07, Rei79, Ros78, SP88a, SG90, SV92, SV94, SV95, SJV96, SJH05, SKB08, SV87, VK06, VS85, VS88a, Wan69b, Wan70a]. Exactly [Che84, Len09]. Example [ADM82, BSSS00, LM91, Mos79, ZS86a]. Examples [EG95, How83, OK09]. Except [JK94]. Exceptional [BK18, Dur22, GFGUMM22, LT23, ADH^{+}21]. Exchange [Bla70, LS75, LS77, ZS86a, CC19]. Excitation [AT07b, Gus86, PT18, Sul20]. Excitations [ABGML72]. Excitatory [Gro69]. exerted [MO21]. Exhaustible [Wan80]. Exhibiting [AO70, BK95, HM93a, SS03b]. Existence [Atk78, BSGC99, BSCS00, BS23, BP14, CS80b, CCK^{+}21, DB96, DFK16, FCGB95, Gaa82, HS82b, KWW05a, Lie77, LY11, Sch75, Tam93, Tho69a, Urb89, Wal01, WPO22, WX22, WW02, Aik19, ABM^{+}22, BS22, dSF20, LR69]. Exotic [LW22]. Expanded [Yan02]. Expansion [BOGK04, EHS18, FR17, KMM82, KT89, KT90, Lou15, ZZ07, BS23]. Expansions [BS99, BS00, Dun01, Dun04, HCB72, HS92, KD17, Kic99, LF02, LT02, LS04, Mac97, Mor86, NR77, NA88, Nem18, PW12, RC05, WL92, Dun22, KHF^{+}20]. Experiment [YMBC04]. Experimental [Bea70, GJ81]. Explanations [CY94]. Explicit [ADM82, CFW87, KL88, Peg88, Ran94c, Ran95a, Zho18, ZZZ10, PP19, SSW22, Ran96b, Ran96a]. Explicitly [MTW16]. Exploratory [STC09]. explore [FCP^{+}21]. Explosion [AMR91a, AMR91b]. Explosive [Mah95, Mah96a]. Exponential [BS16, DL23, Jon80, LO99, Lea16, LM74, Mén96, NAY13, NY14, NAY17, Sta78, WW05, MMO23, VW23]. Exponentially [Kne00, LW95, FH23]. Expression [LNWZ09, Tri99]. Extended [DFG19, Ifi10, Kup06, RD04, Yan01, ZZN12, LSG22]. Extending [YS98]. Extension [FR17, Zen91]. Extensions [ACT87, GKK76]. External [GPS96, RJM78]. Extinction [AT07a, MEHS22]. Extremal [OAVS97]. Extreme [ZZ07, BHN21]. Eye [Ber70].

Facing [Tan85]. Falkner [SKB08]. Families [Bry91, MW98, NY16, YA03, Yan14b, FIM23, Yan21, ZSA22]. Family [PS96, ACS18]. Fast [Fow85a, Isr70, PAS88]. Featuring [Cos97]. Feed [KWW05a]. Feed-Rate [KWW05a]. Feedback [Abl01, KSW14, NPT08, SK23, Wan19]. Fermi [FH23, GP14]. Fermion [Bar84]. fermionic [DIW19]. Ferroelectric [Bax71]. Feynman [Lap87]. fiber [MO21, SZ23]. Fibers [FMM99, Len09, SKP^{+}05, Yan96, Yan97]. Field [CBB87, DKK^{+}14, Dre71, FCL12, Sta76, Stu92, Yan14a, YW92, FDL21, WBZ21, Zhe20, SH76]. Fields [Har98, MM93, Nic21, PCH23]. Fifth [Liu13, AK20, HSS19, LCG21]. Fifth-Order [Liu13, HSS19, LCG21]. Figures [KS02]. Filament [KO13]. filaments [Aik19]. Film [Gal10b]. Films [BB87, Ben06, Gri10, MM90, NLB10]. Filtering [Kli03, KR03]. Finding [Kle72]. Fingers [Pom85]. Finite [BMM87, BFM17, CS80a, ERK^{+}81, FS69, FT19, GR70, GP87, GY93, KM12, LHC04, MC18, Ors71b, Phi98, Ran98, Rei89, Rei73a, Rei73b, Rei75, Tur07, Wan73b, War83, Yu10, ZK13, Zho09, ZS86b, dB75, GMW21, GBSS21, JL19, KDM22, MC21b, PCH23, Sin21, SO20]. Finite-Amplitude [BMM87, GY93, ZS86b, Phi98, MC21b]. Finite-Difference [Ors71b, JL19, KDM22]. Finite-Dimensional [Zho09]. Finite-size [FT19]. Finite-Strain [Rei73a, Rei73b, Rei75]. finite-time [GBSS21]. First [BC89, JK01, Jos15, Kne00, KV06, LW88, LLLZ17, LST06, MC06, NA84a, AE22, KHJK21, KHS22]. First-Kind [MC06]. First-Order [NA84a, AE22, KHJK21, KHS22]. First-Passage [Kne00]. Fisher [GRMOH01, MEHS22, Rod03]. Fisher-type [GRMOH01]. fitness [FSKMP21]. Fixed [CL91, DMPS96, MPB00, Saf73]. Fixed-Volume [MPB00]. Flap [Tan85]. Flat [Lia06, SY06, Tan85]. Flexible [KKS15]. Flexural [MW13, MPG17, TMPVB19]. Flexural-Gravity [MPG17, TMPVB19]. Flights [Nec12]. Floating [Vár09, Vár13, Weg03]. Floquet [BAM98]. Flow [Ack73, AO02, AB03, BM82, BM99, BC89, Ber92, Cam04, CK13, CK14, CLZ91, DPHJ03, DS99, DG05, FLM95, Gaj96, GJ81, Gri90, GY93, GS02, GZC09, Gus86, Gus89, GS90, HS82a, Hen88, Hic85, HS69, How73, How90a, KR78, KM70, Kro77, LPH09, Lan87a, Lin97, Lip06, OCB87, OH95, ÖSV01, PS00, PCZP00, PJ86, Rei74a, Rei79, RS07, RD04, SJH05, Saf74, SS77, Sha91, SY06, Swa18, TWK00, TU05, Tur05a, Tur05b, Tur07, Ulr75, UG83, Wal01, YKH00, YY03, BCCM23, Cai22a, Cai22b, KA19, LBP21, MKP^{+}20, PSG23, UG84]. Flows [AM09, AM03, AM05, BGT07, BG07, BC09, Bar16, BH89, BB69c, BM75, Ben83, Ben84, BZ89, Ber70, BM12b, BDGL02, CHR91, Cho90, CMM^{+}09, CD90, Cro00, DH96a, DS71b, DARB11, EM09, FH98, GT04, Gre69a, Gri81, Hab72, Hal01, Hal06, HS92, HMM87, Her91, Hic83a, Hic83b, Isr70, JMT08, KC05, Lia06, LB81, Lu93, MM93, MBM94, MC02, MDG10, MB95, Mil95, MVBK98, MPD05, NS91, O’N69, Ors71a, PP90, Pas90, Phi98, Ran96c, Ric92, SP88a, SV94, SV95, SVS08, SY15, TG06, VB83, VS96, Vla08, Vla17, VSS98, YK99, ZB92, BV19, BLM21, LYZ20, dMVM19]. Fluctuations [KB98]. Fluid [AKZ16, Ash10, BB69b, BT80, Bry77, BL78, CK13, CV09, CS12, DMPS96, DK05, Fri74, GT04, GOS11, Gre70, GM81, Gri85, Gri90, Gri10, Gub74, KT89, KT90, KR78, Mah96a, MVBK98, MPS98, O’N69, PS99b, PG96, Sch75, SA04, SY15, Swa12, TA07, TEW91, Wal82, Wax84, WNH12, Dya22, Esc23, FMX22, MO21, PSG23, PTDS19]. Fluidized [DS71a]. Fluids [AS80, BI81, BM82, BG82, Gre69b, Isr72b, JTLW13, PP90, Pas90, SS17, TZ14, VS96, KP19a, KP19b]. Flutter [KO13]. Flux [HM94a, MM03, Rat19, ASSV20, WX22]. flux-limited [WX22]. fluxes [KT22]. Foams [DS71a]. Focusing [PV16, Woo84, BO20, BLM21, Slu19, TYZZ21, XF23b]. Fokas [Cau18, Hwa14, zLWjZ22]. Forbidden [WC95]. Force [GPS96, KH81, Rei73a, Rei75, MO21]. Forced [Ben79a, Cam04, CA09, CS15, FH78, FA89, Lan87b, MM01a, MM03, SAM17, TZ14, WK89, WPO22]. forces [WBZ21]. Forchheimer [PS99a]. Forcing [FL73, FIN94, MTVE02, Wan72]. Forest [BBC^{+}10, WA83, Wan85]. Form [BHH^{+}11, CS80b, Ioa89, JTLW13, MR93, PS96, TCK82, WXZ18, ZL17, FT21, TNT21b]. Formal [BDGL02, Fai97, Far82, JG79, JlPG79, JG80, LR69]. Formalism [GHMT08, KR03]. Formation [Kee78, LNVW05, MR84b, MR80, Saf75, WC95]. formations [LWC23]. Forms [Kun86, Mor86, War78]. Formula [CY94, Lap87, Zen91]. Formulae [CFW87, AK20]. Formulas [Abd97, AMT18, FS96, KMM80b, KMM82, Mil01]. Formulation [FMO17, Mil98, YK99, BE22, CF23]. Forward [Tan85]. Forward-Facing [Tan85]. Foundation [BRSV92, Rei70a]. Foundations [And71, CR70, Dou72, DRS74, GR70, LW90]. Four [LP08, ZC10a]. Four-Potential [ZC10a]. Fourier [AKNS74, BP71, FS69, Len16, Mac94, Tri76]. Fourth [Abd97, BCH95, BN98, FC08, Gal08, Gal10a, LR69, WZ07, Lu23]. Fourth-Order [Gal08, Gal10a, Lu23]. Fox [FÖK22]. Fractal [ZHY09]. Fractional [Liu13, PS93, SJB91, ABM^{+}22, BBN22, CCK^{+}21, GSYG21, ZL21]. fractional-order [GSYG21]. Fragmentation [Kne03]. frame [MRE19]. Frames [GM12, GM13, BE22]. Freak [DKZ16]. Fredholm [DZ22a, EHV18, WF12]. Free [Bar84, BGT07, BG07, BC09, Bar16, BC86, BZ92, Cro00, JO89c, JZ86, KL88, MVBK98, O’N69, OCB87, Rei89, Ric92, SV94, SV95, Saf73, SS77, Tam93, CHS21]. Free-Electron [KL88]. Free-Fermion [Bar84]. Frequencies [Sch75, Yu10, FA22]. Frequency [AFN73, Alm90, AMR91a, AMR91b, CHR91, DR77, DIT16, FH78, KA88, MR92, VB76, Vaf20]. Freud [CJK16, FIM23, LDLW20]. Freud-like [LDLW20]. Frobenius [BMD22, Mor80]. Front [BGL97, MMPS13]. Frontal [GV03]. Fronts [BS16, GRMOH01, KS96, MR84b, MKP^{+}20]. Full [CKLS96, Kau80a, HP19]. full-dispersion [HP19]. Fully [Miu74, SK23]. Function [BOGK04, CBB87, CD18a, FG12, GST14, Gre75, Hei71, JTLW13, Kun78, LP11, Nem15, Ulr75, CF23, FLT20, LL22, Min86, SCFFLPA23, Tro21]. Functional [LM74, PS93, CZ21]. Functions [Abr72, BDC97, BRSV92, Bry91, BH91, Cao13, CFS92, CLY95, Dou72, Dun01, Dun04, Dun11, DKZ16, GR70, Har75, HM94a, Hen72, Ioa89, JRS99, KMM80a, KMM80b, KMM82, Lea16, LY11, Loe90, LT99, LF02, LT02, Mil75, MM03, Nem18, Ors69, Rod03, RV18, Sen89, SP88b, Sun93, Tri91, Tri97, Ven92, VS88b, Wan69a, CD18b, DIW19, Dun22, FÖK22, LWAO21, MFIGH22, Min86, Seg21, YLW21]. Fundamental [ACMR01, JRS99, KX18, Sen89]. Further [JG80]. Future [Che69].

G. [BMOT09a]. Galactic [Hun69]. Galaxies [FL73, LL79]. Galaxy [Hun70, Nis79]. Galerkin [FS69, Ors71a]. Galilean [DDM13]. Game [GY95a, YBSM08a, YBSM08b]. Games [FY01]. Gamma [DdY11, Nem15]. Gap [MC18]. Gas [AO02, EY91, Fos70a, Fos70b, Gol04, GS15, Peg88, RSCM05, SA05, SR99, BS22, BS23, KR23]. Gas-Dynamic [EY91]. Gasdynamics [SR12, SAR13]. Gases [Hei71]. Gauge [DFK16, Har98]. Gauss’ [LL22, CG16, GST18, LP11]. Gaussian [LC21, WXZ18, XZ11]. Gaussians [HZ10]. Gegenbauer [GFGUMM22]. Geissinger [Che94]. General [AT07a, BGT07, Ben77, CM09, CLZ91, Gre72c, Gre72a, GK74, Gri90, JO07, Kau06, LFW20, LT07, MS16, OWY11, Pec79, RMMH22, Rei74b, YS98, YL07, YZ20, BS23, BHIS21, WZ22]. generalist [XHW22]. Generalization [Con91, MM03, Mos78, Tri97, Ven92, WF12, RRY22, dSF20]. Generalizations [Tri92]. Generalized [ABT86, ABT87, ACH99, Bax71, BT88, Boy16, CFS92, CJK16, FP12, GZ10, GLL13, Ifi10, Ioa89, JZ10, JW20, JL19, Lin14, LLL12, LF02, MC06, Qu97, RS11, RN15, SJV96, SER^{+}03, Spe74, Sri92, TZ14, Tri97, VA03, VK06, VP10, VJ11, Vas11, WTL^{+}10, Yan12, ZZ13, ZZ08, BHIS21, FMX22, Hol23, KR23, Lu23, LT23, LWC23, NMBD23, YRZ20, KL00, MFIGH22]. generate [BHN21]. Generated [DRS02, MG91, FMN19, dlIJ23]. Generating [Cao13, CFS92, CD18a, GR70, Hen72, KM90, SP88b]. Generation [Aky82, BH89, BMOT09a, BMOT09b, Boy16, EGK05, GY93, GP04, Kau78, Mas72, KA19]. Genetic [WSK10]. Geometric [Bri94, DB94, FMO17, Gre75, Kne00, MW73, DFS22]. Geometrical [DK05, Kne03, KM04, SS03b]. Geometrically [CM13, Kat88, KK90]. Geometries [Bry75, Bry79b, CR70, DA05, RS14]. Geometry [CSS77, Gre76, KS02, RS98, RS12, Vas11]. Geophysical [SG90]. Gerdjikov [MZ23]. Gertler [DH96a, OB94]. GI [YK94a]. GI/M/1 [YK94a]. GI/M/1-K [YK94a]. Gierer [Nec12]. Gilbert [Kal73]. Gillespie [ZGQ12]. Ginibre [BC23, WF17]. Ginzburg [Kee85, Lan87a, Lóp22, SZWM16, War94]. Girth [BBC^{+}10]. Given [CL91, ELY03, Gez79, Gez80, Sch80, PS20b]. Global [BCD21, DD11, DNP17, LY11, LZW23, Mar71, NPT08, Nis79, WWZ10, WQW13, You13, Zho18, MMO23, dSF20]. Global-phase [BCD21]. globally [CCY23]. Goldfish [Cal16]. Gordon [ABT86, ABT87, AFLM18, BT88, BD17, CF22, CMKP16, DKK^{+}14, EFM84, FM83a, Hwa14, JW99, Kau75, KT19, Li06, NS98, PP16, PP08, Sta15, Swa07, Tri88b, bXlZS23]. Gordon/Sinh [AFLM18]. Görtler [BH89, DB96, Wad91]. Governed [CS15, SRJ99, TZHL09]. Governing [Ran94a, Var97]. Gradient [Gal11, KO13, LNVW05]. Gradients [KSW14, LNW04, Wan14]. Grand [Hei71]. Graph [BBC^{+}10, RH07, Wan76a]. graphene [Nic21]. Graphical [Boy16, KS02]. Graphs [Coh91, CKLS96, JO89c, JO89b, JO89a, Wan76b, YY08]. Gravel [Hal06]. Gravitating [Fos70a, Fos70b, Yue81]. Gravitational [Mat89]. Gravity [AM08, AM09, AN13, AM03, AM05, Bau98, Bau99, BH06, Bri90, CS80b, CS80a, CS85, CS12, DMPS96, DMPS97, DB90, Gri77b, Ma82, MW13, MG91, MPG17, MM99b, MM01a, MPS98, Moo00, MP01, MPD05, PMAD07, PR94, RP83, WVBM14, CC19, CW21, GMW21, JW20, TMPVB19]. Gravity-Capillary [AM08, AN13, CS80a, WVBM14, CC19, JW20]. Gravity-Driven [MPD05]. Gravity-Wave [Gri77b]. Gray [LSG22, HM08, KWW05a, KWW06, WW03]. Great [AS80]. Greatest [McL97]. Greedy [Bry91]. Green [DIT16, SCFFLPA23, YLW21]. Grid [Boy16, Pec80c]. grids [Sin21]. Griffiths [Gal71]. Grimshaw [KP19a, KP19b]. Gross [WTW13, Yan14a]. Ground [CM13, GP14, BŠT23, BK23, CI23, PS20a]. ground-based [BK23]. Grounded [Swa18]. Group [JFM21, LLL12, Liu13, MO01, OK10, OK15, PRNA01, PG95, Qu97]. Group-Invariant [PG95]. Group-Multiple [OK10]. Groups [Bry79a, FS96, Rei89]. Growth [BDC97, Ber92, BS93, Gre72d, Hic85, HG01, KS91, OH95, TW03, Wan80, Xu90, Lóp22]. Gubeskys [Ano05b]. Gundelfinger [Kun86]. Gurevich [OF23].

H [BMOT09a, UG84]. Hagen [Wal01]. Hahn [Dur22]. Hairpins [Wat79]. Half [ACMR01, BB12, FP05, GLZ15, Ros10, Vla08, CKS23, CD21b, FvdW21, JGM21, XB22, Zha19]. Half-Line [FP05, GLZ15, CD21b, FvdW21, JGM21, Zha19]. Half-Space [Ros10]. Halos [Wax80]. Hamiltonian [AHP02, AKZ16, BMD22, BDM22, BM22, BK95, CZ20, DH02, GHMT08, Gri10, KS14, Kev82, LN94, Pec80b, RS14, Sat85, SDS12, Swa18, Wes78, YS21, YT14, Zho09, ZS86a]. Hamiltonians [ADM82, Bri90]. Hammerstein [EHV18]. Hamming [KKR71]. Hamming-Adjacent [KKR71]. Hankel [AT90, MC21a, Tél94, ZXZ15]. hard [FT19]. Harmless [NPT08]. Harmonic [BDC97, BC89, DB90, KMM82, Kau78, CI23]. Harmonics [DF09, DDN21]. Harry [BM13]. Hartree [TYZZ21, YRZ20, ZS22]. Harvesting [WA83]. Harvey [CD21a]. Having [Dun01, Dun04, Dun11, HP78, KPS80, Dun20]. Hawk [WWZ10]. hear [BK23]. Heat [How90a, KF10, RS07, SD16, Str11, TW97, GM21]. Heated [O’N69, SS06]. Heavy [DMPS96]. Heeger [MWL23]. Height [BS01, McL97]. Heisenberg [SR12]. Hele [Ric92, YK99]. Helical [STZZ14]. Helicoidal [Wan70c]. Helmholtz [Ma84, YW92]. Herd [ST17]. Hermite [BDS20, CS90, GST18, GUGM18, LY89, RC05]. Hermitean [CI98]. Hermitian [BZZ21, Gal08, IW23, War78]. Hessian [CCMX22]. Heteroclinically [Li06]. heterogeneities [WW21]. Heun [XXZ21]. Heuristic [AH75, CK75, Ulr75]. hexagon [Cha21]. Hexominoes [She82a]. Hierarchical [DGW02]. Hierarchies [BCH95, BM13, EEP03, KA02, MCB97, Sat85, Bła23, MHW20, YY23]. Hierarchy [CDLjZ18, Day80, FC08, GDZ07, GHMT08, KL00, KS02, Nis07, ZC10a, ZC10b, CZjZ21]. Higgs [FCL12]. High [Alm90, ADH^{+}21, Ben84, BGL15, BMB02, FM83b, GLL13, HS82a, Hal12, Mar02, FLT20, LW22, AMR91b]. High-Frequency [Alm90, AMR91b]. High-Order [BGL15, GLL13, Mar02, ADH^{+}21, FLT20, LW22]. Higher [Cos93b, Cos00b, Cos06, Gal06, Hab75, HM08, Lun03, Olv21, RRY22, RANC09, SY03, SDS12, Vag12, CGUGM20, RMMH22, ZTYH20]. Higher-Dimensional [Lun03]. Higher-Order [Cos06, Gal06, Hab75, HM08, RANC09, SY03, SDS12, Cos00b, Olv21, ZTYH20]. Highly [LN13, RS07, FLS23]. Hilbert [CCM12, CCM19, CCF19, CF22, LTY22, MV13, WW05, XZ13]. Hilliard [Hol23, SW00]. Hirota [ALZ21, ACS18, CC17, GRTW85, HM08, JN21, WTL^{+}10]. history [DDAAB20]. HIV [GYZ21, WLS17]. hoc [MS08]. Hodograph [KQ10, QK14, Xu21]. Hodograph-Type [KQ10]. Hohenberg [CNW12]. Hölder [Jin23]. Hole [GZC09]. Holes [HHW94, BHN21]. Holling [XHW22]. Holm [GXW13, GMNP22, GZ10, GZ13, Kau06, SYF22, SS14, dSF20]. Homoclinic [BH98b, DH02, Fow90b, Fow90a, Li02, Tov00, Wri06, YT14]. Homogeneous [BB12, CR86, Hic85, SH76]. Homogenization [KB98]. Homopolar [Lu89, SS95]. Homotopy [Far82]. Honeycomb [CZ15, HHZ20]. honor [CD21a, KP19a, KP19b]. Hooke [LZW23]. Hopf [GM21, Kat81, MML17, RC81, Vag12]. Horizontal [BDC97, Ben06, Hic83a, Hic83b, SS06]. Horseshoes [KH81]. Horton [DGW02]. Host [WLS17, DDMMSYD22, HMBYZ22, WLR22]. Hsieh [FT21]. human [DDMMSYD22]. humoral [HMBYZ22]. hungry [PCH23]. Hybrid [TWK00, YLP95]. Hydraulic [Moo00]. Hydrodynamic [BG81, BC89, Hun70, Wal95, YC15, YM07]. Hydrodynamics [BC85, AP23, Rat19]. Hyper [Car90]. Hyper-rook [Car90]. Hyperbolic [CR86, HM91, HM93a, HM93b, HM94b, HMR86, LT99, MR84a, MM01a, MM03, Rus73, YK02b, ASSV20, CR21, XF23a, Xu21, Cos93a]. Hyperbolicity [GLS23]. hyperconcentrated [FCP^{+}21]. Hyperelastic [Lus80]. hyperelliptic [Wri20]. Hypergeometric [KMM80a, KMM80b, KMM82, LP11, Sri92, DZ22a, FIM23, LL22, MFIGH22]. Hypersonic [YW92]. Hyperspheric [Cla73]. Hypothesis [Vla17]. Hysteresis [FM83b, MKP^{+}20].

i.e [Ano05b]. ice [DDM21]. ice-covered [DDM21]. Ideal [SA05]. Identical [ADM82]. identification [WCYW19]. Identities [DF09, Fre72, Kle75, Mil01, NKUW02, VS91, VS94]. II [AY17b, ABGML72, AMR91b, AN98, BSCS00, BG07, BF16b, BDM22, Cai22b, CCM19, CS80a, CD18a, Cos93b, Cos06, CR70, Fos70b, Hic83b, How86b, HMR86, JlPG79, KK90, KP19b, KS18, Kok20, KL00, LS93, LS77, LR69, LC21, Mah96a, MR84b, MP22, Mil95, MK09, Nem15, Rog14a, RC81, SV95, SR97b, Sta76, SLN16, Sun93, Tam88, TNT21a, Urb90b, VA02, WL92, WXZ18, YKdM^{+}07, YBSM08b, ZC10b]. III [BM22, BMS18, JG80, Mah96b, MW98, ZXZ15]. Ill [LW95]. Ill-Conditioned [LW95]. Illuminated [LK10]. Illustrating [Her82, HMM87, OK09]. image [CCD22]. Imaginary [YL08, AK20]. Imaginary-time [YL08]. immersions [FdST22]. Immiscible [BI81, BM82, LBP21]. immunity [HMBYZ22]. Impact [MMO23]. imperfect [GLH22]. Imperfection [Rei70a]. Implicit [Isr70]. implies [UD20]. Importance [BSSS00]. Imposed [YK99]. Impounding [GJ81]. Improve [EHV14]. Improved [DIT16, GBSS21]. Impurity [Bar83]. In-Core [Wan73b]. in-host [HMBYZ22]. Inaccessible [CW96]. Incidence [GKK18, LWZ23]. Inclined [SY06]. Inclined-Channel [SY06]. including [Pet22]. Inclusion [SRD04]. Incomplete [Gal10a, Nem15]. Incompressible [LCB99, MM93, Ors71a, Ran94a, Ran96c, Ran98, SV94, SY15, Wri85, AN98]. Incompressible-Integrable [AN98]. Incorporating [Mac97]. Increasing [MS08, Wal82]. Increment [Ber72]. Indeed [FZ90]. Independence [LM74]. Independent [CR86, Cos93a, GKM74]. Indeterminacy [War92]. Index [Ano72, Ano73, Ano74, Ano75, Ano76a, Ano76b, Ano77, Ano78, Ano80a, Ano80b, Ano81a, Ano81b, Ano82a, Ano82b, Ano83a, Ano83b, Ano84a, Ano84b, Ano85a, Ano85b, Ano86a, Ano86b, Ano87a, Ano87b, Ano88a, Ano88b, Ano89a, Ano89b, Ano90a, Ano90b, Ano91a, Ano91b, Ano92a, Ano92b, Ano94a, Ano94b, Ano95, Ano96, Ano98, Ano99, Ano00, Ano03, Ano04, Ano05a, Ano06, KS14]. Indices [YY04]. Individual [YBSM08a]. Induced [BC09, FYC05, Kon96, KL00, ST84, Tan85, Wei85, YMBC04, BC23, Lóp22, PSG23]. Inequalities [Car90, Day80, Gal71]. Inequality [Gre75]. Inertial [BB69b, BB69a, Bea70, LPH09, QC96, Tro17]. infection [LP23, LCTZ22, WLR22]. Infiltration [TBKiM18, LSG22]. Infinite [AO02, AO03, Cha71, DWZ21, Dun13, Ifi10, LS04, OC70, She82a, She82b, ZZ07, BCCM23, LW90, Lu23, Wan75]. Infinite-Dimensional [Ifi10]. Infinitesimal [KA02]. Inflow [MPD05, PMAD07]. Influence [Ake12, NA88, Tur07]. Information [Ano16a, Ano16i, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f]. Information-Copyright [Ano16i]. Inhibiting [BC09]. inhibitive [BGMR21]. Inhibitors [Kee78]. Inhomogeneous [HCB72, MS76, WTL^{+}10, WS99, Dun20, Kok20]. Initial [AW97, Ber92, GLZ15, Gri19, Hal01, Hen88, HM15, HS69, JS01, Kau80a, Kau06, LLLZ17, MQ15, MW16, PMAD07, RC05, SD16, XF16, YK02a, ZS22, BO20, FR01]. Initial-Boundary [GLZ15, HM15, MQ15, XF16]. Initial-Boundary-Value [PMAD07]. Initial-to-Interface [SD16]. Initial-Value [Kau80a]. Initial/Boundary [FR01]. Initially [BGL97]. Injection [KC05, Ric92]. Inner [NKUW02]. insecticide [– Dor22]. insecticide-treated [– Dor22]. Instabilities [BR69, Ben79a, BC85, BC86, BC89, CC17, CMKP16, KWW06, MR92, Mor89, PY05, QK94, TW03, YG01]. Instability [BM99, BC09, BG81, BZ89, BZ92, BH14b, BMB02, Cho90, FP05, FH98, GS02, Hag81, HHA86, HMM87, Hic85, HJ15, Joh10, KS14, Mat89, Mes96, OC70, PS96, SDS12, TU05, WSK10, YC15, YW92, ZB92, ZS86b, BHIS21, BBN22, HP19, KLSD23, LYZ20]. Insulators [RS07]. Integer [Bel77, Cla73]. Integrability [Dem23, Fok87, Fok09, GHJ12, HL11, KC92, MR02, Sch01, TZ14, VM11, WTW13, dSF20, BMD22]. Integrable [AM17, AKZ16, AN98, Bła11, BHH^{+}11, CD04, Che11, CWT^{+}15, DLW87, EGP01, EGK05, FIN94, GZ13, HY00, Kon96, KL00, KA02, KM04, Lou15, Lun03, MNW07, NS98, NW07, PEE04, RH07, RS10b, SR12, SAR13, TBKiM18, XF16, XJN21, Yan02, YY18, Zak09, Zho09, ZZZ10, AZ22, ACS18, BF16a, BF16b, BW19, EE19, MP19, MP22, RS21, SYF22, KO13, ZC10b]. Integral [ABCD17, EHV18, LS88, LS93, MC06, SY15, FÖK22, FLS23, VW23, YF21]. Integrals [CFW87, Cao13, KD17, LW88, Lun03, RV18, FLS23, Kok20]. integro [BOPS21]. Integrodifferential [Lan80, MRS88, Wei84]. Interacting [CR86, HMR86, MR84a, GMW21]. Interaction [AN96, AN98, BPS73, BH89, Ben74, BM04, Bri90, FPSA89, GB96, Gri86, GPS96, HS84, Hun92, Kau76, Kau80a, MMPS13, Mar02, MTVE02, SS08, TNM96, TR02, WM94, YK94b, Cai22b, CHS21, Esc23, FMN19, KHJK21, RMMH22]. Interactions [AFN73, Ben76, Ben77, BH98a, BMOT09a, BMOT09b, Cam04, CLS86, DS99, GG84, Gea85, GZ94, Hal01, HM93b, HM94b, Ken74, KL84, LP08, LM16, Mah95, Mah96a, Mah96b, MB95, Mil95, PS96, Ros76b, SB00, STZZ14, TA07, WNH12, ZHY09, CGP23, CC19, KT22, LTY22, Hal12]. Interactions/Self [Hal12]. Interactions/Self-Sustained [Hal12]. Interactive [MS08]. Interest [Lan87a]. Interface [Gro02, SD15, SD16, TW03, CDF^{+}22]. Interfaces [Gal06, SW00]. Interfacial [BM12b, BM14, DB94, NLB10, PTDS19, WVBM14, Nic21]. Interior [AT07b, DF15, KF10]. Intermittency [FM83b]. Intermittent [New02]. Internal [AS80, AKK^{+}09, BC09, BK78b, GG84, Gri77b, Gri81, Gri85, GZ94, Gri16, HCB72, JC02, JC08, KO04, Lam98, Mil76, MG91, PY00, Rub77, RANC09, SDS12, Swa12, TKC81, TCK82, TEW91, WV92, Wit96, Yan97, YKH00, YY03, YGJ18, HW21, KA19, MQ22, Pri23]. Interplay [Cau18, Ste19]. Interpolation [Olv06, VS88b]. Interpolatory [VS93]. Intersectable [Wan76b]. Intersection [Bry79b, ERK^{+}81, Hel74, Wes78]. Interval [AO02, AO03, XF16]. Intracranial [YLP95]. intraspecific [GJLW22]. Intrinsic [RS98]. Intrusive [DMPS97]. invading [FSKMP21]. Invariance [DDM13, Hu98, Kev82, KL84, BL22]. Invariances [Wad78]. Invariant [AHP02, BV19, BJV22, Bri90, DRS74, GM13, JQ13, JS99, LKM14, Oxl82, PG95, VK06, War92, ACS18, Gab21, KDM22, EE19]. Invariants [Gre72a, QJW07, Rei89]. Invasion [CLW22, ELV21, Jin23]. Inverse [AKNS74, AL76, Abl78, AYF83, BT88, BRT89, BN94, BLM21, Cau18, CJM92, DPvdMV13, FA83a, FA83b, FA89, GPFT22, GS16, Kau80b, KM98, KM90, KD92, MHW20, MZ23, Mos78, NA84b, NA84a, OP19, PZ03, PBT11, PV16, POvdMG18, PT18, Vil90, AE22, DvdM23, MN22, NR22, Tro21, XB22]. Inverse-Scattering [NA84b]. Inversion [Coh91, GST14, Mil01, MR11, Zen91]. Inversive [KS02]. Invertible [BM13]. Inviscid [BM99, Ber70, Cho90, CLZ91, GT04, Gre69b, Hen88, Hic83a, Hic83b, SR99, TG01, TR02]. Involving [PS93, Ric92, RTT02, Yeh95, PRAV21]. Irreducible [Hsi73, Kle70]. Isentropic [Wu21, SXZ22]. Ishimori [KM90]. Ising [Abr71, Abr72, BM72, BR73, CBB87, WF12]. Isobe [CI20]. isometric [FdST22]. Isomonodromic [BDM22]. Isomonodromy [XXZ21]. isospectral [GFGUMM22]. Isothermic [DRS02, Sch01]. Isotropic [Ors69]. Isotropically [Gub74]. Issue [Ano16a, Ano16i, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f]. Issues [AY17a, AY17b]. Iterated [Tri88b]. Iterates [Ven92]. Iteration [Isr72a, MC06, YL07]. Ito [Vil04]. Itoh [Wan12]. IV [DK09, GR70, LC21, RF13]. IV. [GJW77, LM91]. Ivanov [MZ23]. IX [Cos00a, DRS74].

J [AAL02, BF16a, BF16b]. Jack [AGJ03]. Jacobi [CZ10, CCF19, Dur22, MMRL23, MC18, MC21a, SW17, SP88b, Wax80, ZXZ15, dlIJ23]. Jaulent [LW22]. Javelin [AB07b, FN05a]. Jet [BMB02, MD99, YW92]. Joint [Gez79, Gez80, KS06, Sch80, BE22]. Jones [BŠT23]. Juggling [YS98]. Jump [MM01a, SS03a, XZ11, LC21]. Justification [Abd97, JW99, JZ10, CM21, CMP22].

K.P.P. [Atk78]. Kac [Lap87]. Kadomtsev [BMOT09a, AYF83, AV91, AW97, BMOT09b, BBN22, DFS97, FA83b, FS86, Fok09, FN22, GM89, GT90, JPS00, KD92, RCMH21, Rat19, VA02]. Kakinuma [CI20]. Kármán [Gao95, ZL17]. Kashin [Haj92]. Kaup [RP11]. KdV [BM13, Bła23, FH10, JW06, KS14, Liu13, LCG21, Lou19, Lu23, NRGO01, Ryb22, TZHL09, VM18]. KdV-Like [KS14]. Keller [PS09, WX22]. Kelvin [CM07, Ma84, YW92]. Kept [CL91]. Kernel [CBKM20, Tél94, Tri88a, DZ22a, FÖK22]. kernels [KHJK21]. Key [MS08, DDMMSYD22]. Kidder [IB15]. Kind [BN98, BH91, Can78, GLN14, KV06, MC06, MMRL23]. Kinematics [Rei73a]. Kinetic [Esc20]. Kinks [BD17, DKK^{+}14, GOS11]. Kissing [BCD21]. Klausmeier [LSG22]. Klein [BD17, CMKP16, DKK^{+}14, JW99, KT19, PP16, PP08, Sta15, Tri88b]. knots [Gab21]. Ko [GKK76]. Kolmogorov [BD70]. Kontorovich [LY13].

Korteweg [AK82, AS77, CKS23, GM93, GP04, GPT^{+}05, Gri15, Gri19, Hab88, HSS19, JZ10, JP22, JPS00, JL19, JN21, Mil79, Mil93, NdJVA22, Ryb23, TZ14, Wad78, XJN21, Yan01]. Kortweg [XF23b]. Kortweg-de [XF23b]. Kowalevski [KC92]. KP [CK09, Dec00, Har11, LGZZ21]. KP-I [LGZZ21]. KPP [Rod03, TNT21b, TNT21a]. Krasovskii [McL97]. Krein [BR19, KS14]. Kuramoto [Hol23, JP22]. Kuznetsov [Joh10, PS10, SB00].

Laboratory [Gre69a, Kau75]. Ladder [DER21]. Ladik [GDZ07, GHMT08, OP19, PV16, ZC10a, ZC10b]. Lagrange [Zen91]. Lagrangian [ASSS12, DFG19, KDM22]. Lagrangians [GM13]. Laguerre [BK18, DK09, FIM23, FT19, GST18, LY89]. Lamellar [RS12]. Laminar [Wor86]. laminated [FCCTR22]. Land [Wan77]. Landau [BGL15, CK13, Kee85, Lan87a, Lóp22, SZWM16, War94]. landscape [FSKMP21]. Langevin [Esc20]. Language [Far82, KK22a]. Laplace [Tri91, Tri99]. Large [AH75, BC09, Ben76, BK78b, Ben79b, BG82, BMOT09a, BMOT09b, BM12b, BS77, CK75, CK04, Fow84, GPT^{+}05, Hic83a, Hic83b, How90a, Jos99, Lea16, LK16, LLLZ17, Mes96, Nau92, NLB10, NMBD23, Nem18, New02, Pom85, RS11, Rei73b, RANC09, SER^{+}03, SRE05, SVS08, Saf71b, SAM17, SLN16, TCK82, Ulr75, VP10, VS85, Wor86, XF23b, BCD21, OP19, BG07]. Large-Amplitude [BMOT09a, BMOT09b, BM12b, GPT^{+}05, Mes96]. Large-Argument [Nem18]. large-degree [BCD21]. Large-Displacement [Rei73b]. Large-Horizontal-Scale [Hic83a, Hic83b]. Large-parameter [Jos99]. Large-Reynolds-Number [CK04]. Large-Scale [How90a]. Large-Time [Lea16, RS11, SER^{+}03, SRE05, SLN16, NMBD23, XF23b]. Laser [Mos14]. Lasers [AHNZ09, KL88, SZ23]. Lateral [SW74, YKdM^{+}07]. Latitude [BN98, Swa18]. Lattice [Bax71, Bel77, CM08, Gre75, Hei71, HK04, JK84, KKR71, LB69, NRGO01, Wan12, WK89, ZZN12, ZZ13, ZZZ10, BŠT23, Luo21, McG23]. Lattices [ACZ12, AZ13, BBC^{+}05, CP12, CMKP16, CZ15, FYC05, GMM05, HAY12, Wag94, YMBC04, HW21, KK22a]. Laurent [AMT18, LT23]. law [LZW23]. Laws [ABCD17, FCGB95, GHMT08, GM12, GM13, HM94a, JMT08, KB98, Miu74, MM01a, MM03, Nou00, Pen17, SS03a, SS04, SDS12, Wad78, YK02a, YK02b, ASSV20, BP20, Hol23, MP19, MP22]. Lax [AZ22, Bar84, BM20, BDM22, Fan11, LY03, SHC06, UD20, XF16]. Layer [AO70, AB03, AKZ16, BMM87, BGT07, BG07, BC09, Bar16, Bry77, BS01, CG07, CM07, CG08, Cam04, CMM^{+}09, Dha94, DIT16, DH96b, FLM95, GB96, GOS11, Gre89, GS02, HHA86, HS84, Her82, IR86, Kat85, KM70, KC05, LM91, Le 04, Lia06, Lin97, MM99a, Mal01b, Mes96, MS76, MM01b, Moo78, OB94, SS06, SA04, SB87, TG01, TU05, Tur05a, Tur05b, Tur07, UG83, Wad91, Wal82, War94, WNH12, WY02, YK99, Cai22a, CW21, MM99b, MPS98, PTDS19, PV23, Pri23, TV22, UG84, dMVM19]. Layered [AT07b, BL78, ZB92, Nic21]. Layers [AW97, Ben83, BZ92, DB96, DB03, DR90, Gaa82, Gaj96, Gre84, Hab72, Hab77, KO04, MBM94, MC02, Wit96, vKD72]. Leading [TU05]. leapfrogging [Aik19]. Learn [Gro70]. Learning [Gro69]. Leaving [OC70]. Lebedev [LY13]. Lebesgue [Lap87]. Lebesgue-Stieltjes [Lap87]. Lectures [Abl78]. Legendre [Dun04, Dun11]. Lemma [DM96, Pec79]. Lenells [zLWjZ22]. Length [KS06]. Lennard [BŠT23]. Lennard-Jones-type [BŠT23]. Leukemia [RKS83]. Level [FR17, WW78, GLH22]. Levi [Tri92]. Lévy [Nec12]. Lie [FS96, JQ13, JQS14, KMM80a, PP19, QJW07, Sat85, Sta85, Van87]. Liénard [Dem23]. life [DDAAB20]. life-history [DDAAB20]. Lifshitz [BGL15]. Lift [ST84, Tan85, Wei85]. Light [YF05]. Lightwave [LK16]. Like [KS14, CLW23, GY95a, LDLW20, VS95a, XF23b]. Limit [AS89, AN96, AN98, BW02, BDGL02, CJN04, Gaa82, GB96, HJ96, Jos99, LW91, MQ15, PP86, Pom85, Stu92, Tum06, TR97, FT21, dMVM19]. limited [SSVS20, WX22]. Limiting [Boo89, How90b, CCY23]. Limits [How90a, Laq81, Vla17, bXlZS23]. Line [AB03, BB12, FP05, GLZ15, NAY13, VA02, WHS17, CKS23, CD21b, FSKMP21, FvdW21, JGM21, MMRL23, RCMH21, Xu21, XB22, Zha19]. Linear [BG81, BL87, Bry91, CR06, Cro98, Dun01, Dun04, DKZ16, GT04, Gre76, HM93a, Hen88, Hic85, HY86, HF17, HG01, HHZ20, HAY12, JT87, JO07, KX18, KLW11, MC21b, NC15a, NY16, PW97, PS18, QJW07, RW69a, Rei70a, SV92, SS81, SS85, Sta15, VK06, VP10, VS94, Wan69b, Wan70a, Wan70b, Wan72, Wan73a, WNH12, WB98, WL92, YF21, YG01, BOPS21, CKS23, CD21b, Cos93a, GMW21, KR22, Lev69, LZW23, NR22, SYF22, YLW21]. Linearizable [EEP03, OF23]. Linearization [KQ10, VK06, ZZN12]. Linearized [CD90, HM15, LLS04, Yan02, Esc23, JGM21]. Linearly [GA79, GA82, Gri86, LKM14, VS95b]. Linkage [Hu98]. linked [CBKM20]. Linking [XHW22]. Liouville [BHN21, Bła11, BN98, HN96, Lan87b]. Lipschitz [Gal06]. Liquid [Ben06, Lan83a, NLB10, Ran94a, Ran96c, Ran98]. Lithography [BO12]. Lived [GPT^{+}05]. LKR [LRS04]. Loaded [Ash10, NA88]. Loads [Wan73a]. Local [ABGML71, ABGML72, ADM82, Atk78, GHMT08, GK74, HM93a, Jos99, PPV08, VS96, YY18, ZL21, FdST22, Hol23, PG96]. Localized [KLW11, LNWZ09, MTW16, PCZP00, WK91, War96, Yan14a, BZZ21, LS23, WW21]. Location [Can78, Hei71]. Locked [AHNZ09]. Locus [HP78, KPS80]. Loewner [SR98]. Logarithmic [GJ84, DWZ21, DZ22b]. logistic [Lóp22]. Lognormal [BKG86]. Lommel [Dun22, Nem18]. Long [AS80, AY95, AM05, Ben73, Ben74, Ben77, BK78b, BM14, CK14, CM03, FM11, GB96, Gri81, Gri85, GS02, GPT^{+}05, GH08, HMM87, KA88, Kup06, Lin14, LCG21, Ma78, Mes96, Miu74, STC09, TA07, TNM96, VB83, WM94, Wri06, YKH00, LG20, PTDS19, RMMH22, RS21, Slu19]. Long-Term [STC09]. Long-Time [AM05, GH08, LCG21, RS21]. Long-Wave [AY95, CK14, GB96, GS02, GPT^{+}05, Ma78, Mes96, Wri06, RMMH22, Slu19]. Longitudinal [PS96, YW92, LZW23, YM07]. Loops [CP12]. Lord [FCCTR22]. Lorentz [GK74]. Lorenz [FM83b, Fow84]. Loss [MK09, YK08]. lossless [Vaf20]. Lot [Bön95]. Lotka [AT07a, CLW22, PCH23, WPO22]. Low [BDGL02, DR77, Haj92, KWW05a, MR92, MWL23, ST84, TWK00, Vaf20, VB76, Wei85, Esc23, LCG21, AMR91a]. Low-Frequency [MR92, VB76, Vaf20, AMR91a]. Low-temperature [MWL23]. Lower [Tur05a, Tur05b]. lozenge [Cha21]. Luke [CS15, Sta15]. Lump [LGZZ21, SB00, VPE09, GWCH23]. lumped [SZ23]. lumps [CCMH21, RCMH21]. Lymphoma [RKS83].

M [BMOT09a, XK93, YK94a, Hab72, Lan73, XK93]. M. [UG84]. Macaulay [Sta75]. Maccari [CCMH21]. Macdonald [NKUW02]. Mach [BDGL02, MR84b]. Mach-Stem [MR84b]. Maclaurin [Mac94, Wax80]. Macroscopic [McG23]. Magic [Bön95]. Magnetic [BM72, BR73, CBB87, YW92, CLW23]. Magneto [SAR13]. Magneto-Gasdynamics [SAR13]. magnetogasdynamics [SXZ22]. Magnetohydrodynamic [Ken76, CLW23]. Magnetohydrodynamics [RS12]. Magnetron [KT89, KT90]. Mahonian [FZ90]. Mahony [ABM^{+}22, CEHS21, GLS23]. Majorizing [EHV18]. malaria [DDAAB20, – Dor22]. Malomed [Ano05b]. Managed [GMM05]. Manakov [Vas11, GR23]. Manifold [EG95]. Manifolds [JS99]. Mantle [Fow86]. Many [Cal16]. Many-Body [Cal16]. Map [BDLP11, FP12, Hwa14, LY13, Lu89, ZHY09]. Mappings [CMM^{+}09]. Maps [APT06, SD16, CGP23, XJN21]. Marcel [Tél94, Tri88a]. Marcum [GST14]. Marginal [SSK82]. Marginally [Hic83a, Hic83b]. Marginals [Gez79, Gez80, Sch80]. Marine [DLW89]. Mark [BF16a, BF16b]. Markov [DK05, PB03]. Markov-Modulated [DK05]. martensitic [WBZ21]. Mass [ACM18, FH23, HW21, LP23]. Mass-in-mass [HW21]. masses [McG23]. massive [CF23]. Master [BF85]. Matched [FR17, Mac97]. Matching [BBC^{+}10, FR17, HK73]. Matchings [GKM74]. material [LLZ23]. Materials [AMR91a, AMR91b, Lus80, RS10b, Wri85]. Math [BMOT09a, UG84]. Mathematical [DARB11, Edw97, Tro17, YBSM08a, YBSM08b, DDAAB20, Pet22, WCYW19]. MATHEMATICS [Ano06]. Mathieu [HG81]. Matrices [AGJ03, CS80c, Kat88, KK90, Pat70, VS95a, WF17, IW23]. Matrix [Abr71, BFM17, CCM12, CI98, GWH11, LB69, PB03, POvdMG18, Str86, VS94, Woo95, ZZ13, CCM19, Cha21, DIW19, DER21, DvdM23, LWAO21, LQYjZ22, XB22]. Matroid [Kun78]. Matroids [Bry79b, KO84, Kun78, Kun80, Oxl82]. Matter [Gre72a]. Max [JO07]. Max-Plus [JO07]. Maxentropic [Gzy00a, PB03]. Maximal [FK77, Wri20]. Maximally [Wan76a]. Maximizing [KKR71]. Maximum [BS01, Can78]. Maxwell [AP23, ALZ21, PS99b, TV22]. Maya [CGUGM20]. Mean [BH89, BP71, BC89, Cam04, CS12, DS99, DM96, Gri77b, Hen88, Lu93, MM93, PS96, Yan14a, Cai22b, KA19]. Mean-Field [Yan14a]. Mean-Flow [Cam04]. Means [Urb89, Urb90a, Urb90b]. measurable [PV23]. Measure [GM75, Gri82, GZ13, Lap87, MN22]. measurement [VMA21]. Measures [GLH22, MMRL23]. Mechanics [Mah96a, Wan73a]. Mechanism [BG81, FL73, HS82a, YC15, LP23]. Mechanisms [KS18]. Media [Amu17, BS16, Dre71, HY86, Lan73, QK94, SS03b, Var97, BO20, HHZ20, ISFH20, LMS19, MKP^{+}20, SA19]. Medium [AO02, CR86, GK95, Mat89, PP90, Pas90, PS00, PT18, QC96, SP88a, Saf71a, Tam89b, TA94, Yue81, KT22]. Meinhardt [Nec12]. Meixner [WQW13]. Mellin [MWL23]. Mel’nikov [LFW20, Li05]. Membrane [GP87, Wax84]. Memory [AY17a, AY17b, Gro73, Kne03, Aou20, SSW22]. Merhasin [Ano05b]. Meridional [BM04]. Meromorphic [CR06, CNW12]. Mesh [MW73]. Mesoscopic [ZGQ12]. Messiter [TNM96]. Metamaterial [MML17]. metamaterials [ACZ23]. Metastable [War94, WS99]. Method [AKS^{+}08, BN94, CFW87, Cau18, Che04, CD21b, CI17, EG95, EHV14, FR17, FS69, GRTW85, HCB72, Hwa14, Isr70, Kau75, Kau78, Kau80a, KK91, LK16, LLC10, LP11, Mac97, MC06, MML17, Mil01, MO01, NA84b, OK10, OK15, OW04, PPV08, RD04, TWK00, TEW91, VS88a, Vas11, Wan73b, Wan73a, Yan15, BW19, CCD22, DS23, KHJK21, SCFFLPA23, bXlZS23]. Methods [Abl72, AH75, CK75, DRS74, GT80, HM77, Lin97, VGW15, YL07, YL08, SO20]. metric [MN22]. Microflora [MMPS13]. Mid [Swa18]. Mid-Latitude [Swa18]. Mikhailov [GY15]. Mills [ACT87, BH16, LQYjZ22]. Minimal [GY95b, Kal73]. Minimizing [Lie77, MRE19]. Minimum [CA09]. Minnaert [FA22]. Miodek [LW22]. Mirakyan [Sun93]. Mirror [HY00, CCD22]. Miscellaneous [Mah96b]. Miscible [HY86]. Mixed [Che94, SS06, SS03b, CCMX22, Xu21]. Mixing [DR90, GB96, JMT08, MM99a, Mal01b, MS76]. Mixture [Gre89, Gre90]. mKdV [DN11, Len16, LS23]. mKP [CZjZ21]. Möbius [Coh91, EE19, Gre75]. Möbius-invariant [EE19]. Modal [FA22, KL84, YM04]. Mode [AHNZ09, Bri90, Gri86, HS84, Sha91]. Mode-Interaction [Bri90]. Mode-Locked [AHNZ09]. Model [ABGML71, Abr71, ABGML72, AM08, AM09, BMM87, BM72, Bar83, Bar84, BF85, Bar16, Bax71, Boa75, BM14, Boo89, BD17, BS93, CBB87, DKZ16, Edw97, FSY78, FMM99, Gal71, GXLX16, GKK18, GPT^{+}05, HMM87, HS82b, HP78, JC08, KPS80, Ken76, KWW05a, KWW06, KS91, LO99, Len09, MMPS13, Mal94, MS76, MPS98, Mor86, MK09, Nec12, Nis79, PS09, RANC09, SG90, Saf74, SF85, ST17, SH76, Sta76, SDS12, Swa12, TBKiM18, VBA97, Vil91, WVBM14, WW02, WW03, WF12, XQZ15, Yan14a, YK08, YLP95, YG01, ZN17, BGMR21, CZ21, CLW22, CF23, CI20, DIW19, DDAAB20, – Dor22, ELV21, FSKMP21, GLH22, GS23, GYZ21, HMBYZ22, HP19, Jin23, JFM21, KHS22, LG20, LSG22, LP23, LCTZ22, LBP21, MWL23, MEHS22, PRAV21, RMMH22, SK23, SSW22, TNT21b, TNT21a]. model [WX22, WCTZ22, WLR22, WW21, WCYW19, WBZ21, XHW22, XYZY22, ZYZ22, Zhe20, BGT07]. Modeling [AO02, AO03, AKS^{+}08, DARB11, MPB00]. Models [Abr72, CJ18, DK05, Gre72d, GŠ06, HSS17, Hun69, Kne03, MSW07, PG96, VSS98, War96, BL22, GKK21, KK22b, KHJK21, SZ23, Slu19, VMA21, YLW21, ZC21]. Moderately [GMW98]. Modes [AB70, Abl72, CM07, FYC05, GB96, Gea85, Isr72b, Le 04, LKM14, NY16, TG01, TU05, Tur05a, Tur05b, Tur07, Sul20, ZSA22]. Modified [DK09, GM89, GT90, GPT^{+}05, KD92, PS10, SRE05, SB00, TZHL09, CCY23, KDM22, LCG21, Lóp22, Luo21, SYF22, XF23b]. Modon [Swa04]. Modulated [Alm90, DK05, Hab88, Yan14a, ACZ23, HHZ20, MC21b]. Modulation [Bau98, Bri94, Bri15, CC17, DH96b, Gri77b, HSS19, JZ10, Rat18, SS03b, BHIS21, BR19, CM21, CMP22, GLS23, LW22]. Modulational [BH14b, EFM84, HJ15, HP19, SDS12, BHIS21]. Modulations [FM83a, Hab91, Ken74, Kli03]. Modules [Rei89]. Moisture [BS16]. Molecular [FCP^{+}21]. Moment [BCKW85, Rei73a, Rei75, YY04, CH18]. Moments [BR73, BW79]. Momentum [Che69, GZ13, How90a]. Monatomic [Fos70a, Fos70b]. Monodromy [DDN21]. Monomer [YY08]. Monopole [TZHL09]. monotone [ELV21]. Monotonicity [KS81]. Morphogen [LNVW05, LNW04, Wan14]. mosquito [DDAAB20]. mosquitoes [ELV21]. Motion [BSGC99, BSCS00, BH14a, BT80, Dha94, Gri77b, JZ86, KO04, Lac82, LO95, LPH09, Lun03, Moo78, NS91, Ran94a, Ran96c, Saf72, Yan95a, LZW23]. Motions [Gub74, MPD05, Ran98, RS12, Urb89, Urb90a, Urb90b]. Motor [Gro69]. Mountain [DM96, XYZY22]. Moutard [DRS02]. Movable [FH09, JK94, Sob04]. Moving [FR08, GM12, GM13, GPS96, Hal12, Kne01]. mucus [FCP^{+}21]. Multi [AB70, Abl72, AKZ16, CMKP16, FMO17, FLT20, GY95b, KR03, Ma79a, MP19, PBT11, LS23]. Multi-breather [FLT20]. Multi-Component [PBT11, AKZ16]. Multi-dark [FMO17]. Multi-dimensional [MP19]. Multi-Duality [GY95b]. Multi-Phase [AB70, Abl72, KR03]. Multi-Site [CMKP16]. Multi-soliton [Ma79a]. Multicommodity [Ulr75]. Multicomponent [Bła23, Mat89, SS14]. Multidimensional [ACT87, AS89, Ben06, FA83a, MR92, MP22, NA84b, NA84a, Vas11, War96, Sin21]. Multidimensions [BLP88]. multifeedback [Wan21]. Multihump [PY05]. Multilayer [BS16]. Multiphase [BLP01, Dec00, Rat18, Ros76b, Rat19]. Multiple [Boa75, KB98, LL22, Mor89, NAY17, OK10, RS80, VW23, WLS17, Yan98, YT14, Wan19, ZTYH20]. Multiple-Scale [NAY17, KB98]. Multiplication [FMRR77]. Multiplications [Che94]. multiplicative [GSYG21]. Multiplicity [PSS75]. Multipoint [PS18]. Multiscale [KHF^{+}20]. Multistep [BHH^{+}11]. Multisymplectic [BLP01]. Multitime [HM77]. Multivaccination [ZN17]. Multivariable [Zen91]. Multivariate [JRS99, Jon80, Olv06, Rei77]. Multiwavelets [Str97]. Mush [LR01].

Naghdi [DIT16]. Nagumo [GRMOH01]. Nagumo- [GRMOH01]. nanoptera [DL23]. nanopterons [FH23]. Narita [Wan12]. Natural [Hal12, SJB91, Yu10]. Navier [AZ22, BW02, CCY23, GG76, LCB99, ÖSV01, Ran94b, Ran94c, Ran95a, Ran95b, Ran96b, Ran96a, VS96]. Near [AB82, ACMR01, BS01, CA09, Edw00, KA88, Ken74, LP11, Ma84, New91, XYZY22, FA22, MWZ22, SZWM16, Xu21]. Near-Critical [Ken74]. Near-optimal [XYZY22]. Nearly [Kev82, DDM21]. Necessarily [LY11]. Necessary [BM99, Mar71]. necessity [Wan19]. Needlelike [Xu90]. Negative [BP71, Gri86, HN96, Yu15, LLZ23]. Negative-Energy [Gri86]. Neighboring [LKK80]. Nekrasov [BS01]. Nematic [LAHM17]. Nerve [FMM99]. nets [– Dor22]. Network [AH75, CK75, Ulr75, GBSS21]. Networks [DGW02, Gro69, Gro70, Gro73, MS08]. Neumann [BDLP11, Che11, CCMX22, DD11, FP12, Hwa14]. Neural [Gro73, GBSS21]. Neutral [Wal01, FSKMP21]. Neutrally [Vár13]. Nevanlinna [CR06]. Newton [Lun03, MRW07]. Newtonian [KR78, Mat89, PP90, Pas90]. Nim [GY95a]. Nim-like [GY95a]. Nizhnik [GWCH23]. NLS [AE22, ACS18, BLP01, CM21, RS98, RP11, Rog14a, VPE09, WZ22, ZTYH20]. No [JK94]. Nodes [GST18]. Noether [GM12, Olv21, Pen17]. noise [CCY23, GSYG21]. Non [Atk78, BPS73, BM12b, BK95, Gal06, Lev69, NY16, PP90, Pas90, PG96, Rei70a, SY15, Tho69a, Yan14b, BM22, BZZ21]. Non- [NY16]. non-autonomous [BM22]. Non-Boussinesq [BM12b]. Non-existence [Tho69a]. Non-Hamiltonian [BK95]. non-Hermitian [BZZ21]. Non-Linear [Rei70a, Lev69]. Non-Lipschitz [Gal06]. Non-Local [Atk78]. Non-Newtonian [PP90, Pas90]. Non-Parity-Time-Symmetric [Yan14b]. Non-Stationary [SY15]. Non-Uniform [BPS73]. Nonautonomous [AT07a, BMD22, BDM22, Bła23]. Nonaxisymmetric [Gus89]. Nonbreaking [Lus80]. Noncanonical [BLP88]. Noncentral [KV06]. Nonclassical [AH95, EG95, HSS19, PS20b]. Nonconstant [LWC23]. Nondecaying [VA02]. Nondispersive [Kau76, RTT02]. nondissipative [SK23]. Nonevolutionary [MNW07, NW07]. Nonexistence [JPS00, ZSA22]. nonflat [PV23]. Nonhomogeneous [RY10, TV22]. Nonhydraulic [MP01]. Nonhydrostatic [AM03, AM05]. Nonhyperbolic [DH02]. Nonideal [SS17]. Nonintegrable [BH97, BH98a, Kic99]. Nonlinear [AB70, Abl71, Abl72, AKNS74, Abl75, AL76, AL77, AS89, Abl01, AM17, AB80, Aky82, ACMR01, AB07a, ALB09, AB10, AH95, BD01, BDLP11, BGT07, BG07, Bar16, BB69c, BL70, Ben73, BM75, Ben79a, BG81, Ben83, BC85, BZ92, BG83, BH97, BH98a, BB12, BP14, BC76, BM14, BGL97, Bry79a, BBC^{+}05, CM07, Cal84, CD04, CD21a, CCD22, CR86, Che11, CM13, Che84, CM03, CMM^{+}09, CEHS19, CMKP16, DLW87, DPvdMV13, DB03, DdY11, DNP17, DFKZ14, DFK16, DR90, DA05, DDM13, DKZ16, ET87, EHV18, FG12, FM11, FLM95, FR08, FA89, Gaj96, Gal11, Gao95, GB96, GWH11, GLZ15, GKT10, GOS11, GKK18, Gri81, Gri85, Gri95, GPT^{+}05, Gri16, GL17, Gro02, GV03, GK95, GLL13, Hab75, Hab77, HjS85, Hab88, Hab91]. Nonlinear [HS82a, HS84, Hal01, Hal06, HM91, HM93b, HM94b, HL11, HG01, How77, HP78, How83, HTWW98, Hu96, HMR86, HAY12, IR86, JQ13, JC02, JC08, JS99, KQ10, KA88, KS96, KC82, KKM87, KPS80, Kee76, Kee82, KM70, Ken74, Kev80, KL84, KL88, KB98, KMP14, KP19a, KP19b, KM90, KA02, KM04, Kon14a, Kon14b, KLW11, LL80, Lan87b, LM91, LM16, Len09, Len16, Li02, LKM14, LK16, LT07, Lie77, LKK80, LB81, LY11, LRS04, LNW04, Ma79a, Ma82, Ma84, MR84a, MRS88, MD99, MM99a, Mas72, MC02, MM90, MQ15, Miu74, MM03, MB84, Mos14, MO01, Nau92, NPT08, NC15b, NAY17, Nou00, Oht09, OWY11, OCB87, ÖM94, OB94, OW04, PP16, PS00, Peg88, PY05, PCZP00, PBT11, POvdMG18, PJ86, QC96, Qu97, QK14, Red80, Rei69, RS10a, RS10b]. Nonlinear [RMA12, Rog14b, Ros78, RS80, Ros76b, Rub77, RD04, SBA05, SKP^{+}05, Sha91, SA19, SR99, SS17, SA04, SB87, Sob04, SY06, Sri91, SS03b, STZZ14, Swa07, SDS12, Swa12, TG06, TKC81, TR97, Var97, VM11, VSS98, Wad91, Wan69a, WK91, WW78, WNH12, Wit96, WB98, Woo84, Wri85, XF16, Yan14a, Yan95b, Yan96, YB96, Yan97, Yan98, YA03, YL07, Yan12, YKH00, Yue81, ZC10b, ZGQ12, Zho09, ACR19, ACR23, BF16a, BF16b, BW19, BO20, BLM21, CBJCN20, CZ20, DFG19, FLT20, Fro19, HHZ20, HS19, Jin23, KHF^{+}20, LTY22, LWZ23, LZW23, Lu23, Luo21, MHW20, RS21, TV22, Wri20, XF23a, Yan21, YZ20, ZSA22, Zha19, FG12]. Nonlinearities [CM13, Gal06, CI23]. Nonlinearity [CZ15, FSY78, SS03b, Yan14a, CBJCN20, DWZ21, DZ22b, PRAV21]. Nonlinearizable [TV22]. Nonlinearly [GMM05, JA16, HSS19]. Nonlocal [AM17, AFLM18, ADM82, Ash10, CDLjZ18, Gri95, HLQ09, MTW16, PS18, RCH17, YY18, Zho18, ALZ21, CCMH21, CLV20, GJLW22, HMBYZ22, ISFH20, KHF^{+}20, LFW20, Lou19, MHW20, MMO23, RS21, WZ22, bXlZS23, XF23b, YZ20]. Nonlocality [GS16, Lou19]. Nonmonotonic [Hag81]. Nonparallel [DB03]. Nonparallelism [HS84]. nonparity [Yan21]. nonparity-time-symmetric [Yan21]. Nonplanar [AB10, RS08, SRJ99]. Nonpolynomial [Lin14, MB09]. Nonreceptors [Wan14]. Nonrepetitive [Li76]. Nonsolvable [MR02]. nonstandard [Lóp22]. Nonstationary [Gaj96, Tur05a, Tur07]. Nonsymmetric [NKUW02, LMS19]. Nonsymmetrical [Bau99]. Nonuniform [FCCTR22, Sin21]. Nonuniqueness [Gal11]. nonvanishing [FH23]. Nonzero [AFLM18, DPvdMV13, PBT11, PV16, CCMH21, MZ23, OP19]. Norm [Haj92, Lu23]. Normal [Le 04, Mor86, Esc23, FCP^{+}21]. Normalized [HK73]. Note [Abl75, AK82, Cao13, Cho90, Cro98, Gre72b, Gre69a, Gre70, Gre90, HK74, Har75, Ioa89, Kle75, KL75, Kne01, Lus80, Rei75, SS03a, VB76, Wan76b, BL22]. Novel [Rog14a, GSYG21, Luo21]. Novikov [GWCH23, FdST22, SS14]. Noyes [SH76, Sta76]. Number [BDGL02, Can78, Cha71, CK04, FM83b, Gro70, HS82a, Hal12, KKR71, KK91, MM01b, Ric84, Saf71b, TWK00, Ven92, WZ08, Wor86, BL22, Dur22]. Numbers [Ben84, BMB02, ELY06, GLN14, KV06, Mén96, Pec80a, SJB91, Tem93]. Numerical [AKS^{+}08, BS01, CS80b, CJ18, DLK16, EY91, FKMS22, GST14, GGS95, GT90, IR86, Isr70, KS18, KS20, LW95, LHC04, MMPS13, MW73, NLB10, Ors71a, PPV08, RF13, RF16, SRJ99, STWW03, Urb90b, Yan15, YLP95, ASSV20, CD21b, DL23, NR22, CS80a, DH19, SR99].

Objective [Bry91]. Objects [Vár13, Wax80]. Oblique [GB96, GZ94, KKS15, KA19]. Observations [Her91]. Obstacle [AT07b, Gri90]. Obstacles [Gzy00b, SCFFLPA23]. Obstructed [Hel74]. Obtain [LT07]. Obtaining [Abl72, VS88a]. occasion [CD21a, KP19a, KP19b]. Occupancy [Dou72]. Ocean [VMA21, DDM21]. Ocean-depth [VMA21]. Oceanic [Gri16]. Oceans [HCB72]. Odd [FG12, Gri78]. Odd-Order [FG12]. ODEs [Bła23, FH09]. Off [BT80, HS19, KA88, TNT21b, TNT21a]. Oikawa [CFiMO18]. Oil [AKS^{+}08]. Okamoto [YY23]. O’Malley [Sri88, WY03]. One [BM72, FYC05, GMM05, HS82b, KWW05a, KWW05b, MW98, Mos79, Nec12, Rei73b, Sta15, SW00, STWW03, Tho69a, TK69, TBKiM18, Tro17, War94, ZGQ12, CCD22, Lu23, NR22, Yan21, YLW21]. One- [BM72, GMM05, Yan21]. One-Dimensional [FYC05, KWW05a, Mos79, Rei73b, Sta15, STWW03, TBKiM18, War94, ZGQ12, Lu23, YLW21]. One-Parameter [MW98]. one-phase [NR22]. One-Space [KWW05b]. Online [HZ10]. Ono [FA83a, KM98, MW16, NS21, RRY22]. Onset [CKLS96, Wal82, YG01]. Open [Bry77]. Operational [CS90, Lap87]. Operator [BP14, CV09, FS86, Fre72, Hoc72, LLS04, Sun93, Tri88b, VS94, Wan12, YL07, XB22]. Operators [ADM82, And71, BLP88, GY15, Jon80, PEE04, BHN21, BZZ21, FÖK22, YS21]. Opposite [Tan86]. Optical [ACZ12, BB87, BO12, BAM98, CP12, DA05, GMM05, Len09, Yan96, Yan97, MWL23]. Optically [FYC05, YMBC04]. Optics [DK05, Kne00, Kne03, KM04, RMA12, SS03b]. Optimal [AB07b, FN05a, Gez79, GKK18, GKK21, Iwa82, Lip06, OH95, Pec80c, SSVS20, Wan80, WA83, Wan85, WSK10, WE17, Slu19, XYZY22, Gez80, Sch80]. Optimization [BK78a, BO12, Urb90b]. Optimum [Bar71, Urb89, Urb90a, Urb90b]. Option [Kne01]. Orbit [BH98b, DH02]. Orbital [AZ13, PP16]. Orbits [Li02]. Order [Bag08, BP71, BGL15, BS88, BS89, CR06, Cos93a, Cos93b, CS93, Cos00a, Cos06, FG12, Gal06, Gal08, Gal10a, Gal11, GLL13, Hab75, Hoc72, HM08, KT89, KT90, LR69, Liu13, LST06, Mal02, Mar02, NA84a, PEE04, RANC09, SR97a, SR97b, SBA05, SY03, Sob04, Sob05, Sob06, SDS12, Var97, WL92, WZ07, YA03, YT14, Yu15, Zau72, AE22, ADH^{+}21, BP20, CGUGM20, Cos00b, FLT20, GSYG21, HSS19, KHJK21, KHS22, LCG21, LW22, Lu23, Mil01, Olv21, RMMH22, YS21, YLW21, ZTYH20]. Ordered [Bel79, Gro69, WA83, Wan85]. Orders [HK73, Hu96, KS81, KS91]. Ordinary [Bag08, BS89, BS93, Dun01, Dun04, LPH03, LR69, SBA05, Sob04, Sob05, Sob06, Woo95, BJV22]. Oregonator [Has76]. Ornstein [GYZ21]. Orr [BSGC99, BSCS00, Rei72, Rei74a, Rei74b, Rei79, Wil72]. Orthogonal [CCM12, CI98, CD18b, DK09, FOX23, GJW77, LY13, WW05, XZ11, Cha21, CCF19, DER21, DFS22, FIM23, IW23, IL20, LL22, LDLW20, MFIGH22, MMRL23, MC21a, VW23]. Orthogonality [Fre87, Str97]. Orthotropic [LW88, LW90]. Oscillating [Mil93, Vla17, Sul20]. Oscillation [Sch75]. Oscillations [BS88, Gri12, Hun69, PY00, SH76, GM21]. Oscillator [FH78, TK69]. Oscillators [GA79, GA82, Gri86, Gri87, Hu96, JA16, KL88, Mil76, MO01, Rub77]. Oscillatory [AY95, EY91, Gal06, Hab88, Hab91, vKD72, FLS23]. Oscilliations [Yan97]. Oseen [Ran95b]. Ostrovsky [GGS95, GH08, GHJ12, HSS17]. Ostrovskyi [PS20a]. outbreaks [SSVS20]. outbursts [BGMR21]. Outflow [Mal94].

P [UG84, Ano05b, Cos06]. P2 [Cos00b]. Pacific [ZC21]. Packet [Cam04, Gri77b, LN13, MM99a, MBM94, Cai22a, Cai22b]. Packets [Ake12, BH06, Ben83, GA11, MB84, OCB87]. Padé [Bak76, DLK16]. page [Ano16a, Ano16i]. Painlevé [Air81, Abd97, Air79, AK20, BCH95, BMD22, BDM22, BM22, Bła23, BMS18, CCM19, CZ10, CD18a, Cla16, CGUGM20, Cos93a, Cos93b, CS93, Cos97, Cos00b, Cos06, DK09, DFS22, FC08, GRTW85, Hab77, HJ96, JK92, JK94, Jos99, JK01, Jos15, Kok20, KC92, LLLZ17, LC21, MW98, MV13, MCB97, NRGO01, RF13, Rog14a, SHC06, Sob06, VA03, WXZ18, XZ11, XZ13, ZXZ15]. Painlevé-III [BMS18]. Painlevé-type [BDM22, BM22]. Pair [Bar84, Fan11, GB96, Gri78, Saf72, Tan86, YK02a, Aik19, AZ22]. Pairs [CSS77, KKR71, LY03, LKM14, SHC06, Wes78, XF16, MMRL23]. Parabolic [Gal06, Hag81, QK14, Rus73, Aou20, CR21, Seg21, Xu21, Cos93b]. Parabolized [HH94]. Parallel [Ash10, BB69c, BM75, BG83, CD90, Dun13, FLM95, FH98, GT04, Hab72]. Parameter [BOGK04, MW98, Mar18, PPV08, Tam93, WQW13, Jos99]. Parameterizations [CA10]. Parameters [ASSS12, Isr72a, Kev80, KL88, OK15, RF13, SW17, Yan14a, Dur22, LZZ23, XXZ21]. Parametric [GKT10, Mal94]. parasite [CR21]. parasitoid [CR21]. Parity [BD17, DFKZ14, Yan14b]. Parity-Time-Symmetric [BD17, DFKZ14, Yan14b]. parliament [BG23]. Part [AY17a, Gri77a, AY17b, BG07, Ben06, BF16a, BF16b, Boy16, Cai22a, Cai22b, DLK16, Fri74, KP19b, LS93, MB95, Mil95, NC15a, NC15b, Rei74a, Rei74b, SR99, Sta71a, YM07, YKdM^{+}07, YBSM08a, YBSM08b, ZC10a, ZC10b]. Partial [AL77, Che84, Cos93a, Cos93b, Fow90a, HK73, KS81, KL81, MS16, PG95, PS20b, Rus73, SV92, VS88a, Wan72, CLW23, Olv21, SO20, ZL21]. Partially [Bel79, BBC^{+}05, Hel74]. Partially-Obstructed [Hel74]. Particle [Bla70, CKL16, LPH09, Moo00, MP01, Esc20]. Particle-Driven [Moo00, MP01]. Particles [AKS^{+}08, FPSA89, NS91, PF89, Stu92]. Partition [CBB87, Hei71, Hol80, DIW19]. Partitioning [BBC^{+}10]. Partitions [Gri82, JK84, Spe74, Sta71a]. Partitions.Part [Sta71b]. Pass [DM96]. Passage [BH98b, DH02, Kev82, Kne00, Rub77]. Passenger [Bar73]. passive [ACM18]. Past [AB03, Che69, Gri90, GY93, Kro77, LK10, TWK00, TW97]. Pasta [FH23]. patches [BS22]. patchy [CLW22, LP23]. Path [Bel79, KS06, LK16].

Path-Based [LK16]. Path-Connected [Bel79]. Paths [ELY03, LMY08]. Pattern [Kee78, MR80, NV17, YS98, LWC23]. Patterns [DdY11, Gal08, Gro69, KH81, MPB00, ST17, War94, WW02, WW03, GJLW22, LW22, WW21, YY23]. Pavlov [GS16]. PDAE [RLW01]. PdE [KT99]. PDEs [FG12, FP05, HF17, JS99, LC21]. Peak [Haj92, LMY08]. Peak- [LMY08]. Peakon [GXW13, XQZ15]. peakons [ACS18]. Peaks [ELY03]. Pearson [MFIGH22]. Penetrative [Str11]. Pentomino [She82b]. Peregrine [KS20]. Period [FM83b, Urb90b]. Periodic [AAA08, AZ13, BH97, BAM98, BDGL02, BBC^{+}05, CBJCN20, DR77, DN11, DB94, DA05, FCL12, GXLX16, GWH11, Gri15, GL17, HSS17, Har11, Has76, JW06, Joh10, Kev82, KH81, Lan80, MA81, MM93, NAY13, NY14, PF89, PP16, PZ03, SER^{+}03, SRE05, SF85, TB02, Urb89, Urb90a, Urb90b, VP10, Ven92, Yan15, YY03, ABM^{+}22, BO20, Cha21, CZ20, FvdW21, Fro19, GMNP22, HCC21, LFW20, Min86, NdJVA22, SCFFLPA23, TMPVB19]. Periodically [Kat81, RC80, RC81, SZ23]. periodicity [MMO23]. Permanence [NPT08, GS23]. Permanent [Bry77, BL78, CS80b, TCK82, Yan98, TNT21b]. Permanent-wave [Yan98]. Permeable [RD04]. permittivity [TV22]. Permutability [BH16]. Permutation [CL91, FZ90, GJ84]. Permutations [RT79]. Persistence [AT07a, JS99]. Persistent [Li02]. Person [PS11]. Perturbation [ABGML71, ABGML72, BD17, CE73, CLJ97, Cro70, FY01, GKT10, JW99, KT89, KT90, KPS80, Lan83b, LS88, LM91, LS93, LST06, LK10, MBT00, MR80, Ros76a, Swa12, TNM96, VA03, Wan77, WM94, TYZZ21]. Perturbations [Abr77, AMT18, CD90, HHW94, KA81, LW91, Mar71, STC09, Sha91, Swa07, WK91, BZZ21, MWZ22, YS21]. Perturbed [BL87, GM93, Gri95, How77, How79, How82, JW06, Kat81, KKM87, Kee85, KO04, KH81, LN94, LW95, Li02, Li05, LLC10, LS04, Ma79b, Mil93, New91, OK09, OK10, OW04, RC80, RC81, Tov00, War92, YT14, ZXZ15, BL22]. pesticide [XYZY22]. Petviashili [VA02]. Petviashvili [BMOT09a, AYF83, AV91, AW97, BMOT09b, BBN22, DFS97, FA83b, FS86, Fok09, FN22, GM89, GT90, JPS00, KD92, RCMH21, Rat19]. Petviashvili-Type [AW97]. Pfaff [Con91]. Phase [AB70, Abl72, CA09, DS71b, Dre71, DFS97, EFM84, Hab88, Hab91, Kat85, KR03, MDG10, NV17, Pas90, Tho69a, TK69, UG83, BCD21, LBP21, MKP^{+}20, NR22, UG84, WBZ21, Zhe20]. phase-field [WBZ21, Zhe20]. Phase-Speed [CA09]. Phenomena [AO03, Kon14a, Kon14b, BOPS21]. Phenomenon [Bla70, Kau76, Tov00]. Phone [Leb73]. Photonic [FYC05, SKP^{+}05, YMBC04, YF05]. Physically [AAA08]. Piecewise [Hen88]. pine [XYZY22]. Pipe [Ber92, Gus89, Wal01]. pipes [ACM18]. Pitaevskii [OF23, WTW13, Yan14a]. Pivot [FA83a]. Planar [BBC^{+}10, Hol80, KT89, KT90, WVBM14, FOX23, LZW23]. Plancherel [IL20]. Plane [Ber71, DdY11, GY93, Gus86, Has76, HS92, KA02, KH73, Len16, LW88, LS04, Ma79b, New91, NC15a, NC15b, Rei79, Sha91, Sta71a, Sta71b, Swa18, Wri06, Rei74a, Sta76]. Plane-Wave [Ma79b]. Planes [Kat85]. Planetary [ET87, Gre69a]. Planform [PR94]. plankton [PRAV21]. Plant [GXLX16]. Plasma [AO03, Boo89]. Plasmodium [DDMMSYD22]. Plasmons [MML17]. Plastic [Ber72]. Plasticity [Ber71]. Plate [KKS15, Lia06, LW90, MPG17, OC70, SS06, Tam93, Tan85, ZL17, DDM21]. Plateau [DM96]. Plates [Ash10, LW88, Rei70a]. Plethystic [Mén96]. Plume [Wor86]. Plus [JO07]. Pochhammer [ZZ08]. Poincaré [CKL16, Lu89]. Point [Abr77, AT07b, CD18a, Dun01, Edw00, Gal11, KC05, Lak72, LT02, LP11, Nou00, OW03, Wax80, Zau72, Dun20, FA22, PS20b]. Points [BH14a, Dun11, EHV14, How77, HP78, HG81, KPS80, KKR71, LW91, ADH^{+}21]. Poiseuille [Ber92, DS99, Gus86, PCZP00, Sha91, Wal01]. Poisson [CLV20, DF15, VM11]. Pol [FH78, JA16]. Polarity [GPT98, Gri15]. Pole [Dun04, LP11]. Poles [BH14a, Ioa89, JK94, CF22, ZTYH20]. Pollaczek [CCF19, HW20, MC21a, RW96, WQW13, ZZ07]. Pollak [Kal73]. Pollutant [AKS^{+}08]. Poly [KC92]. Poly-Painlevé [KC92]. Polygon [DF15]. Polygonal [HF17, ACM18]. Polygons [Sar01]. Polylaplacian [PP19]. Polymerization [GV03]. Polymers [Edw97]. Polynomial [Cos00b, Cos06, OAVS97, Rei78, VS93, VS95a, WTW13, Dem23, YY23]. Polynomiality [MB09]. Polynomials [AGJ03, AMT18, BK18, CCM12, CS90, CFS92, CI98, Cla09, DK09, Di 94, DL94, Fat74, FC08, GJW77, GUGM18, Haj92, JTLW13, LY89, LT99, LY13, Mén96, NKUW02, PW12, PW14, RW96, SW17, SP88b, Sta87, Van87, WW05, WQW13, XZ11, ZZ07, BCD21, BDS20, Cha21, CCF19, CD18b, DIW19, DER21, Dur22, DFS22, FOX23, FIM23, GFGUMM22, HW20, IL20, LL22, LDLW20, LT23, MFIGH22, MMRL23, MC21a, VW23, dlIJ23]. Polytropic [BDGL02]. Popov [GM21]. Population [NPT08, ZN17]. Porous [AO02, BS16, GK95, HY86, KKS15, MPG17, PP90, Pas90, PS00, QK94, QC96, RS07, Saf71a, Str11, Tam89b, TA94, BCCM23, MKP^{+}20]. portrait [BCD21]. posedness [CR21, DZ22b, Lu23, PRAV21]. Poset [Gri82]. Positive [GP87, Gri86]. Positive-Energy [Gri86]. positivity [PRAV21]. Possible [CK13]. Postbuckling [Rei70a]. posteriori [NR22]. Potential [ADM82, Ack73, AO03, CJM92, DFK16, GP14, HN96, Mos79, Yan14a, ZC10a, AP23, CI23, ZL21]. Potentials [BH14a, NAY13, NY14, NY16, War92, Yan14b, DvdM23, FdST22, Yan21, ZSA22]. Powder [DARB11]. Powder-Snow [DARB11]. Power [Haj92, KT22]. Powers [Fat74, SJB91]. Prandtl [Cha71, FM83b, Hal12, Wor86]. Predator [ST17, WWZ10, CZ21, GJLW22, XHW22]. predators [ZYZ22]. predictability [KHS22]. Predictions [MS76]. Preface [AY17a, AY17b, BF16a, BF16b, CD21a, CD18b, KP19a, KP19b, Kon14a, Kon14b]. preferences [DDMMSYD22]. Preliminaries [Gre72c]. Prescribed [DM96, Wan76a]. Presence [CG07, CG08, CR86, FM11, YGJ18]. Pressure [Fow93, RJM78, YLP95, ZL17]. Presymplectic [Bła11]. Pretwisted [Wan75]. prevention [WCTZ22]. Prey [ST17, WWZ10, CZ21, GJLW22, LWC23]. Prey-taxis [ST17, LWC23]. Principle [Bar71, Sri88, Swa18]. Principles [RS10a]. Priori} [Abr77].

Prismatic [Wan93]. Probabilistic [Gzy00b, Kfo70]. Probability [DKZ16, Gzy00a, Hwa97, MC18, NK05, FSKMP21]. Problem [AB07a, AB07b, AH75, BW86, Bri90, BS89, Cal16, CRW16, CJM92, CK75, Cla78, CE73, Cro70, DM96, FR08, GWH11, Hal01, HM91, Hen88, HM15, HS82b, HS69, HP78, IR86, JK92, Kau80b, Kau80a, KPS80, Ken76, KMP14, Kne01, KH81, Leb77, LST06, MBT00, Mon74, Mos78, OAVS97, Oli88, ÖM94, OW03, PZ03, QK14, SKB08, SRD04, Sri88, Tam89a, Ulr75, Ven92, Vil90, Vil91, VA02, Wan75, WSK10, Wil72, WY02, WZ08, XF16, XL15, YY08, BS23, CBJCN20, CCM19, CCMX22, KT19, LTY22, LW22, NdJVA22, NMBD23, OF23, PCH23, PV23, SXZ22, TV22, WCYW19, XF23a, Xu21, XB22, ZS22, Zhe20, CK04]. Problems [Abl72, AKNS74, AO70, AO02, AO03, AH75, BK78a, CCM12, CK75, Cos97, FR17, FA83a, FR01, GY95b, GLZ15, GP87, GK74, HHW94, How77, How82, JT87, JS01, KS14, KKM87, Kau80b, KR78, KLW11, Lak72, Lan83b, LM91, LW95, LLC10, Lip06, LS04, LNW04, MW73, MQ15, MTW16, NAY17, OK10, OW04, PMAD07, PS18, PB03, PT18, Ros76a, RS80, SBA05, SD15, Vár09, Wan73b, Wan73a, Wan77, WK91, War92, YK02a, YK02b, Zau72, AE22, CD21b, CEHS21, DvdM23, KR23, LT23, MN22, MKP^{+}20, NR22]. Procedure [RKS83]. Procesi [CI17]. Process [CD18a, Wal95, ZGQ12, GYZ21]. Processes [GRMOH01, Hal12, HY86]. Processor [YK94a, ZK13]. Produced [DMPS96]. Product [NKUW02]. Products [HK73, WF17]. Profile [DR90, DH96b, Urb89, Urb90a, Urb90b]. Programming [Cla73]. Prohibitions [Lou19]. Prolate [Mil75, RJRTF17, DDN21]. Proliferation [WE17]. Promotion [Bla70]. Proof [Gal71, JK94, Pec79, Plo02, Sen89, Woo95]. Propagation [BMM87, BK78b, CM08, FR17, Gro02, Lan73, Len09, OCB87, Var97, McG23, WLR22, ZC21]. Properties [BCKW85, BGT07, Ben73, BAM98, CMM^{+}09, FYC05, Nem15, SB87, Swa04, SDS12, YB96, ZK13, ACZ23, BHN21, FÖK22, FCP^{+}21, PRAV21, Urb90a]. Property [GRTW85, KS81]. Proportions [Hsi73, Kle70]. Proposal [Str86]. Propulsion [Urb89, Urb90a, Urb90b]. Proteins [STZZ14]. Prototype [LP08]. Pseudo [GUGM18, SW17, Vla17, FdST22]. Pseudo-Jacobi [SW17]. pseudo-potentials [FdST22]. pseudo-spherical [FdST22]. Pseudo-Wronskian [GUGM18]. Pseudospectral [Ors72]. Pseudospherical [BRT89, CT86, Rab89]. Pulse [AHNZ09, FMO17, GXLX16, HSS17, KWW05b, Len09, LLL12, SFO16, SZWM16, CGP23, GPFT22, ML20, SZ23]. Pulse-Splitting [KWW05b]. Pulselike [FMM99]. Pulses [GPT^{+}05, SZ23]. Pump [ÖSV01]. Put [Kne01]. Pycnoclines [LKK80].

Quad [RH07]. Quad-Graph [RH07]. Quadratic [CWT^{+}15, LB69, Lun03, NS91, VGW15]. Quadrature [RF16]. Quadratures [GST18]. Quadruple [SRD04]. Quadrupole [YMBC04]. Qualitative [ZZ08]. Quantification [RC05]. Quantities [MM01b].

Quantization [BM22]. Quantum [Bar77, BG23, Luo21]. quarantine [GKK21]. quarantine-related [GKK21]. Quarter [Len16, LS04]. Quartic [MV13, RS14]. Quasi [Aou20, Boy16, GWH11, KO13, QJW07, WVBM14, CZ20, KR22]. Quasi-Classical [KO13]. Quasi-Linear [QJW07, KR22]. Quasi-Periodic [GWH11, CZ20]. Quasi-Planar [WVBM14]. Quasi-stability [Aou20]. Quasi-Uniform [Boy16]. quasilinear [DZ22b]. Quasiperiodic [FH10, SR99]. Quasisolutions [PG95]. quasistatic [LLZ23]. Quaternion [KX18, LWAO21]. quaternion-matrix [LWAO21]. Queue [YK94a, YK08]. Queues [BK78a, KT99, KM12, XK93, ZK13]. Quicksilver [Jos15]. Quintic [SZWM16]. QUSS [Boy16].

Radial [Mar18, Mos78, Tri91, WX22, ACR19]. radially [JFM21]. Radiation [DDM21, GHO98, HHA86, PY00, PZ03]. Radiative [MG91]. Rado [GKK76]. Radon [MR11]. Random [AGJ03, BM72, BR73, BKG86, BN69, BW79, BFM17, CCY23, Coh91, FY01, Gez79, Gez80, HK74, KO84, KS06, Lan73, Ma82, MS08, MTVE02, RC05, Saf71b, Sch80, Wan72, Wan73a, WF17, FSKMP21, KHJK21, McG23]. range [Ryb22]. Rapid [KB98]. Rapidly [YK02b, Cai22a, Cai22b]. Rare [LK16]. Rate [GKK18, Hic85, KWW05a, Wan80, TNT21b, TNT21a]. Rates [HM86, Hel74, KM12, MB03, FT21]. Ratio [Tho69b]. Rational [Air79, Air81, BMS18, FH09, MW16, RCH17, Ven92, CGUGM20, LFW20, PRAV21, Tro21]. Ratios [Haj92, Seg21]. Ray [HCB72, KK91, KT99, RA10, RMA12]. Rayleigh [Ros10, Wal82, Wal01]. RBCs [DDMMSYD22]. Reacting [AMR91a, AMR91b, MR84b, MR92]. Reaction [Boa75, CW96, Edw00, Gal10a, HM86, HK77, JQS14, Kee76, KO04, KWW05b, KH73, KH81, LO99, MR80, MTW16, SH76, Sta76, STWW03, War96, GSYG21, PRAV21, TNT21b, TNT21a]. reaction-diffusion [TNT21b]. Reactions [ZGQ12]. Reactor [PSS75]. Reactors [LW91, WK91]. Real [CNW12, LLLZ17, SFO16, UD20, MMRL23]. Rearrangements [KL81]. Rebound [Gol04]. Rebours [PS21]. Receptors [LNWZ09]. Reciprocal [BM12a, KS02]. Reciprocals [RSCM05]. Recirculating [LPH09]. Recirculation [DG05]. Recognition [RKS83]. Reconstruction [Gzy00a, HN96, PB03]. Rectangle [Wan73b]. Rectangles [GUGM18]. Rectangular [AGJ03, Tro17]. Rectilinear [FK77, Pec80c]. Recurrence [FS96, GY95a, LLC10, DFS22]. Recursion [BLP88, FS86, GY15, KK91, Laq81, PEE04, Wan12]. Recursive [VS95b]. Reduced [CDLjZ18, GHJ12, How77]. Reducible [JO89a]. Reduction [KMM80b, Qu97, VGW15, WZ22, bXlZS23]. Reductions [CG16, GY15, Liu13, LRS04, LCB99, RMA12, RS14, Rog14a, Wan69b, Wan70a]. Refinable [JRS99]. Refinement [She99]. Reflection [GS15, Mos79, HS19, KA19]. Regime [KWW05a, BGMR21]. Region [BC76, CKLS96, LK10, Ran98]. Regions [WC95]. Registers [Abl01]. Regularity [Str97, BS22, LCG21, ZTYH20]. Regularized [EHS18, Lin14]. Reid [RA10, RMA12]. Related [Bła11, KR78, KLPS18, PT18, SP88b, Zho09, Bła23, DIW19, GKK21, WCYW19]. Relation [DD11, Gre72a, Nis79, DFS22]. Relations [ABCD17, FS96, GY95a, PS93, DER21, MC21a]. Relative [BLP01, Rei89]. Relativity [Gre72c, Gre72a, GK74]. Relaxation [LY11, MDG10, MM03]. Relaxing [Var97]. Release [DMPS96]. Released [MPB00]. Relevance [BSGC99, BSCS00]. Relevant [BHH^{+}11, Ken76]. Remarkable [Yeh95]. Remarks [Bar16, Gzy00b, Ric84]. Remember [Gro70]. remodeling [Jin23]. removal [GSYG21]. Renewal [Sta87]. Renormalization [MO01, OK09, OK10, OK15]. Renormalized [Che04]. Rent [Wan77]. Representable [JO89a, KO84]. Representation [BL87, Bła11, Bry75, CBB87, Ors69, SAR13, TW17, BDM22]. Representations [CS90, Di 94, FS96, HM86, KL00, Ors71a, SV87]. Representing [PMAD07, Ran96c]. Reproduce [Gro70]. reproduction [BL22]. Requirements [Ors71b]. Rescaling [Mos79]. Research [AAL02]. Reservation [MK09]. Reserved [MK09]. Residual [LW88]. Residues [Mil01]. Resolution [Ors71b]. Resolving [STWW03]. Resonance [Abl75, AO70, AB80, AB82, ACMR01, Bar77, BK95, CA09, CM03, CE73, DB90, GKT10, Gri77b, Gri87, Gus89, Kat81, Kev80, Kev82, Ma78, Ma84, Mas85, Mil76, Nou00, Rub77, Sri88, WY03, YM04, CCK^{+}21, FA22, RMMH22]. Resonances [AB00, Amu17, BH98b, Gus86, GS90, SA19, LLZ23]. Resonant [AFN73, BL78, BS88, DA05, FR08, Gri90, GY93, KL84, LKK80, Mah95, Mah96a, MRS88, MB95, Mil95, MTVE02, RP11, RTT02, Str11, TA07, Vas11, CC19, HSS19, HS19, RCMH21]. Resonantly [CR86, HMR86, MR84a, MG91]. Resonating [Peg88]. resonator [ADH^{+}21]. resonators [HW21]. resource [SSW22]. Resources [Wan80, SSVS20]. Respect [DK09, LL22]. response [CZ21]. Rest [BI81]. Restricted [ELY03]. Results [CS80a, ERK^{+}81, How86a, How86b, JPS00, KX18, Mah96b, SR99, Urb90b, Wri85]. Resurgence [Nem15]. Retarded [Tri88b]. Reverberating [Gro73]. Reversal [YF05]. Reverse [AFLM18, MHW20, YZ20]. reverse-space [MHW20]. reverse-time [YZ20]. Reversible [PPV08, You13]. Revisited [Fok09, FH98, GM93, Oli88, Rei81, Ryb22]. revival [BOPS21]. Revolution [RW69b, RW71, Wan69b, Wan70a]. Reynolds [Ben84, BMB02, CK04, HS82a, Saf71b, TWK00]. rheological [FCP^{+}21]. Riccati [Lou15]. Riemann [BFM17, CCM12, CCM19, CCF19, CF22, CEHS21, Fai97, Gre72c, Gre72a, KR23, LTY22, LW22, MV13, NMBD23, SXZ22, WW05, XZ13]. Riesz [BŠT23, Tél94, Tri88a]. Rigid [BT80, BE22, O’N69, Urb89, Urb90a, Urb90b]. Rigorous [CM21, CMP22, JZ10]. Ring [JO07]. Rings [KWW06, Saf70, Saf75, Sta75]. Rip [YC15]. ripples [FH23]. Ritz [FS69]. River [DGW02]. RNA [Wat79]. Robin [BB12, How82]. Robust [KSW14, Wan19, Wan21]. Robustness [CC17, KHS22]. Rod [Wri85, FDL21]. Rods [CLS86]. Roger [KP19a, KP19b]. Rogue [CCMH21, CC17, YY23, FLT20, GWCH23, RMMH22]. Roguing [GXLX16]. Role [Her91]. Roll [YY03]. Rolling [BT80]. Rook [GJW77, Car90]. room [BK23]. Root [EM71]. Roots [OAVS97]. Rossby [Ben79b, BM04, CG07, CG08, Cam04, DG05, KK95, MC02, NC15a, NC15b, War83]. Rota [Spe74]. Rotach [IL20]. Rotating [BM82, BB69b, Ben06, DdY11, Fri74, Gre69b, Gre90, Gri85, Gri90, Gub74, HS69, Isr72b, KR78, O’N69, RA10, Sch75, TA07, TG01, TU05, Tur05a, Tur05b, UG83, VB76, Cai22a, Cai22b, MWZ22, UG84]. Rotating-Disk [TU05, Tur05b]. Rotation [Fos70a, Fos70b, GT90, Wan85, YGJ18, MRE19, Ste19]. Rotation-Modified [GT90]. Rotational [FMN19, GHO98, LB95, Phi98, CW21]. Rotationally [RW69b, RW71, Wan70c]. Rotations [GP87]. Rough [RP83, DDM21]. Roughness [HS82a, PSG23]. Roughness-induced [PSG23]. routes [LCTZ22]. Routing [AH75, CK75]. Rule [PAS88]. Rules [DGW02]. Runaway [Fow86].

s [Can78]. Saalschütz [Con91]. Saddle [DH02, LP11]. Saddle-Center [DH02]. Saffman [Pom85]. Salesman [Cla78]. SAM [Ano05b]. Sample [BCKW85]. sampling [CHSV22]. Sasa [Swa12]. Sasa-Satsuma [Swa12]. Satisfy [ERK^{+}81]. Satsuma [Swa12]. Saturable [CM13]. Saturated [AMR91a, AMR91b, Gri82]. Scalar [CBB87, FCGB95, HM94a, JS01, Li02, MB09, Ors69, ACM18, GM21]. Scale [AY95, AHP02, AH75, Ben76, Che04, CK75, Hic83a, Hic83b, How90a, NAY17, OK10, Str97, Ulr75, KB98]. Scale-Invariant [AHP02]. Scales [New02, LWAO21]. Scaling [MC18, PB03]. scattered [Nic21]. scatterer [FA22]. Scattering [AKNS74, AL76, Abl78, AYF83, ADM82, AT07b, BB69b, BB69a, BT88, BRT89, BN94, Cau18, DPvdMV13, Dun13, FA83a, FA83b, FA89, GS16, Gzy00b, Kau80b, KM98, Kau06, KKS15, MW16, MPG17, Mos79, NA84b, NA84a, PZ03, PBT11, PV16, POvdMG18, Tro21, WK89, AE22, BLM21, GPFT22, KT22, MHW20, MZ23, OP19, SCFFLPA23, Vaf20, XB22]. Scatterings [ZHY09]. Scheduling [Bar73]. Scheme [AL76, KM18, Lin14, MW73, MM03, Nou00]. Schemes [Boy16, Ors71b, ASSV20, BJV22, KDM22]. Schlichting [BH89]. Schnackenberg [You13]. Schnakenberg [WW02, WW21]. Schrieffer [MWL23]. Schrödinger [ACR19, ACR23, AZ22, BG83, BB12, BP14, BLM21, CD04, CI23, CCD22, CM13, CZ20, CEHS19, DPvdMV13, DvdM23, DNP17, DFG19, DFKZ14, DFK16, FLT20, FR08, FA83b, FS86, Fro19, GLZ15, Gri95, GLL13, HL11, KHF^{+}20, Li02, LTY22, Lu23, Luo21, Ma79b, MHW20, MQ15, Mos78, Mos79, Oht09, OWY11, PY05, PBT11, POvdMG18, Rog14b, RS21, Sri91, STZZ14, Swa07, VM11, Woo84, Wri20, XB22, YA03, Yan12, Yan21, YZ20, ZC10b, Zha19, Zho09, MA81]. Schwarzian [EE19, KS02]. Scott [KWW05a, KWW06, LSG22, WW03]. Sea [DPHJ03]. Seas [DLW89]. seasonal [DDAAB20]. Second [BH91, BS88, BS89, Can78, Cla16, Cos93a, Cos93b, CS93, GLN14, Hab77, HJ96, Hoc72, Jos99, Mil01, Olv21, WL92, YT14, Zau72, MMRL23, YLW21]. Second-Order [WL92, YT14, Mil01, YLW21]. Secondary [Cho90, GP04, Kee76]. Section [DPHJ03, HF17, Wal01, SXZ22]. sections [ACM18, KT22]. Secular [Fow85b]. Sediment [EM09, MPS98, MPB00]. Segel [PS09, WX22]. segregation [LWZ23]. Segur [CD21a, Kok20]. SEIR [GKK18]. Selection [CHR91, DH96a, PR94]. selectively [FSKMP21]. Self [ACT87, ACH99, AM05, BH16, Gal10a, KC05, MB03, PP90, Pas90, Pri23, Wal95, WB98, Woo84, Yue81, BO20, LQYjZ22, XB22]. self-adjoint [XB22]. Self-contained [Pri23]. self-dual [LQYjZ22]. Self-Duality [ACH99]. Self-Focusing [Woo84, BO20]. Self-Gravitating [Yue81]. Self-Similar [Gal10a, KC05, MB03, PP90, Pas90, WB98]. Self-Similarity [AM05]. Self-Sustained [Hal12]. Self-Sustaining [Wal95]. Semi [AFN73, AO02, Cos93a, JO07, LW90, OC70, Wan75, BCCM23, LFW20]. semi- [LFW20]. Semi-Infinite [AO02, OC70, Wan75, BCCM23]. Semi-Resonant [AFN73]. Semi-Ring [JO07]. Semiclassical [BO20, MQ15, DFS22]. Semicontinuity [You13, Aou20]. Semidiscrete [TR97, CD21b]. semidiscretization [SYF22]. semigroup [FÖK22]. Semigroups [Li76]. semihyperbolic [BS22]. Semiinfinite [MML17]. Semilinear [Cos93b, Gal06, Gal08, Gal10a, How79, How82]. Semirational [RCH17]. Semisimple [Sat85]. Sense [Vil04]. sensing [ADH^{+}21]. Sensitivity [Rei70a]. Sensory [Gro69]. Separability [Bła11]. Separable [Kau80a, MRW07, RN15, Wan73b]. Separated [DB96]. Separation [Ack73, Gra73, SSK82]. Separatrix [BH94, ZHY09]. Sequences [EHV18, GJW77, JG79, JlPG79, JG80, Kfo70, KL75, Loe90, Rei77, Rei78, Sta87, VS93, VS95b]. Series [BP71, CFW87, KL81, Lia06, LT07, Mac94, RD04, SRD04, Sri92]. Service [BK78a, YK94a]. Set [ELY03, JK84]. Sets [Bel79, Che94, ERK^{+}81, LMY08, XXZ21]. Setting [BRSV92]. Settling [Saf73]. Several [Gez79, Gez80, Sch80, HMR86]. severity [GS23]. Shadowing [Li06]. Shallow [AB07a, ALB09, AB10, Ben92, BM12b, CK09, CP98, DLW89, EGP01, EGK05, GT04, GL17, JC02, LM16, Mar02, Mil95, MM01b, RW69a, Rei69, RA10, SV94, SV95, SF85, SS08, Wan69a, YK94b, BP20, CDF^{+}22, HP19, KDM22, Pri23, VMA21, dMVM19]. Shallow-Water [ALB09, EGP01, MM01b, Pri23, VMA21]. Shannon [YS98]. Shape [AGJ03, AB07b, FN05a, BK23]. Shaped [MPG17, Nis79, WS99]. Shared [ZK13]. Sharing [YK94a]. Sharp [Ack73, BGL97, FDL21, BS23, Seg21]. Shaw [Ric92, YK99]. Shear [BM99, BC09, Bau98, Bau99, BM75, Ben84, BZ89, BZ92, BM04, BM14, CK13, Cho90, DPHJ03, DG05, FLM95, Gri81, GS02, Hal01, HMM87, Hic83a, Hic83b, Hic85, KM70, Lan87a, LB81, MBM94, MC02, Mes96, OCB87, PS96, Phi98, PJ86, RS10b, SP88a, Saf74, SS85, TG06, TKC81, VSS98, War83, YK99, GMW21, LYZ20]. Shear-Strained [RS10b]. Sheared [VB83]. Shearing [Wan70c]. Sheet [MML17, OC70]. Sheets [NV17, Tan86]. Sheffer [Di 94, DL94]. Shell [MW73, NR77, RJM78, RW69a, Rei69, Wan69a]. Shells [Isr72b, JW69, NA88, RW69b, RW71, Wan69b, Wan70a, Wan70b, Wan70c]. Shelves [SA04]. Shift [Abl01, Hab88, Hab91, Wax80]. shifting [WPO22]. shoaling [Ste19]. Shock [Gaa82, GS15, HM93a, HK77, JTLW13, LO95, MR84b, MR92, WC95, CEHS19, HSS19, NS21]. Shock-Wave [JTLW13]. Shocks [AN96, AN98, CR86, EHS18, SA05, SS04, TW03]. Shooting [OW04]. Short [AY95, Ben77, DB03, FMO17, GPT^{+}05, Gro73, HSS17, LLL12, Ma78, Pec79, SFO16, SZWM16, TA07, TNM96, WM94, Wri06, CGP23, GPFT22, LG20, ML20, RMMH22, Ryb22, SZ23]. Short-Lived [GPT^{+}05]. short-pulse [CGP23, GPFT22, SZ23]. short-range [Ryb22]. Short-Scale [AY95]. Short-Wave [Ma78, Wri06, RMMH22]. Shortest [YK08]. Shulman [FCCTR22]. Shunted [Lu89]. sic [VA02]. Sidewalls [Hal12]. Sigma [GŠ06]. Sigmoid [CZ21]. Sign [QJW07, BR19]. Sign-Invariants [QJW07]. Signal [Ben81, Haj92]. Signaling [HM91, KSW14, Wan14, Wan21]. Signals [Bar71]. signature [BR19]. Significant [AAA08, Ben76]. Signless [KV06]. Similar [Gal10a, KC05, MB03, PP90, Pas90, WB98]. Similarity [AM05, BF85, FG12, LCB99, MS08, Red80, SA05, Str97, Tam93, VA03]. Simple [AZ13, Dun01, Dun04, FPSA89, GRMOH01, HS82b, Kau78, Ors71a, SP88a, Sri92, CF22, Pet22, RC80]. Simpler [DL94]. Simplified [Dun11, Nis79, ELV21]. Simulating [LK16]. Simulation [Gre69a, NLB10, Ors71a]. Simulations [MMPS13, FCP^{+}21]. Simultaneous [Cla78, FR01, Sun93]. Sine [ABT86, ABT87, BT88, EFM84, FM83a, Hwa14, Kau75, Li06, NS98, Swa07, CF22, bXlZS23, AFLM18]. Sine-Gordon [ABT86, ABT87, BT88, EFM84, FM83a, Hwa14, Kau75, Li06, NS98, Swa07, CF22, bXlZS23, AFLM18]. Sine-Gordon/Sinh-Gordon [AFLM18]. Single [CR86, Gal11, MR84a, Sha91]. Single-Mode [Sha91]. Singular [Abr77, BHN21, CWT^{+}15, CE73, Cro70, EG95, FG12, Gal10b, HHW94, How77, HP78, KT89, KT90, KPS80, Lan83b, LS88, LM91, LW91, LS93, LST06, LK10, MD99, MR80, Ros76a, Stu92, WF17, YS21]. Singular-Perturbation [KT89]. Singularities [Cos97, FH09, JK94, MVBK98, Mos14, Sob04, TW03, CDF^{+}22, Rat19]. Singularity [CCM12, SS95, SR97a, SR97b, Woo84, WXZ18, PRAV21]. Singularly [Gri95, How77, How79, How82, KKM87, KO04, LW95, Li02, Li05, LLC10, LS04, OK09, OK10, OW04, Tov00, War92]. Sinh [FM83a]. Sinh-Gordon [FM83a, AFLM18]. SIRE [LCTZ22, WCTZ22]. SIS [BL22, LP23, ZN17]. Site [CMKP16, WA83, Wan85]. situ [WCYW19]. SIV [GLH22]. Sivashinsky [Hol23, JP22]. Six [EG95, JK94]. Sixth [MV13]. Size [BOGK04, BFM17, Mar18, FT19]. Skan [SKB08]. skew [IW23]. skew-Hermitian [IW23]. skewness [ACM18]. Skinning [Edw97]. Slab [Mos14]. slender [MO21]. slices [SO20]. Slightly [KM70]. Slip [Lac82]. Slits [Ric92]. Slope [YGJ18, KA19]. Sloping [EM09, MPS98, Swa18, Tro17]. Sloshing [ACMR01, Yan95a]. Slow [BH98b, DH02, LN94, MK09]. Slowly [Abl71, GA79, GA82, Gri86, Gri87, GM93, HK77, Kat85, Kee82, Kev80, KL88, Wal82]. Small [Ben76, Ber92, CM07, DH96a, Gri12, HHW94, Har11, Hun69, Kne00, Mac97, Moo78, NS91, New02, PP86, STC09, Sha91, SY06, TW97, TR97, Urb90a, Urb90b, Wes78, YK94b, YY03, FH23, HW21, JTW22]. Small-Wavelength [DH96a]. Smoluchowski [Sin21]. Smooth [APS98, BB69a, BH06, ACM18, GMNP22]. Snow [DARB11]. Soap [Gri12]. Sobolev [IW23, MMRL23]. Sobolev-orthogonal [IW23]. Soccer [YBSM08a, YBSM08b]. social [BGMR21]. Soft [Gzy00b]. Soil [TBKiM18]. Sojourn [YK94a]. Solar [Ken76]. Solid [Gre72d, Lac82, YG01, Vaf20, Zhe20]. solid-solid [Zhe20]. Solitary [AM09, BG07, BC09, BG82, BH97, CG07, CKL16, Che84, CLS86, CI20, GG84, Gea85, GM93, GZ94, Gri95, GPS96, GPT98, GHO98, GP04, Gri16, GL17, HM08, HAY12, JC02, JTW22, KS14, KA81, Lam98, Luo21, Mar02, PW97, Sta15, SDS12, Swa12, TB02, TEW91, War83, WV92, YL07, YL08, Yan12, ZHY09, GMW21, JW20, JL19, LMS19, Ste19]. Solitary-Wave [Che84, CLS86]. Soliton [BGL15, BMOT09a, BMOT09b, CGP23, CK09, CFiMO18, CM08, DA05, FMO17, FPSA89, Hir09, JTLW13, NY16, PAS88, STZZ14, Swa07, WTL^{+}10, BO20, CCD22, CF23, LFW20, Ma79a, MHW20, Yan21, YZ20, Zha19]. Solitons [APT06, CM13, DLW89, GOS11, GMM05, KA81, Lam98, Mil79, NAY13, NY14, OWY11, PF89, PY00, PY05, POvdMG18, SKP^{+}05, SY03, Yan97, Yan01, YA03, YMBC04, Yan14b, ISFH20, KHF^{+}20, LTY22, MZ23, RCMH21, Ryb22, Slu19, ZTYH20, GMM05, AK82]. Solution [AL77, ABT86, ABT87, AB07b, BB87, BCH95, BGL15, BK95, BS16, BS89, BS01, DLK16, FN05b, FMO17, JW06, Kau78, KMP14, Lea16, Lie77, LW88, Ma78, MW73, MS16, Mil79, MCB97, Rei79, Rod03, SKB08, Sri91, SLN16, Tri88b, VS94, Woo95, Wor86, Yan02, CKS23, HSS19, KT19, KS20, Luo21, MWZ22, NR22]. Solutions [AS77, AK82, ACT87, Abr77, ASSV20, Air79, Air81, AFY15, AKZ16, Atk78, BRT89, BL87, BH97, BMS18, BS88, BS93, CK09, CJM92, CFiMO18, CDLjZ18, Che84, CK14, CR06, CFY17, Cla16, CM09, CNW12, DR77, Dec00, DN11, DNP17, DM96, DFS97, Dun01, Dun04, Dun11, EY91, Fai97, FH10, FCL12, FG12, FCGB95, FMM99, Gal06, Gal10b, GWH11, GXW13, GGS95, GP87, GM93, GH08, GZ10, GLL13, Hag81, Her76, Hir09, How90b, Hu98, IR86, JTLW13, JK94, JK01, Jos15, KPS80, Kee85, KH73, KH81, KLW11, LO99, Lan87a, Lan80, LL80, Lan83b, Lev69, LM74, Lia06, LT07, LR69, LY11, LLL12, LLLZ17, LCB99, Ma79a, Ma79b, MR93, MB03, Nau92, Nec12, OW03, ÖSV01, PS10, PSS75, PMAD07, PG95, Ran94a, Ran94b, Ran94c, Ran95a, Ran95b, Ran96b]. Solutions [Ran96a, RS08, RY10, RS11, RN15, RCH17, Red80, Rei72, Rei74a, Rei74b, Ric84, Ros78, RS80, SG90, SV92, SJV96, SER^{+}03, SJH05, SRE05, SAM17, SA05, SR99, SZWM16, SS04, SH76, Sta76, Swa07, SDS12, Swa12, Tam89a, Tam89b, Tam93, VA03, VK06, VP10, VS85, VS88a, VS96, VPE09, Wan77, WTL^{+}10, War96, WZ07, WZ08, WBZ21, XQZ15, bXlZS23, Yan14a, Yan15, YT14, ZZ13, ZZ08, ZL17, Zho18, ZZZ10, ALZ21, BGMR21, BS22, BS23, BCCM23, CZ20, CCK^{+}21, CF23, CD21b, CGUGM20, CI20, DWZ21, DZ22b, Dun20, Fro19, JFM21, KR22, Kok20, LFW20, LS23, MHW20, MQ22, NdJVA22, PP19, TV22, Wri20, YZ20, Zha19, Zhe20, dSF20, Hic83b]. solvability [Dem23]. Solvable [Cal84, Cal16, Che84, Len09, MTW16]. Solving [AH75, Che11, CK75, Kau75, Kau80a, LLC10, Ulr75, Vas11, YK02b].

Some [Bar73, BK78a, BRT89, Ben73, Ber70, CFS92, CY94, DKK^{+}14, DDM13, ERK^{+}81, ELY06, Gro70, Gzy00a, Hab75, Har75, How86a, How86b, Iwa82, Kle75, KWW05b, KH81, KLW11, LL80, LM74, Ma79a, MW73, MR93, PS11, PP90, Pas90, Peg88, Red80, Ric84, Ros78, Ros76b, SER^{+}03, War96, YY08, YB96]. Sommerfeld [BSGC99, BSCS00, Rei72, Rei74a, Rei74b, Rei79]. Sonine [RV18]. Sound [Cro70, HHA86, SV87, HS19, NMBD23]. Source [AT07b, BO12, JQ13, LNVW05, Qu97, Str11]. Sources [Nou00]. Space [AFLM18, BM75, CR86, Gro70, HMR86, KWW05b, MR84a, Red80, Ros10, SV87, Vla08, Yan14a, ACZ23, MHW20]. Space-Modulated [Yan14a]. Space-Time [AFLM18, Gro70, ACZ23]. Spaces [BD70, GR70, KL00, Sch01, LCG21, MN22, Wu21]. Spacing [BFM17]. Sparse [Fat74, SO20]. Spatial [DB90, KB98, SKP^{+}05, SR12, SS06, SY06, SW00, WCTZ22, WLR22, YW92, LWZ23, SSW22]. Spatially [CHR91, DB94, HM94a, MM03, Wan21, WS99, NMBD23]. Spatiotemporal [GJLW22, Gro69, PSAZ22, SSW22]. Special [AY17a, AY17b, FC08, CD18b, MWZ22]. Specializations [Kun80]. Species [AT07a, CLW22, DDMMSYD22, FSKMP21]. Spectra [Hun69, SZWM16, BZZ21, YS21]. Spectral [AN13, Ben75, BAM98, BD17, Boy16, CV09, CJ18, DKK^{+}14, FMX22, Gal08, GWH11, GT80, HSS17, Har11, Her82, Kli03, KR03, KM90, KD92, Mil01, Mos14, Ors71a, Ors72, Swa04, VM18, YBSM08b, ZK13, LT23, SO20, UD20, GY15]. Spectrum [BG07, Ben75, BP14, HN96, LLS04, Ma82, Saf71b, WXZ18, SZ23, UD20]. Speed [CA09, GPS96, Saf73, ST84, SV87, Wei85, NMBD23]. Spencer [Haj92]. Sperner [Gri77a]. Sphere [Fos70a, Fos70b, Ifi10, KMM80b, Mon74, RF16, AP23, KK22b]. Spheres [Far82]. Spherical [GMW98, Isr72b, PT18, RJM78, SJH05, VP10, FdST22, MQ22]. Spheroidal [BOGK04, HG82, Mar18, Mil75, RJRTF17, DDN21]. Spike [KWW05a, Nec12, STWW03, WW02]. Spike-Type [Nec12]. spiky [WX22]. Spin [Fri74, Gre90, SR12]. Spin-Down [Fri74]. Spinup [BI81]. Spiral [FL73, LL79]. Spline [Lin14]. Splitting [KWW05b]. Spontaneous [JA16, MR84b]. Sports [PS11]. Spot [WW21]. Spotty [WW03]. Spreading [AKS^{+}08, Pec80c]. Spring [CW70, Rei70b, FH23]. Spurious [Lan83b, OW03]. Square [Bax71, POvdMG18]. Squared [CZjZ21, FS86, YL07]. Squared-Operator [YL07]. Squares [Bön95]. Stabilities [LS75, LS77]. Stability [AN13, Ash10, BM12b, BD17, CS85, CMM^{+}09, CJ18, Cro98, DN11, DKK^{+}14, DR90, EFM84, Gaj96, GT04, GMNP22, Gre89, GV03, GK95, GYZ21, HH94, HSS17, HS82a, HS84, Har11, Her82, Her91, HY86, How73, How86a, How86b, HAY12, KS96, Ken74, KWW05a, LP08, Lan80, LNW04, MWZ22, Mar71, MTW16, MSW07, NPT08, O’N69, PP16, PS00, PW97, QC96, RS84, Ros10, SS06, ST17, Sta15, Swa04, TMPVB19, Wal95, WHS17, WW02, Wil72, XB22, Yan95b, ZS86a, vKD72, ABM^{+}22, Aou20, DS23, KS20, LZW23, MMO23, MC21b, PS20a, UD20, WPO22, WX22]. Stabilize [Ben06]. Stabilizing [AB03, JC08]. Stable [Hab91, Hic83a, Hic83b, YRZ20]. Stäckel [BM12a, BM20]. Staggered [MW73]. Stagnation [Edw00, KC05]. Standing [FCL12, LO99]. Starting [EHV14]. State [CM13, GP14, ST17, BS23]. States [Boa75, CW96, MK14, BŠT23, CI23, LWC23, PS20a, WX22]. Static [MW73]. Static-Geometric [MW73]. Statics [KS02, Rei73a]. Stationary [CS12, CNW12, EHS18, GGS95, SY15, Bła23, SZ23, ZSA22]. Statistic [FZ90]. Statistical [RKS83, YKdM^{+}07]. Steadily [Tan86]. Steady [Atk78, BM82, Bau99, Boa75, Boo89, CS80a, HS82b, Ran94a, Ran95a, Ran96a, Ran96c, SS85, ST17, Xu90, LWC23, WX22, MR84b]. Steady-State [ST17]. Steady-Wave [MR84b]. Steepest [CFW87, Lin97]. Steepest-Descent [Lin97]. Stefan [MEHS22, NR22]. Steffensen [EHV14]. Stellar [Hun70, Mar71]. Stem [MR84b]. Step [Ulr75, XF23b]. step-like [XF23b]. Stewartson [FKMS22, ISFH20, KS18, KM90, RCH17, Zho18]. Stieltjes [Lap87, LF02]. Stirling [BH91, Can78, GLN14, KK91, KV06, Mén96, Tem93]. Stochastic [AKS^{+}08, CS80c, PS11, Vil04, VM11, – Dor22, GLH22, KHS22, XYZY22, ZYZ22]. stochastically [BL22]. Stokes [AZ22, BCCM23, BW02, CCY23, DLK16, DH19, DS23, GG76, KLSD23, LPH09, LCB99, McL97, ÖSV01, Plo02, Ran94b, Ran94c, Ran95a, Ran95b, Ran96b, Ran96a, Tov00, VS96, vKD72]. Stopping [YF05]. Storage [YF05]. Strahler [DGW02]. Strain [Ber72, LW88, RW69b, RW71, Rei73a, Rei73b, Rei75]. Strained [RS10b]. Straits [DLW87, DLW89, DPHJ03]. Strange [CJN04]. Strategies [ZN17, GKK21]. Stratified [BG82, CMM^{+}09, CS12, DPHJ03, DS99, Fri74, Gre70, Gri81, GS02, HS69, JMT08, KM70, LB81, Mal01b, PG96, Saf72, SA04, Swa12, TA07, TEW91, VS96, VB76, KT22, Rat19]. Streets [RSCM05]. Stress [RW69b, RW71, Rei73a, Rei75, RS10b, Wan69a, Wax84]. Stretched [RS84]. Stretching [Lia06]. Strictly [KL88]. string [LZW23]. Strings [MS08]. Strip [LS04, Wan75, LBP21]. Stripes [KWW06, MTW16]. Strips [LW90]. Strong [AS89, Ark92, BH06, GG84, Gea85, SA05, WK91, YK94b, DWZ21, DZ22b, Pri23]. Strongly [Bar16, BM14, GOS11, Hab91, JC02, JC08]. Structural [CC17]. Structure [BH14b, LL79, NW07, PRNA01, PZ03, RS84, RS10b, SS95, SR97a, SR97b, SS81, SDS12, Swa18, Wit96, AZ22, CEHS19, FCP^{+}21, ZL21]. Structured [WLS17, GS23, GYZ21, LCTZ22]. Structures [AKZ16, Bry77, BL78, CL91, HK77, Sta78, Yan96, YF05, LLZ23, Nic21]. Struve [Dun22]. Stuart [Ifi10]. Stud [BMOT09a, UG84]. STUDIES [Ano06]. Study [Bea70, EY91, FSY78, GJ81, GT90, KS18, KT99, LHC04, RN15, RF13, RS07, SRJ99, YLP95, FKMS22, GLS23, KS20]. Sturm [BHN21, BN98, HN96, Lan87b]. Subcritical [HMM87, MC02, TYZZ21]. Subharmonic [BH98b, JP22]. submerged [DDM21]. Subset [Kle75]. Subsonic [RSCM05]. Subspaces [JQ13]. subwavelength [ADH^{+}21]. Success [FSKMP21]. Successions [RT79]. Successive [BCKW85]. Suction [DH96b, KC05, Tur05a]. Sudden [DMPS96]. Suddenly [MPB00]. Sufficient [BM99, Mar71]. Suggested [HS82a]. Suitable [GK74]. Sum [SJB91]. Summation [BS99, BS00, Mac94, Mil01]. Sums [BKG86, KL81, Pec80a]. Sundman [BG83]. Super [GXW13]. Superalgebras [Sta85, Van87]. Superantichains [JK84]. Supercritical [DFKZ14, GM21, YRZ20]. Superharmonic [KLSD23, ZS86b]. Superposition [RS10a]. Supersonic [CM08, GB96, Mes96]. Superspace [FH10]. Supersymmetric [FH10, Fan11, XL15, TW17]. Support [GZ13, Yan14b]. Supported [JRS99]. Surface [AB03, BGT07, BG07, BC09, Bar16, BRT89, Ben92, Ben06, BH06, Ben76, BT80, BC86, Cro00, CS15, Edw00, FSY78, FM11, GY95b, Gre77, GM81, HCB72, Lac82, LHC04, LB95, MML17, MPD05, OCB87, Plo02, Pom85, RF16, Ros10, SV94, SV95, Saf71a, TNM96, Wan14, WM94, ZS86b, CHS21, JTW22, LBP21]. Surfaces [CT86, DRS02, GŠ06, Kon96, KL00, MVBK98, NV17, Rab89, Sch01, FdST22]. Surfactant [Ake12]. Surjections [Wag94]. Surroundings [AM05, MP01]. Susceptibility [BM72]. Suspended [Yu10]. Suspensions [MPB00, Saf73]. Sustainable [Wan80]. Sustained [BK95, Hal12]. Sustaining [Wal95]. swarming [KK22b]. swell [ZC21]. Swept [AB03]. Swift [CNW12]. swirl [EWW22]. Swirling [How73]. Symbol [Cos00b, Cos06]. Symbolic [TW17, WTL^{+}10]. Symmetric [BRSV92, BD17, CLY95, CZ15, DKK^{+}14, DFKZ14, Dou72, FC08, GP14, Loe90, MK14, Mos14, NY14, NY16, RW69b, RW71, RV18, SHC06, Sen89, Wan70c, Yan14b, CCMH21, JFM21, Yan21]. Symmetries [AH95, BF85, EG95, Fok87, HLQ09, Ifi10, JQ13, LPH03, LKM14, Pen17, PS21, QJW07, RH07, SVS08, ZC10a, ZC10b, Olv21, ZL21]. Symmetrization [NKUW02]. Symmetry [DB90, DR90, FN05b, Hu98, JQS14, JA16, Kon14a, Kon14b, Liu13, LRS04, MR02, NW07, Qu97, SJH05, Sch98, SW81, Str97, VP10, CZjZ21, EE19, MRE19, PP19, PS20b, ZL21]. symmetry-integrable [EE19]. symplectic [BC23, XJN21]. Synchronizing [Kfo70]. Synthesis [Ben75, Lin02]. Synthetical [XQZ15]. System [BHH^{+}11, CC17, CFiMO18, CD18a, DGW02, DNP17, EGK05, EHS18, FN05b, GZ13, JG79, JlPG79, JG80, JQS14, Ken74, Kev80, LP08, LY11, LRS04, MRS88, Mat89, Mor89, MR80, O’N69, PV16, RP11, RS14, Rog14b, Rog14a, SY03, Tam88, Tum06, TR97, Vil90, WWZ10, WNH12, Woo95, ZS86a, ACR19, ACR23, AE22, CCMH21, CLW23, CR21, DvdM23, Dya20, Esc23, FN22, GSYG21, GR23, GJLW22, GWCH23, JGM21, KHF^{+}20, LZZ23, OP19, Tro21, WPO22, XF23a]. Systematic [BMD22, BDM22, BM22, LP11]. Systems [AT07a, AAA08, AB80, AB00, AKZ16, Bar73, BN94, BL87, BH98a, Bła11, BM12a, BK95, Che11, CP98, DIT16, ERK^{+}81, GY15, GHO98, HM93a, HM77, How79, How83, HY00, KQ10, Kee82, Kev82, KS18, KWW05b, KO13, Lan87b, LS75, LS77, LN94, LK16, Lou15, Lun03, MRW07, Mas85, MK14, MM01a, MM03, MTW16, NA84a, New91, NRGO01, PPV08, QK14, RA10, RC05, SG90, SR97a, SR97b, SS03a, STWW03, VGW15, VS85, Vas11, WC95, XL15, YB96, Yan98, Yan15, YT14, YK02b, Zak09, ZZN12, Zho09, ASSV20, BF16a, BF16b, BW19, CCM19, CZ20, CGUGM20, Dem23, FKMS22, IW23, Lou19, LWC23, LC21, MP19, MP22, MRE19, Olv21, Sul20, YF21]. Szász [Sun93]. Szekeres [Gre72b].

Table [GS90]. Tableaux [CLY95]. Tactics [WWZ10]. Tails [AY95, Mil93]. Tallest [FN05b]. Tandem [KT99]. Tangential [Vla08]. Tank [ACMR01, BS88]. Tanks [Yu10]. Tanner [XHW22]. Target [KH81, WLS17]. Tau [CF23]. Tau-function [CF23]. taxis [LWC23, ST17]. Taylor [LT02, Pom85, SW81]. Technique [LB69]. Techniques [OK09]. Telpaking [AH75, CK75]. Temperature [Fow93, GG76, PS00, QC96, Esc23, MWL23]. Temperature-and-Pressure-Dependent [Fow93]. Temperature-Dependent [PS00, QC96]. Temporal [Tam88, YW92]. Temporally [KH81]. Tennis [NK05]. Tension [Ben06, FSY78, Gre77, Pom85, BGMR21, JTW22, LBP21]. tension-inhibitive [BGMR21]. Tensor [Gre72c, Gre72a]. Term [Gro73, STC09, ZZ08, SYF22]. Terminal [GHO98]. Terms [LT99, Mac97, MV13, DWZ21]. Tests [KC92]. th [EM71]. Their [BSGC99, HLQ09, HAY12, Kon96, KL00, PS96, RSCM05, RW96, SRD04, Sun93, Tan86, Yan97, BSCS00, DFS22, HG81, LS23, MC21a, PS20a, WF12]. Theorem [BD70, Bak76, Bel77, CS80c, Con91, Gre72b, GKK76, Gri77a, Kun86, Olv21, PRNA01, Pen17, Sen89, Tri92, YS98]. Theorems [AT07b, Haj92, JG79, AP23]. Theoretic [Ven92, Wil72]. Theories [Abl71, DKK^{+}14, LW90]. Theory [AO03, Aky82, And71, Ben81, Ben77, BC89, BAM98, Bła11, BC76, BRSV92, BFM17, Bri15, Bry79b, BS88, CJM92, Che69, CR70, DF09, Dou72, DRS74, EG95, FS69, Gal08, Gao95, GY15, GR70, GJW77, Gre69b, Gzy00b, HM91, Her82, Her91, Hol80, How83, JW69, JO07, KMM80a, KS14, Kat85, KX18, Lan80, LR69, LL79, MR84b, Moo00, NR77, PF89, PSS75, Plo02, Rei74b, RW69a, Rei69, Rei73b, RA10, RS10b, Rus73, SRD04, Sta71a, Sta71b, Stu92, Swa04, Swa12, Tam88, TR02, VB76, Wan69b, Wan69a, Wan70a, Wan70b, Wan77, Xu90, YMBC04, ZGQ12, AP23, BHIS21, BR19, CM21, CH18, HSS19, LW22, MO21, Pri23, NK05]. There [Bön95]. Thermal [BS77, Fow86, LMS19]. Thermally [DS99, MPD05]. Thermo [Wei84]. Thermo-Viscoelasticity [Wei84]. Thermoelasticity [ÖM94]. Thermohaline [Mor86]. thermomicropolar [PSG23]. Thermonuclear [HM86]. Thermoviscous [Fow85a]. Thick [GMW98]. Thickness [Moo78]. Thin [BB87, Dha94, Gal10b, GMW98, Hun69, Hun70, Lac82, LB69, MM90, Rei70a, SW74, LBP21, MO21, PSG23]. Thin-Walled [SW74]. Third [Bag08, Cos00a, DB90, Gal11, KT89, KT90, MCB97, PEE04, SR97a, SR97b, SBA05, Sob05, Var97, YA03]. Third-Harmonic [DB90]. Third-Order [Bag08, Cos00a, Gal11, KT89, KT90, PEE04, SR97a, SR97b, SBA05, Var97, YA03]. Thirring [CF23, Vil91]. Thomas [GP14]. Three [BZ89, BE22, BŠT23, Bla70, CLV20, CS85, CMM^{+}09, CLZ91, CS15, DHC00, DFS97, FR01, Gri77a, GS02, HS84, Kau76, Kau80a, LP08, MS16, MK96, MW13, MB09, OB94, Ran95b, STC09, Sha91, STZZ14, TG06, Vár13, dMVM19, WNH12, ZB92, FIM23, JN21]. Three-Coupled [STZZ14]. Three-Dimensional [CS85, CLZ91, CS15, DHC00, Kau80a, MK96, OB94, STC09, Sha91, TG06, BE22, BŠT23, JN21]. Three-Dimensionality [HS84]. Three-Layer [CMM^{+}09, GS02, WNH12, dMVM19]. Three-Phase [DFS97]. Three-Wave [Kau76, Kau80a, MS16]. thresholds [BZZ21, TYZZ21]. Tight [ACZ12]. Tight-Binding [ACZ12]. Tiling [DDNP90, DDNP91]. tilings [Cha21]. Time [AW97, AFLM18, AM05, Bar73, BM75, BW79, BD17, CR86, CW96, DFKZ14, FA83b, FP12, GT80, Isr70, KC82, Lea16, NC15a, NC15b, PPV08, PP08, RS11, Ric92, SER^{+}03, SRE05, SVS08, SAM17, SV87, SB87, SLN16, VP10, WK89, YK94a, Yan14b, Yan15, YF05, ACZ23, DWZ21, FvdW21, Fro19, GBSS21, GH08, Gro70, LWAO21, LCG21, MMO23, NMBD23, PRAV21, RS21, XF23b, YL08, Yan21, YZ20, ZL21]. Time- [SLN16]. Time-Dependent [CW96, FA83b, FP12, Lea16, NC15a, NC15b, PP08, Ric92, SB87]. time-fractional [ZL21]. Time-Periodic [Yan15, FvdW21, Fro19]. Time-Reversible [PPV08]. time-varying [GBSS21]. Times [Tri88b]. Timing [Vla17]. Title [Ano73, Ano76a, Ano76b, Ano78, Ano80b, Ano81b, Ano82b, Ano83b, Ano84b, Ano85b, Ano86b, Ano87b, Ano88b, Ano89b, Ano90b, Ano91b, Ano92b, Ano94b]. Title-Author [Ano76b, Ano80b]. TOC [Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f]. Toda [CM08, HK04, MFIGH22, WK89, YF21]. Toeplitz [Str86]. Tollmien [BSSS00, BH89]. Top [Bar16, JZ86, MB09, O’N69]. Topography [APS98, AB10, BH06, CM03, GY93, KS96, Mil98, MM99b, BP20, CDF^{+}22, FMN19, Ste19]. tor [Mor86]. Torsion [Wan93, SK23]. Total [GG76, Kok20, ZN17]. Tournament [Iwa82]. Trachomatis [WE17]. Tracking [HM93a]. Traffic [MK09, WHS17]. Trailing [Ack73]. Trains [LB81, MB84, Yan98, JP22]. traits [DDAAB20]. Transcendent [Hab77, AK20]. Transcendentally [Mac97]. Transcendents [VA03, XZ13]. Transcritical [GZC09]. Transfer [Abr71, LB69, LKK80, TW97]. Transform [AKNS74, Abl78, AYF83, AT90, BS99, BS00, Cau18, CD21b, DPvdMV13, FA83a, FA89, GS16, KM98, KM90, KD92, LY13, MR11, PBT11, PV16, POvdMG18, Str97, Tél94, Tri76, Tri91, Tri99, BLM21, BM20, FvdW21, GPFT22, MZ23, MWL23, OP19]. Transform-Fourier [AKNS74]. Transformable [VJ11]. Transformation [BT88, Fan11, GK74, GPT98, KMM80b, RS98, Ryb22, Ryb23]. Transformations [BCH95, BRT89, DRS02, GLL13, KQ10, LY03, MV13, MCB97, QK14, RS10a, SR98, Sch01, Vag12, XL15, YY18, Zho18, ALZ21, LT23, ML20, WBZ21]. Transforms [Fre87, Len16, LF02, MHW20]. Transient [HG01, LL80, STC09, SA04, XK93]. Transients [Wal95]. Transition [DHC00, Hab77, Lan87a, Lin02, Tho69a, TK69]. transitions [Zhe20]. Translating [Tan86]. Translation [War92]. Transmission [ACZ23, DDAAB20]. Transport [Bar73, EM09, How90a, MM93, MPS98, ACM18, MWL23]. Transporting [Hal06]. Transversal [Bry75, Kun78]. Transverse [BM72, Har11, Joh10, PP16, BBN22, HCC21, KS20]. transversely [MC21b]. trapeziums [SO20]. Trapped [Isr72b]. Trapping [BN98, Edw97]. Traveling [BGMR21, BBN22, Cla78, DN11, Dya22, Hab91, Har11, MEHS22, PP16, SG90, Tam89a, Tam89b, CW21, HSS19, MMO23, PTDS19, TNT21b, TNT21a, TMPVB19]. Traveling-Wave [SG90]. Travelling [TA94, ZZ08, GMNP22]. treated [– Dor22]. Treatment [MC18, GS23]. Tree [JO89a]. Tree-Representable [JO89a]. Trees [GKM74, HLS03, Kal73, Kle72, KS06]. Trend [Ark92]. Triad [Gri87]. Triads [BL78, Mah95, Mah96a, RTT02]. Triangle [KF10]. Triangles [HK74]. Triangular [MRW07]. tridiagonal [IW23]. Trigonometric [Mac94]. Trigonometrical [SRD04]. Triple [Tum06]. Triple-Deck [Tum06]. Triples [Pec80b]. Triply [Mor89]. Tritronquée} [JK01]. Tropical [NC15a, NC15b]. Trunk [MK09]. Trunks [MK09]. Tsingou [FH23]. tube [FMX22]. Tubes [SW74]. Tubular [LW91, Li06]. Tumor [Gre72d]. Turbulence [Lin02, Mas72, MS76, New02, Ors69, Wal95, Zak09]. Turbulent [Cha71, How90a, MPB00, Saf74]. Turing [ST17]. Turning [Abr77, BH14a, Dun01, Dun11, Lak72, Zau72, Dun20]. twisted [PSAZ22]. Twisting [SW74]. Two [AZ13, Abr71, AN13, AKZ16, BM82, BGT07, BG07, BC09, Bar16, Bry77, Che04, CBB87, CMM^{+}09, Cos93a, DF09, DHC00, DS71b, Dre71, DIT16, Dun11, Dun13, EFM84, Fan11, GOS11, GMM05, GZ10, GZ13, HM93b, HM94b, How90b, KMM80a, KMM80b, KMM82, KM12, KWW06, LP08, Laq81, LAHM17, LT02, LS04, LCB99, MDG10, Mil76, MM99b, MM01b, MPS98, NR77, NAY13, OW03, PS11, Pas90, Ran94a, Ran94b, Ran94c, Ran95a, Ran96b, Ran96a, Red80, Saf71b, ST84, Str97, STWW03, UG83, UG84, Vár09, VM18, Vla17, VSS98, Wei85, XQZ15, YMBC04, ZB92, BP20, CLW22, CW21, ISFH20, KHF^{+}20, LCTZ22, LBP21, LC21, MZ23, MKP^{+}20, NR22, PP19, PTDS19, Pri23, RMMH22, TMPVB19, XF23a, Yan21, YF21]. Two-Boson [Fan11]. Two-Component [GZ10, XQZ15, MZ23, RMMH22]. Two-Dimensional [AZ13, Abr71, AN13, CBB87, DHC00, GMM05, How90b, KWW06, Laq81, LAHM17, LCB99, NR77, NAY13, Ran94a, Ran94b, Ran94c, Ran95a, Ran96b, Ran96a, Saf71b, VM18, YMBC04, VSS98, BP20, ISFH20, KHF^{+}20, PP19, XF23a, Yan21, YF21]. Two-Layer [AKZ16, BGT07, BG07, BC09, Bar16, DIT16, GOS11, MM01b, MM99b, MPS98, CW21, PTDS19, Pri23]. Two-Person [PS11]. Two-Phase [DS71b, Dre71, EFM84, MDG10, Pas90, UG83, UG84, LBP21, MKP^{+}20]. Two-Point [LT02, OW03]. Two-Scale [Che04]. Two-Space-Dimensional [Red80]. two-species [CLW22]. Two-Timing [Vla17]. Two-Variable [KMM80a, KMM80b, KMM82]. Two-Wave [HM93b, HM94b]. Type [AW97, CD04, Che11, CE73, Fai97, FIN94, GY95b, Hab88, Hu96, Joh10, JT87, KQ10, KMP14, Loe90, Nec12, Rei77, Rei81, Sob05, Sob06, Sta87, TZHL09, VS93, ACS18, BŠT23, BMD22, BDM22, BM22, CCF19, CM21, CMP22, FCCTR22, GRMOH01, LG20, Lou19, MC21a, PCH23, XJN21, ZS22, KT19]. Types [CS80b, Liu13]. Tzitzéica [RSCM05].

Uhlenbeck [GYZ21]. Ulam [FH23, Mon74]. Ultradiscrete [Hir09]. Ultrahyperbolic [Tél94, Tri88b, Tri88a]. Umbral [Mor80, She99]. Unbounded [BG82, Fow90a, TA07]. Uncertainty [Bar71, RC05]. Uncomplemented [Kle72]. Unconditional [PS00, YK94a]. Undercooling [PP86]. underdetermined [Olv21]. Underlying [LRS04]. Undular [EGP01, EGK05, YGJ18]. Uneven [Ben92]. Unified [AZ13, Cau18, CD21b, Rus73, FvdW21]. Uniform [BPS73, Boy16, Bry79b, Dha94, Dun22, HW20, Jin23, KD17, KHJK21, Kro77, LY11, LT99, PW12, RJM78, Rei74a, Rei74b, SS81, Seg21, WW05, ZL17]. Uniformly [TR02, Woo95]. Uniformly-Valid [Woo95]. Unigraphic [KL75]. Unimodal [Van87]. Unimodality [Sta85]. Uniquely [JO89a, Wan76b]. Uniqueness [Gri78, Lie77, ABM^{+}22, WPO22]. Unit [Gra73]. Unitarily [ADM82]. Unitary [CZ10, MC18, WXZ18, LC21]. Universal [GM75, RS12, SAR13, CEHS19]. Universally [YL07]. Universally-Convergent [YL07]. Universes [NV17]. Unsolvability [Pat70]. Unstable [Swa07, TG01, Wit96, Wri06]. Unsteady [Lia06]. Upper [FR08, Hic85, KSV01, Sta75, You13, Aou20]. Upper-Semicontinuity [You13]. Urban [Wan77]. Use [BS99, BS00, OK09, MRE19]. Useful [DF09]. Using [DS71a, FN05b, KK91, PS21, RC05, RD04, CD21b, DL23, VMA21].

V [And71, XZ13]. vaccination [GLH22]. vaccine [GLH22]. Valid [TR02, Woo95]. Validity [KMP14, Woo95]. Valley [LMY08]. Valley-Avoiding [LMY08]. Value [AB07a, CRW16, CE73, FR01, GLZ15, Hal01, Hen88, HM15, HS69, How77, JS01, KKM87, Kau80a, Lak72, LM91, LW95, LLC10, MQ15, OW03, OW04, PMAD07, PS18, RS80, SBA05, SKB08, Wan73b, War92, XF16, YK02a, YK02b, Zau72, ZS22]. valued [Cha21, DER21, LWAO21]. Values [WF17, Fro19]. Vandermonde [VS95a]. Vandermonde-like [VS95a]. vanishing [DvdM23, FT21]. Variable [CR86, DLW89, FM11, Gre77, GPS96, GP04, GA11, Gri15, HL11, HCB72, KMM80a, KMM80b, KMM82, KS96, KL84, MR84a, MPD05, PMAD07, RS11, SV92, Sri91, TZ14, VK06, Vag12, VS88a, WTL^{+}10, BP20, CDF^{+}22, PTDS19, SXZ22, Ste19]. Variable-Coefficient [GP04, Gri15, TZ14, WTL^{+}10]. Variable-Inflow [MPD05]. Variables [BKG86, Cos93a, Dec00, Gez79, Gez80, HMR86, Sch80, DS23]. Variants [Cal16, DKK^{+}14]. Variation [OK15, HCC21]. Variational [BH14b, KKM87, RMA12, Sri88, Swa18, BE22, BJV22, MRE19]. Variations [RANC09]. Varies [SV87]. Varying [Abl71, DLW87, GA79, GA82, Gri86, Gri87, GM93, HK77, HZ10, Kat85, Kee82, Kev80, KL88, SW17, WQW13, YK02b, ZN17, GBSS21, Wan21]. Vastly [KM12]. Vector [APT06, AFY15, CS80c, GR23, GR70, HS92, JS01, Li02, PY00, PY05, SKP^{+}05, Yan97, KHF^{+}20, LG20]. vehicle [BK23]. Velocity [DH96b, MM93, MDG10, Saf70, CLW23]. Version [Atk78, YK08]. Versus [DKZ16]. Vertex [Bar84]. Vertical [BM04, Hal12, KKS15, Tam93, LS77]. Very [FG12, Gal10b, GM81, Pom85, Esc23, Seg21]. Veselov [GWCH23]. VI [CZ10, NRGO01]. via [Abl01, AP23, BW19, DRS02, DM96, DFS22, EM09, Fre87, GFGUMM22, KL00, Str97, XZ13, ZZ13]. Viability [AP23]. Vibration [AKK^{+}09, YM04, YM07, YKdM^{+}07, SK23]. Vibrations [Vla08]. VII [Dou72]. violent [GM21]. viral [HMBYZ22, WLR22]. Virasoro [LRS04]. Virus [GXLX16]. Viscid [LO99]. Viscoelastic [LY11, FMX22]. Viscoelasticity [Wei84]. Viscosity [BW02, FSY78, Fow93, PS00, QC96, RS11, VK06, Zhe20]. Viscous [BM82, BT80, BDGL02, CK14, Gre77, GM81, Isr70, Kro77, Lac82, LO95, Le 04, OB94, Phi98, Ran94a, Ran96c, Ran98, Saf70, SB87, SY15, SW00, TR02, Vla08, MO21, PV23]. Vladimirov [Tri97]. Volterra [AT07a, CLW22, PCH23, WPO22]. Volume [Ano75, Ano76a, Ano76b, DMPS96, MPB00, Sin21]. Volumes [Ano77, Ano78, Ano80a, Ano81a, Ano81b, Ano82a, Ano82b, Ano83a, Ano83b, Ano84a, Ano84b, Ano85a, Ano85b, Ano86a, Ano86b, Ano87a, Ano87b, Ano88a, Ano88b, Ano89a, Ano89b, Ano90a, Ano90b, Ano91a, Ano91b, Ano92a, Ano92b, Ano94a, Ano94b, Ano95, Ano96, Ano98, Ano03, Ano04, Ano05a, Ano06, Che94]. Vortex [AKZ16, CM07, CLJ97, Cro98, Hal01, Hal12, KO13, Kro77, Le 04, Moo78, NC15a, NC15b, OC70, PS96, RSCM05, Saf70, Saf72, Saf75, SS77, ST84, SAR13, Tan85, Tan86, Wei85, Aik19, Cai22a, Cai22b]. Vortical [DdY11, MB95, Mil95, SY15]. Vortices [BH89, Cla09, DB96, Gaj96, Ifi10, OB94, RS84, SS81, SW81, Tan86, Wad91, AZ22, DH96a]. Vorticity [BD01, CM07, CS12, Dha94, Her91, Isr72a, PJ86, Saf71b, VSS98, Cai22a, Cai22b, DH19, EWW22]. Vries [AK82, JP22, NdJVA22, AS77, CKS23, GM93, GP04, GPT^{+}05, Gri15, Gri19, Hab88, HSS19, JZ10, JPS00, JL19, JN21, Mil79, Mil93, Ryb23, TZ14, Wad78, XJN21, XF23b, Yan01]. Vries-type [XJN21]. Vries/Kuramoto [JP22].

Waals [BPS73, BS22, GS15]. Waiting [KC82]. Waiting-Time [KC82]. Wake [STC09]. Walk [BW79]. Wall [FLM95, HS82a, Her91]. Walled [SW74]. Wang [Ano05b]. Ward [Vil90]. warning [GLH22]. Water [AM08, Ake12, AN13, Aky82, AB07a, ALB09, AB10, Bau98, Bau99, BMOT09a, BMOT09b, Bry79a, CK09, CS80b, CS80a, CS85, CP98, CS15, DLW87, DHC00, DLK16, EGP01, EGK05, FSY78, GT04, GA11, GL17, GS90, HJ15, JC02, KA88, Kli03, KR03, LM16, LHC04, Ma82, Mar02, MB95, Mil95, MK96, Mil98, MM01b, RA10, SV94, SV95, SF85, SS08, SS85, TBKiM18, VBA97, WVBM14, YK94b, Yu10, ZS86b, BP20, CDF^{+}22, CI20, Dya20, EWW22, GMW21, HP19, KDM22, MC21b, Pri23, VMA21, dMVM19]. Water-Table [GS90]. Watertable [OH95]. Wave [Abl71, Abl72, ABT86, ABT87, AS89, AAA08, Ake12, AB80, AY95, AB00, Atk78, BSGC99, BSCS00, BOGK04, BT88, BN69, BR69, Ben83, BH98a, Bry77, BL78, Bry79a, Cal84, Cam04, Che84, CK14, CLS86, CP98, CHS21, DN11, DS99, Dun13, DDM13, DLK16, EGK05, FH10, FCL12, FR17, GB96, Gea85, Gra73, Gri77b, GM93, GPS96, GPT98, GHO98, GS02, GPT^{+}05, GA11, Gri16, GZ13, Hag81, Hal01, Hal12, HM91, HM93a, HM93b, HM94b, HG82, JTLW13, JC08, Kau76, Kau80a, KL84, KKS15, Kon14a, Kon14b, KH73, LO99, LP08, Lan73, LM16, LL79, Lin14, LB81, LY11, LN13, Ma78, Ma79b, Mah95, Mah96a, Mah96b, MR84b, MM99a, MS16, MBM94, McL97, Mes96, Mil75, MTVE02, Miu74, MPG17, MB84, New02, NAY17, OCB87, PS96, Red80, RJRTF17, RTT02]. Wave [RANC09, SG90, SJV96, SY03, SDS12, Swa12, TA07, Tam89a, Tam89b, Var97, Vas11, WNH12, Wri06, YB96, YL07, Yan12, Yan15, ZZ08, ZS86a, BGMR21, Cai22a, Cai22b, CBJCN20, CI20, CEHS19, DWZ21, DZ22b, HS19, JP22, KR22, LG20, LW22, LZW23, McG23, RS08, RMMH22, Slu19, Ste19, SCFFLPA23, VMA21, XF23a, Yan98, YY23, ZS22, dSF20]. Wave-Mean [DS99, PS96]. wave-short [LG20]. wave-type [LG20]. wave/mean [Cai22b]. Wavefunctions [Mar18]. waveguide [PSAZ22]. Waveguides [BAM98]. Wavelength [DH96a]. Wavelets [JRS99, She99]. Wavenumber [DH96a]. Wavepacket [AM09]. Wavepackets [HHA86]. Waves [AB70, Abl75, AS80, AAA08, APS98, AM08, AM09, AN13, AB82, Aky82, AHP02, Alm90, AMR91a, AMR91b, AN96, AN98, BMM87, BB69b, BB69a, BD01, BGT07, BG07, BC09, BH89, Bau98, Bau99, Bea70, Ben92, BH06, BB69c, Ben73, Ben74, BM75, Ben76, Ben77, BK78b, Ben79a, Ben79b, BG82, BC86, BZ89, BZ92, BM04, BMOT09a, BMOT09b, BM14, Bri90, BBC^{+}05, BS01, CG07, CG08, CKL16, CR86, CK09, CC17, CS80b, CS80a, CS85, CA09, CM03, CS12, CP12, CS15, DLW87, DB03, DG05, DB90, DB94, DHC00, DKZ16, ET87, FL73, FSY78, FM11, GG84, Gea85, GKT10, Gre70, Gri81, Gri85, GY93, GZ94, Gri95, GP04, Gri15, Gri16, GL17, HjS85, Hab88, Hab91, HSS17, Har11, Has76, HM93b, HM94b, HCB72, HMM87, HS82b]. Waves [HM08, HK77, HMR86, Hun92, HJ15, HAY12, JC02, Joh10, KA88, KS14, Ken74, Kli03, KK95, KA81, Kup06, Lam98, LHC04, LKK80, LB95, Lus80, Ma82, Ma84, MR84a, MR92, Mar02, Mas72, MC02, MM90, MB95, Mil95, MK96, Mil98, MW13, MG91, New91, NC15a, NC15b, Nis79, PP16, PR94, Peg88, PW97, PG96, Phi98, Plo02, PJ86, RP83, Ros10, SP88a, SRJ99, SF85, SS08, SV87, SS17, SS85, Sta15, Sta76, SDS12, Swa12, TA94, TB02, TG06, Tro17, TNM96, TKC81, TCK82, TEW91, VBA97, VS85, VB76, VM18, VSS98, Wal01, WVBM14, War83, WV92, WNH12, WM94, Wri85, Wri06, Yan95a, Yan95b, YL07, YL08, YKH00, YY03, Yue81, ZB92, ZHY09, ZS86b, ABM^{+}22, BF16a, BF16b, BBN22, CCMH21, CD21a, CC19, CW21, CI20]. waves [DDM21, DH19, Dya20, Dya22, DS23, EWW22, FLT20, FMN19, GMW21, GMNP22, HCC21, HSS19, HHZ20, HS19, JW20, JTW22, JL19, KA19, KP19a, KP19b, KLSD23, LMS19, Luo21, MMO23, MQ22, MEHS22, MC21b, NS21, PTDS19, Pet22, RMMH22, TNT21b, TNT21a, TMPVB19, WPO22]. Wavetrain [Kee85]. Wavetrains [BLP01, CJ18, EFM84, FM83a, KR03, Rat19]. Weak [CK14, GG84, SS04, War83, YKH00, ZHY09, DWZ21, HS19, AN96]. Weakly [AM03, CR86, GKT10, Gri85, Gri87, Gri95, Gro02, GV03, Hab88, Hal06, HM93b, HM94b, HMR86, Kev80, KB98, KMP14, LL80, Ma82, MR84a, MM99a, Mos14, Peg88, PCZP00, Rub77, SR99, SS17, Sri91, SS03b, TKC81, Yan96, CLV20]. Weber [Dun22]. wedges [RS21]. Weierstrass [KL00]. Weight [CJK16, DK09, Gre76, Pec79, XZ11, ZXZ15]. Weights [GST18, WW05, LDLW20]. Well [CR21, DZ22b, Lu23, Nou00, PRAV21]. Well-Balanced [Nou00]. Well-posedness [CR21, DZ22b, Lu23, PRAV21]. Western [Mal94]. Wetting [BS16, GM81]. Which [AO03, Rab89, LR69]. Whitecapping [DN16]. Whitham [ACR19, ACR23, BHIS21, Bri15, BR19, CJ18, CM21, CMP22, HJ15, JZ10, JW20, JTW22, KLPS18, LW22, RP11]. Whole [YM04, YM07, YKdM^{+}07]. Whole-Body [YM04, YM07, YKdM^{+}07]. Whose [ADM82, How77, Kat88, KK90]. Width [DLW87]. Wiener [Ben75, GM21, MML17, RC05, YY04]. Wightman [Tri97]. Wind [Aky82]. Windowed [SCFFLPA23]. Wing [AB03, SS77]. Wings [ST84, Wei85]. Winning [NK05]. Within [Ors71a, WLS17, Cai22a, Cai22b, DDMMSYD22, WLR22]. Within-Host [WLS17, WLR22]. without [CC19]. WKB [Nis07]. Wolbachia [ELV21]. Wong [Ano05b]. woodpile [DL23]. Woodruff [TNM96]. Wound [SW07]. Wright [FÖK22]. Wronskian [BDS20, GUGM18]. WTC [Kic99].

X [BRSV92]. xfct [MR11]. XI [Cos00a]. XIII [Bag08]. XXXIV [XZ11]. XY [Bar83, ABGML71]. XY-Model [Bar83]. XYh [BF85].

Yajima [CFiMO18]. Yang [ACT87, APT06, BH16, CGP23, LQYjZ22].

Zakharov [Joh10, PS10, SB00, Vas11]. Zero [BW02, CS12, RF13, SZWM16]. Zero-Viscosity [BW02]. zeroing [GBSS21]. Zeros [BFM17, Hei71, Ioa89, RW96, ZZ07, HW20]. Zeroth [BP20]. Zeroth-order [BP20]. Zigzag [KWW06]. Zonal [MM99a, MC02]. Zoomeron [DRS02].

References

Ablowitz:2008:DPW

[AAA08] M. J. Ablowitz, S. A. Ablowitz, and N. Antar. Damping of periodic waves in physically significant wave systems. Studies in Applied Mathematics, 121(3):313–335, October 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00419.x.

Ablowitz:2002:RCD

[AAL02] M. J. Ablowitz, T. R. Akylas, and C. C. Lin. Research contributions of David j. Benney. Studies in Applied Mathematics, 108(1):1–6, January 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
01419.

Ablowitz:1970:EMP

[AB70] M. J. Ablowitz and D. J. Benney. The evolution of multi-phase modes for nonlinear dispersive waves. Studies in Applied Mathematics, 49(3):225–238, September 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1970493225.

Akylas:1980:DRN

[AB80] T. R. Akylas and D. J. Benney. Direct resonance in nonlinear wave systems. Studies in Applied Mathematics, 63(3):209–226, December 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1980633209.

Akylas:1982:EWN

[AB82] T. R. Akylas and D. J. Benney. The evolution of waves near direct-resonance conditions. Studies in Applied Mathematics, 67(2):107–123, October 1982. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1982672107.

Amundsen:2000:RDW

[AB00] D. E. Amundsen and D. J. Benney. Resonances in dispersive wave systems. Studies in Applied Mathematics, 105(3):277–300, October 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00152.

Allen:2003:FPS

[AB03] Leanne Allen and Thomas J. Bridges. Flow past a swept wing with a compliant surface: Stabilizing the attachment-line boundary layer. Studies in Applied Mathematics, 110(4):333– 349, May 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9590.00241.

Antuono:2007:BVP

[AB07a] Matteo Antuono and Maurizio Brocchini. The boundary value problem for the nonlinear shallow water equations. Studies in Applied Mathematics, 119(1):73–93, July 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1365-2966.2007.00378.x.

Atanackovic:2007:MPS

[AB07b] Teodor M. Atanacković and Branislav S. Bačlić. More on the problem and the solution of the optimal shape of a javelin. Studies in Applied Mathematics, 119(2):173–189, August 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2006.00382.x.

Antuono:2010:ANS

[AB10] Matteo Antuono and Maurizio Brocchini. Analysis of the nonlinear shallow water equations over nonplanar topography. Studies in Applied Mathematics, 124(1):85–103, January 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00464.x.

Ankiewicz:2017:CLI

[ABCD17] A. Ankiewicz, A. P. Bassom, P. A. Clarkson, and E. Dowie. Conservation laws and integral relations for the Boussinesq equation. Studies in Applied Mathematics, 139(1):104–128, July 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12174.

Abdullayev:1997:JAF

[Abd97] Alisher S. Abdullayev. Justification of asymptotic formulas for the fourth Painlevé equation. Studies in Applied Mathematics, 99(3):255–283, October 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00061.

Abraham:1971:DLP

[ABGML71] D. B. Abraham, E. Barouch, G. Gallavotti, and A. Martin-Löf. Dynamics of a local perturbation in the XY model. I

— approach to equilibrium. Studies in Applied Mathematics, 50(2):121–131, June 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1971502121.

Abraham:1972:DLP

[ABGML72] D. B. Abraham, E. Barouch, G. Gallavotti, and A. Martin-Löf. Dynamics of a local perturbation in the X–Y model II — excitations. Studies in Applied Mathematics, 51(2):211–218, June 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1972512211.

Ablowitz:1971:ASV

[Abl71] M. J. Ablowitz. Applications of slowly varying nonlinear dispersive wave theories. Studies in Applied Mathematics, 50(4):329–344, December 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1971504329.

Ablowitz:1972:AMO

[Abl72] M. J. Ablowitz. Approximate methods for obtaining multiphase modes in nonlinear dispersive wave problems. Studies in Applied Mathematics, 51(1):17–55, March 1972. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197251117.

Ablowitz:1975:NRN

[Abl75] M. J. Ablowitz. A note on resonance and nonlinear dispersive waves. Studies in Applied Mathematics, 54(1):61–70, March 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197554161.

Ablowitz:1978:LIS

[Abl78] Mark J. Ablowitz. Lectures on the inverse scattering transform. Studies in Applied Mathematics, 58(1):17–94, January 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197858117.

Ablowitz:2001:NFS

[Abl01] M. J. Ablowitz. Nonlinear feedback shift registers via nonlinear evolution. Studies in Applied Mathematics, 107(2):127–136, August 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9590.00181.

Amaral:2022:EUS

[ABM^{+}22] Sabrina Amaral, Handan Borluk, Gulcin M. Muslu, Fábio Natali, and Goksu Oruc. On the existence, uniqueness, and stability of periodic waves for the fractional Benjamin–Bona– Mahony equation. Studies in Applied Mathematics, 148(1):62– 98, January 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12428.

Abraham:1971:TMT

[Abr71] D. B. Abraham. On the transfer matrix for the twodimensional Ising model. Studies in Applied Mathematics, 50(1):71–88, March 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm197150171.

Abraham:1972:CFI

[Abr72] D. B. Abraham. On correlation functions of the Ising and X–Y models. Studies in Applied Mathematics, 51(2):179–209, June 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1972512179.

Abrahamson:1977:PES

[Abr77] Leif R. Abrahamson. A Priori estimates for solutions of singular perturbations with a turning point. Studies in Applied Mathematics, 56(1):51–69, January 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197756151.

Ablowitz:1986:SGW

[ABT86] Mark J. Ablowitz, Richard Beals, and Keti Tenenblat. On the solution of the generalized wave and generalized sine-Gordon equations. Studies in Applied Mathematics, 74(3):177–203, June 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1986743177.

Ablowitz:1987:ACS

[ABT87] Mark J. Ablowitz, Richard Beals, and Keti Tenenblat. Addenda and corrigenda to “On the solution of the generalized wave and generalized sine-Gordon equations”. Studies in Applied Mathematics, 77(1):101, August 1987. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1987771101.

Ablowitz:1999:GCE

[ACH99] M. J. Ablowitz, S. Chakravarty, and R. Halburd. The generalized chazy equation from the self-duality equations. Studies in Applied Mathematics, 103(1):75–88, July 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00121.

Ackerberg:1973:APF

[Ack73] R. C. Ackerberg. Axisymmetric potential flow separation at a sharp trailing edge. Studies in Applied Mathematics, 52(4):377–387, December 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1973524377.

Aminian:2018:MDS

[ACM18] Manuchehr Aminian, Roberto Camassa, and Richard M. McLaughlin. Mass distribution and skewness for passive scalar transport in pipes with polygonal and smooth cross sections. Studies in Applied Mathematics, 141(3):399–417, October 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12230.

Amundsen:2001:ENS

[ACMR01] David E. Amundsen, Edward A. Cox, Michael P. Mortell, and Sandra Reck. Evolution of nonlinear sloshing in a tank near half the fundamental resonance. Studies in Applied Mathematics, 107(1):103–125, July 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
1071180.

Ablowitz:2019:WSR

[ACR19] Mark J. Ablowitz, Justin T. Cole, and Igor Rumanov. On the Whitham system for the radial nonlinear Schrödinger equation. Studies in Applied Mathematics, 142(3):269–313, April 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12254.

Ablowitz:2023:WSD

[ACR23] Mark J. Ablowitz, Justin T. Cole, and Igor Rumanov. On the Whitham system for the (2 + 1)-dimensional nonlinear Schrödinger equation. Studies in Applied Mathematics, 150(2):380–419, February 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12543.

Anco:2018:DCP

[ACS18] Stephen C. Anco, Xiangke Chang, and Jacek Szmigielski. The dynamics of conservative peakons in a family of U(1)-invariant integrable equations of NLS–Hirota type. Studies in Applied Mathematics, 141(4):680–713, November 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12241.

Ablowitz:1987:SME

[ACT87] Mark J. Ablowitz, David G. Costa, and Keti Tenenblat. Solutions of multidimensional extensions of the anti-self-dual Yang–Mills equation. Studies in Applied Mathematics, 77(1):37–46, August 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm198777137.

Ablowitz:2012:TBA

[ACZ12] Mark J. Ablowitz, Christopher W. Curtis, and Yi Zhu. On tight-binding approximations in optical lattices. Studies in Applied Mathematics, 129(4):362–388, November 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00558.x.

Ammari:2023:TPS

[ACZ23] Habib Ammari, Jinghao Cao, and Xinmeng Zeng. Transmission properties of space-time modulated metamaterials. Studies in Applied Mathematics, 150(2):558–581, February 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12549.

Ammari:2021:HOE

[ADH^{+}21] Habib Ammari, Bryn Davies, Erik Orvehed Hiltunen, Hyundae Lee, and Sanghyeon Yu. High-order exceptional points and enhanced sensing in subwavelength resonator arrays. Studies in Applied Mathematics, 146(2):440–462, February 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12349.

Abraham:1982:EEL

[ADM82] P. B. Abraham, B. DeFacio, and H. E. Mose. An explicit example of a local and a nonlocal potential whose Hamiltonians are unitarily equivalent and whose scattering operators are identical. Studies in Applied Mathematics, 66(1):45–51, February 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198266145.

Aktosun:2022:DIS

[AE22] T. Aktosun and R. Ercan. Direct and inverse scattering problems for the first-order discrete system associated with the derivative NLS system. Studies in Applied Mathematics, 148(1):270–339, January 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12441.

Ablowitz:2018:RST

[AFLM18] Mark J. Ablowitz, Bao-Feng Feng, Xu-Dan Luo, and Ziad H. Musslimani. Reverse space-time nonlocal Sine-Gordon/Sinh-Gordon equations with nonzero boundary conditions. Studies in Applied Mathematics, 141(3):267–307, October 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12222.

Ablowitz:1973:SRI

[AFN73] M. J. Ablowitz, B. A. Funk, and A. C. Newell. Semiresonant interactions and frequency dividers. Studies in Applied Mathematics, 52(1):51–74, March 1973. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197352151.

An:2015:CVS

[AFY15] Hongli An, Engui Fan, and Manwai Yuen. The Cartesian vector solutions for the N–dimensional compressible Euler equations. Studies in Applied Mathematics, 134(1):101–119, January 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12056.

Andrews:2003:DRM

[AGJ03] G. E. Andrews, I. P. Goulden, and D. M. Jackson. Determinants of random matrices and Jack polynomials of rectangular shape. Studies in Applied Mathematics, 110(4):377–390, May 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00243.

Atherton:1975:HMS

[AH75] R. W. Atherton and G. M. Homsy. Heuristic methods for solving large scale network routing problems: The telpaking problem. Studies in Applied Mathematics, 54(1):31–60, March 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197554131.

Arrigo:1995:NSN

[AH95] D. J. Arrigo and J. M. Hill. Nonclassical symmetries for nonlinear diffusion and absorption. Studies in Applied Mathematics, 94(1):21–39, January 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm199594121.

Ablowitz:2009:AAP

[AHNZ09] Mark J. Ablowitz, Theodoros P. Horikis, Sean D. Nixon, and Yi Zhu. Asymptotic analysis of pulse dynamics in mode-locked lasers. Studies in Applied Mathematics, 122(4):411–425, May 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00441.x.

Ali:2002:HES

[AHP02] G. Alì, J. K. Hunter, and D. F. Parker. Hamiltonian equations for scale-invariant waves. Studies in Applied Mathematics, 108(3):305–321, April 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
01416.

Aiki:2019:ELP

[Aik19] Masashi Aiki. On the existence of leapfrogging pair of circular vortex filaments. Studies in Applied Mathematics, 143(3):213–243, October 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12273.

Airault:1979:RSP

[Air79] H. Airault. Rational solutions of Painlevé equations. Studies in Applied Mathematics, 61(1):31–53, July 1979. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197961131.

Airault:1981:ACR

[Air81] H. Airault. Addition and correction to “Rational Solutions of Painlevé Equations”. Studies in Applied Mathematics, 64(2): 183, April 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1981642183.

Ablowitz:1982:NAS

[AK82] Mark J. Ablowitz and Yuji Kodama. Note on Asymptotic Solutions of the Korteweg–de Vries Equation with Solitons. Studies in Applied Mathematics, 66(2):159–170, April 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1982662159. See [AS77].

Andreev:2020:CFA

[AK20] Fedor V. Andreev and Alexander V. Kitaev. Connection formulae for asymptotics of the fifth Painlevé transcendent on the imaginary axis: I. Studies in Applied Mathematics, 145(3):397–482, October 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12323.

Akers:2012:SIW

[Ake12] B. F. Akers. Surfactant influence on water wave packets. Studies in Applied Mathematics, 129(1):91–102, July 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00547.x.

Ammari:2009:VAD

[AKK^{+}09] Habib Ammari, Hyeonbae Kang, Eunjoo Kim, Hyundae Lee, and Kaouthar Louati. Vibration analysis for detecting internal corrosion. Studies in Applied Mathematics, 122(1):85–104, January 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2008.00424.x.

Ablowitz:1974:IST

[AKNS74] Mark J. Ablowitz, David J. Kaup, Alan C. Newell, and Harvey Segur. The inverse scattering transform-Fourier analysis for nonlinear problems. Studies in Applied Mathematics, 53(4):249–315, December 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1974534249.

Arkhipov:2008:NMP

[AKS^{+}08] Boris Arkhipov, Vladimir Koterov, Viacheslav Solbakov, Dmitry Shapochkin, and Yulia Yurezanskaya. Numerical modeling of pollutant dispersion and oil spreading by the stochastic discrete particles method. Studies in Applied Mathematics, 120(1):87–104, January 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2007.00394.x.

Akylas:1982:NTG

[Aky82] T. R. Akylas. A nonlinear theory for the generation of water waves by wind. Studies in Applied Mathematics, 67(1):1–24, August 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19826711.

An:2016:MCE

[AKZ16] Hongli An, Man Kam Kwong, and Haixing Zhu. On multicomponent Ermakov systems in a two-layer fluid: Integrable Hamiltonian structures and exact vortex solutions. Studies in Applied Mathematics, 136(2):139–162, February 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12097.

Ablowitz:1976:NDS

[AL76] M. J. Ablowitz and J. F. Ladik. A nonlinear difference scheme and inverse scattering. Studies in Applied Mathematics, 55(3):213–229, September 1976. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1976553213.

Ablowitz:1977:SCN

[AL77] M. J. Ablowitz and J. F. Ladik. On the solution of a class of nonlinear partial difference equations. Studies in Applied Mathematics, 57(1):1–12, July 1977. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19775711.

Antuono:2009:DNS

[ALB09] M. Antuono, V. Liapidevskii, and M. Brocchini. Dispersive nonlinear shallow-water equations. Studies in Applied Mathematics, 122(1):1–28, January 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2008.00422.x.

Almgren:1990:MHF

[Alm90] Robert F. Almgren. Modulated high-frequency waves. Studies in Applied Mathematics, 83(2):159–181, August 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990832159.

An:2021:DTS

[ALZ21] Ling An, Chuanzhong Li, and Lixiang Zhang. Darboux transformations and solutions of nonlocal Hirota and Maxwell– Bloch equations. Studies in Applied Mathematics, 147(1):60– 83, July 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12378.

Antar:2003:WNE

[AM03] N. Antar and T. B. Moodie. Weakly nonhydrostatic effects in compositionally-driven gravity flows. Studies in Applied Mathematics, 111(2):239–267, August 2003. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.t01-1-00233.

Antar:2005:SSE

[AM05] N. Antar and T. B. Moodie. Self-similarity and the evaluation of long-time nonhydrostatic effects in compositionally driven gravity flows in deep surroundings. Studies in Applied Mathematics, 114(3):211–239, April 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2005.01545.x.

Akers:2008:MEG

[AM08] Benjamin Akers and Paul A. Milewski. Model equations for gravity-capillary waves in deep water. Studies in Applied Mathematics, 121(1):49–69, July 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2008.00409.x.

Akers:2009:MEW

[AM09] Benjamin Akers and Paul A. Milewski. A model equation for wavepacket solitary waves arising from capillary-gravity flows. Studies in Applied Mathematics, 122(3):249–274, April 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00432.x.

Ablowitz:2017:INN

[AM17] Mark J. Ablowitz and Ziad H. Musslimani. Integrable nonlocal nonlinear equations. Studies in Applied Mathematics, 139(1): 7–59, July 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12153.

Almgren:1991:AARa

[AMR91a] Robert F. Almgren, Andrew Majda, and Rodolfo R. Rosales. Asymptotic analysis of reacting materials with saturated explosion. I. Low-frequency waves. Studies in Applied Mathematics, 84(4):275–313, May 1991. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991844275.

Almgren:1991:AARb

[AMR91b] Robert F. Almgren, Andrew Majda, and Rodolfo R. Rosales. Asymptotic analysis of reacting materials with saturated explosion. II. High-frequency waves. Studies in Applied Mathematics, 84(4):315–360, May 1991. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991844315.

Ariznabarreta:2018:CBL

[AMT18] Gerardo Ariznabarreta, Manuel Mañas, and Alfredo Toledano. CMV biorthogonal Laurent polynomials: Perturbations and Christoffel formulas. Studies in Applied Mathematics, 140(3):333–400, April 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12202.

Amundsen:2017:RBM

[Amu17] David E. Amundsen. Resonances in bounded media. Studies in Applied Mathematics, 139(2):248–264, August 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12181.

Axel:1996:ISD

[AN96] Ralph M. Axel and Paul K. Newton. The interaction of shocks with dispersive waves I. Weak coupling limit. Studies in Applied Mathematics, 96(2):201–246, February 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996962201.

Axel:1998:ISD

[AN98] Ralph M. Axel and Paul K. Newton. The interaction of shocks with dispersive waves II. Incompressible-integrable limit. Studies in Applied Mathematics, 100(4):311–363, May 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00080.

Akers:2013:SSD

[AN13] B. Akers and D. P. Nicholls. Spectral stability of deep two-dimensional gravity-capillary water waves. Studies in Applied Mathematics, 130(2):81–107, February 2013. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00574.x.

Andrews:1971:FCT

[And71] George E. Andrews. On the foundations of combinatorial theory v, Eulerian differential operators. Studies in Applied Mathematics, 50(4):345–375, December 1971. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1971504345.

Anonymous:1972:I

[Ano72] Anonymous. Index. Studies in Applied Mathematics, 51 (4):417–418, December 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1972514417.

Anonymous:1973:ATI

[Ano73] Anonymous. Author-title index. Studies in Applied Mathematics, 52(4):389–390, December 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1973524389.

Anonymous:1974:I

[Ano74] Anonymous. Index. Studies in Applied Mathematics, 53 (4):337–339, December 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1974534337.

Anonymous:1975:IV

[Ano75]  Anonymous.  Index to volume 54.  Studies in Ap-

plied Mathematics, 54(4):359–361, December 1975. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1975544359.

Anonymous:1976:ATI

[Ano76a] Anonymous. Author-title index to volume 55. Studies in Applied Mathematics, 55(4):365, December 1976. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1976554365.

Anonymous:1976:TAI

[Ano76b] Anonymous. Title-author index to volume 55. Studies in Applied Mathematics, 55(4):367, December 1976. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1976554367.

Anonymous:1977:IV

[Ano77] Anonymous. Index to volumes 56, 57. Studies in Applied Mathematics, 57(3):271–273, December 1977. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1977573271.

Anonymous:1978:ATI

[Ano78] Anonymous. Author-title index to volumes 58, 59. Studies in Applied Mathematics, 59(3):283–284, December 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1978593283.

Anonymous:1980:AIV

[Ano80a] Anonymous. Author index to volumes 60–64. Studies in Applied Mathematics, 63(3):279–281, December 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980633279.

Anonymous:1980:TAI

[Ano80b] Anonymous. Title-author index. Studies in Applied Mathematics, 63(3):282–283, December 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1980633282.

Anonymous:1981:AIV

[Ano81a] Anonymous. Author index to volumes 64 and 65. Studies in Applied Mathematics, 65(3):277–278, December 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1981653277.

Anonymous:1981:TIV

[Ano81b] Anonymous. Title index to volumes 64 and 65. Studies in Applied Mathematics, 65(3):279, December 1981. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1981653279.

Anonymous:1982:AIV

[Ano82a] Anonymous. Author index to volumes 66 and 67. Studies in Applied Mathematics, 67(3):259–260, December 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1982673259.

Anonymous:1982:TIV

[Ano82b] Anonymous. Title index to volumes 66 and 67. Studies in Applied Mathematics, 67(3):261, December 1982. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1982673261.

Anonymous:1983:AIV

[Ano83a] Anonymous. Author index to volumes 68 and 69. Studies in Applied Mathematics, 69(3):265–266, December 1983. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1983693265.

Anonymous:1983:TIV

[Ano83b] Anonymous. Title index to volumes 68 and 69. Studies in Applied Mathematics, 69(3):267, December 1983. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1983693267.

Anonymous:1984:AIV

[Ano84a] Anonymous. Author index to volumes 70 and 71. Studies in Applied Mathematics, 71(3):269–270, December 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1984713269.

Anonymous:1984:TIV

[Ano84b] Anonymous. Title index to volumes 70 and 71. Studies in Applied Mathematics, 71(3):271, December 1984. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1984713271.

Anonymous:1985:AIV

[Ano85a] Anonymous. Author index to volumes 72 and 73. Studies in Applied Mathematics, 73(3):271–272, December 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1985733271.

Anonymous:1985:TIV

[Ano85b] Anonymous. Title index to volumes 72 and 73. Studies in Applied Mathematics, 73(3):273, December 1985. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1985733273.

Anonymous:1986:AIV

[Ano86a] Anonymous. Author index to volumes 74 and 75. Studies in Applied Mathematics, 75(3):273–274, December 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1986753273.

Anonymous:1986:TIV

[Ano86b] Anonymous. Title index to volumes 74 and 75. Studies in Applied Mathematics, 75(3):275–276, December 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1986753275.

Anonymous:1987:AIV

[Ano87a] Anonymous. Author index to volumes 76 and 77. Studies in Applied Mathematics, 77(3):303–304, December 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1987773303.

Anonymous:1987:TIV

[Ano87b] Anonymous. Title index to volumes 76 and 77. Studies in Applied Mathematics, 77(3):305, December 1987. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1987773305.

Anonymous:1988:AIV

[Ano88a] Anonymous. Author index to volumes 78 and 79. Studies in Applied Mathematics, 79(3):291–292, December 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1988793291.

Anonymous:1988:TIV

[Ano88b] Anonymous. Title index to volumes 78 and 79. Studies in Applied Mathematics, 79(3):293, December 1988. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1988793293.

Anonymous:1989:AIV

[Ano89a] Anonymous. Author index to volumes 80 and 81. Studies in Applied Mathematics, 81(3):271–272, December 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1989813271.

Anonymous:1989:TIV

[Ano89b] Anonymous. Title index to volumes 80 and 81. Studies in Applied Mathematics, 81(3):273, December 1989. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1989813273.

Anonymous:1990:AIV

[Ano90a] Anonymous. Author index to volumes 82 and 83. Studies in Applied Mathematics, 83(4):357–358, November 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990834357.

Anonymous:1990:TIV

[Ano90b] Anonymous. Title index to volumes 82 and 83. Studies in Applied Mathematics, 83(4):359–360, November 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990834359.

Anonymous:1991:AIV

[Ano91a] Anonymous. Author index to volumes 84 and 85. Studies in Applied Mathematics, 85(4):373–375, November 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991854373.

Anonymous:1991:TIV

[Ano91b] Anonymous. Title index to volumes 84 and 85. Studies in Applied Mathematics, 85(4):377–378, November 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991854377.

Anonymous:1992:AIV

[Ano92a] Anonymous. Author index to volumes 86 and 87. Studies in Applied Mathematics, 87(4):367–368, November 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1992874367.

Anonymous:1992:TIV

[Ano92b] Anonymous. Title index to volumes 86 and 87. Studies in Applied Mathematics, 87(4):369, November 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1992874369.

Anonymous:1994:AIV

[Ano94a] Anonymous. Author index to volumes 91, 92, and 93. Studies in Applied Mathematics, 93(3):275–277, November 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1994933275.

Anonymous:1994:TIV

[Ano94b] Anonymous. Title index to volumes 91, 92, and 93. Studies in Applied Mathematics, 93(3):279–280, November 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1994933279.

Anonymous:1995:AIV

[Ano95] Anonymous. Author index to volumes 94 and 95. Studies in Applied Mathematics, 95(4):457–459, November 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995954457.

Anonymous:1996:AIV

[Ano96] Anonymous. Author index to volumes 96 and 97. Studies in Applied Mathematics, 97(4):393–395, November 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996974393.

Anonymous:1998:AIV

[Ano98] Anonymous. Author index to volumes 100 and 101: 1998. Studies in Applied Mathematics, 101(4):457–459, November 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00101.

Anonymous:1999:AI

[Ano99] Anonymous. 1999 author index. Studies in Applied Mathematics, 103(4):371–373, November 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
1034138.

Anonymous:2000:AI

[Ano00] Anonymous. 2000 author index. Studies in Applied Mathematics, 105(4):405–407, November 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00158.

Anonymous:2003:IV

[Ano03] Anonymous. Index to volumes 110 and 111. Studies in Applied Mathematics, 111(4):467–469, November 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.t01-1-00001.

Anonymous:2004:AIV

[Ano04] Anonymous. Author index to volumes 112 and 113. Studies in Applied Mathematics, 113(4):413–415, November 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2004.01540.x.

Anonymous:2005:AIV

[Ano05a] Anonymous. Author index to volumes 114 and 115. Studies in Applied Mathematics, 115(4):433–436, November 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2005.115_04_01.x.

Anonymous:2005:EWW

[Ano05b] Anonymous. Errata: Wang and Wong, SAM, 115:1, p. 13 and Gubeskys, Malomed, and Merhasin, SAM, 115:2, p. 266 [i.e., 255]. Studies in Applied Mathematics, 115(3):355, October 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2005.20050826.x. See [WW05,

GMM05].

Anonymous:2006:SAM

[Ano06] Anonymous. STUDIES IN APPLIED MATHEMATICS author index to volumes 116 and 117. Studies in Applied Mathematics, 117(4):391–393, November 2006. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2006.00360.x.

Anonymous:2016:IICa

[Ano16a] Anonymous. Issue information — copyright page. Studies in Applied Mathematics, 136(2):138, February 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12114.

Anonymous:2016:IITa

[Ano16b] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 136(2):137, February 2016. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12115.

Anonymous:2016:IITb

[Ano16c] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 136(3):239, April 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12120.

Anonymous:2016:IITc

[Ano16d] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 136(4):355–356, May 2016. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12129.

Anonymous:2016:IITd

[Ano16e] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 137(1):1–2, July 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12143.

Anonymous:2016:IITe

[Ano16f] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 137(2):155–156, August 2016. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12145.

Anonymous:2016:IITf

[Ano16g] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 137(3):271–272, October 2016. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12146.

Anonymous:2016:IITg

[Ano16h] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 137(4):417–418, November 2016. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12155.

Anonymous:2016:IICb

[Ano16i] Anonymous. Issue information-copyright page. Studies in Applied Mathematics, 136(3):240, April 2016. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12119.

Anonymous:2017:IITa

[Ano17a] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 138(1):1–2, January 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12157.

Anonymous:2017:IITb

[Ano17b] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 138(2):133–134, February 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12162.

Anonymous:2017:IITc

[Ano17c] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 138(3):245–246, April 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12167.

Anonymous:2017:IITd

[Ano17d] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 138(4):369–370, May 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12176.

Anonymous:2017:IITe

[Ano17e] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 139(1):1–2, July 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12182.

Anonymous:2017:IITf

[Ano17f] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 139(2):219–220, August 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12188.

Anonymous:2017:IITg

[Ano17g] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 139(3):369–370, October 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12193.

Anonymous:2017:IITh

[Ano17h] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 139(4):503–504, November 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12196.

Anonymous:2018:IITa

[Ano18a] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 140(1):1–2, January 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12203.

Anonymous:2018:IITb

[Ano18b] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 140(2):131–132, February 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12205.

Anonymous:2018:IITc

[Ano18c] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 140(3):253–254, April 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12213.

Anonymous:2018:IITd

[Ano18d] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 140(4):401–402, May 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12218.

Anonymous:2018:IITe

[Ano18e] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 141(1):1–2, July 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12226.

Anonymous:2018:IITf

[Ano18f] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 141(2):143–144, August 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12228.

Anonymous:2018:IITg

[Ano18g] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 141(3):265–266, October 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12237.

Anonymous:2018:IITh

[Ano18h] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 141(4):419–420, November 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12245.

Anonymous:2019:IITa

[Ano19a] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 142(1):1–2, January 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12251.

Anonymous:2019:IITb

[Ano19b] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 142(2):107–108, February 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12260.

Anonymous:2019:IITc

[Ano19c] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 142(3):213–214, April 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12264.

Anonymous:2019:IITd

[Ano19d] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 142(4):415–416, May 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12269.

Anonymous:2019:IITe

[Ano19e] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 143(1):1–2, July 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12276.

Anonymous:2019:IITf

[Ano19f] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 143(2):121–122, August 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12278.

Anonymous:2019:IITg

[Ano19g] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 143(3):211–212, October 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12286.

Anonymous:2019:IITh

[Ano19h] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 143(4):337–338, November 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12289.

Anonymous:2020:IITa

[Ano20a] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 144(1):1–2, January 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12298.

Anonymous:2020:IITb

[Ano20b] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 144(2):131–132, February 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12301.

Anonymous:2020:IITc

[Ano20c] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 144(3):243–244, April 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12308.

Anonymous:2020:IITd

[Ano20d] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 144(4):387–388, May 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12314.

Anonymous:2020:IITe

[Ano20e] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 145(1):1–2, July 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12324.

Anonymous:2020:IITf

[Ano20f] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 145(2):151–152, August 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12330.

Anonymous:2020:IITg

[Ano20g] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 145(3):323–324, October 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12341.

Anonymous:2020:IITh

[Ano20h] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 145(4):623–624, November 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12350.

Anonymous:2021:IITa

[Ano21a] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 146(1):1–2, January 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12360.

Anonymous:2021:IITb

[Ano21b] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 146(2):299–300, February 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12369.

Anonymous:2021:IITc

[Ano21c] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 146(3):563–564, April 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12377.

Anonymous:2021:IITd

[Ano21d] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 146(4):777–778, May 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12391.

Anonymous:2021:IITe

[Ano21e] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 147(1):1–3, July 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12419.

Anonymous:2021:IITf

[Ano21f] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 147(2):417–419, August 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12431.

Anonymous:2021:IITg

[Ano21g] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 147(3):835–837, October 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12453.

Anonymous:2021:IITh

[Ano21h] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 147(4):1195–1197, November 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12468.

Anonymous:2022:IITa

[Ano22a] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 148(1):1–3, January 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12406.

Anonymous:2022:IITb

[Ano22b] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 148(2):457–459, February 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12408.

Anonymous:2022:IITc

[Ano22c] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 148(3):987–989, April 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12409.

Anonymous:2022:IITd

[Ano22d] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 148(4):1407–1409, May 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12410.

Anonymous:2022:IITe

[Ano22e] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 149(1):1–3, July 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12411.

Anonymous:2022:IITf

[Ano22f] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 149(2):293–295, August 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12412.

Anonymous:2022:IITg

[Ano22g] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 149(3):561–563, October 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12413.

Anonymous:2022:IITh

[Ano22h] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 149(4):839–841, November 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12407.

Anonymous:2023:IITa

[Ano23a] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 150(1):1–3, January 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12502.

Anonymous:2023:IITb

[Ano23b] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 150(2):317–319, February 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12503.

Anonymous:2023:IITc

[Ano23c] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 150(3):603–605, April 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12504.

Anonymous:2023:IITd

[Ano23d] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 150(4):959–961, May 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12505.

Anonymous:2023:IITe

[Ano23e] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 151(1):1–3, July 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12506.

Anonymous:2023:IITf

[Ano23f] Anonymous. Issue information — TOC. Studies in Applied Mathematics, 151(2):407–409, August 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12507.

Ackerberg:1970:BLP

[AO70] R. C. Ackerberg and R. E. O’Malley, Jr. Boundary layer problems exhibiting resonance. Studies in Applied Mathematics, 49(3):277–295, September 1970. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1970493277.

Agarwal:2002:IIP

[AO02] Ravi P. Agarwal and Donal O’Regan. Infinite interval problems modeling the flow of a gas through a semi-infinite porous medium. Studies in Applied Mathematics, 108(3):245–257, April 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.01411.

Agarwal:2003:IIP

[AO03] R. P. Agarwal and D. O’Regan. Infinite interval problems modeling phenomena which arise in the theory of plasma and electrical potential theory. Studies in Applied Mathematics, 111(3):339–358, October 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
t01-1-00237.

Aouadi:2020:QSU

[Aou20] Moncef Aouadi. Quasi-stability and upper semicontinuity for coupled parabolic equations with memory. Studies in Applied Mathematics, 145(3):586–621, October 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12331.

Alexander:2023:VCS

[AP23] Jordan M. Alexander and D. Palaniappan. Viability of circle and sphere theorems in potential theory and hydrodynamics via Maxwell’s conjecture. Studies in Applied Mathematics, 150(2):582–602, February 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12550.

Agnon:1998:DCW

[APS98] Yehuda Agnon, Efim Pelinovsky, and Alexandru Sheremet. Disintegration of cnoidal waves over smooth topography. Studies in Applied Mathematics, 101(1):49–71, July 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00085.

Ablowitz:2006:DVS

[APT06] M. J. Ablowitz, B. Prinari, and A. D. Trubatch. Discrete vector solitons: Composite solitons, Yang–Baxter maps and computation. Studies in Applied Mathematics, 116(1):97–133, January 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2005.00320.x.

Arkeryd:1992:STE

[Ark92] Leif Arkeryd. On the strong L^{1} trend to equilibrium for the Boltzmann equation. Studies in Applied Mathematics, 87(3):283–288, October 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1992873283.

Ablowitz:1977:ASK

[AS77] M. J. Ablowitz and H. Segur. Asymptotic solutions of the Korteweg–de Vries equation. Studies in Applied Mathematics, 57(1):13–44, July 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm197757113.

See note [AK82].

Ablowitz:1980:LIW

[AS80] M. J. Ablowitz and H. Segur. Long internal waves in fluids of great depth. Studies in Applied Mathematics, 62(3):249–262, June 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1980623249.

Ablowitz:1989:SCL

[AS89] Mark J. Ablowitz and Cherie L. Schultz. Strong coupling limit of certain multidimensional nonlinear wave equations. Studies in Applied Mathematics, 80(3):229–238, June 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1989803229.

Ashton:2010:SPF

[Ash10] A. C. L. Ashton. Stability of parallel fluid loaded plates: a nonlocal approach. Studies in Applied Mathematics, 125 (3):301–329, October 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2010.00490.x.

Arkhipov:2012:CAE

[ASSS12] B. Arkhipov, V. Solbakov, M. Solov’ev, and D. Shapochkin. Computation of aqueous environment contamination parameters on the basis of the Lagrangian approach. Studies in Applied Mathematics, 129(4):347–361, November 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00561.x.

Aggarwal:2020:SCC

[ASSV20] Aekta Aggarwal, Manas Ranjan Sahoo, Abhrojyoti Sen, and Ganesh Vaidya. Solutions with concentration for conservation laws with discontinuous flux and its applications to numerical schemes for hyperbolic systems. Studies in Applied Mathematics, 145(2):247–290, August 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12319.

Aguirre:1990:DHT

[AT90] Manuel A. Aguirre and Susana Elena Trione. The distributional Hankel transform of δ^{(k)}(m^{2} + P). Studies in Applied Mathematics, 83(2):111–121, August 1990. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990832111.

Abdurahman:2007:PEG

[AT07a] Xamxinur Abdurahman and Zhidong Teng. Persistence and extinction for general nonautonomous n-species Lotka– Volterra cooperative systems with delays. Studies in Applied Mathematics, 118(1):17–43, January 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2007.00362.x.

Athanasiadis:2007:STA

[AT07b] Christodoulos Athanasiadis and Nikolaos L. Tsitsas. Scattering theorems for acoustic excitation of a layered obstacle by an interior point source. Studies in Applied Mathematics, 118(4):397–418, May 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1365-2966.2007.00375.x.

Atkinson:1978:ESW

[Atk78] C. Atkinson. The existence of steady wave solutions to a non-local version of the K.P.P. equation. Studies in Applied Mathematics, 59(2):95–115, October 1978. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197859295.

Ablowitz:1991:KPE

[AV91] Mark J. Ablowitz and Javier Villarroel. On the Kadomtsev– Petviashvili equation and associated constraints. Studies in Applied Mathematics, 85(3):195–213, October 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991853195.

Ablowitz:1997:ITL

[AW97] Mark J. Ablowitz and Xiao-Ping Wang. Initial time layers and Kadomtsev–Petviashvili-type equations. Studies in Applied Mathematics, 98(2):121–137, February 1997. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00043.

Akylas:1995:SSO

[AY95] T. R. Akylas and T.-S. Yang. On short-scale oscillatory tails of long-wave disturbances. Studies in Applied Mathematics, 94(1):1–20, January 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19959411.

Ablowitz:2017:PSIa

[AY17a] Mark J. Ablowitz and Jianke Yang. Preface: Special issues in memory of David Benney (part i). Studies in Applied Mathematics, 139(1):3–6, July 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12185.

Ablowitz:2017:PSIb

[AY17b] Mark J. Ablowitz and Jianke Yang. Preface: Special issues in memory of David Benney (Part II). Studies in Applied Mathematics, 139(2):221–222, August 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12186.

Ablowitz:1983:IST

[AYF83] M. J. Ablowitz, D. Bar Yaacov, and A. S. Fokas. On the inverse scattering transform for the Kadomtsev–Petviashvili equation. Studies in Applied Mathematics, 69(2):135–143, October 1983. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1983692135.

Ablowitz:2013:UOD

[AZ13] M. J. Ablowitz and Y. Zhu. Unified orbital description of the envelope dynamics in two-dimensional simple periodic lattices. Studies in Applied Mathematics, 131(1):41–71, July 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00575.x.

An:2022:EVI

[AZ22] Hongli An and Haixing Zhu. The elliptical vortices, integrable Ermakov structure, Schrödinger connection, and Lax pair in the compressible Navier–Stokes equation. Studies in Applied Mathematics, 149(4):879–903, November 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12524.

Bagderina:2008:ETO

[Bag08] Yulia Yu. Bagderina. Equivalence of third-order ordinary differential equations to chazy equations I–XIII. Studies in Applied Mathematics, 120(3):293–332, April 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00400.x.

Baker:1976:TCP

[Bak76] George A. Baker, Jr. A theorem on the convergence of Padé approximants. Studies in Applied Mathematics, 55(2):107–117, June 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1976552107.

Besley:1998:POW

[BAM98] J. A. Besley, N. N. Akhmediev, and P. D. Miller. Periodic optical waveguides: Exact Floquet theory and spectral properties. Studies in Applied Mathematics, 101(4):343–355, November 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00097.

Barcilon:1971:OSU

[Bar71] Victor Barcilon. Optimum signals and the uncertainty principle. Studies in Applied Mathematics, 50(2):103–114, June 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1971502103.

Barnett:1973:SST

[Bar73] Arnold Barnett. On scheduling some transport systems with discrete-time passenger arrivals. Studies in Applied Mathematics, 52(1):75–86, March 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm197352175.

Barouch:1977:RQC

[Bar77] Eytan Barouch. Resonance in a quantum chain. Studies in Applied Mathematics, 56(2):169–185, April 1977. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1977562169.

Barouch:1983:IDX

[Bar83] Eytan Barouch. Impurity dynamics of the XY-model. Studies in Applied Mathematics, 69(3):201–209, December 1983. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1983693201.

Barouch:1984:LPF

[Bar84] E. Barouch. Lax pair for the free-fermion eight-vertex model. Studies in Applied Mathematics, 70(2):151–162, April 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1984702151.

Barros:2016:RSN

[Bar16] Ricardo Barros. Remarks on a strongly nonlinear model for two-layer flows with a top free surface. Studies in Applied Mathematics, 136(3):263–287, April 2016. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12104.

Baumstein:1998:MGW

[Bau98] Anatoly I. Baumstein. Modulation of gravity waves with shear in water. Studies in Applied Mathematics, 100(4):365–390, May 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00081.

Baumstein:1999:SNG

[Bau99] Anatoly I. Baumstein. Steady nonsymmetrical gravity waves with shear in water. Studies in Applied Mathematics, 102(3):267–290, April 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00111.

Baxter:1971:GFM

[Bax71] R. J. Baxter. Generalized ferroelectric model on a square lattice. Studies in Applied Mathematics, 50(1):51–69, March 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197150151.

Barcilon:1969:SIWb

[BB69a] Victor Barcilon and Norman Bleistein. Scattering of inertial waves by smooth convex cylinders. Studies in Applied Mathematics, 48(4):351–359, December 1969. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1969484351.

Barcilon:1969:SIWa

[BB69b] Victor Barcilon and Norman Bleistein. Scattering of inertial waves in a rotating fluid. Studies in Applied Mathematics, 48(2):91–104, June 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm196948291.

Benney:1969:NCN

[BB69c] D. J. Benney and R. F. Bergeron, Jr. A new class of nonlinear waves in parallel flows. Studies in Applied Mathematics, 48(3):181–204, September 1969. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1969483181.

Babu:1987:ESO

[BB87] S. V. Babu and E. Barouch. An exact solution for the optical absorbance of thin films. Studies in Applied Mathematics, 77(2):173–182, October 1987. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1987772173.

Biondini:2012:NSE

[BB12] G. Biondini and A. Bui. On the nonlinear Schrödinger equation on the half line with homogeneous Robin boundary conditions. Studies in Applied Mathematics, 129(3):249–271, October 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2012.00553.x.

Buljan:2005:PCW

[BBC^{+}05] H. Buljan, G. Bartal, O. Cohen, T. Schwartz, O. Manela, T. Carmon, M. Segev, J. W. Fleischer, and D. N. Christodoulides.

Partially coherent waves in nonlinear periodic lattices. Studies in Applied Mathematics, 115(2):173–208, August 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2005.00325.x.

Bassa:2010:PPG

[BBC^{+}10] A. Bassa, J. Burns, J. Campbell, A. Deshpande, J. Farley, M. Halsey, S.-Y. Ho, D. Kleitman, S. Michalakis, P.-O. Persson, P. Pylyavskyy, L. Rademacher, A. Riehl, M. Rios, J. Samuel, B. E. Tenner, A. Vijayasarathy, and L. Zhao. Partitioning a planar graph of girth 10 into a forest and a matching. Studies in Applied Mathematics, 124(3):213–228, April 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00468.x.

Borluk:2022:TWT

[BBN22] Handan Borluk, Gabriele Bruell, and Dag Nilsson. Traveling waves and transverse instability for the fractional Kadomtsev– Petviashvili equation. Studies in Applied Mathematics, 149(1): 95–123, July 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12494.

Bobbitt:1976:NTC

[BC76] J. I. Bobbitt and E. Cumberbatch. Nonlinear theory of a caustic region. Studies in Applied Mathematics, 55(3):239–247, September 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1976553239.

Benney:1985:AAN

[BC85] D. J. Benney and K. Chow. An alternative approach to nonlinear instabilities in hydrodynamics. Studies in Applied Mathematics, 73(3):261–267, December 1985. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1985733261.

Benney:1986:IWF

[BC86] D. J. Benney and K. Chow. Instabilities of waves on a free surface. Studies in Applied Mathematics, 74(3):227–243, June 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1986743227.

Benney:1989:MFF

[BC89] D. J. Benney and K. Chow. A mean flow first harmonic theory for hydrodynamic instabilities. Studies in Applied Mathematics, 80(1):37–73, February 1989. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198980137.

Barros:2009:ISI

[BC09] Ricardo Barros and Wooyoung Choi. Inhibiting shear instability induced by large amplitude internal solitary waves in two-layer flows with a free surface. Studies in Applied Mathematics, 122(3):325–346, April 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2009.00436.x.

Byun:2023:ACS

[BC23] Sung-Soo Byun and Christophe Charlier. On the almostcircular symplectic induced Ginibre ensemble. Studies in Applied Mathematics, 150(1):184–217, January 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12537.

Bernardi:2023:SFS

[BCCM23] Francesca Bernardi, Shankararaman Chellam, N. G. Cogan, and M. N. J. Moore. Stokes flow solutions in infinite and semi-infinite porous channels. Studies in Applied Mathematics, 151(1):116–140, July 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12574.

Barhoumi:2021:GPP

[BCD21] A. Barhoumi, A. F. Celsus, and A. Deaño. Global-phase portrait and large-degree asymptotics for the kissing polynomials. Studies in Applied Mathematics, 147(2):448–526, August 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12387.

Bassom:1995:BTS

[BCH95] Andrew P. Bassom, Peter A. Clarkson, and Andrew C. Hicks. Bäcklund transformations and solution hierarchies for the fourth Painlevé equation. Studies in Applied Mathematics, 95(1):1–71, July 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19959511.

Barouch:1985:PSS

[BCKW85] E. Barouch, S. Chow, G. M. Kaufman, and T. H. Wright. Properties of successive sample moment estimators. Studies in Applied Mathematics, 73(3):239–259, December 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1985733239.

Baez-Duarte:1970:KCT

[BD70] Luis Báez-Duarte. C(X)∗ and Kolmogorov’s consistency theorem for Cantor spaces. Studies in Applied Mathematics, 49(4):401–403, December 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1970494401.

Barcilon:2001:NWV

[BD01] A. Barcilon and P. G. Drazin. Nonlinear waves of vorticity. Studies in Applied Mathematics, 106(4):437–479, May 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00174.

Borisov:2017:SSK

[BD17] D. I. Borisov and S. V. Dmitriev. On the spectral stability of kinks in 2D Klein–Gordon model with parity-time-symmetric perturbation. Studies in Applied Mathematics, 138(3):317–342, April 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12156.

Baez-Duarte:1997:HGH

[BDC97] Luis Báez-Duarte and G. A. Cámera. Horizontal growth of harmonic functions. Studies in Applied Mathematics, 98(2):195–202, February 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00045.

Bresch:2002:LMN

[BDGL02] D. Bresch, B. Desjardins, E. Grenier, and C.-K. Lin. Low Mach number limit of viscous polytropic flows: Formal asymptotics in the periodic case. Studies in Applied Mathematics, 109(2):125–149, August 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
01440.

Barone:2011:DNM

[BDLP11] Vincenzo Barone, Silvana De Lillo, Gaia Lupo, and Antonino Polimeno. Dirichlet-to-Neumann map for a nonlinear diffusion equation. Studies in Applied Mathematics, 126(2):145–155, February 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2010.00500.x.

Blaszak:2022:SCNb

[BDM22] Maciej Błaszak, Ziemowit Domański, and Krzysztof Marciniak. Systematic construction of nonautonomous Hamiltonian equations of Painlevé-type. II. Isomonodromic Lax representation. Studies in Applied Mathematics, 149(2):364–415, August 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12495.

Bonneux:2020:CWH

[BDS20] Niels Bonneux, Clare Dunning, and Marco Stevens. Coefficients of Wronskian Hermite polynomials. Studies in Applied Mathematics, 144(3):245–288, April 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12290.

Berkolaiko:2022:TDE

[BE22] Gregory Berkolaiko and Mahmood Ettehad. Three-dimensional elastic beam frames: Rigid joint conditions in variational and differential formulation. Studies in Applied Mathematics, 148(4):1586–1623, May 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12485.

Beardsley:1970:ESI

[Bea70] R. C. Beardsley. An experimental study of inertial waves in a closed cone. Studies in Applied Mathematics, 49(2):187–196, June 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1970492187.

Bell:1977:TCI

[Bel77] David E. Bell. A theorem concerning the integer lattice. Studies in Applied Mathematics, 56(2):187–188, April 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1977562187.

Belding:1979:PCP

[Bel79] Russell Belding. Path-connected partially ordered sets. Studies in Applied Mathematics, 60(2):165–172, April 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1979602165.

Benney:1973:SPL

[Ben73] D. J. Benney. Some properties of long nonlinear waves. Studies in Applied Mathematics, 52(1):45–50, March 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197352145.

Benney:1974:IDL

[Ben74] D. J. Benney. Interaction and dispersion of long waves. Studies in Applied Mathematics, 53(2):137–144, June 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1974532137.

Benedetto:1975:WSS

[Ben75] John J. Benedetto. The Wiener spectrum in spectral synthesis. Studies in Applied Mathematics, 54(2):91–115, June 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197554291.

Benney:1976:SIB

[Ben76] D. J. Benney. Significant interactions between small and large scale surface waves. Studies in Applied Mathematics, 55(2):93– 106, June 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm197655293.

Benney:1977:GTI

[Ben77] D. J. Benney. A general theory for interactions between short and long waves. Studies in Applied Mathematics, 56(1):81–94, January 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm197756181.

Benney:1979:IAF

[Ben79a] D. J. Benney. Instabilities associated with forced nonlinear waves. Studies in Applied Mathematics, 60(1):27–41, February 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197960127.

Benney:1979:LAR

[Ben79b] D. J. Benney. Large amplitude Rossby waves. Studies in Applied Mathematics, 60(1):1–10, February 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19796011.

Benedetto:1981:TCS

[Ben81] John J. Benedetto. The theory of constructive signal analysis. Studies in Applied Mathematics, 65(1):37–80, August 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198165137.

Benney:1983:NWP

[Ben83] D. J. Benney. Nonlinear wave packets in flows with critical layers. Studies in Applied Mathematics, 69(3):177–200, December 1983. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1983693177.

Benney:1984:EDS

[Ben84] D. J. Benney. The evolution of disturbances in shear flows at high Reynolds numbers. Studies in Applied Mathematics, 70(1):1–19, February 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19847011.

Benilov:1992:SWS

[Ben92] Eugene S. Benilov. On the surface waves in a shallow channel with an uneven bottom. Studies in Applied Mathematics, 87(1):1–14, July 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19928711.

Benilov:2006:DST

[Ben06] E. S. Benilov. Does surface tension stabilize liquid films inside a rotating horizontal cylinder?Part 2: Multidimensional disturbances. Studies in Applied Mathematics, 116(1):1–20, January 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2005.00331.x.

Bergeron:1970:SIC

[Ber70] R. F. Bergeron, Jr. Some inviscid “Cat’s eye” flows. Studies in Applied Mathematics, 49(2):197–203, June 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1970492197.

Berg:1971:PDP

[Ber71] C. A. Berg. On plane dilatational plasticity. Studies in Applied Mathematics, 50(3):193–203, September 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1971503193.

Berg:1972:CEP

[Ber72] C. A. Berg. Construction of the equivalent plastic strain increment. Studies in Applied Mathematics, 51(3):311–316, September 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1972513311.

Bergstrom:1992:IAG

[Ber92] Lars Bergström. Initial algebraic growth of small angular dependent disturbances in pipe Poiseuille flow. Studies in Applied Mathematics, 87(1):61–79, July 1992. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199287161.

Barouch:1985:MSS

[BF85] E. Barouch and B. Fuchssteiner. Master symmetries and similarity equations of the XYh model. Studies in Applied Mathematics, 73(3):221–237, December 1985. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1985733221.

Biondini:2016:PMJa

[BF16a] Gino Biondini and Athanassios S. Fokas. Preface: Mark j. Ablowitz, nonlinear waves and integrable systems. Part I. Studies in Applied Mathematics, 137(1):3–9, July 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12140.

Biondini:2016:PMJb

[BF16b] Gino Biondini and Athanassios S. Fokas. Preface: Mark j. Ablowitz, nonlinear waves and integrable systems. Part II. Studies in Applied Mathematics, 137(2):157–158, August 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12141.

Bornemann:2017:FSE

[BFM17] Folkmar Bornemann, Peter J. Forrester, and Anthony Mays. Finite size effects for spacing distributions in random matrix theory: Circular ensembles and Riemann zeros. Studies in Applied Mathematics, 138(4):401–437, May 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12160.

Benney:1981:NML

[BG81] D. J. Benney and L. Håkan Gustavsson. A new mechanism for linear and nonlinear hydrodynamic instability. Studies in Applied Mathematics, 64(3):185–209, June 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1981643185.

Benney:1982:LAS

[BG82] D. J. Benney and R. Grimshaw. Large amplitude solitary waves in unbounded stratified fluids. Studies in Applied Mathematics, 66(2):181–187, April 1982. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1982662181.

Berkshire:1983:CDN

[BG83] F. H. Berkshire and J. D. Gibbon. Collapse in the ndimensional nonlinear Schrödinger equation — a parallel with Sundman’s. Studies in Applied Mathematics, 69(3):229–262, December 1983. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1983693229.

Barros:2007:DNWb

[BG07] Ricardo Barros and Sergey Gavrilyuk. Dispersive nonlinear waves in two-layer flows with free surface Part II. Large amplitude solitary waves embedded into the continuous spectrum. Studies in Applied Mathematics, 119(3):213–251, October 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2007.00384.x.

Bagarello:2023:DQP

[BG23] F. Bagarello and F. Gargano. Dynamics for a quantum parliament. Studies in Applied Mathematics, 150(4):1182–1200, May 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12566.

Broadbridge:1997:DND

[BGL97] P. Broadbridge, J. M. Goard, and M. Lavrentiev, Jr. Degenerate nonlinear diffusion with an initially sharp front. Studies in Applied Mathematics, 99(4):377–391, November 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00064.

Bian:2015:HOS

[BGL15] Dongfen Bian, Boling Guo, and Liming Ling. High-order soliton solution of Landau–Lifshitz equation. Studies in Applied Mathematics, 134(2):181–214, February 2015. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12051.

Bakhshi:2021:TWS

[BGMR21] Marzieh Bakhshi, Anna Ghazaryan, Vahagn Manukian, and Nancy Rodriguez. Traveling wave solutions in a model for social outbursts in a tension-inhibitive regime. Studies in Applied Mathematics, 147(2):650–674, August 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12394.

Barros:2007:DNWa

[BGT07] R. Barros, S. L. Gavrilyuk, and V. M. Teshukov. Dispersive nonlinear waves in two-layer flows with free surface. I. Model derivation and general properties. Studies in Applied Mathematics, 119(3):191–211, October 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2007.00383.x.

Bassom:1989:GMF

[BH89] Andrew P. Bassom and Philip Hall. On the generation of mean flows by the interaction of Görtler vortices and Tollmien–Schlichting waves in curved channel flows. Studies in Applied Mathematics, 81(3):185–219, December 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1989813185.

Butzer:1991:SFS

[BH91] P. L. Butzer and M. Hauss. On Stirling functions of the second kind. Studies in Applied Mathematics, 84(1):71–91, January 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199184171.

Bourland:1994:CAS

[BH94] F. Jay Bourland and Richard Haberman. Connection across a separatrix with dissipation. Studies in Applied Mathematics, 91(2):95–124, February 1994. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199491295.

Berloff:1997:SPS

[BH97] Natalia G. Berloff and Louis N. Howard. Solitary and periodic solutions of nonlinear nonintegrable equations. Studies in Applied Mathematics, 99(1):1–24, July 1997. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00054.

Berloff:1998:NWI

[BH98a] Natalia G. Berloff and Louis N. Howard. Nonlinear wave interactions in nonlinear nonintegrable systems. Studies in Applied Mathematics, 100(3):195–213, April 1998. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00075.

Brothers:1998:SPT

[BH98b] Jerry D. Brothers and Richard Haberman. Slow passage through a homoclinic orbit with subharmonic resonances. Studies in Applied Mathematics, 101(2):211–232, August 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00091.

Benilov:2006:EPS

[BH06] E. S. Benilov and C. P. Howlin. Evolution of packets of surface gravity waves over strong smooth topography. Studies in Applied Mathematics, 116(3):289–301, April 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2006.00344.x.

Bender:2014:CCM

[BH14a] Carl M. Bender and Daniel W. Hook. Complex classical motion in potentials with poles and turning points. Studies in Applied Mathematics, 133(3):318–336, October 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12059.

Bronski:2014:MIV

[BH14b] J. C. Bronski and V. M. Hur. Modulational instability and variational structure. Studies in Applied Mathematics, 132(4):285–331, May 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12029.

Benincasa:2016:BPA

[BH16] Gregorio B. Benincasa and Rod Halburd. Bianchi permutability for the anti-self-dual Yang–Mills equations. Studies in Applied Mathematics, 137(1):110–122, July 2016. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12118.

Brezinski:2011:CFM

[BHH^{+}11] Claude Brezinski, Yi He, Xing-Biao Hu, Jian-Qing Sun, and Hon-Wah Tam. Confluent form of the multistep ϵ-algorithm, and the relevant integrable system. Studies in Applied Mathematics, 127(2):191–209, August 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00518.x.

Binswanger:2021:WMT

[BHIS21] Adam L. Binswanger, Mark A. Hoefer, Boaz Ilan, and Patrick Sprenger. Whitham modulation theory for generalized Whitham equations and a general criterion for modulational instability. Studies in Applied Mathematics, 147(2):724–751, August 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12398.

Belinskiy:2021:SSL

[BHN21] Boris P. Belinskiy, Don B. Hinton, and Roger Nichols. Singular Sturm–Liouville operators with extreme properties that generate black holes. Studies in Applied Mathematics, 147(1):180– 208, July 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12385.

Baker:1981:SRI

[BI81] G. R. Baker and M. Israeli. Spinup from rest of immiscible fluids. Studies in Applied Mathematics, 65(3):249–268, December 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1981653249.

Bihlo:2022:IVS

[BJV22] Alex Bihlo, James Jackaman, and Francis Valiquette. Invariant variational schemes for ordinary differential equations. Studies in Applied Mathematics, 148(3):991–1020, April 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12466.

Barnett:1978:SOP

[BK78a] Arnold Barnett and Daniel J. Kleitman. Some optimization problems with bulk-service queues. Studies in Applied Mathematics, 58(3):277–290, June 1978. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1978583277.

Benney:1978:PLL

[BK78b] D. J. Benney and D. R. S. Ko. The propagation of long large amplitude internal waves. Studies in Applied Mathematics, 59(3):187–199, December 1978. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1978593187.

Bosley:1995:ASN

[BK95] D. L. Bosley and J. Kevorkian. On the asymptotic solution of non-Hamiltonian systems exhibiting sustained resonance. Studies in Applied Mathematics, 94(2):83–130, February 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm199594283.

Bonneux:2018:ELP

[BK18] Niels Bonneux and Arno B. J. Kuijlaars. Exceptional Laguerre polynomials. Studies in Applied Mathematics, 141(4):547–595, November 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12204.

Boutin:2023:CGB

[BK23] Mireille Boutin and Gregor Kemper. Can a groundbased vehicle hear the shape of a room? Studies in Applied Mathematics, 151(1):352–368, July 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12581.

Barouch:1986:SLR

[BKG86] E. Barouch, G. M. Kaufman, and M. L. Glasser. On sums of lognormal random variables. Studies in Applied Mathematics, 75(1):37–55, August 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm198675137.

Benney:1970:AND

[BL70] D. J. Benney and C. G. Lange. The asymptotics of nonlinear diffusion. Studies in Applied Mathematics, 49(1):1–19, March 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19704911.

Bryant:1978:PWS

[BL78] P. J. Bryant and A. K. Laing. Permanent wave structures and resonant triads in a layered fluid. Studies in Applied Mathematics, 59(3):223–243, December 1978. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1978593223.

Benzaid:1987:ARS

[BL87] Z. Benzaid and D. A. Lutz. Asymptotic representation of solutions of perturbed systems of linear difference equations. Studies in Applied Mathematics, 77(3):195–221, December 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1987773195.

Bernardi:2022:NAI

[BL22] Enrico Bernardi and Alberto Lanconelli. A note about the invariance of the basic reproduction number for stochastically perturbed SIS models. Studies in Applied Mathematics, 148(4):1543–1562, May 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12483.

Blau:1970:PPT

[Bla70] Irwin Blau. Promotion phenomenon for three particle exchange. Studies in Applied Mathematics, 49(4):389–399, December 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1970494389.

Blaszak:2011:BPR

[Bła11] Maciej Błaszak. Bi-presymplectic representation of Liouville integrable systems and related separability theory. Studies in Applied Mathematics, 126(4):319–346, May 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00505.x.

Blaszak:2023:MPO

[Bła23] Maciej Błaszak. Multicomponent Painlevé ODEs and related nonautonomous KdV stationary hierarchies. Studies in Applied Mathematics, 151(1):5–34, July 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12571.

Biondini:2021:IST

[BLM21] Gino Biondini, Jonathan Lottes, and Dionyssios Mantzavinos. Inverse scattering transform for the focusing nonlinear Schrödinger equation with counterpropagating flows. Studies in Applied Mathematics, 146(2):371–439, February 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12347.

Boiti:1988:CNR

[BLP88] M. Boiti, J. J-P. Léon, and F. Pempinelli. Canonical and noncanonical recursion operators in multidimensions. Studies in Applied Mathematics, 78(1):1–19, February 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19887811.

Bridges:2001:MRE

[BLP01] Thomas J. Bridges and Fiona E. Laine-Pearson. Multisymplectic relative equilibria, multiphase wavetrains, and coupled NLS equations. Studies in Applied Mathematics, 107(2):137–155, August 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00182.

Barouch:1972:MSR

[BM72] E. Barouch and B. M. McCoy. Magnetic susceptibility of the random one- dimensional transverse Ising model. Studies in Applied Mathematics, 51(1):57–66, March 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197251157.

Benney:1975:EST

[BM75] D. J. Benney and S. A. Maslowe. The evolution in space and time of nonlinear waves in parallel shear flows. Studies in Applied Mathematics, 54(3):181–205, September 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1975543181.

Baker:1982:SVF

[BM82] Gregory R. Baker and Steven J. Mardeusz. The steady viscous flow of two differentially rotating immiscible fluids. Studies in Applied Mathematics, 67(1):63–71, August 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198267163.

Balmforth:1999:NSI

[BM99] N. J. Balmforth and P. J. Morrison. A necessary and sufficient instability condition for inviscid shear flow. Studies in Applied Mathematics, 102(3):309–344, April 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00113.

Biello:2004:EMV

[BM04] Joseph A. Biello and Andrew J. Majda. The effect of meridional and vertical shear on the interaction of equatorial baroclinic and barotropic Rossby waves. Studies in Applied Mathematics, 112(4):341–390, May 2004. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2004.01518.x.

Blaszak:2012:RES

[BM12a] Maciej Błaszak and Krzysztof Marciniak. On reciprocal equivalence of Stäckel systems. Studies in Applied Mathematics, 129(1):26–50, July 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2011.00544.x.

Boonkasame:2012:SLA

[BM12b] Anakewit Boonkasame and Paul Milewski. The stability of large-amplitude shallow interfacial non-Boussinesq flows. Studies in Applied Mathematics, 128(1):40–58, January 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00528.x.

Blaszak:2013:ICK

[BM13] M. Błaszak and K. Marciniak. Invertible coupled KdV and coupled Harry Dym hierarchies. Studies in Applied Mathematics, 131(3):211–228, October 2013. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12008.

Boonkasame:2014:MSN

[BM14] Anakewit Boonkasame and Paul A. Milewski. A model for strongly nonlinear long interfacial waves with background shear. Studies in Applied Mathematics, 133(2):182–213, August 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12036.

Blaszak:2020:STL

[BM20] Maciej Błaszak and Krzysztof Marciniak. Stäckel transform of Lax equations. Studies in Applied Mathematics, 145(2):179– 196, August 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12315.

Blaszak:2022:SCNc

[BM22] Maciej Błaszak and Krzysztof Marciniak. Systematic construction of non-autonomous Hamiltonian equations of Painlevétype. III. Quantization. Studies in Applied Mathematics, 149(2):416–440, August 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12514.

Burns:2002:BIB

[BMB02] A. G. Burns, S. A. Maslowe, and S. N. Brown. Barotropic instability of the Bickley jet at high Reynolds numbers. Studies in Applied Mathematics, 109(4):279–296, November 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00224.

Blaszak:2022:SCNa

[BMD22] Maciej Błaszak, Krzysztof Marciniak, and Ziemowit Domański. Systematic construction of nonautonomous Hamiltonian equations of Painlevé type. I. Frobenius integrability. Studies in Applied Mathematics, 148(3):1208–1250, April 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12473.

Balagondar:1987:PFA

[BMM87] P. M. Balagondar, S. A. Maslowe, and S. Melkonian. The propagation of finite-amplitude waves in a model boundary layer. Studies in Applied Mathematics, 76(2):169–185, April 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1987762169.

Biondini:2009:CGB

[BMOT09a] G. Biondini, K.-I. Maruno, M. Oikawa, and H. Tsuji. Corrigendum: G. Biondini, K.-I. Maruno, M. Oikawa, and H. Tsuji, Soliton Interactions of the Kadomtsev–Petviashvili Equation and Generation of Large-Amplitude Water Waves, Stud. Appl. Math. 122:377–394 (2009). Studies in Applied Mathematics, 123(4):375–376, November 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2009.00459.x. See [BMOT09b].

Biondini:2009:SIK

[BMOT09b] G. Biondini, K.-I. Maruno, M. Oikawa, and H. Tsuji. Soliton interactions of the Kadomtsev–Petviashvili equation and generation of large-amplitude water waves. Studies in Applied Mathematics, 122(4):377–394, May 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00439.x. See corrigen-

dum [BMOT09a].

Bothner:2018:RSP

[BMS18] Thomas Bothner, Peter D. Miller, and Yue Sheng. Rational solutions of the Painlevé-III equation. Studies in Applied Mathematics, 141(4):626–679, November 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12220.

Benney:1969:RWC

[BN69] D. J. Benney and Alan C. Newell. Random wave closures. Studies in Applied Mathematics, 48(1):29–53, March 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm196948129.

Beals:1994:ISM

[BN94] Richard Beals and Roman G. Nouikou. The inverse scattering method and systems of equations q_{xy} = F (q, q_{y}). Studies in Applied Mathematics, 92(3):201–212, July 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1994923201.

Boyd:1998:SLE

[BN98] John P. Boyd and Andrei Natarov. A Sturm–Liouville eigenproblem of the fourth kind: a critical latitude with equatorial trapping. Studies in Applied Mathematics, 101 (4):433–455, November 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00100.

Barouch:2012:SOO

[BO12] E. Barouch and S. A. Orszag. Source optimization in optical lithography. Studies in Applied Mathematics, 128 (2):144–158, February 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2011.00532.x.

Biondini:2020:SDC

[BO20] Gino Biondini and Jeffrey Oregero. Semiclassical dynamics and coherent soliton condensates in self-focusing nonlinear media with periodic initial conditions. Studies in Applied Mathematics, 145(3):325–356, October 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12321.

Boa:1975:MSS

[Boa75] James A. Boa. Multiple steady states in a model biochemical reaction. Studies in Applied Mathematics, 54(1):9–15, March 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19755419.

Barrowes:2004:AES

[BOGK04] Benjamin E. Barrowes, Kevin O’Neill, Tomasz M. Grzegorczyk, and Jin A. Kong. On the asymptotic expansion of the spheroidal wave function and its eigenvalues for complex size parameter. Studies in Applied Mathematics, 113(3):271–301, October 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
0022-2526.2004.01526.x.

Bona:1995:TLM

[Bön95] Miklös Böna. There are a lot of magic squares! Studies in Applied Mathematics, 94(4):415–421, May 1995. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995944415.

Booty:1989:LSC

[Boo89] M. R. Booty. On the limiting steady configurations of a model plasma. Studies in Applied Mathematics, 80(3):203–228, June 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1989803203.

Boulton:2021:NRP

[BOPS21] Lyonell Boulton, Peter J. Olver, Beatrice Pelloni, and David A. Smith. New revival phenomena for linear integro–differential equations. Studies in Applied Mathematics, 147(4):1209–1239, November 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12397.

Boyd:2016:QUS

[Boy16] John P. Boyd. Quasi-uniform spectral schemes (QUSS), Part I: Constructing generalized ellipses for graphical grid generation. Studies in Applied Mathematics, 136(2):189–213, February 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12099.

Benedek:1971:MCF

[BP71] A. Benedek and R. Panzone. On mean convergence of Fourier–Bessel series of negative order. Studies in Applied Mathematics, 50(3):281–292, September 1971. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1971503281.

Biondini:2014:SDO

[BP14] Gino Biondini and Barbara Prinari. On the spectrum of the Dirac operator and the existence of discrete eigenvalues for the defocusing nonlinear Schrödinger equation. Studies in Applied Mathematics, 132(2):138–159, February 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12024.

Bihlo:2020:ZOC

[BP20] Alexander Bihlo and Roman O. Popovych. Zeroth-order conservation laws of two-dimensional shallow water equations with variable bottom topography. Studies in Applied Mathematics, 145(2):291–321, August 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12320.

Barouch:1973:VWI

[BPS73] E. Barouch, J. W. Perram, and E. R. Smith. The van der Waals interaction across non-uniform electrolyte. Studies in Applied Mathematics, 52(2):175–188, June 1973. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1973522175.

Benney:1969:WI

[BR69] D. J. Benney and G. J. Roskes. Wave instabilities. Studies in Applied Mathematics, 48(4):377–385, December 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1969484377.

Barouch:1973:ICR

[BR73] E. Barouch and S. R. Rogers. Ising chain with random magnetic moments. Studies in Applied Mathematics, 52(2):163– 173, June 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1973522163.

Bridges:2019:KSW

[BR19] Thomas J. Bridges and Daniel J. Ratliff. Krein signature and Whitham modulation theory: the sign of characteristics and the “sign characteristic”. Studies in Applied Mathematics, 142(3):314–335, April 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12256.

Bridges:1990:IHM

[Bri90] Thomas J. Bridges. O(2)-invariant Hamiltonians on C^{4} and the (m, n) mode-interaction problem for capillary-gravity waves. Studies in Applied Mathematics, 82(2):93–120, February 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm199082293.

Bridges:1994:GAD

[Bri94] Thomas J. Bridges. Geometric aspects of degenerate modulation equations. Studies in Applied Mathematics, 91 (2):125–151, February 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1994912125.

Bridges:2015:BWM

[Bri15] Thomas J. Bridges. Breakdown of the Whitham modulation theory and the emergence of dispersion. Studies in Applied Mathematics, 135(3):277–294, October 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12086.

Bonetti:1992:FCT

[BRSV92] Flavio Bonetti, Gian-Carlo Rota, Domenico Senato, and Antonietta M. Venezia. On the foundation of combinatorial theory. X. A categorical setting for symmetric functions. Studies in Applied Mathematics, 86(1):1–29, January 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19928611.

Beals:1989:BTI

[BRT89] Richard Beals, Mauro Rabelo, and Keti Tenenblat. Bäcklund transformations and inverse scattering solutions for some pseudospherical surface equations. Studies in Applied Mathematics, 81(2):125–151, October 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1989812125.

Brylawski:1975:ART

[Bry75] Thomas H. Brylawski. An affine representation for transversal geometries. Studies in Applied Mathematics, 54(2):143–160, June 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1975542143.

Bryant:1977:PWS

[Bry77] P. J. Bryant. Permanent wave structures on an open two layer fluid. Studies in Applied Mathematics, 57(3):225–246, December 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1977573225.

Bryant:1979:NWG

[Bry79a] P. J. Bryant. Nonlinear wave groups in deep water. Studies in Applied Mathematics, 61(1):1–30, July 1979. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19796111.

Brylawski:1979:ITE

[Bry79b] Tom Brylawski. Intersection theory for embeddings of matroids into uniform geometries. Studies in Applied Mathematics, 61(3):211–244, December 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1979613211.

Brylawski:1991:GFL

[Bry91] Thomas Brylawski. Greedy families for linear objective functions. Studies in Applied Mathematics, 84(3):221–229, April 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991843221.

Brown:1977:TCL

[BS77] S. N. Brown and K. Stewartson. On thermal convection in a large box. Studies in Applied Mathematics, 57(3):187–204, December 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1977573187.

Byatt-Smith:1988:SSO

[BS88] J. G. Byatt-Smith. On the solutions of a second order differential equation arising in the theory of resonant oscillations in a tank. Studies in Applied Mathematics, 79(2):143–157, October 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1988792143.

Byatt-Smith:1989:ASC

[BS89] J. G. B. Byatt-Smith. The asymptotic solution of a connection problem of a second order ordinary differential equation. Studies in Applied Mathematics, 80(2):109–135, April 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1989802109.

Byatt-Smith:1993:SOD

[BS93] J. G. B. Byatt-Smith. On solutions of ordinary differential equations arising from a model of crystal growth. Studies in Applied Mathematics, 89(2):167–187, July 1993. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1993892167.

Byatt-Smith:1999:BTU

[BS99] J. G. Byatt-Smith. The Borel transform and its use in the summation of asymptotic expansions. Studies in Applied Mathematics, 103(4):339–369, November 1999. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.1034136.

Byatt-Smith:2000:BTU

[BS00] J. G. Byatt-Smith. The Borel transform and its use in the summation of asymptotic expansions. Studies in Applied Mathematics, 105(2):83–113, August 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00143.

Byatt-Smith:2001:NSN

[BS01] J. G. Byatt-Smith. Numerical solution of Nekrasov’s equation in the boundary layer near the crest for waves near the maximum height. Studies in Applied Mathematics, 106(4):393–405, May 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00171.

Budd:2016:MAS

[BS16] Christopher J. Budd and John M. Stockie. Multilayer asymptotic solution for wetting fronts in porous media with exponential moisture diffusivity. Studies in Applied Mathematics, 136(4):424–458, May 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12111.

Barthwal:2022:ERS

[BS22] Rahul Barthwal and T. Raja Sekhar. On the existence and regularity of solutions of semihyperbolic patches to 2-D Euler equations with van der Waals gas. Studies in Applied Mathematics, 148(2):543–576, February 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12449.

Barthwal:2023:ESG

[BS23] R. Barthwal and T. Raja Sekhar. Existence of solutions to gas expansion problem through a sharp corner for 2-D Euler equations with general equation of state. Studies in Applied Mathematics, 151(1):141–170, July 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12575.

Banerjee:2000:EBO

[BSCS00] Mihir B. Banerjee, R. G. Shandil, S. S. Chauhan, and Daleep Sharma. Eigenvalue bounds for the Orr–Sommerfeld equation and their relevance to the existence of backward wave motion — II. Studies in Applied Mathematics, 105(1):31–34, July 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00140.

Banerjee:1999:EBO

[BSGC99] Mihir B. Banerjee, R. G. Shandil, M. G. Gourla, and S. S. Chauhan. Eigenvalue bounds for the Orr–Sommerfeld equation and their relevance to the existence of backward wave motion. Studies in Applied Mathematics, 103(1):43–50, July 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00119.

Banerjee:2000:ITC

[BSSS00] Mihir B. Banerjee, R. G. Shandil, K. S. Shirkot, and Daleep Sharma. Importance of Tollmien’s counter example. Studies in Applied Mathematics, 105(3):191–202, October 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00148.

Betermin:2023:TDL

[BŠT23] Laurent Bétermin, Ladislav Šamaj, and Igor Travěnec. Threedimensional lattice ground states for Riesz and Lennard-Jones-type energies. Studies in Applied Mathematics, 150(1):69–91, January 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12533.

Benney:1980:RMV

[BT80] D. J. Benney and W. J. Timson. The rolling motion of a viscous fluid on and off a rigid surface. Studies in Applied Mathematics, 63(2):93–98, October 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198063293.

Beals:1988:ISB

[BT88] Richard Beals and Keti Tenenblat. Inverse scattering and the Bäcklund transformation for the generalized wave and generalized sine-Gordon equations. Studies in Applied Mathematics, 78(3):227–256, June 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1988783227.

Benson:2019:IDF

[BV19] Joseph Benson and Francis Valiquette. Invariant discrete flows. Studies in Applied Mathematics, 143(1):81–119, July 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12270.

Beyer:1979:MAT

[BW79] W. A. Beyer and M. S. Waterman. Moments of absorption time for a conditioned random walk. Studies in Applied Mathematics, 60(1):83–90, February 1979. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197960183.

Babarsky:1986:EPC

[BW86] Richard Babarsky and Houston G. Wood. On the eigenvalue problem for a certain. Studies in Applied Mathematics, 75(3):249–264, December 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1986753249.

Bona:2002:ZVL

[BW02] Jerry L. Bona and Jiahong Wu. The zero-viscosity limit of the 2D Navier–Stokes equations. Studies in Applied Mathematics, 109(4):265–278, November 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
t01-1-00223.

Biondini:2019:DCC

[BW19] Gino Biondini and Qiao Wang. Discrete and continuous coupled nonlinear integrable systems via the dressing method. Studies in Applied Mathematics, 142(2):139–161, February 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12239.

Xiang:2023:SCL

[bXlZS23] Xiao bo Xiang, Song lin Zhao, and Ying Shi. Solutions and continuum limits to nonlocal discrete sine-Gordon equations: Bilinearization reduction method. Studies in Applied Mathematics, 150(4):1274–1303, May 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12569.

Benney:1989:ITD

[BZ89] D. J. Benney and Y. F. Zhou. Instability of three dimensional waves in shear flows. Studies in Applied Mathematics, 81(1):41–55, August 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm198981141.

Benney:1992:NIC

[BZ92] D. J. Benney and Y. F. Zhou. The nonlinear instability characteristics of waves in free shear layers. Studies in Applied Mathematics, 86(1):53–78, January 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199286153.

Borisov:2021:BTE

[BZZ21] D. I. Borisov, D. A. Zezyulin, and M. Znojil. Bifurcations of thresholds in essential spectra of elliptic operators under localized non-Hermitian perturbations. Studies in Applied Mathematics, 146(4):834–880, May 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12367.

Cho:2009:FWN

[CA09] Yeunwoo Cho and T. R. Akylas. Forced waves near resonance at a phase-speed minimum. Studies in Applied Mathematics, 123(1):1–15, July 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2009.00444.x.

Chakravarty:2010:PCE

[CA10] Sarbarish Chakravarty and Mark J. Ablowitz. Parameterizations of the chazy equation. Studies in Applied Mathematics, 124(2):105–135, February 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2009.00463.x.

Caillol:2022:VWPa

[Cai22a] Philippe Caillol. A vorticity wave packet breaking within a rapidly rotating vortex. Part I: The critical layer flow. Studies in Applied Mathematics, 148(2):825–864, February 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12462.

Caillol:2022:VWPb

[Cai22b] Philippe Caillol. A vorticity wave packet breaking within a rapidly rotating vortex. Part II: wave/mean flow interaction. Studies in Applied Mathematics, 148(2):865–917, February 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12461.

Calogero:1984:SNW

[Cal84] F. Calogero. A solvable nonlinear wave equation. Studies in Applied Mathematics, 70(3):189–199, June 1984. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1984703189.

Calogero:2016:NSV

[Cal16] Francesco Calogero. New solvable variants of the goldfish many-body problem. Studies in Applied Mathematics, 137(1):123–139, July 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12096.

Campbell:2004:WMF

[Cam04] L. J. Campbell. Wave–mean-flow interactions in a forced Rossby wave packet critical layer. Studies in Applied Mathematics, 112(1):39–85, January 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2004.01587.x.

Canfield:1978:LMS

[Can78] E. Rodney Canfield. On the location of the maximum Stirling number(s) of the second kind. Studies in Applied Mathematics, 59(1):83–93, July 1978. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197859183.

Cao:2013:NIC

[Cao13] Jian Cao. A note on q-integrals and certain generating functions. Studies in Applied Mathematics, 131(2):105–118, August 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12002.

Carnielli:1990:HRD

[Car90] Walter A. Carnielli. Hyper-rook domain inequalities. Studies in Applied Mathematics, 82(1):59–69, January 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199082159.

Caudrelier:2018:IBI

[Cau18] Vincent Caudrelier. Interplay between the inverse scattering method and Fokas’s unified transform with an application. Studies in Applied Mathematics, 140(1):3–26, January 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12190.

Chitlaru-Briggs:1987:TDI

[CBB87] Sanda Chitlaru-Briggs and Eytan Barouch. A two-dimensional Ising model in a magnetic field — a scalar representation of the partition function. Studies in Applied Mathematics, 77(1):89– 100, August 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm198777189.

Carreno-Bolanos:2020:PPN

[CBJCN20] Rafael Carreño-Bolaños, Beatriz Juarez-Campos, and Pavel I. Naumkin. Periodic problem for the nonlinear damped wave equation with convective nonlinearity. Studies in Applied Mathematics, 145(1):137–149, July 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12316.

Colbrook:2020:KDE

[CBKM20] Matthew J. Colbrook, Zdravko I. Botev, Karsten Kuritz, and Shev MacNamara. Kernel density estimation with linked boundary conditions. Studies in Applied Mathematics, 145(3):357–396, October 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12322.

Chan:2017:RWA

[CC17] H. N. Chan and K. W. Chow. Rogue waves for an alternative system of coupled Hirota equations: Structural robustness and modulation instabilities. Studies in Applied Mathematics, 139(1):78–103, July 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12170.

Chabane:2019:RIG

[CC19] Malik Chabane and Wooyoung Choi. On resonant interactions of gravity-capillary waves without energy exchange. Studies in Applied Mathematics, 142(4):528–550, May 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12249.

Caudrelier:2022:NMI

[CCD22] Vincent Caudrelier, Nicolas Crampé, and Carlos Mbala Dibaya. Nonlinear mirror image method for nonlinear Schrödinger equation: Absorption/emission of one soliton by a boundary. Studies in Applied Mathematics, 148(2):715–757, February 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12456.

Chen:2019:RHA

[CCF19] Min Chen, Yang Chen, and En-Gui Fan. The Riemann– Hilbert analysis to the Pollaczek–Jacobi type orthogonal polynomials. Studies in Applied Mathematics, 143(1):42–80, July 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12259.

Chen:2021:ESC

[CCK^{+}21] Huiqin Chen, Yaqiong Cui, Shugui Kang, Youmin Lu, and Wenying Feng. Existence of solutions for a class of fractional difference equations at resonance. Studies in Applied Mathematics, 146(4):881–900, May 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12368.

Cassatella-Contra:2012:RHP

[CCM12] G. A. Cassatella-Contra and M. Mañas. Riemann–Hilbert problems, matrix orthogonal polynomials and discrete matrix equations with singularity confinement. Studies in Applied Mathematics, 128(3):252–274, April 2012. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00541.x.

Cassatella-Contra:2019:RHP

[CCM19] Giovanni A. Cassatella-Contra and Manuel Mañas. Riemann– Hilbert problem and matrix discrete Painlevé II systems. Studies in Applied Mathematics, 143(3):272–314, October 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12277.

Cao:2021:RWL

[CCMH21] Yulei Cao, Yi Cheng, Boris A. Malomed, and Jingsong He. Rogue waves and lumps on the nonzero background in the PT -symmetric nonlocal Maccari system. Studies in Applied Mathematics, 147(2):694–723, August 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12396.

Chen:2022:CNP

[CCMX22] Chuanqiang Chen, Li Chen, Xinqun Mei, and Ni Xiang. The classical Neumann problem for a class of mixed Hessian equations. Studies in Applied Mathematics, 148(1):5–26, January 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12429.

Caraballo:2023:RDL

[CCY23] Tomás Caraballo, Zhang Chen, and Dandan Yang. Random dynamics and limiting behaviors for 3D globally modified Navier–Stokes equations driven by colored noise. Studies in Applied Mathematics, 151(1):247–284, July 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12579.

Criminale:1990:ELP

[CD90] W. O. Criminale and P. G. Drazin. The evolution of linearized perturbations of parallel flows. Studies in Applied Mathematics, 83(2):123–157, August 1990. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990832123.

Calogero:2004:NIE

[CD04] F. Calogero and A. Degasperis. New integrable equations of nonlinear Schrödinger type. Studies in Applied Mathematics, 113(1):91–137, July 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2004.01522.x.

Claeys:2018:GFA

[CD18a] Tom Claeys and Antoine Doeraene. The generating function for the Airy point process and a system of coupled Painlevé II equations. Studies in Applied Mathematics, 140(4):403–437, May 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12209.

Clarkson:2018:POP

[CD18b] Peter A. Clarkson and Adri B. Olde Daalhuis. Preface: Orthogonal polynomials, special functions, and applications. Studies in Applied Mathematics, 141(4):421–423, November 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12244.

Carter:2021:PNW

[CD21a] John D. Carter and Bernard Deconinck. Preface: Nonlinear waves in honor of Harvey Segur on the occasion of his 80th birthday. Studies in Applied Mathematics, 147(4):1199–1208, November 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12458.

Cisneros:2021:NSL

[CD21b] Jorge Cisneros and Bernard Deconinck. The numerical solutions of linear semidiscrete evolution problems on the halfline using the unified transform method. Studies in Applied Mathematics, 147(4):1240–1276, November 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12464.

Camassa:2022:EIS

[CDF^{+}22] R. Camassa, R. D’Onofrio, G. Falqui, G. Ortenzi, and M. Pedroni. Evolution of interface singularities in shallow water equations with variable bottom topography. Studies in Applied Mathematics, 148(4):1439–1476, May 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12480.

Chen:2018:SNE

[CDLjZ18] Kui Chen, Xiao Deng, Senyue Lou, and Da jun Zhang. Solutions of nonlocal equations reduced from the AKNS hierarchy. Studies in Applied Mathematics, 141(1):113–141, July 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12215.

Cook:1973:RBV

[CE73] L. Pamela Cook and W. Eckhaus. Resonance in a boundary value problem of singular perturbation type. Studies in Applied Mathematics, 52(2):129–139, June 1973. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1973522129.

Congy:2019:NSE

[CEHS19] T. Congy, G. A. El, M. A. Hoefer, and M. Shearer. Nonlinear Schrödinger equations and the universal description of dispersive shock wave structure. Studies in Applied Mathematics, 142(3):241–268, April 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12247.

Congy:2021:DRP

[CEHS21] T. Congy, G. A. El, M. A. Hoefer, and M. Shearer. Dispersive Riemann problems for the Benjamin–Bona–Mahony equation. Studies in Applied Mathematics, 147(3):1089–1145, October 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12426.

Chen:2022:RHA

[CF22] Meisen Chen and Engui Fan. Riemann–Hilbert approach for discrete sine-Gordon equation with simple and double poles. Studies in Applied Mathematics, 148(3):1180–1207, April 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12472.

Chen:2023:TFF

[CF23] Junchao Chen and Bao-Feng Feng. Tau-function formulation for bright, dark soliton and breather solutions to the massive Thirring model. Studies in Applied Mathematics, 150(1):35– 68, January 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12532.

Chen:2018:DYO

[CFiMO18] Junchao Chen, Bao-Feng Feng, Ken ichi Maruno, and Yasuhiro Ohta. The derivative Yajima–Oikawa system: Bright, dark soliton and breather solutions. Studies in Applied Mathematics, 141(2):145–185, August 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12216.

Chen:1992:SGF

[CFS92] Ming-Po Chen, Chia-Chin Feng, and H. M. Srivastava. Some generating functions for the generalized Bessel polynomials. Studies in Applied Mathematics, 87(4):351–366, November 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1992874351.

Campbell:1987:ESF

[CFW87] John A. Campbell, Per Olof Fröman, and Erik Walles. Explicit series formulae for the evaluation of integrals by the method of steepest descents. Studies in Applied Mathematics, 77(2):151–172, October 1987. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1987772151.

Chow:2017:ASD

[CFY17] KwokWing Chow, EnGui Fan, and ManWai Yuen. The analytical solutions for the N–dimensional damped compressible Euler equations. Studies in Applied Mathematics, 138(3):294–316, April 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12154.

Caillol:2007:RSW

[CG07] P. Caillol and R. H. Grimshaw. Rossby solitary waves in the presence of a critical layer. Studies in Applied Mathematics, 118(4):313–364, May 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2007.00368.x.

Caillol:2008:REW

[CG08] P. Caillol and R. H. Grimshaw. Rossby elevation waves in the presence of a critical layer. Studies in Applied Mathematics, 120(1):35–64, January 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2007.00392.x.

Conte:2016:RGC

[CG16] Robert Conte and A. Michel Grundland. Reductions of Gauss–Codazzi equations. Studies in Applied Mathematics, 137(3):306–327, October 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12121.

Caudrelier:2023:SIY

[CGP23] Vincent Caudrelier, Aikaterini Gkogkou, and Barbara Prinari. Soliton interactions and Yang–Baxter maps for the complex coupled short-pulse equation. Studies in Applied Mathematics, 151(1):285–351, July 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12580.

Clarkson:2020:CMD

[CGUGM20] Peter A. Clarkson, David Gómez-Ullate, Yves Grandati, and Robert Milson. Cyclic Maya diagrams and rational solutions of higher order Painlevé systems. Studies in Applied Mathematics, 144(3):357–385, April 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12300.

Collowald:2018:ACC

[CH18] Mathieu Collowald and Evelyne Hubert. Algorithms for computing cubatures based on moment theory. Studies in Applied Mathematics, 141(4):501–546, November 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12240.

Chan:1971:IPN

[Cha71] Shu-Kwan Chan. Infinite Prandtl number turbulent convection. Studies in Applied Mathematics, 50(1):13–49, March 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197150113.

Charlier:2021:DPL

[Cha21] Christophe Charlier. Doubly periodic lozenge tilings of a hexagon and matrix valued orthogonal polynomials. Studies in Applied Mathematics, 146(1):3–80, January 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12339.

Cheng:1969:TCA

[Che69] Hung Cheng. Theory of complex angular momentum: Past and future. Studies in Applied Mathematics, 48(4):341–349, December 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1969484341.

Cheng:1984:ESN

[Che84] Hung Cheng. An exactly solvable nonlinear partial differential equation with solitary-wave solutions. Studies in Applied Mathematics, 70(3):183–187, June 1984. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1984703183.

Chen:1994:MVG

[Che94] Beifang Chen. The mixed volumes and geissinger multiplications of convex sets. Studies in Applied Mathematics, 91(1):39– 50, January 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm199491139.

Cheng:2004:RTS

[Che04] Hung Cheng. The renormalized two-scale method. Studies in Applied Mathematics, 113(4):381–387, November 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2004.01538.x.

Chen:2011:ANT

[Che11] Jinbing Chen. The application of Neumann type systems for solving integrable nonlinear evolution equations. Studies in Applied Mathematics, 127(2):153–190, August 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00516.x.

Chow:1990:NIS

[Cho90] K. W. Chow. A note on inviscid secondary instability in shear flows. Studies in Applied Mathematics, 83(3):183–192, October 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990833183.

Chomaz:1991:FSC

[CHR91] Jean-Marc Chomaz, Patrick Huerre, and Larry G. Redekopp. A frequency selection criterion in spatially developing flows. Studies in Applied Mathematics, 84(2):119–144, February 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991842119.

Crisan:2021:WCI

[CHS21] Dan Crisan, Darryl D. Holm, and Oliver D. Street. Wave– current interaction on a free surface. Studies in Applied Mathematics, 147(4):1277–1338, November 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12425.

Carrillo:2022:CBS

[CHSV22] J. A. Carrillo, F. Hoffmann, A. M. Stuart, and U. Vaes. Consensus-based sampling. Studies in Applied Mathematics, 148(3):1069–1140, April 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12470.

Chen:1998:HME

[CI98] Yang Chen and Mourad E. H. Ismail. Hermitean matrix ensembles and orthogonal polynomials. Studies in Applied Mathematics, 100(1):33–52, January 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00068.

Constantin:2017:DMD

[CI17] Adrian Constantin and Rossen Ivanov. Dressing method for the Degasperis–Procesi equation. Studies in Applied Mathematics, 138(2):205–226, February 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12149.

Colin:2020:SWS

[CI20] Mathieu Colin and Tatsuo Iguchi. Solitary wave solutions to the Isobe–Kakinuma model for water waves. Studies in Applied Mathematics, 145(1):52–80, July 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12310.

Carles:2023:GSS

[CI23] Rémi Carles and Yavdat Il’yasov. On ground states for the 2D Schrödinger equation with combined nonlinearities and harmonic potential. Studies in Applied Mathematics, 150(1):92–118, January 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12534.

Claassen:2018:NBS

[CJ18] Kyle M. Claassen and Mathew A. Johnson. Numerical bifurcation and spectral stability of wavetrains in bidirectional Whitham models. Studies in Applied Mathematics, 141(2):205–246, August 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12221.

Clarkson:2016:GFW

[CJK16] Peter A. Clarkson, Kerstin Jordaan, and Abey Kelil. A generalized Freud weight. Studies in Applied Mathematics, 136(3):288–320, April 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12105.

Chen:1992:BSI

[CJM92] B. Chen, M. C. Jorge, and A. A. Minzoni. Bifurcation of solutions for an inverse problem in potential theory. Studies in Applied Mathematics, 86(1):31–51, January 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199286131.

Criminale:2004:LCS

[CJN04] W. O. Criminale, T. L. Jackson, and P. W. Nelson. Limit cycle-strange attractor competition. Studies in Applied Mathematics, 112(2):133–160, February 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
0022-2526.2004.01507.x.

Claus:1975:HMS

[CK75] Armin Claus and Daniel J. Kleitman. Heuristic methods for solving large scale network routing problems: The telpaking problem. Studies in Applied Mathematics, 54(1):17–29, March 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197554117.

Cox:2004:LRN

[CK04] S. M. Cox and J. R. King. Large-Reynolds-number asymptotics of the Berman Problem. Studies in Applied Mathematics, 113(3):217–243, October 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
0022-2526.2004.01512.x.

Chakravarty:2009:SSK

[CK09] S. Chakravarty and Y. Kodama. Soliton solutions of the KP equation and application to shallow water waves. Studies in Applied Mathematics, 123(1):83–151, July 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00448.x.

Chesnokov:2013:LDP

[CK13] A. A. Chesnokov and A. K. Khe. Is Landau damping possible in a shear fluid flow? Studies in Applied Mathematics, 131(4):343–358, November 2013. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12013.

Chesnokov:2014:WDS

[CK14] A. A. Chesnokov and P. V. Kovtunenko. Weak discontinuities in solutions of long-wave equations for viscous flow. Studies in Applied Mathematics, 132(1):50–64, January 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12019.

Camassa:2016:SWP

[CKL16] Roberto Camassa, Dongyang Kuang, and Long Lee. Solitary waves and N–particle algorithms for a class of Euler–Poincaré equations. Studies in Applied Mathematics, 137(4):502–546, November 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12132.

Cowen:1996:EFG

[CKLS96] L. J. Cowen, D. J. Kleitman, F. Lasaga, and D. E. Sussman. Enumeration of full graphs: Onset of the asymptotic region. Studies in Applied Mathematics, 96(3):339–350, April 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996963339.

Chatziafratis:2023:BBS

[CKS23] Andreas Chatziafratis, Spyridon Kamvissis, and Ioannis G. Stratis. Boundary behavior of the solution to the linear Korteweg–de Vries equation on the half line. Studies in Applied Mathematics, 150(2):339–379, February 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12542.

Constantineau:1991:CSK

[CL91] Ivan Constantineau and Jacques Labelle. On combinatorial structures kept fixed by the action of a given permutation. Studies in Applied Mathematics, 84(2):105–118, February 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991842105.

Claus:1973:HIP

[Cla73] A. Claus. Hyperspheric integer programming. Studies in Applied Mathematics, 52(2):153–162, June 1973. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1973522153.

Claus:1978:SEA

[Cla78] A. Claus. A simultaneous enumeration approach to the traveling salesman problem. Studies in Applied Mathematics, 58(2):159–163, February 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1978582159.

Clarkson:2009:VP

[Cla09] P. A. Clarkson. Vortices and polynomials. Studies in Applied Mathematics, 123(1):37–62, July 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2009.00446.x.

Clarkson:2016:ASS

[Cla16] Peter A. Clarkson. On Airy solutions of the second Painlevé equation. Studies in Applied Mathematics, 137(1):93–109, July 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12123.

Criminale:1997:VPD

[CLJ97] W. O. Criminale, D. G. Lasseigne, and T. L. Jackson. Vortex perturbation dynamics. Studies in Applied Mathematics, 98(2):99–120, February 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00042.

Clarkson:1986:SWI

[CLS86] Peter A. Clarkson, Randall J. LeVeque, and Ralph Saxton. Solitary-wave interactions in elastic rods. Studies in Applied Mathematics, 75(2):95–121, October 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198675295.

Casati:2020:TCA

[CLV20] Matteo Casati, Paolo Lorenzoni, and Raffaele Vitolo. Three computational approaches to weakly nonlocal Poisson brackets. Studies in Applied Mathematics, 144(4):412–448, May 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12302.

Chen:2022:IAT

[CLW22] Shanshan Chen, Jie Liu, and Yixiang Wu. Invasion analysis of a two-species Lotka–Volterra competition model in an advective patchy environment. Studies in Applied Mathematics, 149(3):762–797, October 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12520.

Chen:2023:MSE

[CLW23] Tiantian Chen, Hongxia Lin, and Jiahong Wu. The 2D magnetohydrodynamic system with Euler-like velocity equation and partial magnetic diffusion. Studies in Applied Mathematics, 150(3):629–649, April 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12551.

Chen:1995:CTS

[CLY95] William Y. C. Chen, Ko-Wei Lih, and Yeong-Nan Yeh. Cyclic tableaux and symmetric functions. Studies in Applied Mathematics, 94(3):327–339, April 1995. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995943327.

Criminale:1991:GTD

[CLZ91] W. O. Criminale, Bruce Long, and Mei Zhu. General threedimensional disturbances to inviscid Couette flow. Studies in Applied Mathematics, 85(3):249–267, October 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991853249.

Choi:2003:LNW

[CM03] Jongho Choi and Paul A. Milewski. Long nonlinear waves in resonance with topography. Studies in Applied Mathematics, 110(1):21–47, January 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00229.

Caillol:2007:SVN

[CM07] P. Caillol and S. A. Maslowe. The small vorticity nonlinear critical layer for Kelvin modes on a vortex. Studies in Applied Mathematics, 118(3):221–254, April 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2007.00372.x.

Cisneros:2008:ASS

[CM08] L. A. Cisneros and A. A. Minzoni. Asymptotics for supersonic soliton propagation in the Toda lattice equation. Studies in Applied Mathematics, 120(4):333–349, May 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00401.x.

Conte:2009:EGA

[CM09] R. Conte and M. Musette. Elliptic general analytic solutions. Studies in Applied Mathematics, 123(1):63–81, July 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00447.x.

Chen:2013:GSG

[CM13] Guanwei Chen and Shiwang Ma. Ground state and geometrically distinct solitons of discrete nonlinear Schrödinger equations with saturable nonlinearities. Studies in Applied Mathematics, 131(4):389–413, November 2013. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12016.

Clarke:2021:RJW

[CM21] W. A. Clarke and R. Marangell. Rigorous justification of the Whitham modulation theory for equations of NLS type. Studies in Applied Mathematics, 147(2):577–621, August 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12390.

Cuevas–Maraver:2016:NIM

[CMKP16] Jesús Cuevas-Maraver, Panayotis G. Kevrekidis, and Dmitry E. Pelinovsky. Nonlinear instabilities of multi-site breathers in Klein–Gordon lattices. Studies in Applied Mathematics, 137(2):214–237, August 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12107.

Chumakova:2009:SPN

[CMM^{+}09] L. Chumakova, F. E. Menzaque, P. A. Milewski, R. R. Rosales, E. G. Tabak, and C. V. Turner. Stability properties and nonlinear mappings of two and three-layer stratified flows. Studies in Applied Mathematics, 122(2):123–137, February 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00426.x.

Clarke:2022:RJW

[CMP22] William A. Clarke, Robert Marangell, and Wesley R. Perkins. Rigorous justification of the Whitham modulation equations for equations of Whitham type. Studies in Applied Mathematics, 149(2):297–323, August 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12500.

Conte:2012:EMS

[CNW12] Robert Conte, Tuen-Wai Ng, and Kwok-Kin Wong. Exact meromorphic stationary solutions of the real cubic Swift–Hohenberg equation. Studies in Applied Mathematics, 129(1):117–131, July 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2012.00546.x.

Cohen:1991:MIR

[Coh91] Joel E. Cohen. Möbius inversion of random acyclic directed graphs. Studies in Applied Mathematics, 84(1):1–6, January 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19918411.

Constantineau:1991:GPS

[Con91] Ivan Constantineau. A generalization of the Pfaff–Saalschütz theorem. Studies in Applied Mathematics, 85(3):243–248, October 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1991853243.

Cosgrove:1993:PCAa

[Cos93a] Christopher M. Cosgrove. Painlevé classification of all semi linear partial differential equations of the second order. I. Hyperbolic equations in two independent variables. Studies in Applied Mathematics, 89(1):1–61, May 1993. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19938911.

Cosgrove:1993:PCAb

[Cos93b] Christopher M. Cosgrove. Painlevé classification of all semilinear partial differential equations of the second order. II. Parabolic and higher dimensional equations. Studies in Applied Mathematics, 89(2):95–151, July 1993. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199389295.

Cosgrove:1997:PCP

[Cos97] Christopher M. Cosgrove. Painlevé classification problems featuring essential singularities. Studies in Applied Mathematics, 98(4):355–433, May 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00053.

Cosgrove:2000:CCI

[Cos00a] Christopher M. Cosgrove. Chazy classes IX–XI of thirdorder differential equations. Studies in Applied Mathematics, 104(3):171–228, April 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00134.

Cosgrove:2000:HOP

[Cos00b] Christopher M. Cosgrove. Higher-order Painlevé equations in the polynomial class I. Bureau symbol P2. Studies in Applied Mathematics, 104(1):1–65, January 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00130.

Cosgrove:2006:HOP

[Cos06] Christopher M. Cosgrove. Higher-order Painlevé equations in the polynomial class II: Bureau symbol P 1. Studies in Applied Mathematics, 116(4):321–413, May 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2006.00346.x.

Clarkson:1998:SWW

[CP98] Peter A. Clarkson and Thomas J. Priestley. Shallow water wave systems. Studies in Applied Mathematics, 101 (4):389–432, November 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00099.

Coles:2012:LEB

[CP12] Matt Coles and Dmitry Pelinovsky. Loops of energy bands for Bloch waves in optical lattices. Studies in Applied Mathematics, 128(3):300–336, April 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2011.00536.x.

Crapo:1970:FCT

[CR70] Henry H. Crapo and Gian-Carlo Rota. On the foundations of combinatorial theory II. Combinatorial geometries. Studies in Applied Mathematics, 49(2):109–133, June 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1970492109.

Cehelsky:1986:RIW

[CR86] Priscilla Cehelsky and Rodolfo R. Rosales. Resonantly interacting weakly nonlinear hyperbolic waves in the presence of shocks: a single space variable in a homogeneous, time independent medium. Studies in Applied Mathematics, 74(2):117– 138, April 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1986742117.

Chiang:2006:NOM

[CR06] Yik-Man Chiang and Simon N. M. Ruijsenaars. On the Nevanlinna order of meromorphic solutions to linear analytic difference equations. Studies in Applied Mathematics, 116(3):257–287, April 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2006.00343.x.

Colombo:2021:WPC

[CR21] Rinaldo M. Colombo and Elena Rossi. Well-posedness and control in a hyperbolic–parabolic parasitoid–parasite system. Studies in Applied Mathematics, 147(3):839–871, October 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12402.

Crow:1970:ASE

[Cro70] S. C. Crow. Aerodynamic sound emission as a singular perturbation problem. Studies in Applied Mathematics, 49(1):21–44, March 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm197049121.

Crowdy:1998:NLS

[Cro98] Darren G. Crowdy. A note on the linear stability of Burgers vortex. Studies in Applied Mathematics, 100(2):107–126, February 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9590.00072.

Crowdy:2000:NAF

[Cro00] Darren G. Crowdy. A new approach to free surface Euler flows with capillarity. Studies in Applied Mathematics, 105(1):35–58, July 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00141.

Chaabi:2016:BVP

[CRW16] S. Chaabi, S. Rigat, and F. Wielonsky. A boundary value problem for conjugate conductivity equations. Studies in Applied Mathematics, 137(3):328–355, October 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12124.

Chen:1980:SGC

[CS80a] B. Chen and P. G. Saffman. Steady gravity-capillary waves on deep water — II. Numerical results for finite amplitude. Studies in Applied Mathematics, 62(2):95–111, April 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198062295.

Chen:1980:NEE

[CS80b] B. Chen and P. G. Sallman. Numerical evidence for the existence of new types of gravity waves of permanent form on deep water. Studies in Applied Mathematics, 62(1):1–21, February 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19806211.

Clapp:1980:BTD

[CS80c] M. H. Clapp and R. C. Shiflett. A Birkhoff theorem for doubly stochastic matrices with vector entries. Studies in Applied Mathematics, 62(3):273–279, June 1980. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980623273.

Chen:1985:TDS

[CS85] B. Chen and P. G. Saffman. Three-dimensional stability and bifurcation of capillary and gravity waves on deep water. Studies in Applied Mathematics, 72(2):125–147, April 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1985722125.

Chatterjea:1990:ORC

[CS90] S. K. Chatterjea and H. M. Srivastava. Operational representations for the classical Hermite polynomials. Studies in Applied Mathematics, 83(4):319–326, November 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990834319.

Cosgrove:1993:PCC

[CS93] Christoper M. Cosgrove and George Scoufis. Painlevé classification of a class of differential equations of the second order and second degree. Studies in Applied Mathematics, 88(1):25– 87, January 1993. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm199388125.

Clamond:2012:SGW

[CS12] D. Clamond and Y. Stepanyants. Stationary gravity waves with the zero mean vorticity in stratified fluid. Studies in Applied Mathematics, 128(1):59–85, January 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00530.x.

Curtis:2015:TDS

[CS15] Christopher W. Curtis and Samuel S. P. Shen. Threedimensional surface water waves governed by the forced Benney–Luke equation. Studies in Applied Mathematics, 135(4):447–465, November 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12091.

Cain:1977:GDP

[CSS77] B. E. Cain, B. D. Saunders, and Hans Schneider. On the geometry of dual pairs. Studies in Applied Mathematics, 56(1):71– 79, January 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm197756171.

Chern:1986:PSE

[CT86] S. S. Chern and K. Tenenblat. Pseudospherical surfaces and evolution equations. Studies in Applied Mathematics, 74(1):55–83, February 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm198674155.

Chugunova:2009:SAO

[CV09] Marina Chugunova and Hans Volkmer. Spectral analysis of an operator arising in fluid dynamics. Studies in Applied Mathematics, 123(3):291–309, October 2009. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00455.x.

Cheng:1970:AS

[CW70] Hung Cheng and Tai Tsun Wu. An aging spring. Studies in Applied Mathematics, 49(2):183–185, June 1970. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1970492183.

Cohen:1996:IST

[CW96] Donald S. Cohen and Thomas P. Witelski. Inaccessible states in time-dependent reaction diffusion. Studies in Applied Mathematics, 97(4):301–319, November 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996974301.

Chu:2021:ART

[CW21] Jifeng Chu and Ling-Jun Wang. Analyticity of rotational traveling gravity two-layer waves. Studies in Applied Mathematics, 146(3):605–634, April 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12358.

Chen:2015:EQS

[CWT^{+}15] Aiyong Chen, Shuangquan Wen, Shengqiang Tang, Wentao Huang, and Zhijun Qiao. Effects of quadratic singular curves in integrable equations. Studies in Applied Mathematics, 134(1):24–61, January 2015. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12060.

Chen:1994:SED

[CY94] Beifang Chen and Yeong-Nan Yeh. Some explanations of Dobinski’s formula. Studies in Applied Mathematics, 92(3):191–199, July 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1994923191.

Chen:2010:PVU

[CZ10] Y. Chen and L. Zhang. Painlevé VI and the unitary Jacobi ensembles. Studies in Applied Mathematics, 125(1):91–112, July 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00483.x.

Curtis:2015:DSH

[CZ15] Christopher W. Curtis and Yi Zhu. Dynamics in PT - symmetric honeycomb lattices with nonlinearity. Studies in Applied Mathematics, 135(2):139–170, August 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12081.

Chen:2020:CHS

[CZ20] Jinbing Chen and Runsu Zhang. The complex Hamiltonian systems and quasi-periodic solutions in the derivative nonlinear Schrödinger equations. Studies in Applied Mathematics, 145(2):153–178, August 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12311.

Chen:2021:DPP

[CZ21] Xianfeng Chen and Xiang Zhang. Dynamics of the predator– prey model with the sigmoid functional response. Studies in Applied Mathematics, 147(1):300–318, July 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12382.

Chen:2021:SES

[CZjZ21] Kui Chen, Cheng Zhang, and Da jun Zhang. Squared eigenfunction symmetry of the DΔ mKP hierarchy and its constraint. Studies in Applied Mathematics, 147(2):752–791, August 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12399.

Dohnal:2005:OSB

[DA05] Tomáš Dohnal and Alejandro B. Aceves. Optical soliton bullets in (2 + 1)D nonlinear Bragg resonant periodic geometries. Studies in Applied Mathematics, 115(2):209–232, August 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2005.00326.x.

Dutykh:2011:MMP

[DARB11] Denys Dutykh, Céline Acary-Robert, and Didier Bresch. Mathematical modeling of powder-snow avalanche flows. Studies in Applied Mathematics, 127(1):38–66, July 2011. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00511.x.

Daykin:1980:HI

[Day80] David E. Daykin. A hierarchy of inequalities. Studies in Applied Mathematics, 63(3):263–274, December 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980633263.

deBraganca:1975:FDB

[dB75] Sérgio Luiz de Bragança. Finite dimensional Baxter algebras. Studies in Applied Mathematics, 54(1):75–89, March 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197554175.

Dias:1990:THR

[DB90] Frédéric Dias and Thomas J. Bridges. The third-harmonic resonance for capillary-gravity waves with O(2) spatial symmetry. Studies in Applied Mathematics, 82(1):13–35, January 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199082113.

Dias:1994:GAS

[DB94] Frédéric Dias and Thomas J. Bridges. Geometric aspects of spatially periodic interfacial waves. Studies in Applied Mathematics, 93(2):93–132, October 1994. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199493293.

Denier:1996:EGV

[DB96] James P. Denier and Andrew P. Bassom. The existence of Görtler vortices in separated boundary layers. Studies in Applied Mathematics, 96(3):247–271, April 1996. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996963247.

Denier:2003:NEN

[DB03] James P. Denier and Andrew P. Bassom. The nonparallel evolution of nonlinear short waves in buoyant boundary layers. Studies in Applied Mathematics, 110(2):139–156, February 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00234.

Dassios:2011:GRD

[DD11] G. Dassios and M. Doschoris. On the global relation and the Dirichlet-to-Neumann correspondence. Studies in Applied Mathematics, 126(1):75–102, January 2011. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00498.x.

Djidjou-Demasse:2020:DAM

[DDAAB20] Ramsés Djidjou-Demasse, Gbenga J. Abiodun, Abiodun M. Adeola, and Joel O. Botai. Development and analysis of a malaria transmission mathematical model with seasonal mosquito life-history traits. Studies in Applied Mathematics, 144(4):389–411, May 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12296.

Duran:2013:GIS

[DDM13] Angel Duran, Denys Dutykh, and Dimitrios Mitsotakis. On the Galilean invariance of some nonlinear dispersive wave equations. Studies in Applied Mathematics, 131(4):359–388, November 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12015.

Das:2021:RWS

[DDM21] Arijit Das, Soumen De, and B. N. Mandal. Radiation of waves by a submerged nearly circular rough plate in ice-covered ocean. Studies in Applied Mathematics, 147(3):935–954, October 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12414.

Djidjou-Demasse:2022:DPR

[DDMMSYD22] Ramsès Djidjou-Demasse, Martin L. Mann-Manyombe, Ousmane Seydi, and Valaire Yatat-Djeumen. Differential preferences for RBCs is key for Plasmodium species evolutionary diversity within human host. Studies in Applied Mathematics, 149(4):1002–1031, November 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12527.

Damiani:1991:CC

[DDN91] E. Damiani, O. D’Antona, and G. Naldi. On the connection constants. Studies in Applied Mathematics, 85(4):289–302, November 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1991854289.

Dawson:2021:MPS

[DDN21] Sean R. Dawson, Holger R. Dullin, and Diana M. H. Nguyen. Monodromy in prolate spheroidal harmonics. Studies in Applied Mathematics, 146(4):953–982, May 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12371.

Damiani:1990:TBB

[DDNP90] E. Damiani, O. D’Antona, G. Naldi, and L. Pavarino. Tiling bricks with bricks. Studies in Applied Mathematics, 83(2):91– 110, August 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm199083291. See erratum

[DDNP91].

Damiani:1991:ETB

[DDNP91] E. Damiani, O. D’Antona, G. Naldi, and L. Pavarino. Erratum: “Tiling bricks with bricks”. Studies in Applied Mathematics, 85(3):287, October 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1991853287.

See [DDNP90].

Derzho:2011:RNV

[DdY11] Oleg G. Derzho and Brad de Young. Rotating nonlinear vortical patterns on a gamma plane. Studies in Applied Mathematics, 126(2):187–201, February 2011. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00502.x.

Deconinck:2000:CVM

[Dec00] Bernard Deconinck. Canonical variables for multiphase solutions of the KP equation. Studies in Applied Mathematics, 104(3):229–292, April 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00135.

Demina:2023:ISP

[Dem23] Maria V. Demina. Integrability and solvability of polynomial Liénard differential systems. Studies in Applied Mathematics, 150(3):755–817, April 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12556.

Deano:2021:LRC

[DER21] Alfredo Deaño, Bruno Eijsvoogel, and Pablo Román. Ladder relations for a class of matrix valued orthogonal polynomials. Studies in Applied Mathematics, 146(2):463–497, February 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12351.

Dassios:2009:TUI

[DF09] G. Dassios and A. S. Fokas. On two useful identities in the theory of ellipsoidal harmonics. Studies in Applied Mathematics, 123(4):361–373, November 2009. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00458.x.

Dimakos:2015:PBE

[DF15] M. Dimakos and A. S. Fokas. The Poisson and the biharmonic equations in the interior of a convex polygon. Studies in Applied Mathematics, 134(4):456–498, May 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12078.

Dhaouadi:2019:ELA

[DFG19] Firas Dhaouadi, Nicolas Favrie, and Sergey Gavrilyuk. Extended Lagrangian approach for the defocusing nonlinear Schrödinger equation. Studies in Applied Mathematics, 142(3):336–358, April 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12238.

Dias:2016:CNS

[DFK16] João-Paulo Dias, Mário Figueira, and Vladimir V. Konotop. Coupled nonlinear Schrödinger equations with a gauge potential: Existence and blowup. Studies in Applied Mathematics, 136(3):241–262, April 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12102.

Dias:2014:SBC

[DFKZ14] João-Paulo Dias, Mário Figueira, Vladimir V. Konotop, and Dmitry A. Zezyulin. Supercritical blowup in coupled parity-time-symmetric nonlinear Schrödinger equations. Studies in Applied Mathematics, 133(4):422–440, November 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12063.

Dubrovin:1997:TPS

[DFS97] B. A. Dubrovin, Ron Flickinger, and Harvey Segur. Threephase solutions of the Kadomtsev–Petviashvili equation. Studies in Applied Mathematics, 99(2):137–203, August 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00059.

Dzhamay:2022:DER

[DFS22] Anton Dzhamay, Galina Filipuk, and Alexander Stokes. Differential equations for the recurrence coefficients of semiclassical orthogonal polynomials and their relation to the Painlevé equations via the geometric approach. Studies in Applied Mathematics, 148(4):1656–1702, May 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12487.

Derzho:2005:RWS

[DG05] Oleg G. Derzho and Roger Grimshaw. Rossby waves on a shear flow with recirculation cores. Studies in Applied Mathematics, 115(4):387–403, November 2005. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2005.00318.x.

DaCosta:2002:HCS

[DGW02] F. P. Da Costa, M. Grinfeld, and J. A. D. Wattis. A hierarchical cluster system based on Horton–Strahler rules for river networks. Studies in Applied Mathematics, 109(3):163–204, October 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00221.

Dando:1996:WSS

[DH96a] Andrew Dando and Philip Hall. Wavenumber selection for small-wavelength gertler Vörtices in curved channel flows. Studies in Applied Mathematics, 97(3):209–234, October 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996973209.

Dwarka:1996:MEA

[DH96b] Pankaj R. Dwarka and Isom H. Herron. The modulation equations for the asymptotic suction velocity profile and the Ekman boundary layer. Studies in Applied Mathematics, 96(2):163–181, February 1996. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996962163.

Diminnie:2002:SPT

[DH02] David C. Diminnie and Richard Haberman. Slow passage through the nonhyperbolic homoclinic orbit of the saddlecenter Hamiltonian bifurcation. Studies in Applied Mathematics, 108(1):65–75, January 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
01423.

Dyachenko:2019:SWC

[DH19] Sergey A. Dyachenko and Vera Mikyoung Hur. Stokes waves with constant vorticity: I. Numerical computation. Studies in Applied Mathematics, 142(2):162–189, February 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12250.

Dhanak:1994:EMT

[Dha94] M. R. Dhanak. On the equation of motion of a thin layer of uniform vorticity. Studies in Applied Mathematics, 92(2):115–125, June 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1994922115.

Dias:2000:TTD

[DHC00] F. Dias and M. Hărăguş-Courcelle. On the transition from two-dimensional to three-dimensional water waves. Studies in Applied Mathematics, 104(2):91–127, February 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00132.

DiBucchianico:1994:RSP

[Di 94] A. Di Bucchianico. Representations of Sheffer polynomials. Studies in Applied Mathematics, 93(1):1–14, September 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19949311.

Duchene:2016:NCT

[DIT16] V. Duchêne, S. Israwi, and R. Talhouk. A new class of twolayer green–Naghdi systems with improved frequency dispersion. Studies in Applied Mathematics, 137(3):356–415, October 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12125.

Dai:2019:APF

[DIW19] Dan Dai, Mourad E. H. Ismail, and Xiang-Sheng Wang. Asymptotics of partition functions in a fermionic matrix model and of related q-polynomials. Studies in Applied Mathematics, 142(1):91–105, January 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12234.

Dominici:2005:GOA

[DK05] Diego Dominici and Charles Knessl. Geometrical optics approach to Markov-modulated fluid models. Studies in Applied Mathematics, 114(1):45–93, January 2005. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2005.01540.x.

Dai:2009:PIA

[DK09] D. Dai and A. B. J. Kuijlaars. Painlevé IV asymptotics for orthogonal polynomials with respect to a modified Laguerre weight. Studies in Applied Mathematics, 122(1):29–83, January 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2008.00423.x.

Demirkaya:2014:SSK

[DKK^{+}14] A. Demirkaya, T. Kapitula, P. G. Kevrekidis, M. Stanislavova, and A. Stefanov. On the spectral stability of kinks in some PT - symmetric variants of the classical Klein–Gordon field theories. Studies in Applied Mathematics, 133(3):298–317, October 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12053.

Dyachenko:2016:PDF

[DKZ16] Alexander I. Dyachenko, Dmitriy I. Kachulin, and Vladimir E. Zakharov. Probability distribution functions of freak waves: Nonlinear versus linear model. Studies in Applied Mathematics, 137(2):189–198, August 2016. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12116.

DiBucchianico:1994:SCS

[DL94] A. Di Bucchianico and D. Loeb. A simpler characterization of Sheffer polynomials. Studies in Applied Mathematics, 92(1):1– 15, May 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19949211.

Deng:2023:EAW

[DL23] Guo Deng and Christopher J. Lustri. Exponential asymptotics of woodpile chain nanoptera using numerical analytic continuation. Studies in Applied Mathematics, 150(2):520–557, February 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12548.

delaIglesia:2023:BBD

[dlIJ23] Manuel D. de la Iglesia and Claudia Juarez. The bilateral birth–death chain generated by the associated Jacobi polynomials. Studies in Applied Mathematics, 151(2):616–642, August 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12605.

Dyachenko:2016:BCS

[DLK16] S. A. Dyachenko, P. M. Lushnikov, and A. O. Korotkevich. Branch cuts of Stokes wave on deep water. Part I: Numerical solution and Padé approximation. Studies in Applied Mathematics, 137(4):419–472, November 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12128.

David:1987:INE

[DLW87] D. David, D. Levi, and P. Winternitz. Integrable nonlinear equations for water waves in straits of varying depth and width. Studies in Applied Mathematics, 76(2):133–168, April 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1987762133.

David:1989:SSS

[DLW89] D. David, D. Levi, and P. Winternitz. Solitons in shallow seas of variable depth and in marine straits. Studies in Applied Mathematics, 80(1):1–23, February 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19898011.

Dozo:1996:SPP

[DM96] E. Lami Dozo and M. C. Mariani. Solutions to the plateau problem for the prescribed mean curvature equation via the mountain pass lemma. Studies in Applied Mathematics, 96(3):351–358, April 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1996963351.

DAlessio:1996:GCP

[DMPS96] S. J. D. D’Alessio, T. Bryant Moodie, J. P. Pascal, and G. E. Swaters. Gravity currents produced by sudden release of a fixed volume of heavy fluid. Studies in Applied Mathematics, 96(4):359–385, May 1996. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996964359.

DAlessio:1997:IGC

[DMPS97] S. J. D. D’Alessio, T. B. Moodie, J. P. Pascal, and G. E. Swaters. Intrusive gravity currents. Studies in Applied Mathematics, 98(1):19–46, January 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00039.

Virissimo:2019:TLF

[dMVM19] Francisco de Melo Viríssimo and Paul A. Milewski. Three-layer flows in the shallow water limit. Studies in Applied Mathematics, 142(4):487–512, May 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12266.

Deconinck:2011:SAP

[DN11] B. Deconinck and M. Nivala. The stability analysis of the periodic traveling wave solutions of the mKdV equation. Studies in Applied Mathematics, 126(1):17–48, January 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00496.x.

Dyachenko:2016:W

[DN16] S. Dyachenko and A. C. Newell. Whitecapping. Studies in Applied Mathematics, 137(2):199–213, August 2016.

CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12126.

Destyl:2017:GBS

[DNP17] Edès Destyl, Sylvere Paul Nuiro, and Pascal Poullet. On the global behavior of solutions of a coupled system of nonlinear Schrödinger equation. Studies in Applied Mathematics, 138(2):227–244, February 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12150.

Doubilet:1972:FCT

[Dou72] Peter Doubilet. On the foundations of combinatorial theory. VII: Symmetric functions through the theory of distribution and occupancy. Studies in Applied Mathematics, 51(4):377–396, December 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1972514377.

Deng:2003:SSF

[DPHJ03] Jian Deng, Larry Pratt, Lou Howard, and Chris Jones. On stratified shear flow in sea straits of arbitrary cross section. Studies in Applied Mathematics, 111(4):409–434, November 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.t01-1-00040.

Demontis:2013:IST

[DPvdMV13] F. Demontis, B. Prinari, C. van der Mee, and F. Vitale. The inverse scattering transform for the defocusing nonlinear Schrödinger equations with nonzero boundary conditions. Studies in Applied Mathematics, 131(1):1–40, July 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00572.x.

Davis:1977:BPS

[DR77] Stephen H. Davis and S. Rosenblat. On bifurcating periodic solutions at low frequency. Studies in Applied Mathematics, 57(1):59–76, July 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm197757159.

Djordjevic:1990:EPS

[DR90] V. D. Djordjevic and L. G. Redekopp. The effect of profile symmetry on the nonlinear stability of mixing layers. Studies in Applied Mathematics, 83(4):287–317, November 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990834287.

Drew:1971:AFE

[Dre71] Donald A. Drew. Averaged field equations for two-phase media. Studies in Applied Mathematics, 50(2):133–166, June 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1971502133.

Doubilet:1974:FCT

[DRS74] Peter Doubilet, Gian-Carlo Rota, and Joel Stein. On the foundations of combinatorial theory: IX combinatorial methods in invariant theory. Studies in Applied Mathematics, 53(3):185–216, September 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1974533185.

Degasperis:2002:ISG

[DRS02] A. Degasperis, C. Rogers, and W. K. Schief. Isothermic surfaces generated via Bäcklund and Moutard transformations: Boomeron and zoomeron connections. Studies in Applied Mathematics, 109(1):39–65, July 2002. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00003.

Drew:1971:AFB

[DS71a] Donald A. Drew and Lee A. Segel. Analysis of fluidized beds and foams using averaged equations. Studies in Applied Mathematics, 50(3):233–257, September 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1971503233.

Drew:1971:AET

[DS71b] Donald A. Drew and Lee A. Segel. Averaged equations for twophase flows. Studies in Applied Mathematics, 50(3):205–231, September 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1971503205.

Denier:1999:WMF

[DS99] James P. Denier and Jillian A. K. Stott. Wave-mean flow interactions in thermally stratified Poiseuille flow. Studies in Applied Mathematics, 102(2):121–136, February 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00106.

Dyachenko:2023:CCV

[DS23] S. A. Dyachenko and A. Semenova. Canonical conformal variables based method for stability of Stokes waves. Studies in Applied Mathematics, 150(3):705–715, April 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12554.

daSilva:2020:IEG

[dSF20] Priscila Leal da Silva and Igor Leite Freire. Integrability, existence of global solutions, and wave breaking criteria for a generalization of the Camassa–Holm equation. Studies in Applied Mathematics, 145(3):537–562, October 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12327.

Dunster:2001:CES

[Dun01] T. M. Dunster. Convergent expansions for solutions of linear ordinary differential equations having a simple turning point, with an application to Bessel functions. Studies in Applied Mathematics, 107(3):293–323, October 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00188.

Dunster:2004:CES

[Dun04] T. M. Dunster. Convergent expansions for solutions of linear ordinary differential equations having a simple pole, with an application to associated Legendre functions. Studies in Applied Mathematics, 113(3):245–270, October 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2004.01525.x.

Dunster:2011:SAS

[Dun11] T. M. Dunster. Simplified asymptotic solutions of differential equations having two turning points, with an application to Legendre functions. Studies in Applied Mathematics, 127(3):250–283, October 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2011.00519.x.

Dunster:2013:EWS

[Dun13]  T. M. Dunster.  Electromagnetic wave scattering by

two parallel infinite dielectric cylinders. Studies in Applied Mathematics, 131(3):302–316, October 2013. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12014.

Dunster:2020:ASI

[Dun20] T. M. Dunster. Asymptotic solutions of inhomogeneous differential equations having a turning point. Studies in Applied Mathematics, 145(3):500–536, October 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12326.

Dunster:2022:UAE

[Dun22] T. M. Dunster. Uniform asymptotic expansions for Lommel, Anger–Weber, and Struve functions. Studies in Applied Mathematics, 148(1):340–372, January 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12442.

Duran:2022:EHJ

[Dur22] Antonio J. Durán. Exceptional Hahn and Jacobi polynomials with an arbitrary number of continuous parameters. Studies in Applied Mathematics, 148(2):606–650, February 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12451.

Demontis:2023:ASM

[DvdM23] Francesco Demontis and Cornelis van der Mee. From the AKNS system to the matrix Schrödinger equation with vanishing potentials: Direct and inverse problems. Studies in Applied Mathematics, 150(2):481–519, February 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12547.

Ding:2021:ITB

[DWZ21] Hang Ding, Renhai Wang, and Jun Zhou. Infinite time blow-up of solutions to a class of wave equations with weak and strong damping terms and logarithmic nonlinearity. Studies in Applied Mathematics, 147(3):914–934, October 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12405.

Dyachenko:2020:CSE

[Dya20] Alexander I. Dyachenko. Canonical system of equations for 1D water waves. Studies in Applied Mathematics, 144(4):493–503, May 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12305.

Dyachenko:2022:TCW

[Dya22] Sergey A. Dyachenko. Traveling capillary waves on the boundary of a fluid disc. Studies in Applied Mathematics, 148(1):125–140, January 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12435.

Dai:2022:ADF

[DZ22a] Dan Dai and Yu Zhai. Asymptotics of the deformed Fredholm determinant of the confluent hypergeometric kernel.

Studies in Applied Mathematics, 149(4):1032–1085, November 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12528.

Ding:2022:WPS

[DZ22b] Hang Ding and Jun Zhou. Well-posedness of solutions for a class of quasilinear wave equations with strong damping and logarithmic nonlinearity. Studies in Applied Mathematics, 149(2):441–486, August 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12498.

Edwards:1997:MMT

[Edw97] David A. Edwards. A mathematical model for trapping skinning in polymers. Studies in Applied Mathematics, 99(1):49– 80, July 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9590.00056.

Edwards:2000:SRN

[Edw00] David A. Edwards. Surface reaction near a stagnation point. Studies in Applied Mathematics, 105(1):1–29, July 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00139.

Euler:2019:MIS

[EE19] Marianna Euler and Norbert Euler. On Möbius-invariant and symmetry-integrable evolution equations and the Schwarzian derivative. Studies in Applied Mathematics, 143(2):139–156, August 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12268.

Euler:2003:LHE

[EEP03] M. Euler, N. Euler, and N. Petersson. Linearizable hierarchies of evolution equations in (1+1) dimensions. Studies in Applied Mathematics, 111(3):315–337, October 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.t01-1-00236.

Ercolani:1984:MST

[EFM84] Nicholas Ercolani, M. Gregory Forest, and David W. McLaughlin. Modulational stability of two-phase sine-Gordon wavetrains. Studies in Applied Mathematics, 71(2):97–101, October 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm198471291.

Estevez:1995:NSS

[EG95] P. G. Estévez and P. R. Gordoa. Nonclassical symmetries and the singular manifold method: Theory and six examples. Studies in Applied Mathematics, 95(1):73–113, July 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199595173.

El:2005:WBG

[EGK05] G. A. El, R. H. J. Grimshaw, and A. M. Kamchatnov. Wave breaking and the generation of undular bores in an integrable shallow water system. Studies in Applied Mathematics, 114(4):395–411, May 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
0022-2526.2005.01560.x.

El:2001:ISW

[EGP01] G. A. El, R. H. J. Grimshaw, and M. V. Pavlov. Integrable shallow-water equations and undular bores. Studies in Applied Mathematics, 106(2):157–186, February 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00163.

El:2018:SES

[EHS18] Gennady A. El, Mark A. Hoefer, and Michael Shearer. Stationary expansion shocks for a regularized Boussinesq system. Studies in Applied Mathematics, 140(1):27–47, January 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12191.

Ezquerro:2014:HID

[EHV14] J. A. Ezquerro and M. A. Hernández-Verón. How to improve the domain of starting points for Steffensen’s method. Studies in Applied Mathematics, 132(4):354–380, May 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12033.

Ezquerro:2018:MSN

[EHV18] José A. Ezquerro and Miguel A. Hernández-Verón. Majorizing sequences for nonlinear Fredholm–Hammerstein integral equations. Studies in Applied Mathematics, 140(3):270–297, April 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12200.

Escobar-Lasso:2021:SMM

[ELV21] Oscar Eduardo Escobar-Lasso and Olga Vasilieva. A simplified monotone model of Wolbachia invasion encompassing Aedes aegypti mosquitoes. Studies in Applied Mathematics, 146(3):565–585, April 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12356.

Eu:2003:DPP

[ELY03] Sen-Peng Eu, Shu-Chung Liu, and Yeong-Nan Yeh. Dyck paths with peaks avoiding or restricted to a given set. Studies in Applied Mathematics, 111(4):453–465, November 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.t01-1-00042.

Eu:2006:CSC

[ELY06] Sen-Peng Eu, Shu-Chung Liu, and Yeong-Nan Yeh. On the congruences of some combinatorial numbers. Studies in Applied Mathematics, 116(2):135–144, February 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2006.00337.x.

Everett:1971:AR

[EM71] C. J. Everett and N. Metropolis. Approximation of the v-th root of N. Studies in Applied Mathematics, 50(2):189–191, June 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1971502189.

Emmett:2009:STD

[EM09] M. Emmett and T. B. Moodie. Sediment transport via dam-break flows over sloping erodible beds. Studies in Applied Mathematics, 123(3):257–290, October 2009. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00453.x.

Ein:1981:SRS

[ERK^{+}81] Lawrence M. H. Ein, David Ross Richman, Daniel J. Kleitman, James Shearer, and Dean Sturtevant. Some results on systems of finite sets that satisfy a certain intersection condition. Studies in Applied Mathematics, 65(3):269–274, December 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1981653269.

Escudero:2020:KEL

[Esc20] Carlos Escudero. Kinetic energy of the Langevin particle. Studies in Applied Mathematics, 145(4):719–738, November 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12333.

Escobedo:2023:LSE

[Esc23] M. Escobedo. On the linearized system of equations for the condensate–normal fluid interaction at very low temperature. Studies in Applied Mathematics, 150(2):448–456, February 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12545.

Eydeland:1987:CNP

[ET87] Alexander Eydeland and Bruce Turkington. On the computation of nonlinear planetary waves. Studies in Applied Mathematics, 76(1):37–67, February 1987. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198776137.

Erhardt:2022:BAA

[EWW22] André H. Erhardt, Erik Wahlén, and Jörg Weber. Bifurcation analysis for axisymmetric capillary water waves with vorticity and swirl. Studies in Applied Mathematics, 149(4):904–942, November 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12525.

E:1991:NSO

[EY91] Weinan E and Huanan Yang. Numerical study of oscillatory solutions of the gas-dynamic equations. Studies in Applied Mathematics, 85(1):29–52, July 1991. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199185129.

Fokas:1983:ISTa

[FA83a] A. S. Fokas and M. J. Ablowitz. The inverse scattering transform for the Benjamin–Ono equation — a pivot to multidimensional problems. Studies in Applied Mathematics, 68(1):1–10, February 1983. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19836811.

Fokas:1983:ISTb

[FA83b] A. S. Fokas and M. J. Ablowitz. On the inverse scattering of the time-dependent Schrödinger equation and the associated Kadomtsev–Petviashvili (i) equation. Studies in Applied Mathematics, 69(3):211–228, December 1983. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1983693211.

Fokas:1989:FNE

[FA89] A. S. Fokas and M. J. Ablowitz. Forced nonlinear evolution equations and the inverse scattering transform. Studies in Applied Mathematics, 80(3):253–272, June 1989. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1989803253.

Feppon:2022:MDP

[FA22] Florian Feppon and Habib Ammari. Modal decompositions and point scatterer approximations near the Minnaert resonance frequencies. Studies in Applied Mathematics, 149(1):164–229, July 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12493.

Fairlie:1997:FSE

[Fai97] D. B. Fairlie. Formal solutions of an evolution equation of Riemann type. Studies in Applied Mathematics, 98 (2):203–205, February 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00046.

Fan:2011:NBB

[Fan11] E. Fan. New bilinear Bäcklund transformation and Lax pair for the supersymmetric two-Boson equation. Studies in Applied Mathematics, 127(3):284–301, October 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00520.x.

Farmer:1982:HSF

[Far82] Frank D. Farmer. Homotopy spheres in formal language. Studies in Applied Mathematics, 66(2):171–179, April 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1982662171.

Fateman:1974:CPS

[Fat74] Richard J. Fateman. On the computation of powers of sparse polynomials. Studies in Applied Mathematics, 53(2):145–155, June 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1974532145.

Filipuk:2008:SFP

[FC08] Galina V. Filipuk and Peter A. Clarkson. The symmetric fourth Painlevé hierarchy and associated special polynomials. Studies in Applied Mathematics, 121(2):157–188, August 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00410.x.

Feng:2022:NLB

[FCCTR22] Baowei Feng, Victor R. Cabanillas, Edson A. Coayla-Teran, and Carlos A. Raposo. Nonuniform laminated beam of Lord–Shulman type. Studies in Applied Mathematics, 149 (4):1123–1154, November 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12530.

Fernandez-Cara:1995:ESC

[FCGB95] E. Fernández-Cara and M. González-Burgos. On the existence of solutions and the convergence of approximations to scalar conservation laws. Studies in Applied Mathematics, 94(4):377– 391, May 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1995944377.

Fan:2012:DPS

[FCL12] E.-G. Fan, K.-W. Chow, and J.-H. Li. On doubly periodic standing wave solutions of the coupled Higgs field equation. Studies in Applied Mathematics, 128(1):86–105, January 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00531.x.

Ford:2021:MDS

[FCP^{+}21] Andrew G. Ford, Xue-Zheng Cao, Micah J. Papanikolas, Takafumi Kato, Richard C. Boucher, Matthew R. Markovetz, David B. Hill, Ronit Freeman, and Mark Gregory Forest. Molecular dynamics simulations to explore the structure and rheological properties of normal and hyperconcentrated airway mucus. Studies in Applied Mathematics, 147(4):1369–1387, November 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12433.

Fang:2021:SEE

[FDL21] Xiaoping Fang, Youjun Deng, and Hongyu Liu. Sharp estimate of electric field from a conductive rod and application. Studies in Applied Mathematics, 146(1):279–297, January 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12348.

Freire:2022:NED

[FdST22] Igor Leite Freire and Rodrigo da Silva Tito. A Novikov equation describing pseudo-spherical surfaces, its pseudopotentials, and local isometric immersions. Studies in Applied Mathematics, 148(2):758–772, February 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12457.

Fernandes:2012:EVS

[FG12] R. S. Fernandes and V. A. Galaktionov. Eigenfunctions and very singular similarity solutions of odd-order nonlinear dispersion PDEs: Toward a “Nonlinear Airy function” and others. Studies in Applied Mathematics, 129(2):163–219, August 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00554.x.

Flaherty:1978:FEF

[FH78] J. E. Flaherty and F. C. Hoppensteadt. Frequency entrainment of a forced van der Pol oscillator. Studies in Applied Mathematics, 58(1):5–15, January 1978. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19785815.

Friedlander:1998:IPF

[FH98] Susan Friedlander and Louis Howard. Instability in parallel flows revisited. Studies in Applied Mathematics, 101(1):1–21, July 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00083.

Filipuk:2009:ROM

[FH09] G. Filipuk and R. G. Halburd. Rational ODEs with movable algebraic singularities. Studies in Applied Mathematics, 123(1):17–36, July 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2009.00445.x.

Fan:2010:QWS

[FH10] E. G. Fan and Y. C. Hon. Quasiperiodic wave solutions of N = 2 supersymmetric KdV equation in superspace. Studies in Applied Mathematics, 125(4):343–371, November 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00491.x.

Faver:2023:MSD

[FH23] Timothy E. Faver and Hermen Jan Hupkes. Mass and spring dimer Fermi–Pasta–Ulam–Tsingou nanopterons with exponentially small, nonvanishing ripples. Studies in Applied Mathematics, 150(4):1046–1153, May 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12564.

Fernandez-Irisarri:2023:LFE

[FIM23] Itsaso Fernández-Irisarri and Manuel Mañas. Laguerre–Freud equations for three families of hypergeometric discrete orthogonal polynomials. Studies in Applied Mathematics, 151(2):509– 535, August 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12601.

Fokas:1994:IEF

[FIN94] A. S. Fokas, A. R. Its, and V. Yu. Novokshenov. Integrable equations with a forcing of a distribution type. Studies in Applied Mathematics, 92(2):97–114, June 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199492297.

Furry:1977:MRC

[FK77] W. H. Furry and D. J. Kleitman. Maximal rectilinear crossing of cycles. Studies in Applied Mathematics, 56(2):159–167, April 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1977562159.

Frauendiener:2022:NSD

[FKMS22] Jörg Frauendiener, Christian Klein, Umar Muhammad, and Nikola Stoilov. Numerical study of Davey–Stewartson I systems. Studies in Applied Mathematics, 149(1):76–94, July 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12491.

Feldman:1973:FMS

[FL73] S. I. Feldman and C. C. Lin. A forcing mechanism for spiral density waves in galaxies. Studies in Applied Mathematics, 52(1):1–20, March 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19735211.

Finlay:1995:NCW

[FLM95] W. H. Finlay, Chonghui Liu, and S. A. Maslowe. The nonlinear critical-wall layer in a parallel shear flow. Studies in Applied Mathematics, 94(1):41–55, January 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199594141.

Ferreira:2023:AAH

[FLS23] Chelo Ferreira, José L. López, and Ester Pérez Sinusía. Asymptotic approximation of a highly oscillatory integral with application to the canonical catastrophe integrals. Studies in Applied Mathematics, 150(1):254–276, January 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12539.

Feng:2020:MBH

[FLT20] Bao-Feng Feng, Liming Ling, and Daisuke A. Takahashi. Multi-breather and high-order rogue waves for the nonlinear Schrödinger equation on the elliptic function background. Studies in Applied Mathematics, 144(1):46–101, January 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12287.

Forest:1983:MSG

[FM83a] M. Gregory Forest and David W. McLaughlin. Modulations of sinh-Gordon and sine-Gordon wavetrains. Studies in Applied Mathematics, 68(1):11–59, February 1983. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198368111.

Fowler:1983:HPD

[FM83b] A. C. Fowler and M. J. McGuinness. Hysteresis, period doubling, and intermittency at high Prandtl number in the Lorenz equations. Studies in Applied Mathematics, 69(2):99–126, October 1983. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm198369299.

Feys:2011:LNS

[FM11] Jan Feys and Sherwin A. Maslowe. Long nonlinear surface waves in the presence of a variable current. Studies in Applied Mathematics, 126(3):303–318, April 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00506.x.

Fusco:1999:PES

[FMM99] D. Fusco, N. Manganaro, and M. Migliardo. Pulselike exact solutions for a model describing nerve fibers. Studies in Applied Mathematics, 103(1):25–42, July 1999. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00118.

Flamarion:2019:RWG

[FMN19] Marcelo V. Flamarion, Paul A. Milewski, and André Nachbin. Rotational waves generated by current-topography interaction. Studies in Applied Mathematics, 142(4):433–464, May 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12253.

Feng:2017:GFM

[FMO17] Bao-Feng Feng, Ken-Ichi Maruno, and Yasuhiro Ohta. Geometric formulation and multi-dark soliton solution to the defocusing complex short pulse equation. Studies in Applied Mathematics, 138(3):343–367, April 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12159.

Faltin:1977:BAA

[FMRR77] F. Faltin, N. Metropolis, B. Ross, and G.-C. Rota. A Boolean analysis of addition and multiplication. Studies in Applied Mathematics, 56(2):147–158, April 1977. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1977562147.

Feng:2022:SAV

[FMX22] Xiao Xuan Feng, Mahyar Mahinzaeim, and Gen Qi Xu. Spectral analysis of a viscoelastic tube conveying fluid with generalized boundary conditions. Studies in Applied Mathematics, 149(3):657–677, October 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12516.

Farjoun:2005:OSJ

[FN05a] Yossi Farjoun and John Neu. The optimal shape of a javelin. Studies in Applied Mathematics, 115(3):339–354, October 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2005.00317.x.

Farjoun:2005:TCD

[FN05b] Yossi Farjoun and John Neu. The tallest column — a dynamical system approach using a symmetry solution. Studies in Applied Mathematics, 115(3):319–337, October 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2005.00316.x.

Fu:2022:CKP

[FN22] Wei Fu and Frank W. Nijhoff. On a coupled Kadomtsev– Petviashvili system associated with an elliptic curve. Studies in Applied Mathematics, 149(4):1086–1122, November 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12529.

Fokas:1987:SI

[Fok87] A. S. Fokas. Symmetries and integrability. Studies in Applied Mathematics, 77(3):253–299, December 1987. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1987773253.

Fokas:2009:KPE

[Fok09] A. S. Fokas. Kadomtsev–Petviashvili equation revisited and integrability in 4 + 2 and 3 + 1. Studies in Applied Mathematics, 122(4):347–359, May 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2009.00437.x.

Fernandez:2022:CSP

[FÖK22] Arran Fernandez, Mehmet Ali Özarslan, and Cemaliye Kürt. A catalogue of semigroup properties for integral operators with Fox–Wright kernel functions. Studies in Applied Mathematics, 148(4):1477–1518, May 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12481.

Foster:1970:RGSa

[Fos70a] M. R. Foster. The rotation of a gravitating sphere in a monatomic gas, I. Studies in Applied Mathematics, 49(1):85– 99, March 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm197049185.

Foster:1970:RGSb

[Fos70b] M. R. Foster. The rotation of a gravitating sphere in a monatomic gas, II. Studies in Applied Mathematics, 49 (3):259–276, September 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1970493259.

Fowler:1984:ALE

[Fow84] A. C. Fowler. Analysis of the Lorenz equations for large r. Studies in Applied Mathematics, 70(3):215–233, June 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1984703215.

Fowler:1985:FTC

[Fow85a] A. C. Fowler. Fast thermoviscous convection. Studies in Applied Mathematics, 72(3):189–219, June 1985. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1985723189.

Fowler:1985:SCC

[Fow85b] A. C. Fowler. Secular cooling in convection. Studies in Applied Mathematics, 72(2):161–171, April 1985. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1985722161.

Fowler:1986:TRE

[Fow86] A. C. Fowler. Thermal runaway in the Earth’s mantle. Studies in Applied Mathematics, 74(1):1–34, February 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19867411.

Fowler:1990:HBP

[Fow90a] A. C. Fowler. Homoclinic bifurcations for partial differential equations in unbounded domains. Studies in Applied Mathematics, 83(4):329–353, November 1990. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990834329.

Fowler:1990:HBD

[Fow90b] A. C. Fowler. Homoclinic bifurcations in n dimensions. Studies in Applied Mathematics, 83(3):193–209, October 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990833193.

Fowler:1993:TDC

[Fow93] A. C. Fowler. Toward a description of convection with temperature-and-pressure-dependent viscosity. Studies in Applied Mathematics, 88(2):113–139, February 1993. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1993882113.

Fasondini:2023:OPC

[FOX23] Marco Fasondini, Sheehan Olver, and Yuan Xu. Orthogonal polynomials on a class of planar algebraic curves. Studies in Applied Mathematics, 151(1):369–405, July 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12582.

Fokas:2005:ACI

[FP05] A. S. Fokas and D. T. Papageorgiou. Absolute and convective instability for evolution PDEs on the half-line. Studies in Applied Mathematics, 114(1):95–114, January 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2005.01541.x.

Fokas:2012:GDN

[FP12] A. S. Fokas and B. Pelloni. Generalized Dirichlet-to-Neumann map in time-dependent domains. Studies in Applied Mathematics, 129(1):51–90, July 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2011.00545.x.

Fokas:1989:ISP

[FPSA89] A. S. Fokas, E. P. Papadopoulou, Y. G. Saridakis, and M. J. Ablowitz. Interaction of simple particles in soliton cellular automata. Studies in Applied Mathematics, 81(2):153–180, October 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1989812153.

Fokas:2001:IBV

[FR01] A. S. Fokas and C. Rogers. Initial/Boundary value problems for simultaneous evolution equations in three dimensions.

Studies in Applied Mathematics, 107(4):391–401, November 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.1074192.

Flavin:2008:UEM

[FR08] J. N. Flavin and C. Rogers. Upper estimates for a moving boundary problem for resonant nonlinear Schrödinger equations. Studies in Applied Mathematics, 121(2):189–198, August 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00417.x.

Faria:2017:ELM

[FR17] L. M. Faria and R. R. Rosales. Equation level matching: an extension of the method of matched asymptotic expansion for problems of wave propagation. Studies in Applied Mathematics, 139(2):265–287, August 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12183.

Freeman:1972:COI

[Fre72] J. M. Freeman. On the classification of operator identities. Studies in Applied Mathematics, 51(1):73–84, March 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197251173.

Freeman:1987:OT

[Fre87] J. M. Freeman. Orthogonality via transforms. Studies in Applied Mathematics, 77(2):119–127, October 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1987772119.

Friedlander:1974:SRS

[Fri74] Susan Friedlander. Spin-down in a rotating stratified fluid: Part I. Studies in Applied Mathematics, 53(2):111–136, June 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1974532111.

Fromm:2019:CSD

[Fro19] Samuel Fromm. Construction of solutions of the defocusing nonlinear Schrödinger equation with asymptotically time-periodic boundary values. Studies in Applied Mathematics, 143(4):404–448, November 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12285.

Fix:1969:FAF

[FS69] George Fix and Gilbert Strang. Fourier analysis of the finite element method in Ritz–Galerkin theory. Studies in Applied Mathematics, 48(3):265–273, September 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1969483265.

Fokas:1986:ROK

[FS86] A. S. Fokas and P. M. Santini. The recursion operator of the Kadomtsev–Petviashvili equation and the squared eigenfunctions of the Schrödinger operator. Studies in Applied Mathematics, 75(2):179–185, October 1986. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1986752179.

Feinsiluer:1996:DRR

[FS96] P. Feinsiluer and R. Schott. Differential relations and recurrence formulas for representations of Lie groups. Studies in Applied Mathematics, 96(4):387–406, May 1996. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996964387.

Farhang-Sardroodi:2021:SPS

[FSKMP21] Suzan Farhang-Sardroodi, Natalia L. Komarova, Marcus Michelen, and Robin Pemantle. Success probability for selectively neutral invading species in the line model with a random fitness landscape. Studies in Applied Mathematics, 146(4):1023–1049, May 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12373.

Ferguson:1978:MES

[FSY78] Warren Ferguson, Philip Saffman, and Henry Yuen. A model equation to study the effects of nonlinearity, surface tension and viscosity in water waves. Studies in Applied Mathematics, 58(2):165–185, February 1978. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1978582165.

Forrester:2019:FSC

[FT19] Peter J. Forrester and Allan K. Trinh. Finite-size corrections at the hard edge for the Laguerre β ensemble. Studies in Applied Mathematics, 143(3):315–336, October 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12279.

Fan:2021:CRV

[FT21] Long Fan and Nafissa Toureche Trouba. Convergence rates of vanishing diffusion limit on conservative form of Hsieh’s equation. Studies in Applied Mathematics, 146(3):753–776, April 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12366.

Fokas:2021:UTE

[FvdW21] A. S. Fokas and M. C. van der Weele. The unified transform for evolution equations on the half-line with time-periodic boundary conditions. Studies in Applied Mathematics, 147(4):1339–1368, November 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12452.

Fuh:2001:RPG

[FY01] Cheng-Der Fuh and Yeong-Nan Yeh. Random perturbation in games of chance. Studies in Applied Mathematics, 107(2):207–215, August 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00185.

Fedele:2005:PDM

[FYC05] Francesco Fedele, Jianke Yang, and Zhigang Chen. Properties of defect modes in one-dimensional optically induced photonic lattices. Studies in Applied Mathematics, 115(2):279–301, August 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2005.00329.x.

Foata:1990:DPS

[FZ90] Dominique Foata and Doron Zeilberger. Denert’s permutation statistic is indeed Euler-Mahonian. Studies in Applied Mathematics, 83(1):31–59, July 1990. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199083131.

Grimshaw:1979:LCS

[GA79] R. Grimshaw and J. S. Allen. Linearly coupled, slowly varying oscillators. Studies in Applied Mathematics, 61(1):55–71, July 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197961155.

Grimshaw:1982:EDL

[GA82] R. Grimshaw and J. S. Allen. The effect of dissipation on linearly coupled, slowly varying oscillators. Studies in Applied Mathematics, 67(3):169–198, December 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1982673169.

Grimshaw:2011:WWP

[GA11] R. H. J. Grimshaw and S. Y. Annenkov. Water wave packets over variable depth. Studies in Applied Mathematics, 126(4):409–427, May 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2010.00508.x.

Gaalswyk:1982:LBE

[Gaa82] Arie Gaalswyk. Limit behavior and the existence of combustion shock layers. Studies in Applied Mathematics, 67 (2):141–168, October 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1982672141.

Gabrovsek:2021:ICB

[Gab21] Boštjan Gabrovšek. An invariant for colored bonded knots. Studies in Applied Mathematics, 146(3):586–604, April 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12357.

Gajjar:1996:NSN

[Gaj96] J. S. B. Gajjar. Nonlinear stability of nonstationary crossflow vortices in compressible boundary layers. Studies in Applied Mathematics, 96(1):53–84, January 1996. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199696153.

Gallavotti:1971:PGI

[Gal71] G. Gallavotti. A proof of the Griffiths inequalities for the X– Y model. Studies in Applied Mathematics, 50(1):89–92, March 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197150189.

Galaktionov:2006:IOS

[Gal06] V. A. Galaktionov. On interfaces and oscillatory solutions of higher-order semilinear parabolic equations with non-Lipschitz nonlinearities. Studies in Applied Mathematics, 117(4):353–389, November 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2006.00359.x.

Galaktionov:2008:HST

[Gal08] V. A. Galaktionov. Hermitian spectral theory and blow-up patterns for a fourth-order semilinear Boussinesq equation. Studies in Applied Mathematics, 121(4):395–431, November 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00421.x.

Galaktionov:2010:ISS

[Gal10a] V. A. Galaktionov. Incomplete self-similar blow-up in a semilinear fourth-order reaction–diffusion equation. Studies in Applied Mathematics, 124(4):347–381, May 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00474.x.

Galaktionov:2010:VSS

[Gal10b] V. A. Galaktionov. Very singular solutions for thin film equations with absorption. Studies in Applied Mathematics, 124(1):39–63, January 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2009.00461.x.

Galaktionov:2011:SPG

[Gal11] Victor A. Galaktionov. Single point gradient blow-up and nonuniqueness for a third-order nonlinear dispersion equation. Studies in Applied Mathematics, 126(2):103–143, February 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2010.00499.x.

Gao:1995:DTN

[Gao95] David Yang Gao. Duality theory in nonlinear buckling analysis for von Kármán equations. Studies in Applied Mathematics, 94(4):423–444, May 1995. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995944423.

Gartside:1996:NIB

[GB96] James Gartside and Thomas F. Balsa. Nonlinear interaction between a pair of oblique modes in a supersonic mixing layer: Long-wave limit. Studies in Applied Mathematics, 97(1):53–98, July 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm199697153.

Gerontitis:2021:IFT

[GBSS21] Dimitris Gerontitis, Ratikanta Behera, Jajati Keshari Sahoo, and Predrag S. Stanimirović. Improved finite-time zeroing neural network for time-varying division. Studies in Applied Mathematics, 146(2):526–549, February 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12354.

Geng:2007:DDA

[GDZ07] Xianguo Geng, H. H. Dai, and Junyi Zhu. Decomposition of the discrete Ablowitz–Ladik hierarchy. Studies in Applied Mathematics, 118(3):281–312, April 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2007.00374.x.

Gear:1985:SIB

[Gea85] J. A. Gear. Strong interactions between solitary waves belonging to different wave modes. Studies in Applied Mathematics, 72(2):95–124, April 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm198572295.

Geza:1979:OJD

[Gez79] Schay Geza. Optimal joint distributions of several random variables with given marginals. Studies in Applied Mathematics, 61(2):179–183, October 1979. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1979612179. See errata [Gez80, Sch80].

Geza:1980:EOJ

[Gez80] Schay Geza. Erratum: Optimal joint distributions of several random variables with given marginals. Studies in Applied Mathematics, 62(1):93, February 1980. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198062193. See [Gez79].

Garcia-Ferrero:2022:EGP

[GFGUMM22] María Ángeles García-Ferrero, David Gómez-Ullate, Robert Milson, and James Munday. Exceptional Gegenbauer polynomials via isospectral deformation. Studies in Applied Mathematics, 149(2):324–363, August 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12510.

Gottlieb:1976:NSE

[GG76] David Gottlieb and Berti Gustafsson. On the Navier– Stokes equations with constant total temperature. Studies in Applied Mathematics, 55(2):167–185, June 1976. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1976552167.

Gear:1984:WSI

[GG84] J. A. Gear and R. Grimshaw. Weak and strong interactions between internal solitary waves. Studies in Applied Mathematics, 70(3):235–258, June 1984. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1984703235.

Gilman:1995:AAN

[GGS95] O. A. Gilman, R. Grimshaw, and Yu. A. Stepanyants. Approximate analytical and numerical solutions of the stationary Ostrovsky equation. Studies in Applied Mathematics, 95(1):115– 126, July 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1995951115.

Grimshaw:2008:LTS

[GH08] Roger Grimshaw and Karl Helfrich. Long-time solutions of the Ostrovsky equation. Studies in Applied Mathematics, 121(1):71–88, July 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2008.00412.x.

Grimshaw:2012:ROE

[GHJ12] R. H. J. Grimshaw, Karl Helfrich, and E. R. Johnson. The reduced Ostrovsky equation: Integrability and breaking. Studies in Applied Mathematics, 129(4):414–436, November 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00560.x.

Gesztesy:2008:LCL

[GHMT08] Fritz Gesztesy, Helge Holden, Johanna Michor, and Gerald Teschl. Local conservation laws and the Hamiltonian formalism for the Ablowitz–Ladik hierarchy. Studies in Applied Mathematics, 120(4):361–423, May 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00405.x.

Grimshaw:1998:TDS

[GHO98] R. H. J. Grimshaw, J.-M. He, and L. A. Ostrovsky. Terminal damping of a solitary wave due to radiation in rotational systems. Studies in Applied Mathematics, 101(2):197–210, August 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00090.

Greenspan:1981:ESF

[GJ81] H. P. Greenspan and Arne V. Johansson. An experimental study of flow over an impounding dike. Studies in Applied Mathematics, 64(3):211–223, June 1981. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1981643211.

Goulden:1984:LCC

[GJ84] I. P. Goulden and D. M. Jackson. A logarithmic connection for circular permutation enumeration. Studies in Applied Mathematics, 70(2):121–139, April 1984. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1984702121.

Geng:2022:SPD

[GJLW22] Dongxu Geng, Weihua Jiang, Yuan Lou, and Hongbin Wang. Spatiotemporal patterns in a diffusive predator–prey system with nonlocal intraspecific prey competition. Studies in Applied Mathematics, 148(1):396–432, January 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12444.

Goldman:1977:RTI

[GJW77] Jay R. Goldman, J. T. Joichi, and Dennis E. White. Rook theory — IV. orthogonal sequences of rook polynomials. Studies in Applied Mathematics, 56(3):267–272, June 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1977563267.

Greenberg:1974:LTS

[GK74] Philip J. Greenberg and James P. Knauer. On a Lorentz transformation suitable for local problems in general relativity. Studies in Applied Mathematics, 53(2):165–183, June 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1974532165.

Guo:1995:DDC

[GK95] Jianlin Guo and P. N. Kaloni. Double-diffusive convection in a porous medium, nonlinear stability, and the Brinkman effect. Studies in Applied Mathematics, 94(3):341–358, April 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995943341.

Greene:1976:EEK

[GKK76] C. Greene, G. O. H. Katona, and D. J. Kleitman. Extensions of the Erdös-Ko–Rado theorem. Studies in Applied Mathematics, 55(1):1–8, March 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19765511.

Grigorieva:2018:OCS

[GKK18] Ellina Grigorieva, Evgenii Khailov, and Andrei Korobeinikov. Optimal control for an SEIR epidemic model with nonlinear incidence rate. Studies in Applied Mathematics, 141(3):353–398, October 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12227.

Grigorieva:2021:OQR

[GKK21] Ellina V. Grigorieva, Evgenii N. Khailov, and Andrei Korobeinikov. Optimal quarantine-related strategies for COVID-19 control models. Studies in Applied Mathematics, 147 (2):622–649, August 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12393.

Greene:1974:CTI

[GKM74] Curtis Greene, Daniel J. Kleitman, and T. L. Magnantit. Complementary trees and independent matchings. Studies in Applied Mathematics, 53(1):57–64, March 1974. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197453157.

Glebov:2010:WND

[GKT10] Sergei Glebov, Oleg Kiselev, and Nikolai Tarkhanov. Weakly nonlinear dispersive waves under parametric resonance perturbation. Studies in Applied Mathematics, 124(1):19–37, January 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00460.x.

Grimshaw:2017:NPS

[GL17] Roger Grimshaw and Zihua Liu. Nonlinear periodic and solitary water waves on currents in shallow water. Studies in Applied Mathematics, 139(1):60–77, July 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12164.

Gamboa:2022:MAW

[GLH22]  Maria Gamboa and Mariajesus Lopez-Herrero.  Mea-

sures to assess a warning vaccination level in a stochastic SIV model with imperfect vaccine. Studies in Applied Mathematics, 148(4):1411–1438, May 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12479.

Guo:2013:HOS

[GLL13] Boling Guo, Liming Ling, and Q. P. Liu. High-order solutions and generalized Darboux transformations of derivative nonlinear Schrödinger equations. Studies in Applied Mathematics, 130(4):317–344, May 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2012.00568.x.

Gawronski:2014:ASC

[GLN14] Wolfgang Gawronski, Lance L. Littlejohn, and Thorsten Neuschel. Asymptotics of Stirling and Chebyshev–Stirling numbers of the second kind. Studies in Applied Mathematics, 133(1):1–17, July 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12037.

Gavrilyuk:2023:HSM

[GLS23] Sergey Gavrilyuk, Yu-Hsi Lin, and Keh-Ming Shyue. Hyperbolicity study of the modulation equations for the Benjamin– Bona–Mahony equation. Studies in Applied Mathematics, 151(2):536–554, August 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12602.

Geng:2015:IBV

[GLZ15] Xianguo Geng, Huan Liu, and Junyi Zhu. Initial-boundary value problems for the coupled nonlinear Schrödinger equation on the half-line. Studies in Applied Mathematics, 135(3):310–346, October 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12088.

Graves:1975:UM

[GM75] William H. Graves and Suzanne M. Molnar. A universal measure. Studies in Applied Mathematics, 54(2):117–134, June 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1975542117.

Greenspan:1981:WSV

[GM81] H. P. Greenspan and B. M. McCay. On the wetting of a surface by a very viscous fluid. Studies in Applied Mathematics, 64(2):94–112, April 1981. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198164295.

Grimshaw:1989:DMK

[GM89] R. Grimshaw and W. K. Melville. On the derivation of the modified Kadomtsev–Petviashvili equation. Studies in Applied Mathematics, 80(3):183–202, June 1989. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1989803183.

Grimshaw:1993:SVS

[GM93] R. Grimshaw and H. Mitsudera. Slowly varying solitary wave solutions of the perturbed Korteweg–de Vries equation revisited. Studies in Applied Mathematics, 90(1):75–86, October 1993. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199390175.

Goncalves:2012:MFN

[GM12] T. M. N. Gonçalves and E. L. Mansfield. On moving frames and Noether’s conservation laws. Studies in Applied Mathematics, 128(1):1–29, January 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2011.00522.x.

Goncalves:2013:MFC

[GM13] T. M. N. Gonçalves and E. L. Mansfield. Moving frames and conservation laws for Euclidean invariant Lagrangians. Studies in Applied Mathematics, 130(2):134–166, February 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00566.x.

Guidotti:2021:WVO

[GM21] Patrick Guidotti and Sandro Merino. On Wiener’s violent oscillations, Popov’s curves, and Hopf’s supercritical bifurcation for a scalar heat equation. Studies in Applied Mathematics, 146(3):677–729, April 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12364.

Gubeskys:2005:ASN

[GMM05] Arthur Gubeskys, Boris A. Malomed, and Ilya M. Merhasin. Alternate Solitons: Nonlinearly managed one- and two-dimensional solitons in optical lattices. Studies in Applied Mathematics, 115(2):255–277, August 2005. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2005.00328.x. See [Ano05b].

Geyer:2022:SSP

[GMNP22] Anna Geyer, Renan H. Martins, Fábio Natali, and Dmitry E. Pelinovsky. Stability of smooth periodic travelling waves in the Camassa–Holm equation. Studies in Applied Mathematics, 148(1):27–61, January 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12430.

Gregory:1998:ADT

[GMW98] R. Douglas Gregory, Thomas I. Milac, and Frederic Y. M. Wan. The axisymmetric deformation of a thin, or moderately thick, elastic spherical cap. Studies in Applied Mathematics, 100(1):67–94, January 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00070.

Gao:2021:CGS

[GMW21] T. Gao, P. A. Milewski, and Z. Wang. Capillary-gravity solitary waves on water of finite depth interacting with a linear shear current. Studies in Applied Mathematics, 147 (3):1036–1057, October 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12422.

Goldsztein:2004:CRG

[Gol04] Guillermo H. Goldsztein. Collapse and rebound of a gas bubble. Studies in Applied Mathematics, 112(2):101–132, February 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2004.01497.x.

Gorshkov:2011:DSN

[GOS11] K. Gorshkov, L. A. Ostrovsky, and I. Soustova. Dynamics of strongly nonlinear kinks and solitons in a two-layer fluid. Studies in Applied Mathematics, 126(1):49–73, January 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00497.x.

Grabmuller:1987:PSA

[GP87] Hans Grabmüller and Robert Pirner. Positive solutions of annular elastic membrane problems with finite rotations. Studies in Applied Mathematics, 77(3):223–252, December 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1987773223.

Grimshaw:2004:GSS

[GP04] R. H. J. Grimshaw and S. R. Pudjaprasetya. Generation of secondary solitary waves in the variable-coefficient Korteweg–de Vries equation. Studies in Applied Mathematics, 112(3):271–279, April 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
0022-2526.2004.01521.x.

Gallo:2014:TFA

[GP14] Clément Gallo and Dmitry Pelinovsky. On the Thomas–Fermi approximation of the ground state in a PT -symmetric confining potential. Studies in Applied Mathematics, 133(4):398–421, November 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12064.

Gkogkou:2022:IST

[GPFT22] Aikaterini Gkogkou, Barbara Prinari, Bao-Feng Feng, and A. David Trubatch. Inverse scattering transform for the complex coupled short-pulse equation. Studies in Applied Mathematics, 148(2):918–963, February 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12463.

Grimshaw:1996:ISW

[GPS96] Roger Grimshaw, Efim Pelinovsky, and Pavel Sakov. Interaction of a solitary wave with an external force moving with variable speed. Studies in Applied Mathematics, 97(3):235–276, October 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1996973235.

Grimshaw:1998:SWT

[GPT98] Roger Grimshaw, Efim Pelinovsky, and Tatjana Talipova. Solitary wave transformation due to a change in polarity. Studies in Applied Mathematics, 101(4):357–388, November 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00098.

Grimshaw:2005:SLL

[GPT^{+}05] Roger Grimshaw, Efim Pelinovsky, Tatiana Talipova, Michael Ruderman, and Róbert Erdélyi. Short-lived large-amplitude pulses in the nonlinear long-wave model described by the modified Korteweg–de Vries equation. Studies in Applied Mathematics, 114(2):189–210, February 2005. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2005.01544.x.

Goldman:1970:FCT

[GR70] Jay Goldman and Gian-Carlo Rota. On the foundations of combinatorial theory IV finite vector spaces and Eulerian generating functions. Studies in Applied Mathematics, 49(3):239–258, September 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1970493239.

Gelash:2023:VBM

[GR23] Andrey Gelash and Anton Raskovalov. Vector breathers in the Manakov system. Studies in Applied Mathematics, 150(3):841– 882, April 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12558.

Graham:1973:SEW

[Gra73]  Keith D. Graham.  Separation of eigenvalues of the

wave equation for the unit ball in R^{N}. Studies in Applied Mathematics, 52(4):329–343, December 1973. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1973524329.

Greenspan:1969:NLS

[Gre69a] H. P. Greenspan. A note on the laboratory simulation of planetary flows. Studies in Applied Mathematics, 48(2):147–152, June 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1969482147.

Greenspan:1969:ITR

[Gre69b] H. P. Greenspan. On the inviscid theory of rotating fluids. Studies in Applied Mathematics, 48(1):19–28, March 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm196948119.

Greenspan:1970:NEW

[Gre70] H. P. Greenspan. A note on edge waves in a stratified fluid. Studies in Applied Mathematics, 49(4):381–388, December 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1970494381.

Greenberg:1972:ARCb

[Gre72a] Phillip Greenberg. The algebra of the Riemann curvature tensor in general relativity: The relation of the invariants of the Einstein curvature tensor to the invariants describing matter. Studies in Applied Mathematics, 51(4):369–376, December 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1972514369.

Greenberg:1972:NTS

[Gre72b] Phillip Greenberg. A note on a theorem of Szekeres. Studies in Applied Mathematics, 51(4):415–416, December 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1972514415.

Greenberg:1972:ARCa

[Gre72c] Phillip J. Greenberg. The algebra of the Riemann curvature tensor in general relativity: Preliminaries. Studies in Applied Mathematics, 51(3):277–308, September 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1972513277.

Greenspan:1972:MGS

[Gre72d] H. P. Greenspan. Models for the growth of a solid tumor by diffusion. Studies in Applied Mathematics, 51(4):317–340, December 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1972514317.

Greenspan:1974:D

[Gre74] H. P. Greenspan. On α-dynamos. Studies in Applied Mathematics, 53(1):35–43, March 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm197453135.

Greene:1975:IMF

[Gre75] Curtis Greene. An inequality for the Möbius function of a geometric lattice. Studies in Applied Mathematics, 54(1):71–74, March 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm197554171.

Greene:1976:WEG

[Gre76] Curtis Greene. Weight enumeration and the geometry of linear codes. Studies in Applied Mathematics, 55(2):119–128, June 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1976552119.

Greenspan:1977:DVD

[Gre77] H. P. Greenspan. On the deformation of a viscous droplet caused by variable surface tension. Studies in Applied Mathematics, 57(1):45–58, July 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm197757145.

Greenspan:1984:CEL

[Gre84] H. P. Greenspan. Compressible Ekman layers on curved boundaries. Studies in Applied Mathematics, 70(2):141–149, April 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1984702141.

Greenspan:1989:SEL

[Gre89] H. P. Greenspan. Stability of the Ekman layer in a mixture. Studies in Applied Mathematics, 81(1):21–32, August 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198981121.

Greenspan:1990:NSE

[Gre90] H. P. Greenspan. A note on spin up effects in a rotating mixture. Studies in Applied Mathematics, 82(1):49–58, January 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199082149.

Griggs:1977:ATP

[Gri77a] J. R. Griggs. Another three part Sperner theorem. Studies in Applied Mathematics, 57(2):181–184, October 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1977572181.

Grimshaw:1977:MIG

[Gri77b] R. H. J. Grimshaw. The modulation of an internal gravitywave packet, and the resonance with the mean motion. Studies in Applied Mathematics, 56(3):241–266, June 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1977563241.

Griggs:1978:UOP

[Gri78] J. R. Griggs. The uniqueness of odd pair designs. Studies in Applied Mathematics, 58(1):1–4, January 1978. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19785811.

Grimshaw:1981:EEL

[Gri81] R. Grimshaw. Evolution equations for long, nonlinear internal waves in stratified shear flows. Studies in Applied Mathematics, 65(2):159–188, October 1981. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1981652159.

Griggs:1982:PMS

[Gri82] Jerrold R. Griggs. Poset measure and saturated partitions. Studies in Applied Mathematics, 66(1):91–93, February 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198266191.

Grimshaw:1985:EEW

[Gri85] R. Grimshaw. Evolution equations for weakly nonlinear, long internal waves in a rotating fluid. Studies in Applied Mathematics, 73(1):1–33, August 1985. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19857311.

Grimshaw:1986:LCS

[Gri86] R. Grimshaw. Linearly coupled, slowly varying oscillators: The interaction of a positive-energy mode with a negative-energy mode. Studies in Applied Mathematics, 74(3):205–226, June 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1986743205.

Grimshaw:1987:TRW

[Gri87] R. Grimshaw. Triad resonance for weakly coupled, slowly varying oscillators. Studies in Applied Mathematics, 77(1):1–35, August 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19877711.

Grimshaw:1990:RFR

[Gri90] R. Grimshaw. Resonant flow of a rotating fluid past an obstacle: The general case. Studies in Applied Mathematics, 83(3):249–269, October 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1990833249.

Grimshaw:1995:WNS

[Gri95] R. Grimshaw. Weakly nonlocal solitary waves in a singularly perturbed nonlinear Schrödinger equation. Studies in Applied Mathematics, 94(3):257–270, April 1995. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995943257.

Grinfeld:2010:HDE

[Gri10] P. Grinfeld. Hamiltonian dynamic equations for fluid films. Studies in Applied Mathematics, 125(3):223–264, October 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00485.x.

Grinfeld:2012:SOS

[Gri12] P. Grinfeld. Small oscillations of a soap bubble. Studies in Applied Mathematics, 128(1):30–39, January 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00523.x.

Grimshaw:2015:CPP

[Gri15] Roger Grimshaw. Change of polarity for periodic waves in the variable-coefficient Korteweg–de Vries equation. Studies in Applied Mathematics, 134(3):363–371, April 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12067.

Grimshaw:2016:NWE

[Gri16] Roger Grimshaw. Nonlinear wave equations for oceanic internal solitary waves. Studies in Applied Mathematics, 136 (2):214–237, February 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12100.

Grimshaw:2019:ICC

[Gri19] Roger Grimshaw. Initial conditions for the cylindrical Korteweg–de Vries equation. Studies in Applied Mathematics, 143(2):176–191, August 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12272.

Garcia-Reimbert:2001:SDP

[GRMOH01] C. García-Reimbert, A. A. Minzoni, R. Ondarza, and J. J. E. Herrera. Simple diffraction processes for Nagumo- and Fishertype diffusion fronts. Studies in Applied Mathematics, 107 (4):367–389, November 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
1074191.

Grossberg:1969:LSP

[Gro69] Stephen Grossberg. On learning of spatiotemporal patterns by networks with ordered sensory and motor components 1. Excitatory components of the cerebellum. Studies in Applied Mathematics, 48(2):105–132, June 1969. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1969482105.

Grossberg:1970:SNC

[Gro70] Stephen Grossberg. Some networks that can learn, remember, and reproduce any number of complicated spacetime. Studies in Applied Mathematics, 49(2):135–166, June 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1970492135.

Grossberg:1973:CES

[Gro73] Stephen Grossberg. Contour enhancement, short term memory, and constancies in reverberating neural networks. Studies in Applied Mathematics, 52(3):213–257, September 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1973523213.

Gross:2002:WND

[Gro02] L. K. Gross. Weakly nonlinear dynamics of interface propagation. Studies in Applied Mathematics, 108(4):323–350, May 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.01415.

Gibbon:1985:PPH

[GRTW85] J. D. Gibbon, P. Radmore, M. Tabor, and D. Wood. The Painlevé property and Hirota’s method. Studies in Applied Mathematics, 72(1):39–63, February 1985. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198572139.

Gustavsson:1990:DDR

[GS90] L. Håkan Gustavsson and R. Shanthini. Degeneracies and direct resonances in water-table flow. Studies in Applied Mathematics, 83(1):61–89, July 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm199083161.

Grimshaw:2002:LWI

[GS02] Roger Grimshaw and Yuri Skyrnnikov. Long-wave instability in a three-layer stratified shear flow. Studies in Applied Mathematics, 108(1):77–88, January 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
01424.

Grundland:2006:SAS

[GŠ06] A. M. Grundland and L. Šnobl. Surfaces associated with sigma models. Studies in Applied Mathematics, 117(4):335–351, November 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9590.2006.00358.
x.

Gupta:2015:SRD

[GS15] Neelam Gupta and V. D. Sharma. On shock reflection– diffraction in a van der Waals gas. Studies in Applied Mathematics, 135(2):171–195, August 2015. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12084.

Grinevich:2016:NIS

[GS16] P. G. Grinevich and P. M. Santini. Nonlocality and the inverse scattering transform for the Pavlov equation. Studies in Applied Mathematics, 137(1):10–27, July 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12127.

Gion:2023:BBP

[GS23] Hiromu Gion and Yasuhisa Saito. Backward bifurcation and permanence of a disease-severity-structured epidemic model with treatment. Studies in Applied Mathematics, 150(4):1026– 1045, May 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12563.

Gil:2014:ANI

[GST14] Amparo Gil, Javier Segura, and Nico M. Temme. The asymptotic and numerical inversion of the Marcum Q-function. Studies in Applied Mathematics, 133(2):257–278, August 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12050.

Gil:2018:AAN

[GST18] Amparo Gil, Javier Segura, and Nico M. Temme. Asymptotic approximations to the nodes and weights of Gauss– Hermite and Gauss–Laguerre quadratures. Studies in Applied Mathematics, 140(3):298–332, April 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12201.

Gao:2021:NFO

[GSYG21] Juanjuan Gao, Jiebao Sun, Wenjuan Yao, and Zhichang Guo. A novel fractional-order reaction diffusion system for the multiplicative noise removal. Studies in Applied Mathematics, 147(2):421–447, August 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12386.

Gottlieb:1980:TDS

[GT80] David Gottlieb and Eli Turkel. On time discretizations for spectral methods. Studies in Applied Mathematics, 63(1):68– 86, August 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm198063167.

Grimshaw:1990:RMK

[GT90] R. Grimshaw and S. Tang. The rotation-modified Kadomtsev– Petviashvili equation: an analytical and numerical study. Studies in Applied Mathematics, 83(3):223–248, October 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990833223.

Gavrilyuk:2004:LSP

[GT04] S. L. Gavrilyuk and V. M. Teshukov. Linear stability of parallel inviscid flows of shallow water and bubbly fluid. Studies in Applied Mathematics, 113(1):1–29, July 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2004.01486.x.

Gubbins:1974:DAI

[Gub74] David Gubbins. Dynamo action of isotropically driven motions of a rotating fluid. Studies in Applied Mathematics, 53(2):157– 164, June 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1974532157.

Gomez-Ullate:2018:DRP

[GUGM18] David Gómez-Ullate, Yves Grandati, and Robert Milson. Durfee rectangles and pseudo-Wronskian equivalences for Hermite polynomials. Studies in Applied Mathematics, 141(4):596–625, November 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12225.

Gustavsson:1986:EDR

[Gus86] L. Håkan Gustavsson. Excitation of direct resonances in plane Poiseuille flow. Studies in Applied Mathematics, 75 (3):227–248, December 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1986753227.

Gustavsson:1989:DRN

[Gus89] L. Håkan Gustavsson. Direct resonance of nonaxisymmetric disturbances in pipe flow. Studies in Applied Mathematics, 80(2):95–108, April 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm198980295.

Gross:2003:WNS

[GV03] L. K. Gross and V. A. Volpert. Weakly nonlinear stability analysis of frontal polymerization. Studies in Applied Mathematics, 110(4):351–375, May 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00242.

Guo:2023:DRL

[GWCH23] Lijuan Guo, Lihong Wang, Lei Chen, and Jingsong He. Dynamics of the rogue lump in the asymmetric Nizhnik–Novikov– Veselov system. Studies in Applied Mathematics, 151(1):35–59, July 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12572.

Geng:2011:QPS

[GWH11] Xianguo Geng, Lihua Wu, and Guoliang He. Quasiperiodic solutions of nonlinear evolution equations associated with a 3 × 3 matrix spectral problem. Studies in Applied Mathematics, 127(2):107–140, August 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00513.x.

Gao:2016:PVD

[GXLX16] Shujing Gao, Lijun Xia, Yang Liu, and Dehui Xie. A plant virus disease model with periodic environment and pulse roguing. Studies in Applied Mathematics, 136(4):357–381, May 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12109.

Geng:2013:SCH

[GXW13] Xianguo Geng, Bo Xue, and Lihua Wu. A super Camassa– Holm equation with N–peakon solutions. Studies in Applied Mathematics, 130(1):1–16, January 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2012.00555.x.

Grimshaw:1993:RGF

[GY93] R. Grimshaw and Z. Yi. Resonant generation of finiteamplitude waves by flow past topography on a β-plane. Studies in Applied Mathematics, 88(2):89–112, February 1993. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199388289.

Gan:1995:NLG

[GY95a] Boon-Beng Gan and Yeong-Nan Yeh. A nim-like game and dynamic recurrence relations. Studies in Applied Mathematics, 95(2):213–228, August 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1995952213.

Gao:1995:MDM

[GY95b] David Yang Gao and Wei H. Yang. Multi-duality in minimal surface — type problems. Studies in Applied Mathematics, 95(2):127–146, August 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1995952127.

Gerdjikov:2015:CSM

[GY15] V. S. Gerdjikov and A. B. Yanovski. CBC systems with Mikhailov reductions by Coxeter automorphism: I. Spectral theory of the recursion operators. Studies in Applied Mathematics, 134(2):145–180, February 2015. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12065.

Guo:2021:SDA

[GYZ21] Wenjuan Guo, Ming Ye, and Qimin Zhang. Stability in distribution for age-structured HIV model with delay and driven by Ornstein–Uhlenbeck process. Studies in Applied Mathematics, 147(2):792–815, August 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12400.

Grimshaw:1994:OIB

[GZ94] R. Grimshaw and Y. Zhu. Oblique interactions between internal solitary waves. Studies in Applied Mathematics, 92(3):249– 270, July 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1994923249.

Guo:2010:STC

[GZ10] Zhengguang Guo and Yong Zhou. On solutions to a two-component generalized Camassa–Holm equation. Studies in Applied Mathematics, 124(3):307–322, April 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00472.x.

Guo:2013:WBM

[GZ13] Zhengguang Guo and Mingxuan Zhu. Wave breaking and measure of momentum support for an integrable Camassa–Holm system with two components. Studies in Applied Mathematics, 130(4):417–430, May 2013. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12004.

Grimshaw:2009:TFH

[GZC09] R. H. J. Grimshaw, D.-H. Zhang, and K. W. Chow. Transcritical flow over a hole. Studies in Applied Mathematics, 122(3):235–248, April 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2009.00431.x.

Gzyl:2000:MRS

[Gzy00a] Henryk Gzyl. Maxentropic reconstruction of some probability distributions. Studies in Applied Mathematics, 105 (3):235–243, October 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00150.

Gzyl:2000:PRT

[Gzy00b] Henryk Gzyl. Probabilistic remarks in the theory of scattering by soft obstacles. Studies in Applied Mathematics, 105(4):375–384, November 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00156.

Haberman:1972:CLM

[Hab72] Richard Haberman. Critical layers m parallel flows. Studies in Applied Mathematics, 51(2):139–161, June 1972. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1972512139.

Haberman:1975:SHO

[Hab75] Richard Haberman. Some higher-order nonlinear evolution equations. Studies in Applied Mathematics, 54(4):275–282, December 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1975544275.

Haberman:1977:NTL

[Hab77] Richard Haberman. Nonlinear transition layers — the second Painlevé transcendent. Studies in Applied Mathematics, 57(3):247–270, December 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1977573247.

Haberman:1988:MPS

[Hab88] Richard Haberman. The modulated phase shift for weakly dissipated nonlinear oscillatory waves of the Korteweg–de Vries type. Studies in Applied Mathematics, 78(1):73–90, February 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198878173.

Haberman:1991:PSM

[Hab91] Richard Haberman. Phase shift modulations for stable, oscillatory, traveling, strongly nonlinear waves. Studies in Applied Mathematics, 84(1):57–69, January 1991. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199184157.

Hagan:1981:INW

[Hag81] Patrick S. Hagan. The instability of nonmonotonic wave solutions of parabolic equations. Studies in Applied Mathematics, 64(1):57–88, February 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm198164157.

Hajela:1992:PLP

[Haj92] D. J. Hajela. On polynomials with low peak signal to power (L_{∞} to L_{2} norm) ratios and theorems of Kashin and Spencer. Studies in Applied Mathematics, 87(1):81–94, July 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199287181.

Hall:2001:NIV

[Hal01] Philip Hall. On the nonlinear initial value problem for vortex– wave interactions in shear flows. Studies in Applied Mathematics, 106(1):69–83, January 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00161.

Hall:2006:NEE

[Hal06] Philip Hall. Nonlinear evolution equations and the braiding of weakly transporting flows over gravel beds. Studies in Applied Mathematics, 117(1):27–69, July 2006. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2006.00341.x.

Hall:2012:VWI

[Hal12] Philip Hall. Vortex–wave Interactions/Self-Sustained processes in high Prandtl number natural convection in a vertical channel with moving sidewalls. Studies in Applied Mathematics, 129(1):1–25, July 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2011.00543.x.

Harper:1975:NSC

[Har75] L. H. Harper. A note on some classes of Boolean functions. Studies in Applied Mathematics, 54(2):161–164, June 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1975542161.

Harris:1998:ACC

[Har98] Adam Harris. Analytic continuation of complex gauge fields. Studies in Applied Mathematics, 101(3):309–319, October 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00095.

Haragus:2011:TSS

[Har11] Mariana Haragus. Transverse spectral stability of small periodic traveling waves for the KP equation. Studies in Applied Mathematics, 126(2):157–185, February 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00501.x.

Hastings:1976:PPW

[Has76] S. P. Hastings. Periodic plane waves for the Oregonator. Studies in Applied Mathematics, 55(4):293–299, December 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1976554293.

Hwang:2012:SWT

[HAY12] G. Hwang, T. R. Akylas, and J. Yang. Solitary waves and their linear stability in nonlinear lattices. Studies in Applied Mathematics, 128(3):275–298, April 2012. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00538.x.

Hector:1972:RME

[HCB72] D. Hector, J. Cohen, and N. Bleistein. Ray method expansions for surface and internal waves in inhomogeneous oceans of variable depth. Studies in Applied Mathematics, 51(2):121–137, June 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1972512121.

Henderson:2021:VBP

[HCC21] Diane M. Henderson, John D. Carter, and Megan E. Catalano. On the variation of bi-periodic waves in the transverse direction. Studies in Applied Mathematics, 147(4):1388–1408, November 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12446.

Heilmann:1971:LZG

[Hei71] Ole J. Heilmann. Location of the zeros of the grand partition function of certain classes of lattice gases. Studies in Applied Mathematics, 50(4):385–390, December 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1971504385.

Hellinger:1974:EAR

[Hel74] Steven J. Hellinger. Equilibrium arrival rates at a partiallyobstructed intersection. Studies in Applied Mathematics, 53(1):65–71, March 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm197453165.

Henle:1972:DGF

[Hen72] Michael Henle. Dissection of generating functions. Studies in Applied Mathematics, 51(4):397–410, December 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1972514397.

Henningson:1988:IIV

[Hen88] Dan S. Henningson. The inviscid initial value problem for a piecewise linear mean flow. Studies in Applied Mathematics, 78(1):31–56, February 1988. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198878131.

Hershenov:1976:SDE

[Her76] Joseph Hershenov. Solutions of the differential equation u′′′ + λ^{2}zu′ + (α − 1)λ^{2}u = 0. Studies in Applied Mathematics, 55(4):301–314, December 1976. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1976554301.

Herron:1982:DEI

[Her82] Isom H. Herron. A diffusion equation illustrating spectral theory for boundary layer stability. Studies in Applied Mathematics, 67(3):231–241, December 1982. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1982673231.

Herron:1991:ORV

[Her91] Isom H. Herron. Observations on the role of vorticity in the stability theory of wall bounded flows. Studies in Applied Mathematics, 85(3):269–286, October 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991853269.

Hitzazis:2017:LEP

[HF17] I. Hitzazis and A. S. Fokas. Linear elliptic PDEs in a cylindrical domain with a polygonal cross-section. Studies in Applied Mathematics, 139(2):288–321, August 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12187.

Hunter:1981:EME

[HG81] C. Hunter and B. Guerrieri. The eigenvalues of Mathieu’s equation and their branch points. Studies in Applied Mathematics, 64(2):113–141, April 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1981642113.

Hunter:1982:EAS

[HG82] C. Hunter and B. Guerrieri. The eigenvalues of the angular spheroidal wave equation. Studies in Applied Mathematics, 66(3):217–240, June 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1982663217.

Hooper:2001:TLG

[HG01] A. P. Hooper and R. Grimshaw. Transient linear growth and nonlinear effects. Studies in Applied Mathematics, 106(1):47–68, January 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00160.

Haj-Hariri:1994:CAP

[HH94] Hossein Haj-Hariri. Characteristics analysis of the parabolized stability equations. Studies in Applied Mathematics, 92(1):41– 53, May 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm199492141.

Haj-Hariri:1986:SRI

[HHA86] H. Haj-Hariri and T. R. Akylas. Sound radiation by instability wavepackets in a boundary layer. Studies in Applied Mathematics, 75(1):57–76, August 1986. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198675157.

Haj-Hariri:1994:SPE

[HHW94] Hossein Haj-Hariri and Hubertus J. Weinitschke. Singular perturbations of elliptic problems on domains with small holes. Studies in Applied Mathematics, 92(1):55–93, May 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199492155.

Hu:2020:LNE

[HHZ20] Pipi Hu, Liu Hong, and Yi Zhu. Linear and nonlinear electromagnetic waves in modulated honeycomb media. Studies in Applied Mathematics, 144(1):18–45, January 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12284.

Hickernell:1983:ELHa

[Hic83a] Fred J. Hickernell. The evolution of large-horizontal-scale disturbances in marginally stable, inviscid, shear flows. I. Derivation of the amplitude evolution equations. Studies in Applied Mathematics, 69(1):1–21, August 1983. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19836911.

Hickernell:1983:ELHb

[Hic83b] Fred J. Hickernell. The evolution of large-horizontal-scale disturbances in marginally stable, inviscid, shear flows. II. Solutions of the Boussinesq equation. Studies in Applied Mathematics, 69(1):23–49, August 1983. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198369123.

Hickernell:1985:UBG

[Hic85] Fred J. Hickernell. An upper bound on the growth rate of a linear instability in a homogeneous shear flow. Studies in Applied Mathematics, 72(1):87–93, February 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198572187.

Hirota:2009:NSU

[Hir09] Ryogo Hirota. New solutions to the ultradiscrete soliton equations. Studies in Applied Mathematics, 122(4):361–376, May 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00438.x.

Halburd:1996:CLS

[HJ96] Rod Halburd and Nalini Joshi. The coalescence limit of the second Painlevé equation. Studies in Applied Mathematics, 97(1): 1–15, July 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19969711.

Hur:2015:MIW

[HJ15] Vera Mikyoung Hur and Mathew A. Johnson. Modulational instability in the Whitham equation for water waves. Studies in Applied Mathematics, 134(1):120–143, January 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12061.

Haberman:1985:NCC

[HjS85] Richard Haberman and Ren ji Sun. Nonlinear cusped caustics for dispersive waves. Studies in Applied Mathematics, 72(1):1– 37, February 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19857211.

Hsieh:1973:NMD

[HK73] W. N. Hsieh and D. J. Kleitman. Normalized matching in direct products of partial orders. Studies in Applied Mathematics, 52(3):285–289, September 1973. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1973523285.

Halasz:1974:NRT

[HK74] Sylvia Halász and D. J. Kleitman. A note on random triangles. Studies in Applied Mathematics, 53(3):225–237, September 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1974533225.

Howard:1977:SVW

[HK77] L. N. Howard and N. Kopell. Slowly varying waves and shock structures in reaction–diffusion equations. Studies in Applied Mathematics, 56(2):95–145, April 1977. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197756295.

Hoffmann:2004:DCT

[HK04] Tim Hoffmann and Nadja Kutz. Discrete curves in CP and the Toda lattice. Studies in Applied Mathematics, 113(1):31–55, July 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2004.01493.x.

He:2011:DIV

[HL11] J. He and Y. Li. Designable integrability of the variable coefficient nonlinear Schrödinger equations. Studies in Applied Mathematics, 126(1):1–15, January 2011. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00495.x.

Hu:2009:NSB

[HLQ09] Xing-Biao Hu, Sen-Yue Lou, and Xian-Min Qian. Nonlocal symmetries for bilinear equations and their applications. Studies in Applied Mathematics, 122(3):305–324, April 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00435.x.

He:2003:TCE

[HLS03] Li He, Xiangwei Liu, and Gilbert Strang. Trees with Cantor eigenvalue distribution. Studies in Applied Mathematics, 110(2):123–138, February 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00233.

Hoppensteadt:1977:MMS

[HM77] Frank C. Hoppensteadt and Willard L. Miranker. Multitime methods for systems of difference equations. Studies in Applied Mathematics, 56(3):273–289, June 1977. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1977563273.

Haubold:1986:ART

[HM86] H. J. Haubold and A. M. Mathai. Analytic representations of thermonuclear reaction rates. Studies in Applied Mathematics, 75(2):123–137, October 1986. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1986752123.

He:1991:SPN

[HM91] Yuanping He and T. B. Moodie. The signaling problem in nonlinear hyperbolic wave theory. Studies in Applied Mathematics, 85(4):343–372, November 1991. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991854343.

He:1993:SWT

[HM93a] Yuanping He and T. Bryant Moodie. Shock wave tracking for hyperbolic systems exhibiting local linear degeneracy. Studies in Applied Mathematics, 89(3):195–232, August 1993. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1993893195.

He:1993:TWI

[HM93b] Yuanping He and T. Bryant Moodie. Two-wave interactions for weakly nonlinear hyperbolic waves. Studies in Applied Mathematics, 88(3):241–268, April 1993. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1993883241.

He:1994:SCL

[HM94a] Yuanping He and T. Bryant Moodie. Scalar conservation laws and spatially dependent flux functions. Studies in Applied Mathematics, 91(3):215–245, March 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1994913215.

He:1994:TWI

[HM94b] Yuanping He and T. Bryant Moodie. Two-wave interactions for weakly nonlinear hyperbolic waves. Studies in Applied Mathematics, 91(3):275–277, March 1994. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1994913275.

Hoseini:2008:EHO

[HM08] S. M. Hoseini and T. R. Marchant. Evolution of higherorder Gray Hirota solitary waves. Studies in Applied Mathematics, 121(2):117–139, August 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2008.00407.x.

Himonas:2015:IBV

[HM15] A. Alexandrou Himonas and Dionyssios Mantzavinos. On the initial-boundary value problem for the linearized Boussinesq equation. Studies in Applied Mathematics, 134(1):62–100, January 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12055.

Hu:2022:DND

[HMBYZ22] Jing Hu, Anke Meyer-Baese, Ming Ye, and Qimin Zhang. Dynamics of a nonlocal dispersal in-host viral model with humoral immunity. Studies in Applied Mathematics, 149(3):705–728, October 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12518.

Herron:1987:MEI

[HMM87] I. H. Herron, S. A. Maslowe, and S. Melkonian. A model equation illustrating subcritical instability to long waves in shear flows. Studies in Applied Mathematics, 76(3):265–275, June 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1987763265.

Hunter:1986:RIW

[HMR86] J. K. Hunter, A. Majda, and R. Rosales. Resonantly interacting, weakly nonlinear hyperbolic waves. II. Several space variables. Studies in Applied Mathematics, 75(3):187–226, December 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1986753187.

Henkin:1996:RAP

[HN96] G. M. Henkin and N. N. Novikova. The reconstruction of the attracting potential in the Sturm–Liouville equation through characteristics of negative discrete spectrum. Studies in Applied Mathematics, 97(1):17–52, July 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199697119.

Hochstadt:1972:SOD

[Hoc72] Harry Hochstadt. A second order differential operator with eigenvalue degeneracy. Studies in Applied Mathematics, 51(1):99–105, March 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm197251199.

Holladay:1980:PTP

[Hol80] Kenneth Holladay. A partition theory of planar animals. Studies in Applied Mathematics, 63(2):169–183, October 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980632169.

Holba:2023:CCL

[Hol23] Pavel Holba. Complete classification of local conservation laws for generalized Cahn–Hilliard–Kuramoto–Sivashinsky equation. Studies in Applied Mathematics, 151(1):171–182, July 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12576.

Howard:1973:SCS

[How73] Louis N. Howard. On the stability of compressible swirling flow. Studies in Applied Mathematics, 52(1):39–43, March 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197352139.

Howes:1977:SPN

[How77] F. A. Howes. Singularly perturbed nonlinear boundary-value problems whose reduced equations have singular points. Studies in Applied Mathematics, 57(2):135–180, October 1977.

CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1977572135.

Howes:1979:SPS

[How79] F. A. Howes. Singularly perturbed semilinear systems. Studies in Applied Mathematics, 61(3):185–209, December 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1979613185.

Howes:1982:SPS

[How82] F. A. Howes. Singularly perturbed semilinear Robin problems. Studies in Applied Mathematics, 67(2):125–139, October 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1982672125.

Howes:1983:NDS

[How83] F. A. Howes. Nonlinear dispersive systems: Theory and examples. Studies in Applied Mathematics, 69(1):75–97, August 1983. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198369175.

Howes:1986:SSRa

[How86a] F. A. Howes. Some stability results for advection–diffusion equations. Studies in Applied Mathematics, 74(1):35–53, February 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm198674135.

Howes:1986:SSRb

[How86b] F. A. Howes. Some stability results for advection–diffusion equations. II. Studies in Applied Mathematics, 75(2):153–162, October 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1986752153.

Howard:1990:LTH

[How90a] Louis N. Howard. Limits on the transport of heat and momentum by turbulent convection with large-scale flow. Studies in Applied Mathematics, 83(4):273–285, November 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990834273.

Howes:1990:TDA

[How90b] F. A. Howes. Two-dimensional advection–diffusion equations with constant limiting solutions. Studies in Applied Mathematics, 82(3):245–265, April 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1990823245.

Howes:1978:MNP

[HP78] F. A. Howes and S. V. Parter. A model nonlinear problem having a continuous locus of singular points. Studies in Applied Mathematics, 58(3):249–262, June 1978. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1978583249.

Hur:2019:MIF

[HP19] Vera Mikyoung Hur and Ashish Kumar Pandey. Modulational instability in a full-dispersion shallow water model. Studies in Applied Mathematics, 142(1):3–47, January 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12231.

Howard:1969:IVP

[HS69] Louis N. Howard and William L. Siegmann. On the initial value problem for rotating stratified flow. Studies in Applied Mathematics, 48(2):153–169, June 1969. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1969482153.

Hall:1982:SMN

[HS82a] P. Hall and F. T. Smith. A suggested mechanism for nonlinear wall roughness effects on high Reynolds number flow stability. Studies in Applied Mathematics, 66(3):241–265, June 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1982663241.

Holmes:1982:EOD

[HS82b] Philip Holmes and D. S. Stewart. The existence of one dimensional steady detonation waves in a simple model problem. Studies in Applied Mathematics, 66(2):121–143, April 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1982662121.

Hall:1984:ENT

[HS84] P. Hall and F. T. Smith. On the effects of nonparallelism, three-dimensionality, and mode interaction in nonlinear boundary-layer stability. Studies in Applied Mathematics, 70(2):91–120, April 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm198470291.

Henningson:1992:VEE

[HS92] Dan S. Henningson and Peter J. Schmid. Vector eigenfunction expansions for plane channel flows. Studies in Applied Mathematics, 87(1):15–43, July 1992. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199287115.

Hunter:2019:RRW

[HS19] John K. Hunter and Evan B. Smothers. On the resonant reflection of weak, nonlinear sound waves off an entropy wave. Studies in Applied Mathematics, 143(2):157–175, August 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12271.

Hsieh:1973:PID

[Hsi73] W. N. Hsieh. Proportions of irreducible diagrams. Studies in Applied Mathematics, 52(3):277–283, September 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1973523277.

Hakkaev:2017:SSC

[HSS17] S. Hakkaev, M. Stanislavova, and A. Stefanov. Spectral stability for classical periodic waves of the Ostrovsky and short pulse models. Studies in Applied Mathematics, 139(3):405–433, October 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12166.

Hoefer:2019:MTS

[HSS19] Mark A. Hoefer, Noel F. Smyth, and Patrick Sprenger. Modulation theory solution for nonlinearly resonant, fifth-order Korteweg–de Vries, nonclassical, traveling dispersive shock waves. Studies in Applied Mathematics, 142(3):219–240, April 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12246.

Hsieh:1998:DNE

[HTWW98] D. Y. Hsieh, S. Q. Tang, X. P. Wang, and L. X. Wu. Dissipative nonlinear evolution equations and chaos. Studies in Applied Mathematics, 101(3):233–266, October 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00092.

Hu:1996:ABA

[Hu96] Jishan Hu. Asymptotics beyond all orders for a certain type of nonlinear oscillators. Studies in Applied Mathematics, 96(1):85–109, January 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm199696185.

Hu:1998:LBA

[Hu98] Jishan Hu. Linkage between adiabatic invariance and symmetry of solutions. Studies in Applied Mathematics, 101(2):99–126, August 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00087.

Hunter:1969:SSO

[Hun69] C. Hunter. The spectra of small oscillations of thin disk galactic models. Studies in Applied Mathematics, 48(1):55–76, March 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm196948155.

Hunter:1970:SHE

[Hun70] C. Hunter. Stellar hydrodynamic equations for a thin disk galaxy. Studies in Applied Mathematics, 49(1):59–68, March 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197049159.

Hunter:1992:IEW

[Hun92] John K. Hunter. Interaction of elastic waves. Studies in Applied Mathematics, 86(4):281–314, May 1992. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1992864281.

Huang:2020:UAZ

[HW20] Xiao-Min Huang and R. Wong. Uniform asymptotics and zeros of the associated Pollaczek polynomials. Studies in Applied Mathematics, 145(4):625–646, November 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12292.

Hadadifard:2021:MML

[HW21] Fazel Hadadifard and J. Douglas Wright. Mass-in-mass lattices with small internal resonators. Studies in Applied Mathematics, 146(1):81–98, January 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12340.

Hwang:1997:AEE

[Hwa97] Hsien-Kuei Hwang. Asymptotic estimates of elementary probability distributions. Studies in Applied Mathematics, 99 (4):393–417, November 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00065.

Hwang:2014:FMD

[Hwa14] Guenbo Hwang. The Fokas method: The Dirichlet to Neumann map for the sine-Gordon equation. Studies in Applied Mathematics, 132(4):381–406, May 2014. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12035.

Hickernell:1986:LSM

[HY86] F. J. Hickernell and Y. C. Yortsos. Linear stability of miscible displacement processes in porous media in the absence of dispersion. Studies in Applied Mathematics, 74(2):93–115, April 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198674293.

Hu:2000:MSI

[HY00] Jishan Hu and Min Yan. The mirror systems of integrable equations. Studies in Applied Mathematics, 104(1):67–90, January 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00131.

Hu:2010:OCV

[HZ10] Ting Hu and Ding-Xuan Zhou. Online classification with varying Gaussians. Studies in Applied Mathematics, 124(1):65–83, January 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9590.2009.00462.
x.

Iacono:2015:KEE

[IB15] Roberto Iacono and John P. Boyd. The Kidder equation: u_{xx}+ 2xu_{x}/\sqrt{1 − αu} = 0. Studies in Applied Mathematics, 135(1): 63–85, July 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12073.

Ifidon:2010:SVE

[Ifi10] E. O. Ifidon. Stuart vortices extended to a sphere admits infinite-dimensional generalized symmetries. Studies in Applied Mathematics, 125(3):265–273, October 2010. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00487.x.

Ismail:2020:DEA

[IL20] Mourad Ismail and Chun-Kong Law. A difference equation approach to Plancherel–Rotach asymptotics for q-orthogonal polynomials. Studies in Applied Mathematics, 145(4):696–718, November 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12332.

Ioakimidis:1989:NCF

[Ioa89] N. I. Ioakimidis. A note on the closed-form determination of zeros and poles of generalized analytic functions. Studies in Applied Mathematics, 81(3):265–269, December 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1989813265.

Ierley:1986:ANS

[IR86] Glenn R. Ierley and Otto G. Ruehr. Analytic and numerical solutions of a nonlinear boundary-layer problem. Studies in Applied Mathematics, 75(1):1–36, August 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19867511.

Ioannou-Sougleridis:2020:DSD

[ISFH20] Ioannis Ioannou-Sougleridis, Dimitrios J. Frantzeskakis, and Theodoros P. Horikis. A Davey–Stewartson description of two-dimensional solitons in nonlocal media. Studies in Applied Mathematics, 144(1):3–17, January 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12283.

Israeli:1970:FIN

[Isr70] M. Israeli. A fast implicit numerical method for time dependent viscous flows. Studies in Applied Mathematics, 49 (4):327–349, December 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1970494327.

Israeli:1972:EIP

[Isr72a] M. Israeli. On the evaluation of iteration parameters for the boundary vorticity. Studies in Applied Mathematics, 51(1):67– 71, March 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm197251167.

Israeli:1972:TMR

[Isr72b] Moshe Israeli. On trapped modes of rotating fluids in spherical shells. Studies in Applied Mathematics, 51(3):219–237, September 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1972513219.

Iserles:2023:SOS

[IW23] Arieh Iserles and Marcus Webb. Sobolev-orthogonal systems with tridiagonal skew-Hermitian differentiation matrices. Studies in Applied Mathematics, 150(2):420–447, February 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12544.

Iwahori:1982:SOC

[Iwa82] Nobuko Iwahori. On some optimal constructions of chess tournament combinations. Studies in Applied Mathematics, 66(1):53–67, February 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm198266153.

Jutley:2016:SSB

[JA16] Jessica J. Jutley and Alejandro B. Aceves. Spontaneous PT - symmetry breaking in nonlinearly coupled van der Pol oscillators. Studies in Applied Mathematics, 137(2):256–270, August 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12139.

Jo:2002:DSN

[JC02] Tae-Chang Jo and Wooyoung Choi. Dynamics of strongly nonlinear internal solitary waves in shallow water. Studies in Applied Mathematics, 109(3):205–227, October 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00222.

Jo:2008:SSN

[JC08] Tae-Chang Jo and Wooyoung Choi. On stabilizing the strongly nonlinear internal wave model. Studies in Applied Mathematics, 120(1):65–85, January 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2007.00393.x.

Junior:2021:GCA

[JFM21] Altemir Bortuli Junior, Igor Leite Freire, and Norberto Anibal Maidana. Group classification and analytical solutions of a radially symmetric avascular cancer model. Studies in Applied Mathematics, 147(3):978–1006, October 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12416.

Jackson:1979:FCEa

[JG79] D. M. Jackson and I. P. Goulden. A formal calculus for the enumerative system of sequences — I. Combinatorial theorems. Studies in Applied Mathematics, 61(2):141–178, October 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1979612141.

Jackson:1980:FCE

[JG80] D. M. Jackson and I. P. Goulden. A formal calculus for the enumerative system of sequences — III. Further developments. Studies in Applied Mathematics, 62(2):113–141, April 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980622113.

Johnston:2021:LCB

[JGM21] C. M. Johnston, Clarence T. Gartman, and Dionyssios Mantzavinos. The linearized classical Boussinesq system on the half-line. Studies in Applied Mathematics, 146(3):635–657, April 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12359.

Jin:2023:UBE

[Jin23] Chunhua Jin. Uniform boundedness and eventual Hölder continuity to a cancer invasion model with remodeling of ECM and nonlinear diffusion. Studies in Applied Mathematics, 151(1):183–207, July 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12577.

Jichang:1984:SLP

[JK84] Sha Jichang and D. J. Kleitman. Superantichains in the lattice of partitions of a set. Studies in Applied Mathematics, 71(3):207–241, December 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1984713207.

Joshi:1992:PCP

[JK92] Nalini Joshi and Martin D. Kruskal. The Painlevé connection problem: an asymptotic approach. I. Studies in Applied Mathematics, 86(4):315–376, May 1992. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1992864315.

Joshi:1994:DPS

[JK94] Nalini Joshi and Martin D. Kruskal. A direct proof that solutions of the six Painlevé equations have no movable singularities except poles. Studies in Applied Mathematics, 93 (3):187–207, November 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1994933187.

Joshi:2001:BTS

[JK01] N. Joshi and A. V. Kitaev. On Boutroux’s Tritronquée solutions of the first Painlevé equation. Studies in Applied Mathematics, 107(3):253–291, October 2001. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00187.

Joshi:2019:GSW

[JL19] N. Joshi and C. J. Lustri. Generalized solitary waves in a finite-difference Korteweg–de Vries equation. Studies in Applied Mathematics, 142(3):359–384, April 2019.

CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12252.

Jackson:1979:FCEb

[JlPG79] D. M. Jackson and l. P. Goulden. A formal calculus for the enumerative system of sequences — II. Applications. Studies in Applied Mathematics, 61(3):245–277, December 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1979613245.

Jacobson:2008:MCC

[JMT08] Tivon Jacobson, Paul A. Milewski, and Esteban G. Tabak. Mixing closures for conservation laws in stratified flows. Studies in Applied Mathematics, 121(1):89–116, July 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00403.x.

Joshi:2021:TDC

[JN21] Nalini Joshi and Nobutaka Nakazono. On the threedimensional consistency of Hirota’s discrete Korteweg–de Vries equation. Studies in Applied Mathematics, 147(4):1409–1424, November 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12421.

Jamison:1989:RGC

[JO89a] B. Jamison and S. Olariu. P_{4}-reducible graphs — class of uniquely tree-representable graphs. Studies in Applied Mathematics, 81(1):79–87, August 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm198981179.

Jamison:1989:NCB

[JO89b] B. Jamison and S. Olariu. A new class of brittle graphs. Studies in Applied Mathematics, 81(1):89–92, August 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198981189.

Jamison:1989:CFG

[JO89c] B. Jamison and S. Olariu. On a class of P_{5}-free graphs. Studies in Applied Mathematics, 81(1):33–39, August 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198981133.

Joshi:2007:GTL

[JO07] Nalini Joshi and Chris Ormerod. The general theory of linear difference equations over the max-plus semi-ring. Studies in Applied Mathematics, 118(1):85–97, January 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2007.00364.x.

Johnson:2010:TIP

[Joh10] Mathew A. Johnson. The transverse instability of periodic waves in Zakharov–Kuznetsov type equations. Studies in Applied Mathematics, 124(4):323–345, May 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00473.x.

Joni:1980:MEO

[Jon80] S. A. Joni. Multivariate exponential operators. Studies in Applied Mathematics, 62(2):175–182, April 1980. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980622175.

Joshi:1999:SPE

[Jos99] Nalini Joshi. The second Painlevé equation in the largeparameter limit i: Local asymptotic analysis. Studies in Applied Mathematics, 102(4):345–373, May 1999. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00114.

Joshi:2015:QSD

[Jos15] Nalini Joshi. Quicksilver solutions of a q-difference first Painlevé equation. Studies in Applied Mathematics, 134 (2):233–251, February 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12066.

Johnson:2022:SDW

[JP22] Mathew A. Johnson and Wesley R. Perkins. Subharmonic dynamics of wave trains in the Korteweg–de Vries/Kuramoto– Sivashinsky equation. Studies in Applied Mathematics, 148 (3):1274–1302, April 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12475.

Joshi:2000:NRK

[JPS00] Nalini Joshi, Johannes Petersen, and Luke Schubert. Nonexistence results for the Korteweg–de Vries and Kadomtsev– Petviashvili equations. Studies in Applied Mathematics, 105 (4):361–374, November 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00155.

Ji:2013:CLB

[JQ13] Lina Ji and Changzheng Qu. Conditional Lie–Bäcklund symmetries and invariant subspaces to nonlinear diffusion equations with convection and source. Studies in Applied Mathematics, 131(3):266–301, October 2013. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12010.

Ji:2014:CLB

[JQS14] Lina Ji, Changzheng Qu, and Shoufeng Shen. Conditional Lie–Bäcklund symmetry of evolution system and application for reaction–diffusion system. Studies in Applied Mathematics, 133(1):118–149, July 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12042.

Joni:1979:CBC

[JR79] S. A. Joni and G.-C. Rota. Coalgebras and bialgebras in combinatorics. Studies in Applied Mathematics, 61(2):93–139, October 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm197961293.

Ji:1999:MCS

[JRS99] Hui Ji, Sherman D. Riemenschneider, and Zuowei Shen. Multivariate compactly supported fundamental refinable functions, duals, and biorthogonal wavelets. Studies in Applied Mathematics, 102(2):173–204, February 1999. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00108.

Jones:1999:PIM

[JS99] Don A. Jones and Steve Shkoller. Persistence of invariant manifolds for nonlinear PDEs. Studies in Applied Mathematics, 102(1):27–67, January 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00103.

Joseph:2001:IBV

[JS01] K. T. Joseph and P. L. Sachdev. Initial boundary value problems for scalar and vector Burgers equations. Studies in Applied Mathematics, 106(4):481–505, May 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00175.

Jorge:1987:LPA

[JT87] L. P. Jorge and K. Tenenblat. Linear problems associated with evolution equations of the type u_{tt} = F (u, u_{x}, u_{xx}, u_{t}). Studies in Applied Mathematics, 77(2):103–117, October 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1987772103.

Jiang:2013:BFS

[JTLW13] Yan Jiang, Bo Tian, Min Li, and Pan Wang. Bilinear form and N–shock-wave solutions for a (2 + 1)- dimensional breaking soliton equation in certain fluids with the Bell polynomials and auxiliary function. Studies in Applied Mathematics, 131(4):331–342, November 2013. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12012.

Johnson:2022:SWW

[JTW22] Mathew A. Johnson, Tien Truong, and Miles H. Wheeler. Solitary waves in a Whitham equation with small surface tension. Studies in Applied Mathematics, 148(2):773–812, February 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12459.

Johnson:1969:ADT

[JW69] M. W. Johnson and O. E. Widera. An asymptotic dynamic theory for cylindrical shells. Studies in Applied Mathematics, 48(3):205–226, September 1969. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1969483205.

Jiang:1999:JPA

[JW99] X.-H. Jiang and R. Wong. Justification of a perturbation approximation of the Klein–Gordon equation. Studies in Applied Mathematics, 102(4):375–417, May 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00115.

Jiang:2006:AAP

[JW06] X. H. Jiang and R. Wong. Asymptotic analysis of a perturbed periodic solution for the KdV equation. Studies in Applied Mathematics, 116(1):21–33, January 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2005.00332.x.

Johnson:2020:GSW

[JW20] Mathew A. Johnson and J. Douglas Wright. Generalized solitary waves in the gravity-capillary Whitham equation. Studies in Applied Mathematics, 144(1):102–130, January 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12288.

Julian:1986:MFA

[JZ86] William H. Julian and Joseph D. Zund. The motion of a free asymmetric top in n dimensions. Studies in Applied Mathematics, 75(2):171–177, October 1986. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1986752171.

Johnson:2010:RJW

[JZ10] M. A. Johnson and K. Zumbrun. Rigorous justification of the Whitham modulation equations for the generalized Korteweg–de Vries equation. Studies in Applied Mathematics, 125(1):69–89, July 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2010.00482.x.

Kodama:1981:PSS

[KA81] Yuji Kodama and Mark J. Ablowitz. Perturbations of solitons and solitary waves. Studies in Applied Mathematics, 64(3):225– 245, June 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1981643225.

Kantzios:1988:LNW

[KA88] Yorgos D. Kantzios and T. R. Akylas. Long nonlinear water waves in a channel near a cut-off frequency. Studies in Applied Mathematics, 78(1):57–72, February 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198878157.

Konopelchenko:2002:NDP

[KA02] B. Konopelchenko and Martínez L. Alonso. Nonlinear dynamics on the plane and integrable hierarchies of infinitesimal deformations. Studies in Applied Mathematics, 109 (4):313–336, November 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00226.

Kataoka:2019:MFG

[KA19] Takeshi Kataoka and T. R. Akylas. On mean flow generation due to oblique reflection of internal waves at a slope. Studies in Applied Mathematics, 142(4):419–432, May 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12257.

Kallman:1973:CGP

[Kal73] Robert R. Kallman. A conjecture of Gilbert and Pollak on minimal trees. Studies in Applied Mathematics, 52(2):141–151, June 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1973522141.

Kath:1981:RPP

[Kat81] William L. Kath. Resonance in periodically perturbed Hopf bifurcation. Studies in Applied Mathematics, 65(2):95–112, October 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm198165295.

Kath:1985:SVP

[Kat85] William L. Kath. Slowly varying phase planes and boundarylayer theory. Studies in Applied Mathematics, 72(3):221–239, June 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1985723221.

Kato:1988:DMW

[Kat88] Yoshifumi Kato. The determinants of matrices whose elements decrease geometrically in diagonal direction. Studies in Applied Mathematics, 78(3):167–176, June 1988. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1988783167.

Kaup:1975:MSS

[Kau75] D. J. Kaup. Method for solving the sine-Gordon equation in laboratory coordinates. Studies in Applied Mathematics, 54(2):165–179, June 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1975542165.

Kaup:1976:TWI

[Kau76] D. J. Kaup. The three-wave interaction — a nondispersive phenomenon. Studies in Applied Mathematics, 55(1):9–44, March 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19765519.

Kaup:1978:SHG

[Kau78] D. J. Kaup. Simple harmonic generation: an exact method of solution. Studies in Applied Mathematics, 59(1):25–35, July 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197859125.

Kaup:1980:MSS

[Kau80a] D. J. Kaup. A method for solving the separable initialvalue problem of the full three-dimensional three-wave interaction. Studies in Applied Mathematics, 62(1):75–83, February 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198062175.

Kaup:1980:ISP

[Kau80b] D. J. Kaup. On the inverse scattering problem for cubic eigenvalue problems of the class ψ_{xxx} + 6Qψ_{x} + 6R_{ψ} = λψ. Studies in Applied Mathematics, 62(3):189–216, June 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980623189.

Kaup:2006:ESC

[Kau06] D. J. Kaup. Evolution of the scattering coefficients of the Camassa–Holm equation, for general initial data. Studies in Applied Mathematics, 117(2):149–164, August 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2006.00350.x.

Kevorkian:1998:MSH

[KB98] J. Kevorkian and D. L. Bosley. Multiple-scale homogenization for weakly nonlinear conservation laws with rapid spatial fluctuations. Studies in Applied Mathematics, 101(2):127–183, August 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00088.

Kath:1982:WTB

[KC82] William L. Kath and Donald S. Cohen. Waiting-time behavior in a nonlinear diffusion equation. Studies in Applied Mathematics, 67(2):79–105, October 1982. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198267279.

Kruskal:1992:PKP

[KC92] Martin D. Kruskal and Peter A. Clarkson. The Painlevé– Kowalevski and poly-Painlevé tests for integrability. Studies in Applied Mathematics, 86(2):87–165, February 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199286287.

King:2005:SSS

[KC05] J. R. King and S. M. Cox. Self-similar “Stagnation point” boundary layer flows with suction or injection. Studies in Applied Mathematics, 115(1):73–107, July 2005. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2005.01563.

Konopelchenko:1992:IST

[KD92] B. G. Konopelchenko and V. G. Dubrovsky. Inverse spectral transform for the modified Kadomtsev–Petviashvili equation. Studies in Applied Mathematics, 86(3):219–268, April 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1992863219.

Khwaja:2017:CCA

[KD17] S. Farid Khwaja and A. B. Olde Daalhuis. Computation of the coefficients appearing in the uniform asymptotic expansions of integrals. Studies in Applied Mathematics, 139(4):551–567, November 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12172.

Kaptsov:2022:CIF

[KDM22] Evgeniy I. Kaptsov, Vladimir A. Dorodnitsyn, and Sergey V. Meleshko. Conservative invariant finite-difference schemes for the modified shallow water equations in Lagrangian coordinates. Studies in Applied Mathematics, 149(3):729–761, October 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12519.

Keener:1976:SBN

[Kee76] J. P. Keener. Secondary bifurcation in nonlinear diffusion reaction equations. Studies in Applied Mathematics, 55(3):187–211, September 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1976553187.

Keener:1978:AIP

[Kee78] James P. Keener. Activators and inhibitors in pattern formation. Studies in Applied Mathematics, 59(1):1–23, July 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19785911.

Keener:1982:CBS

[Kee82] James P. Keener. Chaotic behavior in slowly varying systems of nonlinear differential equations. Studies in Applied Mathematics, 67(1):25–44, August 1982. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198267125.

Keefe:1985:DPW

[Kee85] Laurence R. Keefe. Dynamics of perturbed wavetrain solutions to the Ginzburg–Landau equation. Studies in Applied Mathematics, 73(2):91–153, October 1985. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198573291.

Kennett:1974:MNI

[Ken74] Rosemary G. Kennett. On the modulations and nonlinear interactions of waves and the nonlinear stability of a near-critical system. Studies in Applied Mathematics, 53(4):317–336, December 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1974534317.

Kennett:1976:MMC

[Ken76] Rosemary G. Kennett. A model for magnetohydrodynamic convection relevant to the solar dynamo problem. Studies in Applied Mathematics, 55(1):65–81, March 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197655165.

Kevorkian:1980:RWN

[Kev80] J. Kevorkian. Resonance in a weakly nonlinear system with slowly varying parameters. Studies in Applied Mathematics, 62(1):23–67, February 1980. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198062123.

Kevorkian:1982:AIP

[Kev82] J. Kevorkian. Adiabatic invariance and passage through resonance for nearly periodic Hamiltonian systems. Studies in Applied Mathematics, 66(2):95–119, April 1982. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198266295.

Kalimeris:2010:HEI

[KF10] K. Kalimeris and A. S. Fokas. The heat equation in the interior of an equilateral triangle. Studies in Applied Mathematics, 124(3):283–305, April 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2009.00471.x.

Kfoury:1970:SSP

[Kfo70] Denis J. Kfoury. Synchronizing sequences for probabilistic automata. Studies in Applied Mathematics, 49(1):101–103, March 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1970491101.

Kopell:1973:PWS

[KH73] N. Kopell and L. N. Howard. Plane wave solutions to reaction–diffusion equations. Studies in Applied Mathematics, 52(4):291–328, December 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1973524291.

Kopell:1981:TPH

[KH81] N. Kopell and L. N. Howard. Target patterns and horseshoes from a perturbed central-force problem: Some temporally periodic solutions to reaction–diffusion equations. Studies in Applied Mathematics, 64(1):1–56, February 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19816411.

Koutsokostas:2020:MEV

[KHF^{+}20] Georgios N. Koutsokostas, Theodoros P. Horikis, Dimitrios J. Frantzeskakis, Barbara Prinari, and Gino Biondini. Multiscale expansions and vector solitons of a two-dimensional nonlocal nonlinear Schrödinger system. Studies in Applied Mathematics, 145(4):739–764, November 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12334.

Ko:2021:UEE

[KHJK21] Dongnam Ko, Seung-Yeal Ha, Shi Jin, and Doheon Kim. Uniform error estimates for the random batch method to the firstorder consensus models with antisymmetric interaction kernels. Studies in Applied Mathematics, 146(4):983–1022, May 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12372.

Ko:2022:SRC

[KHS22] Dongnam Ko, Seung-Yeal Ha, and Woojoo Shim. On the stochastic robustness of complete clustering predictability for a first-order consensus model. Studies in Applied Mathematics, 148(3):1364–1406, April 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12478.

Kichenassamy:1999:WEN

[Kic99] S. Kichenassamy. WTC expansions and nonintegrable equations. Studies in Applied Mathematics, 102(1):1–26, January 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00102.

Kato:1990:DMW

[KK90] Yoshifumi Kato and Masao Koike. The determinants of matrices whose elements decrease geometrically in the diagonal direction. II. Studies in Applied Mathematics, 82(4):291–304, May 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1990824291.

Knessl:1991:SNA

[KK91] Charles Knessl and Joseph B. Keller. Stirling number asymptotics from recursion equations using the ray method. Studies in Applied Mathematics, 84(1):43–56, January 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199184143.

Knessl:1995:RW

[KK95] Charles Knessl and Joseph B. Keller. Rossby waves. Studies in Applied Mathematics, 94(4):359–376, May 1995. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995944359.

Kapitula:2022:LCL

[KK22a] Todd Kapitula and Panayotis G. Kevrekidis. Language competition on lattices. Studies in Applied Mathematics, 148(1):219–247, January 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12439.

Kim:2022:EBS

[KK22b] Dohyun Kim and Jeongho Kim. On the emergent behavior of the swarming models on the complex sphere. Studies in Applied Mathematics, 148(3):1303–1338, April 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12476.

Kath:1987:VAN

[KKM87] W. L. Kath, C. Knessl, and B. J. Matkowsky. A variational approach to nonlinear singularly perturbed boundary-value problems. Studies in Applied Mathematics, 77(1):61–88, August 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198777161.

Kleitman:1971:CMN

[KKR71] D. J. Kleitman, M. M. Krieger, and B. L. Rothschild. Configurations maximizing the number of pairs of Hamming-adjacent lattice points. Studies in Applied Mathematics, 50(2):115–119, June 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1971502115.

Koley:2015:OWS

[KKS15] S. Koley, R. B. Kaligatla, and T. Sahoo. Oblique wave scattering by a vertical flexible porous plate. Studies in Applied Mathematics, 135(1):1–34, July 2015. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12076.

Kleitman:1975:NUS

[KL75] Daniel J. Kleitman and S.-Y. Li. A note on unigraphic sequences. Studies in Applied Mathematics, 54(4):283–287, December 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1975544283.

Kleitman:1981:PSR

[KL81] D. J. Kleitman and Shuo-Yen Robert Li. On the partial sums of the rearrangements of a complex series. Studies in Applied Mathematics, 64(2):177–182, April 1981. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1981642177.

Kevorkian:1984:RMI

[KL84] J. Kevorkian and H. K. Li. Resonant modal interactions and adiabatic invariance for a nonlinear wave equation in a variable domain. Studies in Applied Mathematics, 71(1):1–64, August 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19847111.

Kevorkian:1988:EAS

[KL88] J. Kevorkian and Y. P. Li. Explicit approximations for strictly nonlinear oscillators with slowly varying parameters with applications to free-electron lasers. Studies in Applied Mathematics, 78(2):111–165, April 1988. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1988782111.

Konopelchenko:2000:IST

[KL00] B. G. Konopelchenko and G. Landolfi. Induced surfaces and their integrable dynamics II. Generalized Weierstrass representations in 4-d spaces and deformations via Ds hierarchy. Studies in Applied Mathematics, 104(2):129–169, February 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00133.

Kleitman:1970:PID

[Kle70] Daniel J. Kleitman. Proportions of irreducible diagrams. Studies in Applied Mathematics, 49(3):297–299, September 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1970493297.

Kleitman:1972:FUT

[Kle72] D. J. Kleitman. Finding uncomplemented trees. Studies in Applied Mathematics, 51(3):309–310, September 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1972513309.

Kleitman:1975:NSS

[Kle75] Daniel J. Kleitman. A note on some subset identities. Studies in Applied Mathematics, 54(4):289–292, December 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1975544289.

Kliakhandler:2003:MDW

[Kli03] I. L. Kliakhandler. Modulations of deep water waves and spectral filtering. Studies in Applied Mathematics, 111 (3):301–313, October 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
t01-1-00235.

Klein:2018:WRE

[KLPS18] Christian Klein, Felipe Linares, Didier Pilod, and Jean-Claude Saut. On Whitham and related equations. Studies in Applied Mathematics, 140(2):133–177, February 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12194.

Korotkevich:2023:SIS

[KLSD23] A. O. Korotkevich, P. M. Lushnikov, A. Semenova, and S. A. Dyachenko. Superharmonic instability of Stokes waves. Studies in Applied Mathematics, 150(1):119–134, January 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12535.

Kropinski:2011:AAL

[KLW11] M. C. Kropinski, A. E. Lindsay, and M. J. Ward. Asymptotic analysis of localized solutions to some linear and nonlinear biharmonic eigenvalue problems. Studies in Applied Mathematics, 126(4):347–408, May 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2010.00507.x.

Kelly:1970:NCL

[KM70] R. E. Kelly and S. A. Maslowe. The nonlinear critical layer in a slightly stratified shear flow. Studies in Applied Mathematics, 49(4):301–326, December 1970. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1970494301.

Konopelchenko:1990:IST

[KM90] B. G. Konopelchenko and B. T. Matkarimov. Inverse spectral transform for the nonlinear evolution equation generating the Davey–Stewartson and Ishimori equations. Studies in Applied Mathematics, 82(4):319–359, May 1990. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990824319.

Kaup:1998:IST

[KM98] D. J. Kaup and Y. Matsuno. The inverse scattering transform for the Benjamin–Ono equation. Studies in Applied Mathematics, 101(1):73–98, July 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00086.

Konopelchenko:2004:IEN

[KM04] B. Konopelchenko and A. Moro. Integrable equations in nonlinear geometrical optics. Studies in Applied Mathematics, 113(4):325–352, November 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
0022-2526.2004.01536.x.

Knessl:2012:AAT

[KM12] Charles Knessl and John A. Morrison. Asymptotic analysis of two coupled queues with vastly different arrival rates and finite customer capacities. Studies in Applied Mathematics, 128(2):107–143, February 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2011.00529.x.

Koornwinder:2018:DAS

[KM18] Tom H. Koornwinder and Marta Mazzocco. Dualities in the q-Askey scheme and degenerate DAHA. Studies in Applied Mathematics, 141(4):424–473, November 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12229.

Kalnins:1980:LTT

[KMM80a] E. G. Kalnins, H. L. Manocha, and Willard Miller, Jr. The Lie theory of two-variable hypergeometric functions. Studies in Applied Mathematics, 62(2):143–173, April 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980622143.

Kalnins:1980:TRF

[KMM80b] E. G. Kalnins, H. L. Manocha, and Willard Miller, Jr. Transformation and reduction formulas for two-variable hypergeometric functions on the sphere S_{2}. Studies in Applied Mathematics, 63(2):155–167, October 1980. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980632155.

Kalnins:1982:HAE

[KMM82] E. G. Kalnins, H. L. Manocha, and Willard Miller, Jr. Harmonic analysis and expansion formulas for two-variable hypergeometric functions. Studies in Applied Mathematics, 66(1):69–89, February 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm198266169.

Khusnutdinova:2014:VWN

[KMP14] K. R. Khusnutdinova, K. R. Moore, and D. E. Pelinovsky. Validity of the weakly nonlinear solution of the Cauchy problem for the Boussinesq-type equation. Studies in Applied Mathematics, 133(1):52–83, July 2014. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12034.

Knessl:2000:GOA

[Kne00] Charles Knessl. Geometric optics approach to first-passage distributions: Caustic boundaries and exponentially small eigenvalues. Studies in Applied Mathematics, 105(4):301–332, November 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9590.00153.

Knessl:2001:NMB

[Kne01] Charles Knessl. A note on a moving boundary problem arising in the American put option. Studies in Applied Mathematics, 107(2):157–183, August 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00183.

Knessl:2003:GOM

[Kne03] Charles Knessl. Geometrical optics and models of computer memory fragmentation. Studies in Applied Mathematics, 111(2):185–238, August 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
t01-1-00232.

Kelly:1984:RRM

[KO84] D. G. Kelly and J. G. Oxley. On random representable matroids. Studies in Applied Mathematics, 71(3):181–205, December 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1984713181.

Knaub:2004:MIL

[KO04] Karl R. Knaub and Robert E. O’Malley, Jr. The motion of internal layers in singularly perturbed advection– diffusion-reaction equations. Studies in Applied Mathematics, 112(1):1–15, January 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2004.01502.x.

Konopelchenko:2013:QCA

[KO13] B. G. Konopelchenko and G. Ortenzi. Quasi-classical approximation in vortex filament dynamics. Integrable systems, gradient catastrophe, and flutter. Studies in Applied Mathematics, 130(2):167–199, February 2013. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00563.x.

Kokocki:2020:TIA

[Kok20] Piotr Kokocki. Total integrals of Ablowitz-Segur solutions for the inhomogeneous Painlevé II equation. Studies in Applied Mathematics, 144(4):504–547, May 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12307.

Konopelchenko:1996:IST

[Kon96] B. G. Konopelchenko. Induced surfaces and their integrable dynamics. Studies in Applied Mathematics, 96(1):9–51, January 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm19969619.

Konotop:2014:PNWa

[Kon14a] Vladimir V. Konotop. Preface: Nonlinear wave phenomena and PT -symmetry. Studies in Applied Mathematics, 133(3):279–280, October 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12068.

Konotop:2014:PNWb

[Kon14b] Vladimir V. Konotop. Preface: Nonlinear wave phenomena and PT -symmetry. Studies in Applied Mathematics, 133(4):351–352, November 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12069.

Khusnutdinova:2019:PNWa

[KP19a] Karima R. Khusnutdinova and Dmitry E. Pelinovsky. Preface: Nonlinear waves in fluids in honor of Roger Grimshaw on the occasion of his 80th birthday. Studies in Applied Mathematics, 142(3):215–218, April 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12262.

Khusnutdinova:2019:PNWb

[KP19b] Karima R. Khusnutdinova and Dmitry E. Pelinovsky. Preface: Nonlinear waves in fluids in honor of Roger Grimshaw on the occasion of his 80th birthday: Part II. Studies in Applied Mathematics, 142(4):417–418, May 2019.

CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12267.

Kedem:1980:SMN

[KPS80] Gershon Kedem, Seymour V. Parter, and Michael Steuerwalt. The solutions of a model nonlinear singular perturbation problem having a continuous locus of singular points. Studies in Applied Mathematics, 63(2):119–146, October 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1980632119.

Kang:2010:HTT

[KQ10] Jing Kang and Changzheng Qu. Hodograph-type transformations for linearization of systems of nonlinear diffusion equations. Studies in Applied Mathematics, 124(3):247–281, April 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00470.x.

Kazaki:1978:FNF

[KR78] J. Y. Kazaki and R. S. Rivlin. Flow of a Newtonian fluid between eccentric rotating cylinders and related problems. Studies in Applied Mathematics, 58(3):209–247, June 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1978583209.

Kliakhandler:2003:SFF

[KR03] I. L. Kliakhandler and B. Y. Rubinstein. Spectral filtering formalism and its application for multi-phase deep-water wavetrains. Studies in Applied Mathematics, 110(3):205–219, April 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00237.

Kohler:2022:BSQ

[KR22] Simon Kohler and Wolfgang Reichel. Breather solutions for a quasi-linear 1 + 1-dimensional wave equation. Studies in Applied Mathematics, 148(2):689–714, February 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12455.

Kumar:2023:RPG

[KR23] Akshay Kumar and R. Radha. Riemann problems for generalized gas dynamics. Studies in Applied Mathematics, 150(4):1154–1181, May 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12565.

Kroll:1977:UFP

[Kro77] John Kroll. Uniform flow past an axially uniform viscous vortex. Studies in Applied Mathematics, 57(3):205–223, December 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1977573205.

Kleitman:1981:MPP

[KS81] D. J. Kleitman and J. B. Shearer. A monotonicity property of partial orders. Studies in Applied Mathematics, 65(1):81–83, August 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm198165181.

Kruskal:1991:ABA

[KS91] Martin D. Kruskal and Harvey Segur. Asymptotics beyond all orders in a model of crystal growth. Studies in Applied Mathematics, 85(2):129–181, August 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991852129.

Karsten:1996:NSB

[KS96] Richard H. Karsten and Gordon E. Swaters. Nonlinear stability of baroclinic fronts in a channel with variable topography. Studies in Applied Mathematics, 96(2):183–199, February 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996962183.

Konopelchenko:2002:RFG

[KS02] B. G. Konopelchenko and W. K. Schief. Reciprocal figures, graphical statics, and inversive geometry of the Schwarzian BKP hierarchy. Studies in Applied Mathematics, 109(2):89– 124, August 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9590.00402.

Knessl:2006:JPL

[KS06] Charles Knessl and Wojciech Szpankowski. On the joint path length distribution in random binary trees. Studies in Applied Mathematics, 117(2):109–147, August 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2006.00349.x.

Kapitula:2014:HKI

[KS14] Todd Kapitula and Atanas Stefanov. A Hamiltonian–Krein (instability) index theory for solitary waves to KdV-like eigenvalue problems. Studies in Applied Mathematics, 132(3):183– 211, April 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12031.

Klein:2018:NSB

[KS18] Christian Klein and Nikola Stoilov. Numerical study of blow-up mechanisms for Davey–Stewartson II systems. Studies in Applied Mathematics, 141(1):89–112, July 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12214.

Klein:2020:NST

[KS20] Christian Klein and Nikola Stoilov. Numerical study of the transverse stability of the peregrine solution. Studies in Applied Mathematics, 145(1):36–51, July 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12306.

Kirousis:2001:UBA

[KSV01] Lefteris M. Kirousis, Yannis C. Stamatiou, and Malvina Vamvakari. Upper bounds and asymptotics for the q-binomial coefficients. Studies in Applied Mathematics, 107(1):43–62, July 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.1071177.

Kushner:2014:NAF

[KSW14] T. Kushner, A. Simonyan, and F. Y. M. Wan. A new approach to feedback for robust signaling gradients. Studies in Applied Mathematics, 133(1):18–51, July 2014. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12041.

Kaup:1989:TOS

[KT89] D. J. Kaup and Gary E. Thomas. The third-order singularperturbation expansion of the planar cold-fluid magnetron equations. Studies in Applied Mathematics, 81(1):57–78, August 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm198981157. See errata [KT90].

Kaup:1990:ETO

[KT90] D. J. Kaup and Gary E. Thomas. Errata for: The thirdorder singular perturbation expansion of the planar coldfluid magnetron equations. Studies in Applied Mathematics, 83(3):271–272, October 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1990833271.

See [KT89].

Knessl:1999:RAP

[KT99] Charles Knessl and Charles Tier. A ray approximation to a PdE that arises in the study of tandem queues. Studies in Applied Mathematics, 102(1):87–120, January 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00105.

Khusnutdinova:2019:DTS

[KT19] Karima R. Khusnutdinova and Matthew R. Tranter. D’Alemberttype solution of the Cauchy problem for the Boussinesq-Klein– Gordon equation. Studies in Applied Mathematics, 142(4):551– 585, May 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12263.

Kalogeropoulos:2022:EID

[KT22] Andreas Kalogeropoulos and Nikolaos L. Tsitsas. Electromagnetic interactions of dipole distributions with a stratified medium: Power fluxes and scattering cross sections. Studies in Applied Mathematics, 148(3):1040–1068, April 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12469.

Kung:1978:AFM

[Kun78] Joseph. P. S. Kung. The alpha function of a matroid — I. Transversal matroids. Studies in Applied Mathematics, 58(3):263–275, June 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1978583263.

Kung:1980:SM

[Kun80] Joseph P. S. Kung. On specializations of matroids. Studies in Applied Mathematics, 62(2):183–187, April 1980. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980622183.

Kung:1986:GTB

[Kun86] Joseph P. S. Kung. Gundelfinger’s theorem on binary forms. Studies in Applied Mathematics, 75(2):163–169, October 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1986752163.

Kupershmidt:2006:EEL

[Kup06] Boris A. Kupershmidt. Extended equations of long waves. Studies in Applied Mathematics, 116(4):415–434, May 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2006.00347.x.

Kyriakoussis:2006:ASN

[KV06] A. Kyriakoussis and M. G. Vamvakari. On asymptotics for the signless noncentral q-Stirling numbers of the first kind. Studies in Applied Mathematics, 117(3):191–213, October 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2006.00352.x.

Kolokolnikov:2005:ESS

[KWW05a] Theodore Kolokolnikov, Michael J. Ward, and Juncheng Wei. The existence and stability of spike equilibria in the one-dimensional Gray–Scott model: The low feed-rate regime. Studies in Applied Mathematics, 115(1):21–71, July 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2005.01554.

Kolokolnikov:2005:PSS

[KWW05b] Theodore Kolokolnikov, Michael J. Ward, and Juncheng Wei. Pulse-splitting for some reaction–diffusion systems in one-space dimension. Studies in Applied Mathematics, 114 (2):115–165, February 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.0022-
2526.2005.01542.x.

Kolokolnikov:2006:ZBI

[KWW06] T. Kolokolnikov, M. J. Ward, and J. Wei. Zigzag and breakup instabilities of stripes and rings in the two-dimensional Gray– Scott model. Studies in Applied Mathematics, 116(1):35–95, January 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1365-2966.2005.0333.x.

Kou:2018:LQD

[KX18] Kit Ian Kou and Yong-Hui Xia. Linear quaternion differential equations: Basic theory and fundamental results. Studies in Applied Mathematics, 141(1):3–45, July 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12211.

Lacey:1982:MST

[Lac82] A. A. Lacey. The motion with slip of a thin viscous droplet over a solid surface. Studies in Applied Mathematics, 67(3):217–230, December 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1982673217.

Lewis:2017:NET

[LAHM17] A. H. Lewis, D. G. A. L. Aarts, P. D. Howell, and A. Majumdar. Nematic equilibria on a two-dimensional annulus. Studies in Applied Mathematics, 138(4):438–466, May 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12161.

Lakin:1972:BVP

[Lak72] W. D. Lakin. Boundary value problems with a turning point. Studies in Applied Mathematics, 51(3):261–275, September 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1972513261.

Lamb:1998:SIW

[Lam98] Kevin G. Lamb. Are solitary internal waves solitons? Studies in Applied Mathematics, 101(3):289–308, October 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00094.

Lange:1973:WPM

[Lan73] Charles G. Lange. Wave propagation m continuous random media. Studies in Applied Mathematics, 52(4):359–376, December 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1973524359.

Landman:1980:BST

[Lan80] Kerry A. Landman. Bifurcation and stability theory of periodic solutions for integrodifferential equations. Studies in Applied Mathematics, 62(3):217–248, June 1980. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980623217.

Landman:1983:CCL

[Lan83a] K. A. Landman. On the crenation of a compound liquid droplet. Studies in Applied Mathematics, 69(1):51–73, August 1983. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198369151.

Lange:1983:SSS

[Lan83b] Charles G. Lange. On spurious solutions of singular perturbation problems. Studies in Applied Mathematics, 68(3):227–257, June 1983. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1983683227.

Landamn:1987:SGL

[Lan87a] Michael J. Landamn. Solutions of the Ginzburg–Landau equation of interest in shear flow transition. Studies in Applied Mathematics, 76(3):187–237, June 1987. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1987763187.

Lange:1987:AAF

[Lan87b] Charles G. Lange. Asymptotic analysis of forced nonlinear Sturm–Liouville systems. Studies in Applied Mathematics, 76(3):239–263, June 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1987763239.

Lapidus:1987:FKF

[Lap87] Michel L. Lapidus. The Feynman–Kac formula with a Lebesgue-Stieltjes measure and Feynman’s operational calculus. Studies in Applied Mathematics, 76(2):93–132, April 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198776293.

Laquer:1981:ALT

[Laq81] H. Turner Laquer. Asymptotic limits for a two-dimensional recursion. Studies in Applied Mathematics, 64(3):271–277, June 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1981643271.

Lieb:1969:CTQ

[LB69] Elliott H. Lieb and W. A. Beyer. Clusters on a thin quadratic lattice (transfer matrix technique). Studies in Applied Mathematics, 48(1):77–90, March 1969. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm196948177.

Liu:1981:ENW

[LB81] A. K. Liu and D. J. Benney. The evolution of nonlinear wave trains in stratified shear flows. Studies in Applied Mathematics, 64(3):247–269, June 1981. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1981643247.

Lu:1995:RES

[LB95] Qingyuan Lu and D. J. Benney. Rotational effects on surface waves. Studies in Applied Mathematics, 94(1):77–82, January 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199594177.

Lunowa:2021:AMI

[LBP21] Stephan B. Lunowa, Carina Bringedal, and Iuliu Sorin Pop. On an averaged model for immiscible two-phase flow with surface tension and dynamic contact angle in a thin strip. Studies in Applied Mathematics, 147(1):84–126, July 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12376.

Lyu:2021:GUE

[LC21] Shulin Lyu and Yang Chen. Gaussian unitary ensembles with two jump discontinuities, PDEs, and the coupled Painlevé II and IV systems. Studies in Applied Mathematics, 146(1):118–138, January 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12343.

Ludlow:1999:SRE

[LCB99] David K. Ludlow, Peter A. Clarkson, and Andrew P. Bassom. Similarity reductions and exact solutions for the twodimensional incompressible Navier–Stokes equations. Studies in Applied Mathematics, 103(3):183–240, October 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00125.

Liu:2021:LTA

[LCG21] Nan Liu, Mingjuan Chen, and Boling Guo. Long-time asymptotic behavior of the fifth-order modified KdV equation in low regularity spaces. Studies in Applied Mathematics, 147(1):230– 299, July 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12379.

Lu:2022:DAA

[LCTZ22] Jingjing Lu, Qiaoling Chen, Zhidong Teng, and Tingting Zheng. Dynamical analysis of an age-structured SIRE epidemic model with two routes of infection in environment. Studies in Applied Mathematics, 148(2):461–497, February 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12447.

Long:2020:AOP

[LDLW20] Wen-Gao Long, Dan Dai, Yu-Tian Li, and Xiang-Sheng Wang. Asymptotics of orthogonal polynomials with asymptotic Freud-like weights. Studies in Applied Mathematics, 144(2):133–163, February 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12291.

LeDizes:2004:VCL

[Le 04] Stéphane Le Dizès. Viscous critical-layer analysis of vortex normal modes. Studies in Applied Mathematics, 112(4):315–332, May 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.0022-2526.2004.01514.x.

Leach:2016:LTS

[Lea16] J. A. Leach. The large-time solution of Burgers’ equation with time-dependent coefficients. I. The coefficients are exponential functions. Studies in Applied Mathematics, 136(2):163–188, February 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12098.

Lebensold:1973:ECP

[Leb73] Kenneth Lebensold. Efficient communication by phone calls. Studies in Applied Mathematics, 52(4):345–358, December 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1973524345.

Lebensold:1977:DP

[Leb77] Kenneth Lebensold. A divisibility problem. Studies in Applied Mathematics, 56(3):291–294, June 1977. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1977563291.

Lenells:2009:ESM

[Len09] Jonatan Lenells. Exactly solvable model for nonlinear pulse propagation in optical fibers. Studies in Applied Mathematics, 123(2):215–232, August 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2009.00454.x.

Lenells:2016:NFT

[Len16] Jonatan Lenells. Nonlinear Fourier transforms and the mKdV equation in the quarter plane. Studies in Applied Mathematics, 136(1):3–63, January 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12089.

Levinson:1969:ABS

[Lev69] N. Levinson. Asymptotic behavior of solutions of non-linear differential equations. Studies in Applied Mathematics, 48 (4):285–297, December 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1969484285.

Lopez:2002:AEG

[LF02] José L. López and Chelo Ferreira. Asymptotic expansions of generalized Stieltjes transforms of algebraically decaying functions. Studies in Applied Mathematics, 108(2):187–215, February 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.01404.

Li:2020:GSS

[LFW20] Ming Li, Heming Fu, and Chengfa Wu. General soliton and (semi-)rational solutions to the nonlocal Mel’nikov equation on the periodic background. Studies in Applied Mathematics, 145(1):97–136, July 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12313.

Landman:1982:CCD

[LG82] K. A. Landman and H. P. Greenspan. On the crenation of coated droplets. Studies in Applied Mathematics, 66(3):189– 216, June 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1982663189.

Li:2020:VLW

[LG20] Ruomeng Li and Xianguo Geng. On a vector long waveshort wave-type model. Studies in Applied Mathematics, 144(2):164–184, February 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12293.

Lester:2021:LCK

[LGZZ21] Charles Lester, Andrey Gelash, Dmitry Zakharov, and Vladimir Zakharov. Lump chains in the KP-I equation. Studies in Applied Mathematics, 147(4):1425–1442, November 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12420.

Li:2004:NSE

[LHC04] Yi A. Li, James M. Hyman, and Wooyoung Choi. A numerical study of the exact evolution equations for surface waves in water of finite depth. Studies in Applied Mathematics, 113(3):303–324, October 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
0022-2526.2004.01534.x.

Li:1976:ANS

[Li76] Shuo-Yen R. Li. Annihilators in nonrepetitive semigroups. Studies in Applied Mathematics, 55(1):83–85, March 1976.

CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197655183.

Li:2002:SPV

[Li02] Yanguang (Charles) Li. Singularly perturbed vector and scalar nonlinear Schrödinger equations with persistent homoclinic orbits. Studies in Applied Mathematics, 109(1):19–38, July 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00002.

Li:2005:MAS

[Li05] Y. Charles Li. Melnikov analysis for a singularly perturbed DSII equation. Studies in Applied Mathematics, 114(3):285–306, April 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.0022-
2526.2005.01531.x.

Li:2006:CSA

[Li06] Y. Charles Li. Chaos and shadowing around a heteroclinically tubular cycle with an application to sine-Gordon equation. Studies in Applied Mathematics, 116(2):145–171, February 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2006.00336.x.

Liao:2006:SSU

[Lia06] Shijun Liao. Series solutions of unsteady boundary-layer flows over a stretching flat plate. Studies in Applied Mathematics, 117(3):239–263, October 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2006.00354.x.

Lieb:1977:EUM

[Lie77] Elliott H. Lieb. Existence and uniqueness of the minimizing solution of Choquard’s nonlinear equation. Studies in Applied Mathematics, 57(2):93–105, October 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197757293.

Lingwood:1997:ABS

[Lin97] R. J. Lingwood. On the application of the Briggs’ and steepest-descent methods to a boundary-layer flow. Studies in Applied Mathematics, 98(3):213–254, April 1997. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00048.

Lin:2002:TTS

[Lin02] C. C. Lin. On transition to turbulence and synthesis of alternative approaches. Studies in Applied Mathematics, 108(1):7–28, January 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
01420.

Lin:2014:NSS

[Lin14] Bin Lin. A nonpolynomial spline scheme for the generalized regularized long wave equation. Studies in Applied Mathematics, 132(2):160–182, February 2014. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12022.

Lippert:2006:DAC

[Lip06] Ross A. Lippert. Discrete approximations to continuum optimal flow problems. Studies in Applied Mathematics, 117 (4):321–333, November 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2006.00357.x.

Liu:2013:CGC

[Liu13] Hanze Liu. Complete group classifications and symmetry reductions of the fractional fifth-order KdV types of equations. Studies in Applied Mathematics, 131(4):317–330, November 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12011.

Lynch:2010:DDP

[LK10] S. Lynch and C. Knessl. Drift–diffusion past an ellipse: Singular perturbation analysis of the illuminated region. Studies in Applied Mathematics, 125(2):113–151, August 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00477.x.

Li:2016:PBM

[LK16] J. Li and W. L. Kath. A path-based method for simulating large deviations and rare events in nonlinear lightwave systems. Studies in Applied Mathematics, 137(2):159–173, August 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12122.

Liu:1980:RTE

[LKK80] A. K. Liu, T. Kubota, and D. R. S. Ko. Resonant transfer of energy between nonlinear waves in neighboring pycnoclines. Studies in Applied Mathematics, 63(1):25–45, August 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198063125.

Li:2014:NMS

[LKM14] K. Li, P. G. Kevrekidis, and B. A. Malomed. Nonlinear modes and symmetries in linearly coupled pairs of PT -invariant dimers. Studies in Applied Mathematics, 133(3):281–297, October 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12038.

Lin:1979:DWT

[LL79] C. C. Lin and Y. Y. Lau. Density wave theory of spiral structure of galaxies. Studies in Applied Mathematics, 60(2):97–163, April 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm197960297.

Lange:1980:TSS

[LL80] C. G. Lange and D. A. Larson. Transient solutions to some weakly diffusive nonlinear diffusion equations. Studies in Applied Mathematics, 63(3):249–262, December 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980633249.

Lima:2022:MOP

[LL22] Hélder Lima and Ana Loureiro. Multiple orthogonal polynomials with respect to Gauss’ hypergeometric function. Studies in Applied Mathematics, 148(1):154–185, January 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12437.

Lin:2010:RMS

[LLC10] B. Lin, K. Li, and Z. Cheng. Recurrence method for solving singularly perturbed boundary value problems. Studies in Applied Mathematics, 125(4):331–342, November 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00489.x.

Liu:2012:CGC

[LLL12] H. Liu, J. Li, and L. Liu. Complete group classification and exact solutions to the generalized short pulse equation. Studies in Applied Mathematics, 129(1):103–116, July 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00548.x.

Long:2017:RSF

[LLLZ17] W.-G. Long, Y.-T. Li, S.-Y. Liu, and Y.-Q. Zhao. Real solutions of the first Painlevé equation with large initial data. Studies in Applied Mathematics, 139(4):505–532, November 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12171.

Latushkin:2004:SLE

[LLS04] Y. Latushkin, Y. C. Li, and M. Stanislavova. The spectrum of a linearized 2D Euler operator. Studies in Applied Mathematics, 112(3):259–270, April 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
0022-2526.2004.01510.x.

Li:2023:ERC

[LLZ23] Hongjie Li, Hongyu Liu, and Jun Zou. Elastodynamical resonances and cloaking of negative material structures beyond quasistatic approximation. Studies in Applied Mathematics, 150(3):716–754, April 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12555.

Levinson:1974:CIE

[LM74] Norman Levinson and Clement McCalla. Completeness and independence of the exponential solutions of some functional differential equations. Studies in Applied Mathematics, 53(1): 1–15, March 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm19745311.

Lange:1991:SPA

[LM91] Charles G. Lange and Robert M. Miura. Singular perturbation analysis of boundary value problems for differentialdifference equations. IV. a nonlinear example with layer behavior. Studies in Applied Mathematics, 84(3):231–273, April 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991843231.

Lannes:2016:NWC

[LM16] David Lannes and Fabien Marche. Nonlinear wave–current interactions in shallow water. Studies in Applied Mathematics, 136(4):382–423, May 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12110.

Louis:2019:DSW

[LMS19] Simon A. Louis, Tim R. Marchant, and Noel F. Smyth. 2-d solitary waves in thermal media with nonsymmetric boundary conditions. Studies in Applied Mathematics, 142(4):586–607, May 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12243.

Liu:2008:DPP

[LMY08] Shu-Chung Liu, Jun Ma, and Yeong-Nan Yeh. Dyck paths with peak- and valley-avoiding sets. Studies in Applied Mathematics, 121(3):263–289, October 2008. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00415.x.

Lebovitz:1994:SEP

[LN94] Norman R. Lebovitz and Anatoly Neishtadt. Slow evolution in perturbed Hamiltonian systems. Studies in Applied Mathematics, 92(2):127–144, June 1994. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1994922127.

Luz:2013:WPD

[LN13] Ana Maria Luz and André Nachbin. Wave packet defocusing due to a highly disordered bathymetry. Studies in Applied Mathematics, 130(4):393–416, May 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00571.x.

Lander:2005:ADS

[LNVW05] A. D. Lander, Q. Nie, B. Vargas, and F. Y. M. Wan. Aggregation of a distributed source in morphogen gradient formation. Studies in Applied Mathematics, 114(4):343–374, May 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2005.01556.x.

Lou:2004:NEP

[LNW04] Y. Lou, Q. Nie, and F. Y. M. Wan. Nonlinear eigenvalue problems in the stability analysis of morphogen gradients. Studies in Applied Mathematics, 113(2):183–215, August 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2004.01530.x.

Lander:2009:LEE

[LNWZ09] A. D. Lander, Q. Nie, F. Y. M. Wan, and Y.-T. Zhang. Localized ectopic expression of Dpp receptors in a Drosophila embryo. Studies in Applied Mathematics, 123(2):175–214, August 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00450.x.

Laforgue:1995:VSM

[LO95] Jacques G. L. Laforgue and Robert E. O’Malley, Jr. Viscous shock motion for advection–diffusion equations. Studies in Applied Mathematics, 95(2):147–170, August 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995952147.

Laforgue:1999:EAV

[LO99] Jacques G. L. Laforgue and Robert E. O’Malley, Jr. Exponential asymptotics, the viscid Burgers’ equation, and standing wave solutions for a reaction–advection–diffusion model. Studies in Applied Mathematics, 102(2):137–172, February 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00107.

Loeb:1990:SSF

[Loe90] Daniel E. Loeb. Sequences of symmetric functions of binomial type. Studies in Applied Mathematics, 83(1):1–30, July 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19908311.

Lopez:2022:NCD

[Lóp22] José Luis López. On nonstandard chemotactic dynamics with logistic growth induced by a modified complex Ginzburg–Landau equation. Studies in Applied Mathematics, 148(1):248–269, January 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12440.

Lou:2015:CRE

[Lou15] S. Y. Lou. Consistent Riccati expansion for integrable systems. Studies in Applied Mathematics, 134(3):372–402, April 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12072.

Lou:2019:PCN

[Lou19] S. Y. Lou. Prohibitions caused by nonlocality for nonlocal Boussinesq-KdV type systems. Studies in Applied Mathematics, 143(2):123–138, August 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12265.

Laine-Pearson:2008:STT

[LP08] F. E. Laine-Pearson. The stability of two, three, and four wave interactions of a prototype system. Studies in Applied Mathematics, 120(1):17–34, January 2008. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2007.00391.x.

Lopez:2011:SSP

[LP11] José L. López and Pedro J. Pagola. A systematic “Saddle point near a pole” asymptotic method with application to the Gauss hypergeometric function. Studies in Applied Mathematics, 127(1):24–37, July 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2010.00510.x.

Li:2023:SEM

[LP23] Huicong Li and Rui Peng. An SIS epidemic model with mass action infection mechanism in a patchy environment. Studies in Applied Mathematics, 150(3):650–704, April 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12553.

Laine-Pearson:2003:CDS

[LPH03] F. E. Laine-Pearson and P. E. Hydon. Classification of discrete symmetries of ordinary differential equations. Studies in Applied Mathematics, 111(3):269–299, October 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.t01-1-00234.

Laine-Pearson:2009:IPM

[LPH09] F. E. Laine-Pearson and P. E. Hydon. Inertial particle motion in recirculating Stokes flow. Studies in Applied Mathematics, 122(2):139–152, February 2009. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00427.x.

Li:2022:CMA

[LQYjZ22] Shangshuai Li, Changzheng Qu, Xiangxuan Yi, and Da jun Zhang. Cauchy matrix approach to the SU(2) self-dual Yang– Mills equation. Studies in Applied Mathematics, 148(4):1703– 1721, May 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12488.

Lin:1969:ATC

[LR69] C. C. Lin and A. L. Rabenstein. On the asymptotic theory of a class of ordinary differential equations of fourth order II. Existence of solutions which are approximated by the formal solutions. Studies in Applied Mathematics, 48(4):311–340, December 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1969484311.

Loper:2001:MCC

[LR01] David E. Loper and Paul H. Roberts. Mush–Chimney convection. Studies in Applied Mathematics, 106(2):187–227, February 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00165.

Lou:2004:NDL

[LRS04] S. Y. Lou, C. Rogers, and W. K. Schief. On a nonlinear discretized LKR system: Reductions and underlying Virasoro symmetry algebras. Studies in Applied Mathematics, 113(4):353–380, November 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
0022-2526.2004.01537.x.

Lebovitz:1975:ESA

[LS75] N. R. Lebovitz and R. J. Schaar. Exchange of stabilities in autonomous systems. Studies in Applied Mathematics, 54(3):229–260, September 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1975543229.

Lebovitz:1977:ESA

[LS77] N. R. Lebovitz and R. J. Schaar. Exchange of stabilities in autonomous systems — II. Vertical bifurcation.

Studies in Applied Mathematics, 56(1):1–50, January 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19775611.

Lange:1988:SPA

[LS88] Charles G. Lange and Donald R. Smith. Singular perturbation analysis of integral equations. Studies in Applied Mathematics, 79(1):1–63, August 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19887911.

Lange:1993:SPA

[LS93] Charles G. Lange and Donald R. Smith. Singular perturbation analysis of integral equations: Part II. Studies in Applied Mathematics, 90(1):1–74, October 1993. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19939011.

Lopez:2004:AET

[LS04] José L. López and Ester Pérez Sinusía. Asymptotic expansions for two singularly perturbed convection—diffusion problems with discontinuous data: The quarter plane and the infinite strip. Studies in Applied Mathematics, 113(1):57–89, July 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2004.01508.x.

Ling:2023:MEL

[LS23] Liming Ling and Xuan Sun. The multi elliptic-localized solutions and their asymptotic behaviors for the mKdV equation. Studies in Applied Mathematics, 150(1):135–183, January 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12536.

Li:2022:BAE

[LSG22] Jing Li, Gui-Quan Sun, and Zun-Guang Guo. Bifurcation analysis of an extended Klausmeier–Gray–Scott model with infiltration delay. Studies in Applied Mathematics, 148(4):1519– 1542, May 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12482.

Lopez:2006:FOA

[LST06] José L. López, Ester Pérez Sinusía, and Nico M. Temme. First order approximation of an elliptic 3D singular perturbation problem. Studies in Applied Mathematics, 116(3):303–319, April 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2006.00345.x.

Lopez:1999:UAB

[LT99] José L. López and Nico M. Temme. Uniform approximations of Bernoulli and Euler polynomials in terms of hyperbolic functions. Studies in Applied Mathematics, 103(3):241–258, October 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00126.

Lopez:2002:TPT

[LT02] José L. López and Nico M. Temme. Two-point Taylor expansions of analytic functions. Studies in Applied Mathematics, 109(4):297–311, November 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00225.

Liao:2007:GAO

[LT07] S. Liao and Y. Tan. A general approach to obtain series solutions of nonlinear differential equations. Studies in Applied Mathematics, 119(4):297–354, November 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2007.00387.x.

Luo:2023:ELB

[LT23] Yu Luo and Satoshi Tsujimoto. Exceptional Laurent biorthogonal polynomials through spectral transformations of generalized eigenvalue problems. Studies in Applied Mathematics, 151(2):585–615, August 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12603.

Li:2022:RHP

[LTY22] Yuan Li, Shou-Fu Tian, and Jin-Jie Yang. Riemann– Hilbert problem and interactions of solitons in the ncomponent nonlinear Schrödinger equations. Studies in Applied Mathematics, 148(2):577–605, February 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12450.

Lu:1989:SHD

[Lu89] Ya Yan Lu. The shunted homopolar dynamo — an analytic approach to a Poincaré map. Studies in Applied Mathematics, 80(3):239–251, June 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1989803239.

Lu:1993:BMF

[Lu93] Ya Yan Lu. Bifurcation to mean flows in convection. Studies in Applied Mathematics, 88(1):1–23, January 1993. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19938811.

Lu:2023:WPG

[Lu23] Yufeng Lu. Well-posedness of generalized KdV and onedimensional fourth-order derivative nonlinear Schrödinger equations for data with an infinite L^{2} norm. Studies in Applied Mathematics, 150(3):883–898, April 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12559.

Lundmark:2003:HDI

[Lun03] Hans Lundmark. Higher-dimensional integrable Newton systems with quadratic integrals of motion. Studies in Applied Mathematics, 110(3):257–296, April 2003. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00239.

Luo:2021:EAS

[Luo21] Jingxi Luo. Exact analytical solution of a novel modified nonlinear Schrödinger equation: Solitary quantum waves on a lattice. Studies in Applied Mathematics, 146(2):550–562, February 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12355.

Lustman:1980:NNW

[Lus80] Liviu Lustman. A note on nonbreaking waves in hyperelastic materials. Studies in Applied Mathematics, 63(2):147–154, October 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1980632147.

Lin:1988:FIR

[LW88] Yihan Lin and Frederic Y. M. Wan. First integrals and the residual solution for orthotropic plates in plane strain or axisymmetric deformations. Studies in Applied Mathematics, 79(2):93–125, October 1988. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198879293.

Lin:1990:SIO

[LW90] Yihan Lin and Frederic Y. M. Wan. Semi infinite orthotropic cantilevered strips and the foundations of plate theories. Studies in Applied Mathematics, 82(3):217–244, April 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990823217.

Lange:1991:SPL

[LW91] Charles G. Lange and Hubertus J. Weinitschke. Singular perturbations of limit points with application to tubular chemical reactors. Studies in Applied Mathematics, 84(1):7–42, January 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19918417.

Lee:1995:ANA

[LW95] June-Yub Lee and Michael J. Ward. On the asymptotic and numerical analyses of exponentially ill-conditioned singularly perturbed boundary value problems. Studies in Applied Mathematics, 94(3):271–326, April 1995. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995943271.

Liu:2022:EWP

[LW22] Yaqing Liu and Deng-Shan Wang. Exotic wave patterns in Riemann problem of the high-order Jaulent–Miodek equation: Whitham modulation theory. Studies in Applied Mathematics, 149(3):588–630, October 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12513.

Li:2021:CQM

[LWAO21] Zhien Li, Chao Wang, Ravi P. Agarwal, and Donal O’Regan. Commutativity of quaternion-matrix–valued functions and quaternion matrix dynamic equations on time scales. Studies in Applied Mathematics, 146(1):139–210, January 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12344.

Luo:2023:NSS

[LWC23] Demou Luo, Qiru Wang, and Li Chen. Nonconstant steady states and pattern formations of generalized 1D cross-diffusion systems with prey-taxis. Studies in Applied Mathematics, 150(3):899–958, April 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12560.

Liu:2023:EDS

[LWZ23] Guodong Liu, Hao Wang, and Xiaoyan Zhang. Epidemic dynamics and spatial segregation driven by cognitive diffusion and nonlinear incidence. Studies in Applied Mathematics, 151(2):643–675, August 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12606.

Labelle:1989:CLC

[LY89] Jacques Labelle and Yeong Nan Yeh. The combinatorics of Laguerre, Charlier, and Hermite polynomials. Studies in Applied Mathematics, 80(1):25–36, February 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198980125.

Li:2003:LPD

[LY03] Yanguang (Charles) Li and Artyom V. Yurov. Lax pairs and Darboux transformations for Euler equations. Studies in Applied Mathematics, 111(1):101–113, July 2003. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.t01-1-00229.

Liu:2011:GEU

[LY11] W. Liu and J. Yu. Global existence and uniform decay of solutions for a coupled system of nonlinear viscoelastic wave equations with not necessarily differentiable relaxation functions. Studies in Applied Mathematics, 127(4):315–344, November 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00524.x.

Loureiro:2013:KLT

[LY13] Ana F. Loureiro and S. Yakubovich. The Kontorovich– Lebedev transform as a map between d-orthogonal polynomials. Studies in Applied Mathematics, 131(3):229–265, October 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12009.

Lin:2020:BIS

[LYZ20] Zhiwu Lin, Jincheng Yang, and Hao Zhu. Barotropic instability of shear flows. Studies in Applied Mathematics, 144(3):289– 326, April 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12297.

Liu:2023:GNS

[LZW23] Jianli Liu, Xinyu Zhao, and Kunlai Wu. Global nonlinear stability of longitudinal wave for the planar motion of elastic string with linear Hooke’s law. Studies in Applied Mathematics, 150(2):321–338, February 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12541.

Li:2023:DED

[LZZ23] Kun Li, Maozhu Zhang, and Zhaowen Zheng. Dependence of eigenvalues of Dirac system on the parameters. Studies in Applied Mathematics, 150(4):1201–1216, May 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12567.

Ma:1978:CSL

[Ma78] Yan-Chow Ma. The complete solution of the long-wave–short-wave resonance equations. Studies in Applied Mathematics, 59(3):201–221, December 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1978593201.

Ma:1979:MSS

[Ma79a] Yan-Chow Ma. On the multi-soliton solutions of some nonlinear evolution equations. Studies in Applied Mathematics, 60(1):73–82, February 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm197960173.

Ma:1979:PPW

[Ma79b] Yan-Chow Ma. The perturbed plane-wave solutions of the cubic Schrödinger equation. Studies in Applied Mathematics, 60(1):43–58, February 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm197960143.

Ma:1981:PCS

[MA81] Yan-Chow Ma and Mark J. Ablowitz. The periodic cubic Schrõdinger equation. Studies in Applied Mathematics, 65(2):113–158, October 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1981652113.

Ma:1982:BRS

[Ma82] Yan-Chow Ma. On the bifurcation of a random spectrum of weakly nonlinear deep-water gravity waves. Studies in Applied Mathematics, 66(2):145–158, April 1982. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1982662145.

Ma:1984:NKH

[Ma84] Yan-Chow Ma. On nonlinear Kelvin–Helmholtz waves near direct resonance. Studies in Applied Mathematics, 70(3):201– 213, June 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1984703201.

MacDonald:1994:EMS

[Mac94] D. A. MacDonald. Euler–Maclaurin summation of trigonometric Fourier series. Studies in Applied Mathematics, 92(3):213– 248, July 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1994923213.

MacGillivray:1997:MIT

[Mac97] A. D. MacGillivray. A method for incorporating transcendentally small terms into the method of matched asymptotic expansions. Studies in Applied Mathematics, 99(3):285–310, October 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00062.

Mahoney:1995:WIE

[Mah95] Daniel J. Mahoney. On wave interactions, i: Explosive resonant triads. Studies in Applied Mathematics, 95(4):381–417, November 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1995954381.

Mahoney:1996:WIIa

[Mah96a] D. J. Mahoney. On wave interactions II. Explosive resonant triads in fluid mechanics. Studies in Applied Mathematics, 96(4):407–454, May 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1996964407.

Mahoney:1996:WIIb

[Mah96b] D. J. Mahoney. On wave interactions III. Miscellaneous results. Studies in Applied Mathematics, 96(4):455–465, May 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996964455.

Mallier:1994:PMW

[Mal94] Roland Mallier. On the parametric model of western boundary outflow. Studies in Applied Mathematics, 91(1):17–25, January 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199491117.

Malkus:2001:BD

[Mal01a] Willem V. R. Malkus. Borders of disorder. Studies in Applied Mathematics, 107(4):325–336, November 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.1074189.

Mallier:2001:DSM

[Mal01b] Roland Mallier. On disturbances to a stratified mixing layer. Studies in Applied Mathematics, 107(2):185–206, August 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00184.

Malkus:2002:DOA

[Mal02] Willem V. R. Malkus. From disorder to order August 2000. Studies in Applied Mathematics, 108(1):29–37, January 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.01421.

Mark:1971:NSG

[Mar71] James W-K. Mark. On necessary and sufficient global stability criteria for axisymmetric perturbations of the stellar dynamic disk. Studies in Applied Mathematics, 50(1):1–12, March 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm19715011.

Marchant:2002:HOI

[Mar02] T. R. Marchant. High-order interaction of solitary waves on shallow water. Studies in Applied Mathematics, 109(1):1–17, July 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00001.

Martin:2018:AAR

[Mar18] P. A. Martin. Asymptotic approximations for radial spheroidal wavefunctions with complex size parameter. Studies in Applied Mathematics, 140(3):255–269, April 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12199.

Maslowe:1972:GCA

[Mas72] S. A. Maslowe. The generation of clear air turbulence by nonlinear waves. Studies in Applied Mathematics, 51(1):1–16, March 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm19725111.

Maslowe:1985:DRD

[Mas85] S. A. Maslowe. Direct resonance in double-diffusive systems. Studies in Applied Mathematics, 73(1):59–74, August 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198573159.

Mathai:1989:SDE

[Mat89] A. M. Mathai. On a system of differential equations connected with the gravitational instability of a multicomponent medium in Newtonian cosmology. Studies in Applied Mathematics, 80(1):75–93, February 1989. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198980175.

Moroz:1984:NAE

[MB84] I. M. Moroz and J. Brindley. Nonlinear amplitude evolution of baroclinic wave trains and wave packets. Studies in Applied Mathematics, 70(1):21–61, February 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198470121.

Milewski:1995:RIBa

[MB95] P. A. Milewski and D. J. Benney. Resonant interactions between vortical flows and water waves. Part I: Deep water. Studies in Applied Mathematics, 94(2):131–167, February 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995942131.

Miller:2003:RCS

[MB03] Joel C. Miller and Andrew J. Bernoff. Rates of convergence to self-similar solutions of Burgers’ equation. Studies in Applied Mathematics, 111(1):29–40, July 2003. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.t01-2-00226.

Mizrahi:2009:TCS

[MB09] Eti Mizrahi and Ayşe Hümeyra Bilge. Toward the classification of scalar nonpolynomial evolution equations: Polynomiality in top three derivatives. Studies in Applied Mathematics, 123(3):233–255, October 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2009.00451.x.

Maslowe:1994:WPC

[MBM94] S. A. Maslowe, D. J. Benney, and D. J. Mahoney. Wave packet critical layers in shear flows. Studies in Applied Mathematics, 91(1):1–16, January 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19949111.

MacGillivray:2000:PAP

[MBT00] A. D. MacGillivray, R. J. Braun, and G. Tanoğlu. Perturbation analysis of a problem of Carrier’s. Studies in Applied Mathematics, 104(4):293–311, May 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00136.

Maslowe:2002:SRW

[MC02] S. A. Maslowe and S. R. Clarke. Subcritical Rossby waves in zonal shear flows with nonlinear critical layers. Studies in Applied Mathematics, 108(1):89–103, January 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.01425.

Margetis:2006:GIM

[MC06] Dionisios Margetis and Jaehyuk Choi. Generalized iteration method for first-kind integral equations. Studies in Applied Mathematics, 117(1):1–25, July 2006. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2006.00340.x.

Min:2018:GPD

[MC18] Chao Min and Yang Chen. Gap probability distribution of the Jacobi unitary ensemble: an elementary treatment, from finite n to double scaling. Studies in Applied Mathematics, 140(2):202–220, February 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12198.

Min:2021:DDA

[MC21a] Chao Min and Yang Chen. Differential, difference, and asymptotic relations for Pollaczek–Jacobi type orthogonal polynomials and their Hankel determinants. Studies in Applied Mathematics, 147(1):390–416, July 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12392.

Murashige:2021:LST

[MC21b] Sunao Murashige and Wooyoung Choi. Linear stability of transversely modulated finite-amplitude capillary waves on deep water. Studies in Applied Mathematics, 146(2):498–525, February 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12353.

Milne:1997:BTS

[MCB97] Alice E. Milne, Peter A. Clarkson, and Andrew P. Bassom. Bäcklund transformations and solution hierarchies for the third Painlevé equation. Studies in Applied Mathematics, 98(2):139–194, February 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00044.

McGinnis:2023:MWP

[McG23] Joshua A. McGinnis. Macroscopic wave propagation for 2D lattice with random masses. Studies in Applied Mathematics, 151(2):752–790, August 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12612.

McLeod:1997:SKC

[McL97] J. B. McLeod. The Stokes and Krasovskii conjectures for the wave of greatest height. Studies in Applied Mathematics, 98(4):311–333, May 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00051.

Mallier:1999:NES

[MD99] R. Mallier and A. M. J. Davis. Nonlinear evolution of singular disturbances to the Bickley jet. Studies in Applied Mathematics, 103(2):149–181, August 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00124.

Meyapin:2010:VER

[MDG10] Y. Meyapin, D. Dutykh, and M. Gisclon. Velocity and energy relaxation in two-phase flows. Studies in Applied Mathematics, 125(2):179–212, August 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2010.00484.x.

McCue:2022:TWB

[MEHS22] Scott W. McCue, Maud El-Hachem, and Matthew J. Simpson. Traveling waves, blow-up, and extinction in the Fisher–Stefan model. Studies in Applied Mathematics, 148(2):964–986, February 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12465.

Mendez:1996:PEP

[Mén96] Miguel Méndez. Plethystic exponential polynomials and plethystic Stirling numbers. Studies in Applied Mathematics, 96(1):1–8, January 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19969611.

Messiter:1996:LAL

[Mes96] A. F. Messiter. Large-amplitude long-wave instability of a supersonic shear layer. Studies in Applied Mathematics, 97(2):167–207, August 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1996972167.

Manas:2022:PED

[MFIGH22] Manuel Mañas, Itsaso Fernández-Irisarri, and Omar F. González-Hernández. Pearson equations for discrete orthogonal polynomials: I. Generalized hypergeometric functions and Toda equations. Studies in Applied Mathematics, 148(3):1141– 1179, April 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12471.

Mitsudera:1991:ERD

[MG91] H. Mitsudera and R. Grimshaw. Effects of radiative damping on resonantly generated internal gravity waves. Studies in Applied Mathematics, 84(3):183–206, April 1991. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991843183.

Ma:2020:IST

[MHW20] Wen-Xiu Ma, Yehui Huang, and Fudong Wang. Inverse scattering transforms and soliton solutions of nonlocal reversespace nonlinear Schrödinger hierarchies. Studies in Applied Mathematics, 145(3):563–585, October 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12329.

Miles:1975:AAP

[Mil75] John W. Miles. Asymptotic approximations for prolate spheroidal wave functions. Studies in Applied Mathematics, 54(4):315–349, December 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1975544315.

Miles:1976:IRT

[Mil76] John W. Miles. On internal resonance of two damped oscillators. Studies in Applied Mathematics, 55(4):351–359, December 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1976554351.

Miles:1979:ASK

[Mil79] John W. Miles. The asymptotic solution of the Korteweg– de Vries equation in the absence of solitons. Studies in Applied Mathematics, 60(1):59–72, February 1979. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197960159.

Milewski:1993:OTP

[Mil93] Paul Milewski. Oscillating tails in the perturbed Korteweg–de Vries equation. Studies in Applied Mathematics, 90(1):87–90, October 1993. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm199390187.

Milewski:1995:RIBb

[Mil95] P. A. Milewski. Resonant interactions between vortical flows and water waves. Part II: Shallow water. Studies in Applied Mathematics, 94(3):225–256, April 1995. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995943225.

Milewski:1998:FWW

[Mil98] Paul A. Milewski. A formulation for water waves over topography. Studies in Applied Mathematics, 100(1):95–106, January 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00071.

Milson:2001:SRS

[Mil01] R. Milson. Spectral residues of second-order differential equations: a new method for summation identities and inversion formulas. Studies in Applied Mathematics, 107 (4):337–366, November 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
1074190.

Minzoni:1986:CFP

[Min86] A. A. Minzoni. On the completeness of the functions e^{−np(x)} cos nx, e^{−np(x)} sin nx for n ≥ 0 and p(x) a 2π periodic function. Studies in Applied Mathematics, 75(3):265–269, December 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1986753265.

Miura:1974:CLF

[Miu74] Robert M. Miura. Conservation laws for the fully nonlinear long wave equations. Studies in Applied Mathematics, 53(1):45–56, March 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm197453145.

Milewski:1996:TDW

[MK96] Paul A. Milewski and Joseph B. Keller. Three-dimensional water waves. Studies in Applied Mathematics, 97(2):149–166, August 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1996972149.

Morrison:2009:AAL

[MK09] John A. Morrison and Charles Knessl. Asymptotic analysis of a loss model with trunk reservation II: Trunks reserved for slow traffic. Studies in Applied Mathematics, 122(2):153–193, February 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2008.00428.x.

Molina:2014:ESC

[MK14] Mario I. Molina and Yuri S. Kivshar. Embedded states in the continuum for PT -symmetric systems. Studies in Applied Mathematics, 133(3):337–350, October 2014. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12058.

Mitra:2020:FTP

[MKP^{+}20] K. Mitra, T. Köppl, I. S. Pop, C. J. van Duijn, and R. Helmig. Fronts in two-phase porous media flow problems: The effects of hysteresis and dynamic capillarity. Studies in Applied Mathematics, 144(4):449–492, May 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12304.

Mao:2020:SPE

[ML20] Hui Mao and Q. P. Liu. The short pulse equation: Bäcklund transformations and applications. Studies in Applied Mathematics, 145(4):791–811, November 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12336.

Melkonian:1990:AND

[MM90] S. Melkonian and S. A. Maslowe. Analysis of a nonlinear diffusive amplitude equation for waves on thin films. Studies in Applied Mathematics, 82(1):37–48, January 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199082137.

Majda:1993:EMF

[MM93] Andrew J. Majda and Richard M. McLaughlin. The effect of mean flows on enhanced diffusivity in transport by incompressible periodic velocity fields. Studies in Applied Mathematics, 89(3):245–279, August 1993. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1993893245.

Mallier:1999:WNE

[MM99a] R. Mallier and S. A. Maslowe. Weakly nonlinear evolution of a wave packet in a zonal mixing layer. Studies in Applied Mathematics, 102(1):69–85, January 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00104.

Montgomery:1999:TLG

[MM99b] P. J. Montgomery and T. B. Moodie. Two-layer gravity currents with topography. Studies in Applied Mathematics, 102(3):221–266, April 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00110.

Montgomery:2001:JCH

[MM01a] P. J. Montgomery and T. B. Moodie. Jump conditions for hyperbolic systems of forced conservation laws with an application to gravity currents. Studies in Applied Mathematics, 106(3):367–392, April 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00170.

Montgomery:2001:NCQ

[MM01b] P. J. Montgomery and T. B. Moodie. On the number of conserved quantities for the two-layer shallow-water equations. Studies in Applied Mathematics, 106(2):229–259, February 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00166.

Montgomery:2003:GRS

[MM03] Patrick J. Montgomery and T. Bryant Moodie. Generalization of a relaxation scheme for systems of forced nonlinear hyperbolic conservation laws with spatially dependent flux functions. Studies in Applied Mathematics, 110(1):1–19, January 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00228.

Margetis:2017:WHM

[MML17] Dionisios Margetis, Matthias Maier, and Mitchell Luskin. On the Wiener–Hopf method for surface plasmons: Diffraction from semiinfinite metamaterial sheet. Studies in Applied Mathematics, 139(4):599–625, November 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12180.

Ma:2023:IND

[MMO23] Manjun Ma, Wentao Meng, and Chunhua Ou. Impact of nonlocal dispersal and time periodicity on the global exponential stability of bistable traveling waves. Studies in Applied Mathematics, 150(3):818–840, April 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12557.

Malaga:2013:CMI

[MMPS13] C. Málaga, A. Minzoni, R. G. Plaza, and C. Simeoni. A chemotactic model for interaction of antagonistic microflora colonies: Front asymptotics and numerical simulations. Studies in Applied Mathematics, 130(3):264–294, April 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12000.

Marcato:2023:CPM

[MMRL23] G. A. Marcato, F. Marcellán, A. Sri Ranga, and Yen Chi Lun. Coherent pairs of measures of the second kind on the real line and Sobolev orthogonal polynomials. An application to a Jacobi case. Studies in Applied Mathematics, 151(2):475–508, August 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12583.

Memoli:2022:DDI

[MN22] Facundo Mémoli and Tom Needham. Distance distributions and inverse problems for metric measure spaces. Studies in Applied Mathematics, 149(4):943–1001, November 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12526.

Mikhailov:2007:CIN

[MNW07] Alexander V. Mikhailov, Vladimir S. Novikov, and Jing Ping Wang. On classification of integrable nonevolutionary equations. Studies in Applied Mathematics, 118(4):419–457, May 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2007.00376.x.

Mudavanhu:2001:RGM

[MO01] B. Mudavanhu and R. E. O’Malley, Jr. A renormalization group method for nonlinear oscillators. Studies in Applied Mathematics, 107(1):63–79, July 2001. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.1071178.

Mori:2021:ASB

[MO21] Yoichiro Mori and Laurel Ohm. Accuracy of slender body theory in approximating force exerted by thin fiber on viscous fluid. Studies in Applied Mathematics, 147(1):127–179, July 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12380.

Montejano:1974:PUC

[Mon74] Luis Montejano. On a problem of Ulam concerning a characterization of the sphere. Studies in Applied Mathematics, 53(3):243–248, September 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1974533243.

Moore:1978:EMV

[Moo78] D. W. Moore. The equation of motion of a vortex layer of small thickness. Studies in Applied Mathematics, 58(2):119–140, February 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1978582119.

Moodie:2000:HTP

[Moo00] T. B. Moodie. Hydraulic theory and particle-driven gravity currents. Studies in Applied Mathematics, 105(2):115–120, August 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00144.

Morris:1980:FEU

[Mor80] Robert A. Morris. Frobenius endomorphisms in the Umbral calculus. Studies in Applied Mathematics, 62(1):85–92, February 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm198062185.

Moroz:1986:AEN

[Mor86] Irene M. Moroz. Amplitude expansions and normal forms in a model tor thermohaline convection. Studies in Applied Mathematics, 74(2):155–170, April 1986. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1986742155.

Moroz:1989:MIT

[Mor89] Irene M. Moroz. Multiple instabilities in a triply diffusive system. Studies in Applied Mathematics, 80(2):137–164, April 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1989802137.

Moses:1978:GDI

[Mos78] Harry E. Moses. A generalization of the direct and inverse problem for the radial Schrödinger equation. Studies in Applied Mathematics, 58(3):187–207, June 1978. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1978583187.

Moses:1979:EER

[Mos79] H. E. Moses. An example of the effect of rescaling of the reflection coefficient on the scattering potential for the one-dimensional Schrödinger equation. Studies in Applied Mathematics, 60(2):177–181, April 1979. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1979602177.

Mostafazadeh:2014:SSC

[Mos14] Ali Mostafazadeh. Spectral singularities and CPA-laser action in a weakly nonlinear PT -symmetric bilayer slab. Studies in Applied Mathematics, 133(4):353–372, November 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12054.

Moodie:2001:NEP

[MP01] T. B. Moodie and J. P. Pascal. Nonhydraulic effects in particle-driven gravity currents in deep surroundings. Studies in Applied Mathematics, 107(3):217–251, October 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00186.

Makridin:2019:MDC

[MP19] Zakhar V. Makridin and Maxim V. Pavlov. Multi-dimensional conservation laws and integrable systems. Studies in Applied Mathematics, 143(4):339–355, November 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12280.

Makridin:2022:MCL

[MP22] Zakhar V. Makridin and Maxim V. Pavlov. Multidimensional conservation laws and integrable systems II. Studies in Applied Mathematics, 148(2):813–824, February 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12460.

Moodie:2000:MSD

[MPB00] T. B. Moodie, J. P. Pascal, and John C. Bowman. Modeling sediment deposition patterns arising from suddenly released fixed-volume turbulent suspensions. Studies in Applied Mathematics, 105(4):333–359, November 2000. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00154.

Moodie:2005:TEM

[MPD05] T. B. Moodie, J. P. Pascal, and S. J. D. D’Alessio. Thermally enhanced motions of variable-inflow surface gravitydriven flows. Studies in Applied Mathematics, 115(4):405–432, November 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9590.2005.00319.
x.

Mondal:2017:FGW

[MPG17] Arpita Mondal, Srikumar Panda, and R. Gayen. Flexuralgravity wave scattering by a circular-arc-shaped porous plate. Studies in Applied Mathematics, 138(1):77–102, January 2017.

CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12137.

Moodie:1998:STD

[MPS98] T. B. Moodie, J. P. Pascal, and G. E. Swaters. Sediment transport and deposition from a two-layer fluid model of gravity currents on sloping bottoms. Studies in Applied Mathematics, 100(3):215–244, April 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00076.

Miller:2015:IBV

[MQ15] Peter D. Miller and Zhenyun Qin. Initial-boundary value problems for the defocusing nonlinear Schrödinger equation in the semiclassical limit. Studies in Applied Mathematics, 134(3):276–362, April 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12075.

Martin:2022:ESI

[MQ22] Calin Iulian Martin and Ronald Quirchmayr. Exact solutions and internal waves for the Antarctic circumpolar current in spherical coordinates. Studies in Applied Mathematics, 148(3):1021–1039, April 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12467.

Mottoni:1980:SPA

[MR80] P. Mottoni and F. Rothe. A singular perturbation analysis for a reaction–diffusion system describing pattern formation. Studies in Applied Mathematics, 63(3):227–247, December 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1980633227.

Majda:1984:RIW

[MR84a] Andrew Majda and Rodolfo Rosales. Resonantly interacting weakly nonlinear hyperbolic waves. I. A single space variable. Studies in Applied Mathematics, 71(2):149–179, October 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1984712149.

Majda:1984:TSM

[MR84b] Andrew Majda and Rodolfo Rosales. A theory for spontaneous Mach-stem formation in reacting shock fronts. II. Steady-wave bifurcations and the evidence for breakdown. Studies in Applied Mathematics, 71(2):117–148, October 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1984712117.

Majda:1992:LFM

[MR92] A. Majda and V. Roytburd. Low-frequency multidimensional instabilities for reacting shock waves. Studies in Applied Mathematics, 87(2):135–174, August 1992. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1992872135.

McAsey:1993:SCF

[MR93] M. McAsey and L. A. Rubel. Some closed-form solutions of Burgers’ equation. Studies in Applied Mathematics, 88(3):173– 190, April 1993. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1993883173.

Muriel:2002:IEA

[MR02] C. Muriel and J. L. Romero. Integrability of equations admitting the nonsolvable symmetry algebra so(3, R). Studies in Applied Mathematics, 109(4):337–352, November 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00227.

Miqueles:2011:IXR

[MR11] Eduardo X. Miqueles and Álvaro Rodolfo De Pierro. On the inversion of the xfct Radon transform. Studies in Applied Mathematics, 127(4):394–419, November 2011. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00527.x.

Mansfield:2019:URM

[MRE19] E. L. Mansfield and A. Rojo-Echeburúa. On the use of the rotation minimizing frame for variational systems with Euclidean symmetry. Studies in Applied Mathematics, 143(3):244–271, October 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12275.

Majda:1988:CSI

[MRS88] Andrew Majda, Rodolfo Rosales, and Maria Schonbek. A canonical system of integrodifferential equations arising in resonant nonlinear acoustics. Studies in Applied Mathematics, 79(3):205–262, December 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1988793205.

Marciniak:2007:SSC

[MRW07] Krzysztof Marciniak and Stefan Rauch-Wojciechowski. Separable systems of coordinates for triangular Newton equations q̈ = M(q, …, q). Studies in Applied Mathematics, 118(1):45–84, January 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2007.00363.x.

Milinazzo:1976:TMP

[MS76] F. Milinazzo and P. G. Saffman. Turbulence model predictions for the inhomogeneous mixing layer. Studies in Applied Mathematics, 55(1):45–63, March 1976. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197655145.

Makri:2008:ISI

[MS08] Effie Makri and Yannis C. Stamatiou. An interactive, similarity increasing algorithm for random strings with applications to key agreement in ad hoc networks. Studies in Applied Mathematics, 121(2):141–155, August 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00408.x.

Martin:2016:TGS

[MS16] Ruth A. Martin and Harvey Segur. Toward a general solution of the three-wave partial differential equations. Studies in Applied Mathematics, 137(1):70–92, July 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12133.

Mulone:2007:SEM

[MSW07] G. Mulone, B. Straughan, and Wendi Wang. Stability of epidemic models with evolution. Studies in Applied Mathematics, 118(2):117–132, February 2007. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2007.00367.x.

Milewski:2002:RWI

[MTVE02] Paul A. Milewski, Esteban G. Tabak, and Eric Vanden-Eijnden. Resonant wave interaction with random forcing and dissipation. Studies in Applied Mathematics, 108 (1):123–144, January 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
01427.

Moyles:2016:ESN

[MTW16] I. Moyles, W. H. Tse, and M. J. Ward. Explicitly solvable nonlocal eigenvalue problems and the stability of localized stripes in reaction–diffusion systems. Studies in Applied Mathematics, 136(1):89–136, January 2016. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12093.

Mazzocco:2013:CQT

[MV13] Marta Mazzocco and Raimundas Vidunas. Cubic and quartic transformations of the sixth Painlevé equation in terms of Riemann–Hilbert correspondence. Studies in Applied Mathematics, 130(1):17–48, January 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2012.00562.x.

Milewski:1998:SFS

[MVBK98] P. Milewski, J.-M. Vanden-Broeck, and Joseph B. Keller. Singularities on free surfaces of fluid flows. Studies in Applied Mathematics, 100(3):245–267, April 1998. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00077.

Mallett:1973:SGD

[MW73] R. L. Mallett and F. Y. M. Wan. The static-geometric duality and a staggered mesh scheme in the numerical solution of some Shell problems. Studies in Applied Mathematics, 52(1):21–38, March 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm197352121.

Mansfield:1998:OPF

[MW98] Elizabeth L. Mansfield and Helen N. Webster. On oneparameter families of Painlevé III. Studies in Applied Mathematics, 101(3):321–341, October 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00096.

Milewski:2013:TDF

[MW13] P. A. Milewski and Z. Wang. Three dimensional flexural– gravity waves. Studies in Applied Mathematics, 131(2):135– 148, August 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12005.

Miller:2016:DSB

[MW16] Peter D. Miller and Alfredo N. Wetzel. Direct scattering for the Benjamin–Ono equation with rational initial data. Studies in Applied Mathematics, 137(1):53–69, July 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12101.

Margetis:2023:SHM

[MWL23] Dionisios Margetis, Alexander B. Watson, and Mitchell Luskin. On the Su–Schrieffer–Heeger model of electron transport:

Low-temperature optical conductivity by the Mellin transform. Studies in Applied Mathematics, 151(2):555–584, August 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12604.

Ma:2022:SPN

[MWZ22] Liangliang Ma, Jiahong Wu, and Qian Zhang. Stability of 3D perturbations near a special 2D solution to the rotating Boussinesq equations. Studies in Applied Mathematics, 148(4):1624– 1655, May 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12486.

Ma:2023:IST

[MZ23] Xinxin Ma and Junyi Zhu. Inverse scattering transform for the two-component Gerdjikov–Ivanov equation with nonzero boundary conditions: Dark–dark solitons. Studies in Applied Mathematics, 151(2):676–715, August 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12607.

Nachman:1984:MISb

[NA84a] Adrian I. Nachman and Mark J. Ablowitz. Multidimensional inverse scattering for first-order systems. Studies in Applied Mathematics, 71(3):251–262, December 1984. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1984713251.

Nachman:1984:MISa

[NA84b] Adrian I. Nachman and Mark J. Ablowitz. A multidimensional inverse-scattering method. Studies in Applied Mathematics, 71(3):243–250, December 1984. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1984713243.

Nair:1988:AEI

[NA88] S. Nair and A. M. Ahmadshahi. Asymptotic expansions and influence coefficients for edge-loaded conical shells. Studies in Applied Mathematics, 78(2):91–110, April 1988. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198878291.

Naumkin:1992:ABS

[Nau92] P. I. Naumkin. Asymptotic behavior of solutions to nonlinear equations with dissipation for large x and t. Studies in Applied Mathematics, 87(1):45–60, July 1992. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199287145.

Nixon:2013:EAL

[NAY13] Sean D. Nixon, T. R. Akylas, and Jianke Yang. Exponential asymptotics for line solitons in two-dimensional periodic potentials. Studies in Applied Mathematics, 131(2):149–178, August 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12006.

Nixon:2017:NAE

[NAY17] Sean D. Nixon, T. R. Akylas, and Jianke Yang. New aspects of exponential asymptotics in multiple-scale nonlinear wave problems. Studies in Applied Mathematics, 139(2):223–247, August 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12179.

Nikitina:2015:DVRa

[NC15a] L. V. Nikitina and L. J. Campbell. Dynamics of vortex Rossby waves in tropical cyclones, Part 1: Linear time-dependent evolution on an f-plane. Studies in Applied Mathematics, 135(4):377–421, November 2015. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12094.

Nikitina:2015:DVRb

[NC15b] L. V. Nikitina and L. J. Campbell. Dynamics of vortex Rossby waves in tropical cyclones, Part 2: Nonlinear time-dependent asymptotic analysis on a β-plane. Studies in Applied Mathematics, 135(4):422–446, November 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12095.

Naumkin:2022:ASP

[NdJVA22] Pavel I. Naumkin and José de Jesús Villela-Aguilar. Asymptotics of solutions to periodic problem for the Korteweg– de Vries–Burgers equation. Studies in Applied Mathematics, 149(2):523–536, August 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12501.

Nec:2012:STS

[Nec12] Y. Nec. Spike-type solutions to one dimensional Gierer– Meinhardt model with Lévy flights. Studies in Applied Mathematics, 129(3):272–299, October 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2012.00550.x.

Nemes:2015:RPI

[Nem15] Gergő Nemes. The resurgence properties of the incomplete gamma function II. Studies in Applied Mathematics, 135(1): 86–116, July 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12077.

Nemes:2018:EBL

[Nem18] Gergő Nemes. Error bounds for the large-argument asymptotic expansions of the Lommel and Allied functions. Studies in Applied Mathematics, 140(4):508–541, May 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12210.

Newton:1991:BNP

[New91] Paul K. Newton. Branching near plane waves in perturbed dispersive systems. Studies in Applied Mathematics, 85(1):75–92, July 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm199185175.

Newell:2002:WTA

[New02] Alan C. Newell. Wave turbulence is almost always intermittent at either small or large scales. Studies in Applied Mathematics, 108(1):39–64, January 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
01422.

Nicholls:2021:ASF

[Nic21] David P. Nicholls. On analyticity of scattered fields in layered structures with interfacial graphene. Studies in Applied Mathematics, 147(2):527–576, August 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12389.

Nishimoto:1979:GDR

[Nis79] Toshihiko Nishimoto. Global dispersion relation for density waves in a certain simplified model of a disk shaped galaxy. Studies in Applied Mathematics, 60(1):11–26, February 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197960111.

Nishikawa:2007:TEW

[Nis07] Yukihiro Nishikawa. Toward the exact WKB analysis of the P_{II} hierarchy. Studies in Applied Mathematics, 119(1):1–71, July 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2007.00377.x.

Newton:2005:PWT

[NK05] Paul K. Newton and Joseph B. Keller. Probability of winning at tennis I. Theory and data. Studies in Applied Mathematics, 114(3):241–269, April 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
0022-2526.2005.01547.x.

Nishino:2002:SNM

[NKUW02] Akinori Nishino, Yasushi Komori, Hideaki Ujino, and Miki Wadati. Symmetrization of nonsymmetric Macdonald polynomials and Macdonald’s inner product identities. Studies in Applied Mathematics, 108(4):399–425, May 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.01431.

Nave:2010:DNS

[NLB10] J.-C. Nave, X. D. Liu, and S. Banerjee. Direct numerical simulation of liquid films with large interfacial Deformation^{†}. Studies in Applied Mathematics, 125(2):153–177, August 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00479.x.

Needham:2023:RPG

[NMBD23] David J. Needham, John C. Meyer, John Billingham, and Catherine Drysdale. The Riemann problem for a generalized Burgers equation with spatially decaying sound speed. I large-time asymptotics. Studies in Applied Mathematics, 150(4):963–995, May 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12561.

Noussair:2000:ANR

[Nou00] Ahmed Noussair. Analysis of nonlinear resonance in conservation laws with point sources and well-balanced scheme. Studies in Applied Mathematics, 104(4):313–352, May 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00137.

Nie:2008:HFC

[NPT08] Linfei Nie, Jigen Peng, and Zhidong Teng. Harmless feedback control for permanence and global asymptotic stability in nonlinear delay population equation. Studies in Applied Mathematics, 120(3):247–263, April 2008. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00396.x.

Nair:1977:AEE

[NR77] S. Nair and E. Reissner. On asymptotic expansions and error bounds in the derivation of two-dimensional Shell theory. Studies in Applied Mathematics, 56(3):189–217, June 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1977563189.

Nanda:2022:ENS

[NR22] Phalguni Nanda and Gujji Murali Mohan Reddy. Efficient numerical solution of one-phase linear inverse Stefan and Cauchy–Stefan problems in two dimensions: a posteriori error control. Studies in Applied Mathematics, 148(4):1563–1585, May 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12484.

Nijhoff:2001:DPE

[NRGO01] F. W. Nijhoff, A. Ramani, B. Grammaticos, and Y. Ohta. On discrete Painlevé equations associated with the lattice KdV systems and the Painlevé VI equation. Studies in Applied Mathematics, 106(3):261–314, April 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00167.

Nadim:1991:MSP

[NS91] Ali Nadim and Howard A. Stone. The motion of small particles and droplets in quadratic flows. Studies in Applied Mathematics, 85(1):53–73, July 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm199185153.

Nimmo:1998:IDD

[NS98] J. J. C. Nimmo and W. K. Schief. An integrable discretization of a 2+1-dimensional sine-Gordon equation. Studies in Applied Mathematics, 100(3):295–309, April 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00079.

Nguyen:2021:DSW

[NS21] Lu Trong Khiem Nguyen and Noel Frederick Smyth. Dispersive shock waves for the Boussinesq Benjamin–Ono equation. Studies in Applied Mathematics, 147(1):32–59, July 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12381.

Newell:2017:ESP

[NV17] Alan C. Newell and Shankar C. Venkataramani. Elastic sheets, phase surfaces, and pattern universes. Studies in Applied Mathematics, 139(2):322–368, August 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12184.

Novikov:2007:SSI

[NW07] Vladimir S. Novikov and Jing Ping Wang. Symmetry structure of integrable nonevolutionary equations. Studies in Applied Mathematics, 119(4):393–428, November 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2007.00390.x.

Nixon:2014:EAS

[NY14] Sean D. Nixon and Jianke Yang. Exponential asymptotics for solitons in PT -symmetric periodic potentials. Studies in Applied Mathematics, 133(4):373–397, November 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12057.

Nixon:2016:BSF

[NY16] Sean D. Nixon and Jianke Yang. Bifurcation of soliton families from linear modes in non-PT -symmetric complex potentials. Studies in Applied Mathematics, 136(4):459–483, May 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12117.

Oaxaca-Adams:1997:EPC

[OAVS97] Guillermo Oaxaca-Adams and Luis Verde-Star. An extremal problem concerning the derivatives of a polynomial at its roots. Studies in Applied Mathematics, 98(1):47–59, January 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00040.

Otto:1994:NDV

[OB94] S. R. Otto and Andrew P. Bassom. Nonlinear development of viscous gertler vortices in a three-dimensional boundary layer. Studies in Applied Mathematics, 92(1):17–39, May 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199492117.

Orszag:1970:IVS

[OC70] S. A. Orszag and S. C. Crow. Instability of a vortex sheet leaving a semi-infinite plate. Studies in Applied Mathematics, 49(2):167–181, June 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1970492167.

Oikawa:1987:PNW

[OCB87] M. Oikawa, K. Chow, and D. J. Benney. The propagation of nonlinear wave packets in a shear flow with a free surface. Studies in Applied Mathematics, 76(1):69–92, February 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198776169.

Opanasenko:2023:LAE

[OF23] Stanislav Opanasenko and Evgeny V. Ferapontov. Linearizable Abel equations and the gurevich–Pitaevskii problem. Studies in Applied Mathematics, 150(3):607–628, April 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12552.

Olsson:1995:ODG

[OH95] P. J. Olsson and D. S. Henningson. Optimal disturbance growth in watertable flow. Studies in Applied Mathematics, 94(2):183–210, February 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1995942183.

Ohta:2009:DCN

[Oht09] Yasuhiro Ohta. Discretization of coupled nonlinear Schrödinger equations. Studies in Applied Mathematics, 122(4):427–447, May 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2009.00443.x.

OMalley:2009:EIU

[OK09] Jr. Robert E. O’Malley and Eleftherios Kirkinis. Examples illustrating the use of renormalization techniques for singularly perturbed differential equations. Studies in Applied Mathematics, 122(2):105–122, February 2009. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00425.x.

OMalley:2010:CRG

[OK10] Robert E. O’Malley, Jr. and Eleftherios Kirkinis. A combined renormalization group-multiple scale method for singularly perturbed problems. Studies in Applied Mathematics, 124(4):383–410, May 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2009.00475.x.

OMalley:2015:VPR

[OK15] Robert E. O’Malley, Jr. and Eleftherios Kirkinis. Variation of parameters and the renormalization group method. Studies in Applied Mathematics, 134(2):215–232, February 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12052.

Olive:1988:BPR

[Oli88] Gloria Olive. The ballot problem revisited. Studies in Applied Mathematics, 78(1):21–30, February 1988. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198878121.

Olver:2006:MI

[Olv06] Peter J. Olver. On multivariate interpolation. Studies in Applied Mathematics, 116(2):201–240, February 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2006.00335.x.

Olver:2021:HOS

[Olv21] Peter J. Olver. Higher-order symmetries of underdetermined systems of partial differential equations and Noether’s second theorem. Studies in Applied Mathematics, 147(3):904–913, October 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12404.

Oncu:1994:AAN

[ÖM94] T. Sabri Öncü and T. Bryant Moodie. Asymptotic analysis of a nonlinear problem in thermoelasticity. Studies in Applied Mathematics, 93(2):163–186, October 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1994932163.

ONeil:1969:SFD

[O’N69] Elizabeth J. O’Neil. The stability of flows in a differentially heated rotating fluid system with rigid bottom and free top. Studies in Applied Mathematics, 48(3):227–256, September 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1969483227.

Ortiz:2019:IST

[OP19] Alyssa K. Ortiz and Barbara Prinari. Inverse scattering transform for the defocusing Ablowitz–Ladik system with arbitrarily large nonzero background. Studies in Applied Mathematics, 143(4):373–403, November 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12282.

Dordevic:2022:SMM

[– Dor22] Jasmina – Dor– dević. A stochastic model for malaria and its behavior under insecticide-treated nets. Studies in Applied Mathematics, 149(3):631–656, October 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12515.

Orszag:1969:RIT

[Ors69] Steven A. Orszag. Representation of isotropic turbulence by scalar functions. Studies in Applied Mathematics, 48 (3):275–279, September 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1969483275.

Orszag:1971:NSI

[Ors71a] Steven A. Orszag. Numerical simulation of incompressible flows within simple boundaries. I. Galerkin (spectral) representations. Studies in Applied Mathematics, 50(4):293–327, December 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1971504293.

Orszag:1971:RRF

[Ors71b] Steven A. Orszag. On the resolution requirements of finitedifference schemes. Studies in Applied Mathematics, 50 (4):395–397, December 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1971504395.

Orszag:1972:CPS

[Ors72] Steven A. Orszag. Comparison of pseudospectral and spectral approximation. Studies in Applied Mathematics, 51 (3):253–259, September 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1972513253.

Oztekin:2001:PFS

[ÖSV01] A. Öztekin, B. R. Seymour, and E. Varley. Pump flow solutions of the Navier–Stokes equations. Studies in Applied Mathematics, 107(1):1–41, July 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
1071176.

Ou:2003:TPB

[OW03] C. H. Ou and R. Wong. On a two-point boundary-value problem with spurious solutions. Studies in Applied Mathematics, 111(4):377–408, November 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
t01-1-00039.

Ou:2004:SMN

[OW04] C. H. Ou and R. Wong. Shooting method for nonlinear singularly perturbed boundary-value problems. Studies in Applied Mathematics, 112(2):161–200, February 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2004.01509.x.

Ohta:2011:GDD

[OWY11] Yasuhiro Ohta, Deng-Shan Wang, and Jianke Yang. General N–dark–dark solitons in the coupled nonlinear Schrödinger equations. Studies in Applied Mathematics, 127(4):345–371, November 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9590.2011.00525.
x.

Oxley:1982:CBI

[Oxl82] James G. Oxley. On Crapo’s beta invariant for matroids. Studies in Applied Mathematics, 66(3):267–277, June 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1982663267.

Papatheodorou:1988:RFC

[PAS88] T. S. Papatheodorou, M. J. Ablowitz, and Y. G. Saridakis. A rule for fast computation and analysis of soliton automata. Studies in Applied Mathematics, 79(2):173–184, October 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1988792173.

Pascal:1990:SSSb

[Pas90] H. Pascal. Some self-similar two-phase flows of non-Newtonian fluids through a porous medium. Studies in Applied Mathematics, 82(4):305–318, May 1990. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990824305.

Paterson:1970:UM

[Pat70] Michael S. Paterson. Unsolvability in 3 × 3 matrices. Studies in Applied Mathematics, 49(1):105–107, March 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1970491105.

Preda:2003:MRC

[PB03] Vasile Preda and Costel Bălcău. On maxentropic reconstruction of countable Markov chains and matrix scaling problems. Studies in Applied Mathematics, 111(1):85–100, July 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.t01-1-00228.

Prinari:2011:IST

[PBT11] Barbara Prinari, Gino Biondini, and A. David Trubatch. Inverse scattering transform for the multi-component nonlinear Schrödinger equation with nonzero boundary conditions. Studies in Applied Mathematics, 126(3):245–302, April 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00504.x.

Pan:2023:ATD

[PCH23] Yan Pan, Xiang-Ke Chang, and Xing-Biao Hu. An application of a qd-type discrete hungry Lotka–Volterra equation over finite fields to a decoding problem. Studies in Applied Mathematics, 151(2):450–474, August 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12584.

Ponziani:2000:WNA

[PCZP00] D. Ponziani, C. M. Casciola, F. Zirilli, and R. Piva. Weakly nonlinear analysis of a localized disturbance in Poiseuille flow. Studies in Applied Mathematics, 105(2):121–141, August 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00145.

Peck:1979:SPG

[Pec79] G. W. Peck. Short proof of a general weight Burnside lemma. Studies in Applied Mathematics, 60(2):173–176, April 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1979602173.

Peck:1980:ECS

[Pec80a] G. W. Peck. Erdös conjecture on sums of distinct numbers. Studies in Applied Mathematics, 63(1):87–92, August 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198063187.

Peck:1980:HCA

[Pec80b] G. W. Peck. Hamiltonian cycles of adjacent triples. Studies in Applied Mathematics, 63(3):275–278, December 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980633275.

Peck:1980:OSD

[Pec80c] G. W. Peck. Optimal spreading in an n dimensional rectilinear grid. Studies in Applied Mathematics, 62(1):69–74, February 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198062169.

Petersson:2004:ROC

[PEE04] Niclas Petersson, Norbert Euler, and Marianna Euler. Recursion operators for a class of integrable third-order evolution equations. Studies in Applied Mathematics, 112 (2):201–225, February 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.0022-
2526.2004.01511.x.

Pego:1988:SER

[Peg88] Robert L. Pego. Some explicit resonating waves in weakly nonlinear gas dynamics. Studies in Applied Mathematics, 79(3):263–270, December 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1988793263.

Peng:2017:SCL

[Pen17] Linyu Peng. Symmetries, conservation laws, and Noether’s theorem for differential-difference equations. Studies in Applied Mathematics, 139(3):457–502, October 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12168.

Petkov:2022:NSM

[Pet22] Pavel Petkov. New and simple mathematical description of epidemics including consecutive waves. Studies in Applied Mathematics, 148(3):1339–1363, April 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12477.

Papatheodorou:1989:ETP

[PF89] T. S. Papatheodorou and A. S. Fokas. Evolution theory, periodic particles, and solitons in cellular automata. Studies in Applied Mathematics, 80(2):165–182, April 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1989802165.

Pucci:1995:QGI

[PG95] Edvige Pucci and Giuseppe Giuseppe. Quasisolutions as group-invariant solutions for partial differential equations. Studies in Applied Mathematics, 94(3):211–223, April 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995943211.

Pelinovsky:1996:NLM

[PG96] Dmitry E. Pelinovsky and Roger H. J. Grimshaw. Non local models for envelope waves in a stratified fluid. Studies in Applied Mathematics, 97(4):369–391, November 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996974369.

Phillips:1998:FAR

[Phi98] W. R. C. Phillips. Finite-amplitude rotational waves in viscous shear flows. Studies in Applied Mathematics, 101(1):23–47, July 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00084.

Pullin:1986:NWS

[PJ86] D. I. Pullin and P. A. Jacobs. Nonlinear waves in a shear flow with a vorticity discontinuity. Studies in Applied Mathematics, 75(1):77–93, August 1986. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198675177.

Plotnikov:2002:PSC

[Plo02] P. I. Plotnikov. A proof of the Stokes conjecture in the theory of surface waves. Studies in Applied Mathematics, 108(2):217–244, February 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
01408.

Pascal:2007:SIB

[PMAD07] J. P. Pascal, T. B. Moodie, N. Antar, and S. J. D. D’Alessio. Solutions for initial-boundary-value problems representing gravity currents arising from variable inflow. Studies in Applied Mathematics, 119(2):127–171, August 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1365-2966.2007.00381.x.

Pomeau:1985:STF

[Pom85] Y. Pomeau. The Saffman–Taylor fingers in the limit of very large surface tension. Studies in Applied Mathematics, 73(1):75–89, August 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm198573175.

Prinari:2018:IST

[POvdMG18] Barbara Prinari, Alyssa K. Ortiz, Cornelis van der Mee, and Marek Grabowski. Inverse scattering transform and solitons for square matrix nonlinear Schrödinger equations. Studies in Applied Mathematics, 141(3):308–352, October 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12223.

Pelce:1986:DSU

[PP86] P. Pelce and Y. Pomeau. Dendrites in the small undercooling limit. Studies in Applied Mathematics, 74(3):245–258, June 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1986743245.

Pascal:1990:SSSa

[PP90] H. Pascal and F. Pascal. On some self-similar flows of non-Newtonian fluids through a porous medium. Studies in Applied Mathematics, 82(1):1–12, January 1990. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19908211.

Pelloni:2008:KGE

[PP08] B. Pelloni and D. A. Pinotsis. The Klein–Gordon equation in a domain with time-dependent boundary. Studies in Applied Mathematics, 121(3):291–312, October 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00416.x.

Pava:2016:TOS

[PP16] Jaime Angulo Pava and Ramón G. Plaza. Transverse orbital stability of periodic traveling waves for nonlinear Klein– Gordon equations. Studies in Applied Mathematics, 137 (4):473–501, November 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12131.

Papamikos:2019:LSA

[PP19] Georgios Papamikos and Tristan Pryer. A Lie symmetry analysis and explicit solutions of the two-dimensional ∞- polylaplacian. Studies in Applied Mathematics, 142(1):48–64, January 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12232.

Pedregal:2008:NML

[PPV08] Pablo Pedregal, Francisco Periago, and Jorge Villena. A numerical method of local energy decay for the boundary controllability of time-reversible distributed parameter systems. Studies in Applied Mathematics, 121(1):27–47, July 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00406.x.

Pavithran:1994:CGW

[PR94] S. Pavithran and L. G. Redekopp. The coupling of gravity waves and convection: Amplitude equations and planform selection. Studies in Applied Mathematics, 93(3):209–250, November 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1994933209.

Perasso:2021:WPP

[PRAV21] Antoine Perasso, Quentin Richard, Irene Azzali, and Ezio Venturino. Well-posedness, positivity, and time asymptotics properties for a reaction–diffusion model of plankton communities, involving a rational nonlinearity with singularity. Studies in Applied Mathematics, 146(1):211–232, January 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12345.

Priede:2023:SCT

[Pri23] Janis Priede. Self-contained two-layer shallow-water theory of strong internal bores. Studies in Applied Mathematics, 150(2):457–480, February 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12546.

Padmavati:2001:GSC

[PRNA01] B. Sri Padmavati, G. P. Rajasekhar, S. D. Nigam, and T. Amaranath. Group structure in circle theorem. Studies in Applied Mathematics, 106(4):407–417, May 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00172.

Pandey:1993:FCA

[PS93] R. N. Pandey and H. M. Srivastava. Fractional calculus and its applications involving certain classes of functional relations. Studies in Applied Mathematics, 89(2):153–165, July 1993. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1993892153.

Phillips:1996:FWM

[PS96] W. R. C. Phillips and Q. Shen. A family of wave-mean shear interactions and their instability to longitudinal vortex form. Studies in Applied Mathematics, 96(2):143–161, February 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1996962143.

Payne:1999:CCD

[PS99a] L. E. Payne and B. Straughan. Convergence and continuous dependence for the Brinkman–Forchheimer equations. Studies in Applied Mathematics, 102(4):419–439, May 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00116.

Payne:1999:CEM

[PS99b] L. E. Payne and B. Straughan. Convergence of the equations for a Maxwell fluid. Studies in Applied Mathematics, 103(3):267–278, October 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00128.

Payne:2000:UNS

[PS00] L. E. Payne and B. Sstraughan. Unconditional nonlinear stability in temperature-dependent viscosity flow in a porous medium. Studies in Applied Mathematics, 105(1):59–81, July 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00142.

Payne:2009:DKS

[PS09] L. E. Payne and B. Straughan. Decay for a Keller–Segel chemotaxis model. Studies in Applied Mathematics, 123 (4):337–360, November 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2009.00457.x.

Panthee:2010:ABC

[PS10] M. Panthee and M. Scialom. Asymptotic behavior for a class of solutions to the critical modified Zakharov–Kuznetsov equation. Studies in Applied Mathematics, 124(3):229–245, April 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00469.x.

Paindaveine:2011:SAS

[PS11] Davy Paindaveine and Yvik Swan. A stochastic analysis of some two-person sports. Studies in Applied Mathematics, 127(3):221–249, October 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2011.00517.x.

Pelloni:2018:NMB

[PS18] Beatrice Pelloni and David Andrew Smith. Nonlocal and multipoint boundary value problems for linear evolution equations. Studies in Applied Mathematics, 141(1):46–88, July 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12212.

Posukhovskyi:2020:GSO

[PS20a] Iurii Posukhovskyi and Atanas Stefanov. On the ground states of the ostrovskyi equation and their stability. Studies in Applied Mathematics, 144(4):548–575, May 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12309.

Pucci:2020:PDE

[PS20b] Edvige Pucci and Giuseppe Saccomandi. Partial differential equations admitting a given nonclassical point symmetry. Studies in Applied Mathematics, 145(1):81–96, July 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12312.

Pucci:2021:USR

[PS21] Edvige Pucci and Giuseppe Saccomandi. Using symmetries à Rebours. Studies in Applied Mathematics, 146(1):99–117, January 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12342.

Parker:2022:SDT

[PSAZ22] Ross Parker, Yannan Shen, Alejandro Aceves, and John Zweck. Spatiotemporal dynamics in a twisted, circular waveguide array. Studies in Applied Mathematics, 149(2):537–560, August 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12511.

Pazanin:2023:RIE

[PSG23] Igor Pažanin and Francisco J. Suárez-Grau. Roughnessinduced effects on the thermomicropolar fluid flow through a thin domain. Studies in Applied Mathematics, 151(2):716–751, August 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12611.

Parter:1975:MSD

[PSS75] Seymour V. Parter, Myron L. Stein, and Paul R. Stein. On the multiplicity of solutions of a differential equation arising in chemical reactor theory. Studies in Applied Mathematics, 54(4):293–314, December 1975. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1975544293.

Prokopiou:2018:EES

[PT18] Prokopios Prokopiou and Nikolaos L. Tsitsas. Electromagnetic excitation of a spherical medium by an arbitrary dipole and related inverse problems. Studies in Applied Mathematics, 140(4):438–464, May 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12206.

Pelinovsky:2019:ILT

[PTDS19] Efim Pelinovsky, Tatiana Talipova, Ira Didenkulova, and Ekaterina Didenkulova (Shurgalina). Interfacial long traveling waves in a two-layer fluid with variable depth. Studies in Applied Mathematics, 142(4):513–527, May 2019.

CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12235.

Prinari:2016:IST

[PV16] Barbara Prinari and Federica Vitale. Inverse scattering transform for the focusing Ablowitz–Ladik system with nonzero boundary conditions. Studies in Applied Mathematics, 137(1): 28–52, July 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12103.

Phan:2023:BLP

[PV23] Tuoc Phan and Darío A. Valdebenito. A boundary layer problem in nonflat domains with measurable viscous coefficients. Studies in Applied Mathematics, 150(1):5–34, January 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12531.

Pego:1997:CLS

[PW97] Robert L. Pego and Michael I. Weinstein. Convective linear stability of solitary waves for Boussinesq equations. Studies in Applied Mathematics, 99(4):311–375, November 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00063.

Pan:2012:UAE

[PW12] J. H. Pan and R. Wong. Uniform asymptotic expansions for the discrete Chebyshev polynomials. Studies in Applied Mathematics, 128(4):337–384, May 2012. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00537.x.

Pan:2014:ADC

[PW14] J. H. Pan and R. Wong. Asymptotics of the discrete Chebyshev polynomials. Studies in Applied Mathematics, 132(1):13–49, January 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12018.

Pelinovsky:2000:IOR

[PY00] Dmitry E. Pelinovsky and Jianke Yang. Internal oscillations and radiation damping of vector solitons. Studies in Applied Mathematics, 105(3):245–276, October 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00151.

Pelinovsky:2005:IMV

[PY05] Dmitry E. Pelinovsky and Jianke Yang. Instabilities of multihump vector solitons in coupled nonlinear Schrödinger equations. Studies in Applied Mathematics, 115(1):109–137, July 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2005.01565.

Perez:2003:IPS

[PZ03] L. O. Castellano Pérez and Jorge P. Zubelli. On the inverse problem for scattering of electromagnetic radiation by a periodic structure. Studies in Applied Mathematics, 111(2):115–166, August 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
t01-1-00230.

Qin:1996:NCS

[QC96] Y. Qin and J. Chadam. Nonlinear convective stability in a porous medium with temperature-dependent viscosity and inertial drag. Studies in Applied Mathematics, 96(3):273–288, April 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1996963273.

Qu:2007:CLB

[QJW07] Changzheng Qu, Lina Ji, and Lizhen Wang. Conditional Lie Bäcklund symmetries and sign-invariants to quasi-linear diffusion equations. Studies in Applied Mathematics, 119 (4):355–391, November 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2007.00388.x.

Qin:1994:CIA

[QK94] Y. Qin and P. N. Kaloni. Convective instabilities in anisotropic porous media. Studies in Applied Mathematics, 91(3):189–204, March 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1994913189.

Qu:2014:HTC

[QK14] Changzheng Qu and Jing Kang. Hodograph transformations and Cauchy problem to systems of nonlinear parabolic equations. Studies in Applied Mathematics, 132(2):81–111, February 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12025.

Qu:1997:GCG

[Qu97] Changzheng Qu. Group classification and generalized conditional symmetry reduction of the nonlinear diffusion– convection equation with a nonlinear source. Studies in Applied Mathematics, 99(2):107–136, August 1997. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00058.

Rogers:2010:ERR

[RA10] Colin Rogers and Hongli An. Ermakov–Ray–Reid systems in (2 + 1)-dimensional rotating shallow water theory. Studies in Applied Mathematics, 125(3):275–299, October 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00488.x.

Rabelo:1989:EWD

[Rab89] Mauro L. Rabelo. On equations which describe pseudospherical surfaces. Studies in Applied Mathematics, 81(3):221–248, December 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1989813221.

Ranger:1994:CSE

[Ran94a] K. B. Ranger. A class of solutions for the equations governing the steady two-dimensional motion of a viscous incompressible liquid. Studies in Applied Mathematics, 92(2):145–158, June 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1994922145.

Ranger:1994:CST

[Ran94b] K. B. Ranger. The construction of solutions to the twodimensional Navier Stokes equations. Studies in Applied Mathematics, 91(3):205–214, March 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1994913205.

Ranger:1994:EST

[Ran94c] K. B. Ranger. Explicit solutions of the two-dimensional Navier Stokes equations. Studies in Applied Mathematics, 91(1):27– 37, January 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm199491127.

Ranger:1995:ESS

[Ran95a] K. B. Ranger. Explicit solutions of the steady twodimensional Navier–Stokes equations. Studies in Applied Mathematics, 94(2):169–181, February 1995. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995942169.

Ranger:1995:SSO

[Ran95b] K. B. Ranger. Solutions of the Stokes, Oseen, and Navier–Stokes equations in three dimensions. Studies in Applied Mathematics, 95(2):193–212, August 1995. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995952193.

Ranger:1996:AESb

[Ran96a] K. B. Ranger. Addendum to “Explicit solutions of the steady two-dimensional Navier–Stokes equations”. Studies in Applied Mathematics, 97(1):101–102, July 1996. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996971101.

Ranger:1996:AESa

[Ran96b]  K. B. Ranger.  Addendum to “Explicit solutions of

the two-dimensional Navier–Stokes equations”. Studies in Applied Mathematics, 97(1):99–100, July 1996. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199697199.

Ranger:1996:CAA

[Ran96c] K. B. Ranger. A constructive analytical approach to the calculation of flows representing the steady motion of a viscous incompressible liquid. Studies in Applied Mathematics, 96(3):289–305, April 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1996963289.

Ranger:1998:AMV

[Ran98] K. B. Ranger. Arbitrary motions of a viscous incompressible liquid in a finite region. Studies in Applied Mathematics, 101(2):185–195, August 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00089.

RuizDeZarate:2009:HOI

[RANC09] Ailín Ruiz De Zárate, Daniel G. Alfaro Vigo, André Nachbin, and Wooyoung Choi. A higher-order internal wave model accounting for large bathymetric variations. Studies in Applied Mathematics, 122(3):275–294, April 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00433.x.

Ratliff:2018:DDM

[Rat18] Daniel J. Ratliff. Double degeneracy in multiphase modulation and the emergence of the Boussinesq equation. Studies in Applied Mathematics, 140(1):48–77, January 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12189.

Ratliff:2019:FSM

[Rat19] Daniel J. Ratliff. Flux singularities in multiphase wavetrains and the Kadomtsev–Petviashvili equation with applications to stratified hydrodynamics. Studies in Applied Mathematics, 142(2):109–138, February 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12242.

Rosenblat:1980:PPB

[RC80] S. Rosenblat and Donald S. Cohen. Periodically perturbed bifurcation — 1. Simple bifurcation. Studies in Applied Mathematics, 63(1):1–23, August 1980. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19806311.

Rosenblat:1981:PPB

[RC81] S. Rosenblat and Donals S. Cohen. Periodically perturbed bifurcation. II. Hopf bifurcation. Studies in Applied Mathematics, 64(2):143–175, April 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1981642143.

Rubinstein:2005:UQS

[RC05] Robert Rubinstein and Meelan Choudhari. Uncertainty quantification for systems with random initial conditions using Wiener–Hermite expansions. Studies in Applied Mathematics, 114(2):167–188, February 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
0022-2526.2005.01543.x.

Rao:2017:RSS

[RCH17] J. Rao, Y. Cheng, and J. He. Rational and semirational solutions of the nonlocal Davey–Stewartson equations. Studies in Applied Mathematics, 139(4):568–598, November 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12178.

Rao:2021:CRC

[RCMH21] Jiguang Rao, Kwok Wing Chow, Dumitru Mihalache, and Jingsong He. Completely resonant collision of lumps and line solitons in the Kadomtsev–Petviashvili I equation. Studies in Applied Mathematics, 147(3):1007–1035, October 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12417.

Rudraiah:2004:NFB

[RD04] N. Rudraiah and P. A. Dinesh. Nonlinear flow between permeable disks using computer-extended series method. Studies in Applied Mathematics, 113(2):163–182, August 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2004.01528.x.

Redekopp:1980:SSS

[Red80] Larry G. Redekopp. Similarity solutions of some twospace-dimensional nonlinear wave evolution equations. Studies in Applied Mathematics, 63(3):185–207, December 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980633185.

Reissner:1969:ENS

[Rei69] E. Reissner. On the equations of nonlinear shallow Shell theory. Studies in Applied Mathematics, 48(2):171–175, June 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1969482171.

Reissner:1970:PBI

[Rei70a] E. Reissner. On postbuckling behavior and imperfection sensitivity of thin elastic plates on a non-linear elastic foundation. Studies in Applied Mathematics, 49(1):45–57, March 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197049145.

Reissner:1970:AS

[Rei70b] Eric Reissner. An aging spring. Studies in Applied Mathematics, 49(3):205–224, September 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1970493205.

Reid:1972:CAS

[Rei72] W. H. Reid. Composite approximations to the solutions of the Orr–Sommerfeld equation. Studies in Applied Mathematics, 51(4):341–368, December 1972. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1972514341.

Reissner:1973:KSF

[Rei73a] E. Reissner. On kinematics and statics of finite-strain force and moment stress elasticity. Studies in Applied Mathematics, 52(2):97–101, June 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm197352297.

Reissner:1973:ODL

[Rei73b] E. Reissner. On one-dimensional large-displacement finitestrain beam theory. Studies in Applied Mathematics, 52(2): 87–95, June 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm197352287.

Reid:1974:UAAa

[Rei74a] W. H. Reid. Uniform asymptotic approximations to the solutions of the Orr–Sommerfeld equation. Part 1. Plane Couette flow. Studies in Applied Mathematics, 53(2):91–110, June 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197453291.

Reid:1974:UAAb

[Rei74b] W. H. Reid. Uniform asymptotic approximations to the solutions of the Orr–Sommerfeld equation Part 2. The general theory. Studies in Applied Mathematics, 53(3):217–224, September 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1974533217.

Reissner:1975:NEF

[Rei75] E. Reissner. Note on the equations of finite-strain force and moment stress elasticity. Studies in Applied Mathematics, 54(1):1–8, March 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19755411.

Reiner:1977:MSB

[Rei77] David L. Reiner. Multivariate sequences of binomial type. Studies in Applied Mathematics, 57(2):119–133, October 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1977572119.

Reiner:1978:CPS

[Rei78] David L. Reiner. The combinatorics of polynomial sequences. Studies in Applied Mathematics, 58(2):95–117, February 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm197858295.

Reid:1979:ESO

[Rei79] W. H. Reid. An exact solution of the Orr–Sommerfeld equation for plane Couette flow. Studies in Applied Mathematics, 61(1): 83–91, July 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm197961183.

Reiner:1981:EBT

[Rei81] David L. Reiner. Eulerian binomial type revisited. Studies in Applied Mathematics, 64(1):89–93, February 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198164189.

Reiner:1989:FMR

[Rei89] Victor Reiner. Free modules of relative invariants of finite groups. Studies in Applied Mathematics, 81(2):181–184, October 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1989812181.

Reeger:2013:PIB

[RF13] Jonah A. Reeger and Bengt Fornberg. Painlevé IV with both parameters zero: a numerical study. Studies in Applied Mathematics, 130(2):108–133, February 2013. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00564.x.

Reeger:2016:NQS

[RF16] Jonah A. Reeger and Bengt Fornberg. Numerical quadrature over the surface of a sphere. Studies in Applied Mathematics, 137(2):174–188, August 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12106.

Rasin:2007:SID

[RH07] Olexandr G. Rasin and Peter E. Hydon. Symmetries of integrable difference equations on the quad-graph. Studies in Applied Mathematics, 119(3):253–269, October 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2007.00385.x.

Richman:1984:SRN

[Ric84] David R. Richman. Some remarks on the number of solutions to the equation f(X_{1}) + ⋯ + f(X_{n}) = 0. Studies in Applied Mathematics, 71(3):263–266, December 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1984713263.

Richardson:1992:HSF

[Ric92] S. Richardson. Hele Shaw flows with time-dependent free boundaries involving injection through slits. Studies in Applied Mathematics, 87(2):175–194, August 1992. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1992872175.

Rauch:1978:BCS

[RJM78] Harry E. Rauch, Neal H. Jacobs, and Jonathan L. Marz. Buckling of a complete spherical Shell under uniform external pressure. Studies in Applied Mathematics, 58(2):141–158, February 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1978582141.

Richard-Jung:2017:NCE

[RJRTF17] F. Richard-Jung, J.-P. Ramis, J. Thomann, and F. Fauvet. New characterizations for the eigenvalues of the prolate spheroidal wave equation. Studies in Applied Mathematics, 138(1):3–42, January 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12134.

Rosenbaum:1983:SPC

[RKS83] D. M. Rosenbaum, D. J. Kleitman, and M. E. Saks. A statistical procedure for cluster recognition with application to Atlanta leukemia–lymphoma data. Studies in Applied Mathematics, 68(1):61–88, February 1983. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198368161.

Reid:2001:DEC

[RLW01] Gregory J. Reid, Ping Lin, and Allan D. Wittkopf. Differential elimination–completion algorithms for DAE and PDAE. Studies in Applied Mathematics, 106(1):1–45, January 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00159.

Rogers:2012:ERR

[RMA12] Colin Rogers, Boris Malomed, and Hongli An. Ermakov–Ray– Reid reductions of variational approximations in nonlinear optics. Studies in Applied Mathematics, 129(4):389–413, November 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2012.00557.x.

Rao:2022:GHO

[RMMH22] Jiguang Rao, Boris A. Malomed, Dumitru Mihalache, and Jingsong He. General higher — order breathers and rogue waves in the two-component long-wave–short-wave resonance interaction model. Studies in Applied Mathematics, 149 (4):843–878, November 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12523.

Rao:2015:SSS

[RN15] Ch. Srinivasa Rao and Smriti Nath. A study of separable solutions of a generalized Burgers equation. Studies in Applied Mathematics, 134(4):403–419, May 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12079.

Rodrigo:2003:EBF

[Rod03] M. Rodrigo. Evolution of bounding functions for the solution of the KPP–Fisher equation in bounded domains. Studies in Applied Mathematics, 110(1):49–61, January 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00230.

Rogers:2014:NEP

[Rog14a] Colin Rogers. A novel Ermakov–Painlevé II system: N + 1dimensional coupled NLS and elastodynamic reductions. Studies in Applied Mathematics, 133(2):214–231, August 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12039.

Rogers:2014:CNS

[Rog14b] Colin Rogers. On a coupled nonlinear Schrödinger system: a Ermakov connection. Studies in Applied Mathematics, 132(3):247–265, April 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12027.

Rosenblat:1976:AES

[Ros76a] S. Rosenblat. Asymptotically equivalent singular perturbation problems. Studies in Applied Mathematics, 55(3):249–280, September 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1976553249.

Roskes:1976:SNM

[Ros76b] G. J. Roskes. Some nonlinear multiphase interactions. Studies in Applied Mathematics, 55(3):231–238, September 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1976553231.

Rosales:1978:ESS

[Ros78] Rodoljo R. Rosales. Exact solutions of some nonlinear evolution equations. Studies in Applied Mathematics, 59 (2):117–151, October 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1978592117.

Rosini:2010:SSR

[Ros10] M. D. Rosini. Stability of surface Rayleigh waves in an elastic half-space. Studies in Applied Mathematics, 124 (2):179–211, February 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2009.00467.x.

Rosales:1983:GWC

[RP83] Rodolfo R. Rosales and George C. Papanicolaou. Gravity waves in a channel with a rough bottom. Studies in Applied Mathematics, 68(2):89–102, April 1983. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198368289.

Rogers:2011:DWB

[RP11] Colin Rogers and Oktay Pashaev. On a 2+1-dimensional Whitham–Broer–Kaup system: a resonant NLS connection. Studies in Applied Mathematics, 127(2):141–152, August 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00514.x.

Riano:2022:HDG

[RRY22] Oscar Riaño, Svetlana Roudenko, and Kai Yang. Higher dimensional generalization of the Benjamin–Ono equation: 2D case. Studies in Applied Mathematics, 148(2):498–542, February 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12448.

Rosenblat:1980:MSN

[RS80] S. Rosenblat and R. Szeto. Multiple solutions of nonlinear boundary-value problems. Studies in Applied Mathematics, 63(2):99–117, October 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm198063299.

Robinson:1984:SSS

[RS84] A. C. Robinson and P. G. Saffman. Stability and structure of stretched vortices. Studies in Applied Mathematics, 70(2):163– 181, April 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1984702163.

Rogers:1998:IGN

[RS98] C. Rogers and W. K. Schief. Intrinsic geometry of the NLS equation and its Auto–Bäcklund transformation. Studies in Applied Mathematics, 101(3):267–287, October 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00093.

Roy:2007:SHF

[RS07] Samir Roy and S. Sundar. Study of heat flow through highly porous heat insulators. Studies in Applied Mathematics, 118(1):1–15, January 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2007.00361.x.

Rao:2008:AWS

[RS08] Ch. Srinivasa Rao and Engu Satyanarayana. Asymptotic N-wave solutions of the nonplanar Burgers equation. Studies in Applied Mathematics, 121(2):199–221, August 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00411.x.

Rogers:2010:BTS

[RS10a] C. Rogers and W. K. Schief. Bäcklund transformations and superposition principles in nonlinear elastodynamics. Studies in Applied Mathematics, 124(2):137–149, February 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00465.x.

Rogers:2010:IST

[RS10b] C. Rogers and W. K. Schief. Integrable structure in the theory of stress concentration in shear-strained nonlinear elastic materials. Studies in Applied Mathematics, 125(1):39–53, July 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00480.x.

Rao:2011:LTA

[RS11] Ch. Srinivasa Rao and Engu Satyanarayana. Large-time asymptotics for solutions of a generalized Burgers equation with variable viscosity. Studies in Applied Mathematics, 127(1):1–23, July 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2011.00509.x.

Rogers:2012:GCL

[RS12] C. Rogers and A. Szereszewski. On the geometry of complexlamellar magnetohydrodynamics: Universal motions. Studies in Applied Mathematics, 128(3):225–251, April 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00535.x.

Rogers:2014:BCS

[RS14] C. Rogers and W. K. Schief. On a Boussinesq capillarity system: Hamiltonian reductions and associated quartic geometries. Studies in Applied Mathematics, 132(1):1–12, January 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12017.

Rybalko:2021:CWL

[RS21] Yan Rybalko and Dmitry Shepelsky. Curved wedges in the long-time asymptotics for the integrable nonlocal nonlinear Schrödinger equation. Studies in Applied Mathematics, 147(3):872–903, October 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12403.

Rogers:2005:TVS

[RSCM05] C. Rogers, W. K. Schief, K. W. Chow, and C. C. Mak. On Tzitzéica vortex streets and their reciprocals in subsonic gas dynamics. Studies in Applied Mathematics, 114(3):271–283, April 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.0022-2526.2005.01548.x.

Reilly:1979:CSP

[RT79] J. W. Reilly and S. M. Tanny. Counting successions in permutations. Studies in Applied Mathematics, 61(1):73–81, July 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197961173.

Rosales:2002:RTI

[RTT02] Rodolfo R. Rosales, Esteban G. Tabak, and Cristina V. Turner. Resonant triads involving a nondispersive wave. Studies in Applied Mathematics, 108(1):105–122, January 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.01426.

Rubenfeld:1977:PWC

[Rub77] L. A. Rubenfeld. The passage of weakly coupled nonlinear oscillators through internal resonance. Studies in Applied Mathematics, 57(1):77–92, July 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm197757177.

Russell:1973:UBC

[Rus73] David L. Russell. A unified boundary controllability theory for hyperbolic and parabolic partial differential equations. Studies in Applied Mathematics, 52(3):189–211, September 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1973523189.

Rosler:2018:BDS

[RV18] Margit Rösler and Michael Voit. Beta distributions and Sonine integrals for Bessel functions on symmetric cones. Studies in Applied Mathematics, 141(4):474–500, November 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12217.

Reissner:1969:ELS

[RW69a] E. Reissner and F. Y. M. Wan. On the equations of linear shallow Shell theory. Studies in Applied Mathematics, 48(2):133– 145, June 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1969482133.

Reissner:1969:RSS

[RW69b] E. Reissner and F. Y. M. Wan. Rotationally symmetric stress and strain in shells of revolution. Studies in Applied Mathematics, 48(1):1–17, March 1969. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19694811.

Reissner:1971:RSS

[RW71] E. Reissner and F. Y. M. Wan. On rotationally symmetric stress and strain in anisotropic shells of revolution. Studies in Applied Mathematics, 50(4):391–394, December 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1971504391.

Rui:1996:ABP

[RW96] Bo Rui and R. Wong. Asymptotic behavior of the Pollaczek polynomials and their zeros. Studies in Applied Mathematics, 96(3):307–338, April 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1996963307.

Rao:2010:SNB

[RY10] Ch. Srinivasa Rao and Manoj K. Yadav. Solutions of a nonhomogeneous Burgers equation. Studies in Applied Mathematics, 124(4):411–422, May 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2009.00478.x.

Rybkin:2022:BDT

[Ryb22] Alexei Rybkin. The binary Darboux transformation revisited and KdV solitons on arbitrary short-range backgrounds. Studies in Applied Mathematics, 148(1):141–153, January 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12436.

Rybkin:2023:CAB

[Ryb23] Alexei Rybkin. A continuous analog of the binary Darboux transformation for the Korteweg–de Vries equation. Studies in Applied Mathematics, 151(1):208–246, July 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12578.

Skopovi:2004:TND

[SA04] Ivan Skopovi and T. R. Akylas. Transient nonlinear disturbances and shelves in a stratified fluid layer. Studies in Applied Mathematics, 113(2):139–161, August 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2004.01527.x.

Sharma:2005:SSS

[SA05] V. D. Sharma and Rajan Arora. Similarity solutions for strong shocks in an ideal gas. Studies in Applied Mathematics, 114(4):375–394, May 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
0022-2526.2005.01557.x.

Shatnawi:2019:RBM

[SA19] Taqi A. M. Shatnawi and David E. Amundsen. Resonances in bounded media: Nonlinear and boundary effects. Studies in Applied Mathematics, 143(2):192–210, August 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12274.

Saffman:1970:VVV

[Saf70] P. G. Saffman. The velocity of viscous vortex rings. Studies in Applied Mathematics, 49(4):371–380, December 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1970494371.

Saffman:1971:BCS

[Saf71a] P. G. Saffman. On the boundary condition at the surface of a porous medium. Studies in Applied Mathematics, 50(2):93– 101, June 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm197150293.

Saffman:1971:SDR

[Saf71b] P. G. Saffman. On the spectrum and decay of random twodimensional vorticity distributions at large Reynolds number. Studies in Applied Mathematics, 50(4):377–383, December 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1971504377.

Saffman:1972:MVP

[Saf72] P. G. Saffman. The motion of a vortex pair in a stratified atmosphere. Studies in Applied Mathematics, 51(2):107–119, June 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1972512107.

Saffman:1973:SSF

[Saf73] P. G. Saffman. On the settling speed of free and fixed suspensions. Studies in Applied Mathematics, 52(2):115–127, June 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1973522115.

Saffman:1974:MET

[Saf74] P. G. Saffman. Model equations for turbulent shear flow. Studies in Applied Mathematics, 53(1):17–34, March 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197453117.

Saffman:1975:FVR

[Saf75] P. G. Saffman. On the formation of vortex rings. Studies in Applied Mathematics, 54(3):261–268, September 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1975543261.

Satyanarayana:2017:LTA

[SAM17] Engu Satyanarayana, Mohd Ahmed, and V. Murugan. Large time asymptotics with error estimates to solutions of a forced Burgers equation. Studies in Applied Mathematics, 138(2):185–204, February 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12148.

Sarkaria:2001:BP

[Sar01] K. S. Sarkaria. Bénard polygons. Studies in Applied Mathematics, 106(2):139–155, February 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
t01-1-00163.

Schief:2013:UIA

[SAR13] W. K. Schief, H. An, and C. Rogers. Universal and integrable aspects of an elliptic vortex representation in 2+1dimensional magneto-gasdynamics. Studies in Applied Mathematics, 130(1):49–79, January 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2012.00559.x.

Sattinger:1985:HHS

[Sat85] D. H. Sattinger. Hamiltonian hierarchies on semisimple Lie algebras. Studies in Applied Mathematics, 72(1):65–86, February 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198572165.

Satir:1999:CQP

[Sat99] Ahmet Satır. Classification of q_{t} = P (x, t, q, q_{x}, q_{xx}). Studies in Applied Mathematics, 102(2):205–219, February 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00109.

Smith:1987:PTD

[SB87] F. T. Smith and R. J. Bodonyi. Properties of a timedependent nonlinear viscous critical layer. Studies in Applied Mathematics, 77(2):129–150, October 1987. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1987772129.

Sipcic:2000:LIC

[SB00] R. Sipcic and D. J. Benney. Lump interactions and collapse in the modified Zakharov–Kuznetsov equation. Studies in Applied Mathematics, 105(4):385–403, November 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00157.

Sachdev:2005:BVP

[SBA05] P. L. Sachdev, N. M. Bujurke, and V. B. Awati. Boundary value problems for third-order nonlinear ordinary differential equations. Studies in Applied Mathematics, 115 (3):303–318, October 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2005.00310.x.

Strauszer-Caussade:2023:WGF

[SCFFLPA23] Thomas Strauszer-Caussade, Luiz M. Faria, Agustín Fernandez-Lado, and Carlos Pérez-Arancibia. Windowed green function method for wave scattering by periodic arrays of 2D obstacles. Studies in Applied Mathematics, 150(1):277–315, January 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12540.

Schaeffer:1975:EDF

[Sch75] David G. Schaeffer. On the existence of discrete frequencies of oscillation in a rotating fluid. Studies in Applied Mathematics, 54(3):269–274, September 1975. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1975543269.

Schay:1980:EOJ

[Sch80] Geza Schay. Erratum: Optimal joint distributions of several random variables with given marginals. Studies in Applied Mathematics, 62(3):281, June 1980. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980623281. See [Gez79].

Schwarz:1998:SAA

[Sch98] Fritz Schwarz. Symmetry analysis of Abel’s equation. Studies in Applied Mathematics, 100(3):269–294, April 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00078.

Schief:2001:ISS

[Sch01] W. K. Schief. Isothermic surfaces in spaces of arbitrary dimension: Integrability, discretization, and Bäcklund transformations — a discrete Calapso equation. Studies in Applied Mathematics, 106(1):85–137, January 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00162.

Sheils:2015:IPD

[SD15] Natalie E. Sheils and Bernard Deconinck. Interface problems for dispersive equations. Studies in Applied Mathematics, 134(3):253–275, April 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12070.

Sheils:2016:IIM

[SD16] Natalie E. Sheils and Bernard Deconinck. Initial-to-interface maps for the heat equation on composite domains. Studies in Applied Mathematics, 137(1):140–154, July 2016. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12138.

Swaters:2012:CLH

[SDS12] G. E. Swaters, H. V. Dosser, and B. R. Sutherland. Conservation laws, Hamiltonian structure, modulational instability properties and solitary wave solutions for a higherorder model describing nonlinear internal waves. Studies in Applied Mathematics, 128(2):159–182, February 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00533.x.

Segura:2021:UVS

[Seg21] Javier Segura. Uniform (very) sharp bounds for ratios of parabolic cylinder functions. Studies in Applied Mathematics, 147(2):816–833, August 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12401.

Senato:1989:BPF

[Sen89] Domenico Senato. A bijective proof of the fundamental theorem on symmetric functions. Studies in Applied Mathematics, 81(1):1–5, August 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19898111.

Sachdev:2003:LTA

[SER^{+}03] P. L. Sachdev, B. O. Enflo, Ch. Srinivasa Rao, B. Mayil Vaganan, and Poonam Goyal. Large-time asymptotics for periodic solutions of some generalized Burgers equations. Studies in Applied Mathematics, 110(2):181–204, February 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00236.

Segur:1985:AMP

[SF85] Harvey Segur and Allan Finkel. An analytical model of periodic waves in shallow water. Studies in Applied Mathematics, 73(3):183–220, December 1985. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1985733183.

Shen:2016:RCC

[SFO16] Shoufeng Shen, Bao-Feng Feng, and Yasuhiro Ohta. From the real and complex coupled dispersionless equations to the real and complex short pulse equations. Studies in Applied Mathematics, 136(1):64–88, January 2016. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12092.

Sachdev:1990:ETW

[SG90] P. L. Sachdev and Neelam Gupta. Exact travelingwave solutions for model geophysical systems. Studies in Applied Mathematics, 82(4):267–289, May 1990. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990824267.

Stanshine:1976:ASFa

[SH76] J. A. Stanshine and L. N. Howard. Asymptotic solutions of the Field–Noyes model for the Belousov reaction — I. Homogeneous oscillations. Studies in Applied Mathematics, 55(2):129– 165, June 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1976552129.

Shanthini:1991:NDS

[Sha91] Rajaratnam Shanthini. On the nonlinear development of small three-dimensional perturbations in plane Poiseuille flow: Single-mode approach. Studies in Applied Mathematics, 84(2):93–104, February 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm199184293.

Sen:2006:LPS

[SHC06] A. Sen, A. N. W. Hone, and P. A. Clarkson. On the Lax pairs of the symmetric Painlevé equations. Studies in Applied Mathematics, 117(4):299–319, November 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2006.00356.x.

Shearer:1982:BHI

[She82a] James B. Shearer. Barring hexominoes from the infinite checkerboard. Studies in Applied Mathematics, 67(3):243–255, December 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1982673243.

Shearer:1982:BPI

[She82b] James B. Shearer. Barring the Z pentomino from the infinite checkerboard. Studies in Applied Mathematics, 67(1):73–77, August 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm198267173.

Shen:1999:CWU

[She99] Jianhong Shen. Combinatorics for wavelets: The Umbral refinement equation. Studies in Applied Mathematics, 103(2):121–147, August 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00123.

Singh:2021:NFV

[Sin21] Mehakpreet Singh. New finite volume approach for multidimensional Smoluchowski equation on nonuniform grids. Studies in Applied Mathematics, 147(3):955–977, October 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12415.

Srivastava:1991:FCS

[SJB91] H. M. Srivastava, J. M. C. Joshi, and C. S. Bisht. Fractional calculus and the sum of powers of natural numbers. Studies in Applied Mathematics, 85(2):183–193, August 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991852183.

Sachdev:2005:ESC

[SJH05] P. L. Sachdev, K. T. Joseph, and M. Ejanul Haque. Exact solutions of compressible flow equations with spherical symmetry. Studies in Applied Mathematics, 114(4):325–342, May 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2005.01552.x.

Sachdev:1996:EWS

[SJV96] P. L. Sachdev, K. T. Joseph, and B. Mayil Vaganan. Exact N–wave solutions of generalized Burgers equations. Studies in Applied Mathematics, 97(4):349–367, November 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996974349.

Shubov:2023:ADE

[SK23] Marianna A. Shubov and Laszlo P. Kindrat. Asymptotic distribution of the eigenvalues of the bending-torsion vibration model with fully nondissipative boundary feedback. Studies in Applied Mathematics, 150(4):996–1025, May 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12562.

Sachdev:2008:EAS

[SKB08] P. L. Sachdev, Ramesh B. Kudenatti, and N. M. Bujurke. Exact analytic solution of a boundary value problem for the Falkner–Skan equation. Studies in Applied Mathematics, 120(1):1–16, January 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2007.00386.x.

Salgueiro:2005:SVS

[SKP^{+}05] José R. Salgueiro, Yuri S. Kivshar, Dmitry E. Pelinovsky, Verónica Simón, and Humberto Michinel. Spatial vector solitons in nonlinear photonic crystal fibers. Studies in Applied Mathematics, 115(2):157–171, August 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2005.00324.x.

Sulaiman:2016:LTS

[SLN16] F. B. Ali Sulaiman, J. A. Leach, and D. J. Needham. The large-time solution of Burgers’ equation with time- dependent coefficients. II. Algebraic coefficients. Studies in Applied Mathematics, 137(3):273–305, October 2016. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12130.

Slunyaev:2019:OFS

[Slu19] Alexey Slunyaev. On the optimal focusing of solitons and breathers in long-wave models. Studies in Applied Mathematics, 142(3):385–413, April 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12261.

Snowball:2020:SSF

[SO20] Ben Snowball and Sheehan Olver. Sparse spectral and p-finite element methods for partial differential equations on disk slices and trapeziums. Studies in Applied Mathematics, 145(1):3–35, July 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12303.

Sobolevsky:2004:MSC

[Sob04] S. Sobolevsky. Movable singularities of a class of nonlinear ordinary differential equations of arbitrary order. Studies in Applied Mathematics, 112(3):227–234, April 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2004.01472.x.

Sobolevsky:2005:BTO

[Sob05] S. Sobolevsky. Binomial-type ordinary differential equations of the third order. Studies in Applied Mathematics, 114(1):1–15, January 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.0022-
2526.2005.01533.x.

Sobolevsky:2006:PCB

[Sob06] S. Sobolevsky. Painlevé classification of binomial type ordinary differential equations of the arbitrary order. Studies in Applied Mathematics, 117(3):215–237, October 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2006.00353.x.

Sachdev:1988:ESW

[SP88a] P. L. Sachdev and Varughese Philip. Exact simple waves on shear flows in a compressible barotropic medium. Studies in Applied Mathematics, 79(3):193–203, December 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1988793193.

Srivastava:1988:CGF

[SP88b] H. M. Srivastava and Blagoj S. Popov. A class of generating functions for the Jacobi and related polynomials. Studies in Applied Mathematics, 78(3):177–182, June 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1988783177.

Spencer:1974:GRC

[Spe74] J. H. Spencer. A generalized Rota conjecture for partitions. Studies in Applied Mathematics, 53(3):239–241, September 1974. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1974533239.

Sachdev:1997:SSTa

[SR97a] P. L. Sachdev and Sharadha Ramanan. Singularity structure of third-order dynamical systems. I. Studies in Applied Mathematics, 98(3):255–275, April 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00049.

Sachdev:1997:SSTb

[SR97b] P. L. Sachdev and Sharadha Ramanan. Singularity structure of third-order dynamical systems. II. Studies in Applied Mathematics, 98(3):277–310, April 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00050.

Schief:1998:LTA

[SR98] W. K. Schief and C. Rogers. Loewner transformations: Adjoint and binary Darboux connections. Studies in Applied Mathematics, 100(4):391–422, May 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00082.

Shefter:1999:QSW

[SR99] Michael Shefter and Rodolfo R. Rosales. Quasiperiodic solutions in weakly nonlinear gas dynamics. Part I. Numerical results in the inviscid case. Studies in Applied Mathematics, 103(4):279–337, November 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
1034137.

Schief:2012:EES

[SR12] W. K. Schief and C. Rogers. The Euler equations of spatial gasdynamics and the integrable Heisenberg spin equation. Studies in Applied Mathematics, 128(4):407–419, May 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00539.x.

Singh:2004:QTS

[SRD04] B. M. Singh, J. Rokne, and R. S. Dhaliwal. Quadruple trigonometrical series equations and their application to an inclusion problem in the theory of elasticity. Studies in Applied Mathematics, 112(1):17–37, January 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2004.01468.x.

Sachdev:2005:LTA

[SRE05] P. L. Sachdev, Ch. Srinivasa Rao, and B. O. Enflo. Largetime asymptotics for periodic solutions of the modified Burgers equation. Studies in Applied Mathematics, 114(3):307–323, April 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.0022-2526.2005.01551.x.

Srinivasan:1988:VPA

[Sri88] R. Srinivasan. A variational principle for the Ackerberg– O’Malley resonance problem. Studies in Applied Mathematics, 79(3):271–289, December 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1988793271.

Srinivasan:1991:ASW

[Sri91] Radhakrishnan Srinivasan. Asymptotic solution of the weakly nonlinear Schrödinger equation with variable coefficients. Studies in Applied Mathematics, 84(2):145–165, February 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991842145.

Srivastava:1992:SAE

[Sri92] H. M. Srivastava. A simple algorithm for the evaluation of a class of generalized hypergeometric series. Studies in Applied Mathematics, 86(1):79–86, January 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199286179.

Sachdev:1999:ANS

[SRJ99] P. L. Sachdev, Ch. Srinivasa Rao, and K. T. Joseph. Analytic and numerical study of N-waves governed by the nonplanar Burgers equation. Studies in Applied Mathematics, 103(2):89–120, August 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00122.

Saffman:1977:FWA

[SS77] P. G. Saffman and J. S. Sheffield. Flow over a wing with an attached free vortex. Studies in Applied Mathematics, 57(2):107–117, October 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1977572107.

Saffman:1981:SLA

[SS81] P. G. Saffman and R. Szeto. Structure of a linear array of uniform vortices. Studies in Applied Mathematics, 65(3):223–248, December 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1981653223.

Simmen:1985:SDW

[SS85] J. A. Simmen and P. G. Saffman. Steady deep-water waves on a linear shear current. Studies in Applied Mathematics, 73(1):35– 57, August 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm198573135.

Sachdev:1995:SSC

[SS95] P. L. Sachdev and R. Sarathy. Singularity structure and chaotic behavior of the homopolar disk dynamo. Studies in Applied Mathematics, 95(4):345–380, November 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995954345.

Srinivasan:2003:NJC

[SS03a] G. K. Srinivasan and V. D. Sharma. A note on the jump conditions for systems of conservation laws. Studies in Applied Mathematics, 110(4):391–396, May 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00244.

Srinivasan:2003:MEW

[SS03b] Gopala Krishna Srinivasan and V. D. Sharma. Modulation equations of weakly nonlinear geometrical optics in media exhibiting mixed nonlinearity. Studies in Applied Mathematics, 110(2):103–122, February 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00232.

Srinivasan:2004:EDA

[SS04] Gopala Krishna Srinivasan and V. D. Sharma. Energy dissipated across shocks in weak solutions of conservation laws. Studies in Applied Mathematics, 112(3):281–291, April 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2004.01501.x.

Sengupta:2006:SSM

[SS06] T. K. Sengupta and K. Venkata Subbaiah. Spatial stability for mixed convection boundary layer over a heated horizontal plate. Studies in Applied Mathematics, 117(3):265–298, October 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2006.00355.x.

Sekhar:2008:ISW

[SS08] T. Raja Sekhar and V. D. Sharma. Interaction of shallow water waves. Studies in Applied Mathematics, 121(1):1–25, July 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2008.00402.x.

Strachan:2014:NAC

[SS14] Ian A. B. Strachan and Błażej M. Szablikowski. Novikov algebras and a classification of multicomponent Camassa–Holm equations. Studies in Applied Mathematics, 133(1):84–117, July 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12040.

Shukla:2017:WNW

[SS17] Triveni P. Shukla and V. D. Sharma. Weakly nonlinear waves in nonideal fluids. Studies in Applied Mathematics, 138(3):247– 268, April 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12151.

Stewartson:1982:MS

[SSK82] Keith Stewartson, F. T. Smith, and Kalle Kaups. Marginal separation. Studies in Applied Mathematics, 67(1):45–61, August 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm198267145.

Sepulveda-Salcedo:2020:OCD

[SSVS20] Lilian Sofia Sepulveda-Salcedo, Olga Vasilieva, and Mikhail Svinin. Optimal control of dengue epidemic outbreaks under limited resources. Studies in Applied Mathematics, 144 (2):185–212, February 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12295.

Song:2022:SDD

[SSW22] Yongli Song, Junping Shi, and Hao Wang. Spatiotemporal dynamics of a diffusive consumer-resource model with explicit spatial memory. Studies in Applied Mathematics, 148(1):373–395, January 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12443.

Saffman:1984:VIL

[ST84] P. G. Saffman and S. Tanveer. Vortex induced lift on two dimensional low speed wings. Studies in Applied Mathematics, 71(1):65–78, August 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm198471165.

Song:2017:SSS

[ST17] Yongli Song and Xiaosong Tang. Stability, steady-state bifurcations, and Turing patterns in a predator–prey model with herd behavior and prey-taxis. Studies in Applied Mathematics, 139(3):371–404, October 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12165.

Stanley:1971:TAPa

[Sta71a] Richard P. Stanley. Theory and application of plane partitions:Part 1. Studies in Applied Mathematics, 50(2):167–188, June 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1971502167.

Stanley:1971:TAPb

[Sta71b] Richard P. Stanley. Theory and application of plane Partitions.Part 2. Studies in Applied Mathematics, 50(3):259–279, September 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1971503259.

Stanley:1975:UBC

[Sta75] Richard P. Stanley. The upper bound conjecture and Cohen-Macaulay rings. Studies in Applied Mathematics, 54(2):135– 142, June 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1975542135.

Stanshine:1976:ASFb

[Sta76] J. A. Stanshine. Asymptotic solutions of the field–Noyes model for the Belousov reaction — II. Plane waves. Studies in Applied Mathematics, 55(4):327–349, December 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1976554327.

Stanley:1978:ES

[Sta78] Richard P. Stanley. Exponential structures. Studies in Applied Mathematics, 59(1):73–82, July 1978. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197859173.

Stanley:1985:ULS

[Sta85] Richard P. Stanley. Unimodality and Lie superalgebras. Studies in Applied Mathematics, 72(3):263–281, June 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1985723263.

Stam:1987:PBT

[Sta87] A. J. Stam. Polynomials of binomial type and renewal sequences. Studies in Applied Mathematics, 77(3):183–193, December 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1987773183.

Stanislavova:2015:LSS

[Sta15] Milena Stanislavova. Linear stability of solitary waves for the one-dimensional Benney–Luke and Klein–Gordon equations. Studies in Applied Mathematics, 134(1):1–23, January 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12062.

Scarsoglio:2009:EAT

[STC09] S. Scarsoglio, D. Tordella, and W. O. Criminale. An exploratory analysis of the transient and long-term behavior of small three-dimensional perturbations in the circular cylinder wake. Studies in Applied Mathematics, 123(2):153–173, August 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00449.x.

Stepanyants:2019:EIB

[Ste19] Yury Stepanyants. The effects of interplay between the rotation and shoaling for a solitary wave on variable topography. Studies in Applied Mathematics, 142(4):465–486, May 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12255.

Strang:1986:PTM

[Str86] Gilbert Strang. A proposal for Toeplitz matrix calculations. Studies in Applied Mathematics, 74(2):171–176, April 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1986742171.

Strela:1997:MRO

[Str97] Vasily Strela. Multiwavelets: Regularity, orthogonality, and symmetry via two–scale similarity transform. Studies in Applied Mathematics, 98(4):335–354, May 1997. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00052.

Straughan:2011:CDH

[Str11] B. Straughan. Continuous dependence on the heat source in resonant porous penetrative convection. Studies in Applied Mathematics, 127(3):302–314, October 2011. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00521.x.

Stuart:1992:PSA

[Stu92] David M. Stuart. Particles in a singular asymptotic limit of a classical field theory. Studies in Applied Mathematics, 87(3):239–281, October 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1992873239.

Sun:2003:NCR

[STWW03] Wentao Sun, Tao Tang, Michael J. Ward, and Juncheng Wei. Numerical challenges for resolving spike dynamics for two one-dimensional reaction–diffusion systems. Studies in Applied Mathematics, 111(1):41–84, July 2003. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.t01-1-00227.

Sun:2014:SIT

[STZZ14] Wen-Rong Sun, Bo Tian, Hui Zhong, and Hui-Ling Zhen. Soliton interactions for the three-coupled discrete nonlinear Schrödinger equations in the alpha helical proteins. Studies in Applied Mathematics, 132(1):65–80, January 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12021.

Sultanov:2020:BAM

[Sul20] Oskar A. Sultanov. Bifurcations of autoresonant modes in oscillating systems with combined excitation. Studies in Applied Mathematics, 144(2):213–241, February 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12294.

Sun:1993:SAF

[Sun93] Xie-Hua Sun. On the simultaneous approximation of functions and their derivatives by the Szász–Mirakyan operator (II). Studies in Applied Mathematics, 89(3):189–194, August 1993. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1993893189.

Seymour:1987:ERA

[SV87] Brian Seymour and Eric Varley. Exact representations for acoustical waves when the sound speed varies in space and time. Studies in Applied Mathematics, 76(1):1–35, February 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19877611.

Sachdev:1992:ESL

[SV92] P. L. Sachdev and B. Mayil Vaganan. Exact solutions of linear partial differential equations with variable coefficients. Studies in Applied Mathematics, 87(3):213–237, October 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1992873213.

Sachdev:1994:EFS

[SV94] P. L. Sachdev and B. Mayil Vaganan. Exact free surface flows for shallow water equations i: The incompressible case. Studies in Applied Mathematics, 93(3):251–274, November 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1994933251.

Sachdev:1995:EFS

[SV95] P. L. Sachdev and B. Mayil Vaganan. Exact free surface flows for shallow water equations II: The compressible case. Studies in Applied Mathematics, 94(1):57–76, January 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199594157.

Spigler:1999:CDE

[SV99] Renato Spigler and Marco Vianello. On the complex differential equation Y′′ + G(z)Y = 0 in Banach algebras. Studies in Applied Mathematics, 102(3):291–308, April 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00112.

Sachdev:2008:SLT

[SVS08] P. L. Sachdev, B. Mayil Vaganan, and G. Sivagami. Symmetries and large time asymptotics of compressible Euler flows with damping. Studies in Applied Mathematics, 120 (2):105–128, February 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2007.00397.x.

Seaman:1974:LBT

[SW74]  W. J. Seaman and F. Y. M. Wan.  Lateral bend-

ing and twisting of thin-walled curved tubes. Studies in Applied Mathematics, 53(1):73–89, March 1974. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197453173.

Shearer:1981:BSB

[SW81] M. Shearer and I. C. Walton. On bifurcation and symmetry in Benard convection and Taylor vortices. Studies in Applied Mathematics, 65(1):85–93, August 1981. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198165185.

Sun:2000:DCI

[SW00] Xiaodi Sun and Michael J. Ward. Dynamics and coarsening of interfaces for the viscous Cahn — Hilliard equation in one spatial dimension. Studies in Applied Mathematics, 105(3):203–234, October 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00149.

Sadovsky:2007:EEE

[SW07] Alexander Sadovsky and Frederic Y. M. Wan. The elastodynamics of embryonic epidermal wound closure. Studies in Applied Mathematics, 118(4):365–395, May 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1365-2966.2007.00370.x.

Song:2017:APJ

[SW17] Z. Song and R. Wong. Asymptotics of pseudo-Jacobi polynomials with varying parameters. Studies in Applied Mathematics, 139(1):179–217, July 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12177.

Swaters:2004:SPM

[Swa04] Gordon E. Swaters. Spectral properties in modon stability theory. Studies in Applied Mathematics, 112(3):235–258, April 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2004.01506.x.

Swaters:2007:PSS

[Swa07] Gordon E. Swaters. Perturbations of soliton solutions to the unstable nonlinear Schrödinger and sine-Gordon equations. Studies in Applied Mathematics, 118(2):99–116, February 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2007.00365.x.

Swaters:2012:PTS

[Swa12] Gordon E. Swaters. Perturbation theory for the solitary wave solutions to a Sasa-Satsuma model describing nonlinear internal waves in a continuously stratified fluid. Studies in Applied Mathematics, 128(4):420–432, May 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00542.x.

Swaters:2018:HSV

[Swa18] Gordon E. Swaters. Hamiltonian structure and a variational principle for grounded abyssal flow on a sloping bottom in a mid-latitude β-plane. Studies in Applied Mathematics, 141(2):247–263, August 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12224.

Sheng:2022:RPI

[SXZ22] Wancheng Sheng, Tao Xiao, and Qinglong Zhang. Riemann problem for isentropic magnetogasdynamics in a variable crosssection duct. Studies in Applied Mathematics, 149(1):266–292, July 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12496.

Shchesnovich:2003:HOS

[SY03] Valery S. Shchesnovich and Jianke Yang. Higher-order solitons in the N–wave system. Studies in Applied Mathematics, 110(4):297–332, May 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00240.

Spindler:2006:NSE

[SY06] Richard Spindler and Jun Yu. Nonlinear spatial evolution of small disturbances from boundary conditions in flat inclined-channel flow. Studies in Applied Mathematics, 116 (2):173–200, February 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2006.00334.x.

Stepanyants:2015:BIC

[SY15] Y. A. Stepanyants and E. I. Yakubovich. The Bernoulli integral for a certain class of non-stationary viscous vortical flows of incompressible fluid. Studies in Applied Mathematics, 135(3):295–309, October 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12087.

Sheng:2022:ISM

[SYF22] Han-Han Sheng, Guo-Fu Yu, and Bao-Feng Feng. An integrable semidiscretization of the modified Camassa– Holm equation with linear dispersion term. Studies in Applied Mathematics, 149(1):230–265, July 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12497.

Shinglot:2023:ESP

[SZ23] Vrushaly Shinglot and John Zweck. The essential spectrum of periodically stationary pulses in lumped models of short-pulse fiber lasers. Studies in Applied Mathematics, 150(1):218–253, January 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12538.

Shen:2016:SSP

[SZWM16] Yannan Shen, John Zweck, Shaokang Wang, and Curtis R. Menyuk. Spectra of short pulse solutions of the cubic– quintic complex Ginzburg–Landau equation near zero dispersion. Studies in Applied Mathematics, 137(2):238–255, August 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12136.

Tam:1994:CTC

[TA94] K. K. Tam and Andonowati. Computation of travelling combustion waves in a porous medium. Studies in Applied Mathematics, 91(2):179–187, February 1994. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1994912179.

Tabaei:2007:RLS

[TA07] Ali Tabaei and T. R. Akylas. Resonant long–short wave interactions in an unbounded rotating stratified fluid. Studies in Applied Mathematics, 119(3):271–296, October 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2007.00389.x.

Tam:1988:TES

[Tam88] K. K. Tam. The temporal evolution of a system in combustion theory. II. Studies in Applied Mathematics, 79(2):159–171, October 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1988792159.

Tam:1989:TWSa

[Tam89a] K. K. Tam. Traveling wave solutions for a combustion problem. Studies in Applied Mathematics, 81(2):117–124, October 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1989812117.

Tam:1989:TWSb

[Tam89b] K. K. Tam. Traveling wave solutions for combustion in a porous medium. Studies in Applied Mathematics, 81 (3):249–263, December 1989. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1989813249.

Tam:1993:EPD

[Tam93] K. K. Tam. On the existence and parameter dependence of similarity solutions for free convection on a vertical plate. Studies in Applied Mathematics, 88(2):141–152, February 1993. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1993882141.

Tanveer:1985:VIL

[Tan85] S. Tanveer. Vortex induced lift on a flat plate with a curved forward-facing flap. Studies in Applied Mathematics, 72(2):173–187, April 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1985722173.

Tanveer:1986:STP

[Tan86] S. Tanveer. A steadily translating pair of equal and opposite vortices with vortex sheets on their boundaries. Studies in Applied Mathematics, 74(2):139–154, April 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1986742139.

Tan:2002:ESW

[TB02] Benkui Tan and John P. Boyd. Envelope solitary waves and periodic waves in the AB equations. Studies in Applied Mathematics, 109(2):67–87, August 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00401.

Triadis:2018:IDM

[TBKiM18] Dimetre Triadis, Philip Broadbridge, Kenji Kajiwara, and Ken ichi Maruno. Integrable discrete model for one-dimensional soil water infiltration. Studies in Applied Mathematics, 140(4):483– 507, May 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12208.

Tung:1982:LAI

[TCK82] Ka-Kit Tung, Tony F. Chan, and Toshi Kubota. Large amplitude internal waves of permanent form. Studies in Applied Mathematics, 66(1):1–44, February 1982. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19826611.

Tellez:1994:DHT

[Tél94] Manuel A. Aguirre Téllez. The distributional Hankel transform of marcel Riesz’s ultrahyperbolic kernel. Studies in Applied Mathematics, 93(2):133–162, October 1994. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1994932133.

Temme:1993:AES

[Tem93] N. M. Temme. Asymptotic estimates of Stirling numbers. Studies in Applied Mathematics, 89(3):233–243, August 1993. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1993893233.

Turkington:1991:CMS

[TEW91] Bruce Turkington, Alexander Eydeland, and Sheng Wang. A computational method for solitary internal waves in a continuously stratified fluid. Studies in Applied Mathematics, 85(2):93–127, August 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm199185293.

Turkyilmazoglu:2001:AAC

[TG01] M. Turkyilmazoglu and J. S. B. Gajjar. An analytic approach for calculating absolutely unstable inviscid modes of the boundary layer on a rotating disk. Studies in Applied Mathematics, 106(4):419–435, May 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00173.

Teshukov:2006:TDN

[TG06] V. M. Teshukov and S. L. Gavrilyuk. Three-dimensional nonlinear dispersive waves on shear flows. Studies in Applied Mathematics, 116(3):241–255, April 2006. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2006.00342.x.

Thompson:1969:EDE

[Tho69a] Colin J. Thompson. Eigenvalue degeneracy and existence and non-existence of a phase transition in one dimension. Studies in Applied Mathematics, 48(4):299–309, December 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1969484299.

Thompson:1969:RCE

[Tho69b] Colin J. Thompson. On the ratio of consecutive eigenvalues in N–dimensions. Studies in Applied Mathematics, 48 (3):281–283, September 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1969483281.

Thompson:1969:PTE

[TK69] Colin J. Thompson and Mark Kac. Phase transition and eigenvalue degeneracy of a one dimensional anharmonic oscillator. Studies in Applied Mathematics, 48(3):257–264, September 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1969483257.

Tung:1981:WNI

[TKC81] Ka-Kit Tung, Denny R. S. Ko, and Jerry J. Chang. Weakly nonlinear internal waves in shear. Studies in Applied Mathematics, 65(3):189–221, December 1981. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1981653189.

Trichtchenko:2019:SPT

[TMPVB19] Olga Trichtchenko, Paul Milewski, Emilian Părău, and Jean-Marc Vanden-Broeck. Stability of periodic traveling flexural-gravity waves in two dimensions. Studies in Applied Mathematics, 142(1):65–90, January 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12233.

Trulsen:1996:CPA

[TNM96] K. Trulsen, M. Naciri, and C. C. Mei. Comment on “A perturbation analysis of an interaction between long and short surface waves” by Woodruff and Messiter. Studies in Applied Mathematics, 96(1):141–142, January 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996961141. See [WM94].

Tisbury:2021:ETWb

[TNT21a] Alex D. O. Tisbury, David J. Needham, and Alexandra Tzella. The evolution of traveling waves in a KPP reaction–diffusion model with cut-off reaction rate. II. Evolution of traveling waves. Studies in Applied Mathematics, 146(2):330–370, February 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12352.

Tisbury:2021:ETWa

[TNT21b] Alex D. O. Tisbury, David J. Needham, and Alexandra Tzella. The evolution of traveling waves in a KPP reaction-diffusion model with cut-off reaction rate. I. Permanent form traveling waves. Studies in Applied Mathematics, 146(2):301–329, February 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12337.

Tovbis:2000:BHC

[Tov00] Alexander Tovbis. Breaking homoclinic connections for a singularly perturbed differential equation and the Stokes phenomenon. Studies in Applied Mathematics, 104(4):353–386, May 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00138.

Turner:1997:SDL

[TR97] Cristina Vilma Turner and Rodolfo Rubén Rosales. The small dispersion limit for a nonlinear semidiscrete system of equations. Studies in Applied Mathematics, 99(3):205–254, October 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00060.

Turkyilmazoglu:2002:UVA

[TR02] M. Turkyilmazoglu and A. I. Ruban. A uniformly valid asymptotic approach to the inviscid–viscous interaction theory. Studies in Applied Mathematics, 108(2):161–185, February 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.01403.

Trione:1976:FTC

[Tri76] Susana Elena Trione. On the Fourier transform of causal distributions. Studies in Applied Mathematics, 55(4):315–326, December 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1976554315.

Trione:1988:MRU

[Tri88a] Susana Elena Trione. On marcel Riesz’s ultrahyperbolic kernel. Studies in Applied Mathematics, 79(3):185–191, December 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1988793185.

Trione:1988:ERU

[Tri88b] Susana Elena Trione. On the elementary retarded, ultrahyperbolic solution of the Klein–Gordon operator, iterated k times. Studies in Applied Mathematics, 79(2):127–141, October 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1988792127.

Trione:1991:LTR

[Tri91] Susana Elena Trione. On the Laplace transform of radial functions. Studies in Applied Mathematics, 85(4):303–315, November 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1991854303.

Trion:1992:GTB

[Tri92] S. E. Trion. On the generalizations of a theorem of Beppo Levi. Studies in Applied Mathematics, 87(3):195–211, October 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1992873195.

Trione:1997:GVW

[Tri97] Susana Elena Trione. On the generalization of vladimirov and Wightman generalized functions. Studies in Applied Mathematics, 98(3):207–211, April 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00047.

Trione:1999:ELT

[Tri99] Susana Elena Trione. The expression of the Laplace transform of (P ± i0)^{λ} as a K-transform. Studies in Applied Mathematics, 103(3):259–265, October 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00127.

Troitskaya:2017:MAI

[Tro17] Saule Troitskaya. Mathematical analysis of inertial waves in rectangular basins with one sloping boundary. Studies in Applied Mathematics, 139(3):434–456, October 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12169.

Trogdon:2021:SIS

[Tro21] Thomas Trogdon. Scattering and inverse scattering for the AKNS system: a rational function approach. Studies in Applied Mathematics, 147(4):1443–1480, November 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12434.

Turkyilmazoglu:2005:CMR

[TU05] M. Turkyilmazoglu and N. Uygun. Compressible modes of the rotating-disk boundary-layer flow leading to absolute instability. Studies in Applied Mathematics, 115(1):1–20, July 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2005.01549.

Tumin:2006:BES

[Tum06] Anatoli Tumin. Biorthogonal eigenfunction system in the triple-deck limit. Studies in Applied Mathematics, 117(2):165– 190, August 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9590.2006.00351.
x.

Turkyilmazoglu:2005:ESN

[Tur05a] M. Turkyilmazoglu. Effects of suction on the nonstationary lower branch modes of a compressible boundary layer flow due to a rotating disk. Studies in Applied Mathematics, 115(4):357–385, November 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2005.00330.x.

Turkyilmazoglu:2005:LBM

[Tur05b] M. Turkyilmazoglu. Lower branch modes of the compressible boundary layer flow due to a rotating-disk. Studies in Applied Mathematics, 114(1):17–43, January 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2005.01535.x.

Turkyilmazoglu:2007:IFA

[Tur07] M. Turkyilmazoglu. Influence of finite amplitude disturbances on the nonstationary modes of a compressible boundary layer flow. Studies in Applied Mathematics, 118(3):199–220, April 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2007.00366.x.

Tikhov:2022:NSE

[TV22] Stanislav V. Tikhov and Dmitry V. Valovik. Nonlinearizable solutions in an eigenvalue problem for Maxwell’s equations with nonhomogeneous nonlinear permittivity in a layer. Studies in Applied Mathematics, 149(3):565–587, October 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12512.

Titcombe:1997:CHT

[TW97] Michèle S. Titcombe and Michael J. Ward. Convective heat transfer past small cylindrical bodies. Studies in Applied Mathematics, 99(1):81–105, July 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00057.

Tsemekhman:2003:SSI

[TW03] V. Tsemekhman and J. S. Wettlaufer. Singularities, shocks, and instabilities in interface growth. Studies in Applied Mathematics, 110(3):221–256, April 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00238.

Tian:2017:SRC

[TW17] Kai Tian and Jing Ping Wang. Symbolic representation and classification of N = 1Supersymmetric evolutionary equations. Studies in Applied Mathematics, 138(4):467–498, May 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12163.

Titcombe:2000:HML

[TWK00] Michèle S. Titcombe, Michael J. Ward, and Mary Catherine Kropinski. A hybrid method for low Reynolds number flow past an asymmetric cylindrical body. Studies in Applied Mathematics, 105(2):165–190, August 2000. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00147.

Tian:2021:ETB

[TYZZ21] Shuai Tian, Ying Yang, Rui Zhou, and Shihui Zhu. Energy thresholds of blow-up for the Hartree equation with a focusing subcritical perturbation. Studies in Applied Mathematics, 146(3):658–676, April 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12362.

Tian:2014:IGV

[TZ14] Shou-Fu Tian and Hong-Qing Zhang. On the integrability of a generalized variable-coefficient forced Korteweg–de Vries equation in fluids. Studies in Applied Mathematics, 132(3):212–246, April 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12026.

Tang:2009:MBG

[TZHL09] X. Y. Tang, J. Zhao, F. Huang, and S. Y. Lou. Monopole blocking governed by a modified KdV type equation. Studies in Applied Mathematics, 122(3):295–304, April 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00434.x.

Upsal:2020:RLS

[UD20] Jeremy Upsal and Bernard Deconinck. Real Lax spectrum implies spectral stability. Studies in Applied Mathematics, 145(4):765–790, November 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12335.

Ungarish:1983:TPF

[UG83] M. Ungarish and H. P. Greenspan. On two-phase flow in a rotating boundary layer. Studies in Applied Mathematics, 69(2):145–175, October 1983. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1983692145.

See erratum [UG84].

Ungarish:1984:EMU

[UG84] M. Ungarish and H. P. Greenspan. Erratum: M. Ungarish and H. P. Greenspan, Two-phase flow in a rotating layer, Stud. Appl. Math. 69:145–174 (1983). Studies in Applied Mathematics, 70(1):87–90, February 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198470187. See [UG83].

Ulrich:1975:HAS

[Ulr75] Lionel Yves Ulrich. Heuristic algorithms for solving a large scale multicommodity flow problem on a network with a step function cost. Studies in Applied Mathematics, 54(3):207–227, September 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1975543207.

Urbach:1989:EOP

[Urb89] H. P. Urbach. Existence of optimum propulsion by means of periodic motions of a rigid profile. Studies in Applied Mathematics, 81(2):93–116, October 1989. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198981293.

Urbach:1990:OPMa

[Urb90a] H. P. Urbach. On optimum propulsion by means of small periodic motions of a rigid profile. I. Properties of optimum motions. Studies in Applied Mathematics, 82(2):121–180, February 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1990822121.

Urbach:1990:OPMb

[Urb90b] H. P. Urbach. On optimum propulsion by means of small periodic motions of a rigid profile II. Optimization of the period and numerical results. Studies in Applied Mathematics, 82(3):181–215, April 1990. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990823181.

Villarroel:2002:CPK

[VA02] J. Villarroel and M. J. Ablowitz. The Cauchy problem for the Kadomtsev–Petviashili [sic] II equation with nondecaying data along a line. Studies in Applied Mathematics, 109(3):151–162, October 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00220.

Vaganan:2003:DSA

[VA03] B. Mayil Vaganan and R. Asokan. Direct similarity analysis of generalized Burgers equations and perturbation solutions of Euler–Painlevé transcendents. Studies in Applied Mathematics, 111(4):435–451, November 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
t01-1-00041.

Vafeas:2020:LFD

[Vaf20] Panayiotis Vafeas. Low-frequency dipolar electromagnetic scattering by a solid ellipsoid in lossless environment. Studies in Applied Mathematics, 145(2):217–246, August 2020.

CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12318.

Vaganan:2012:CHT

[Vag12] B. Mayil Vaganan. Cole–Hopf transformations for higher dimensional Burgers equations with variable coefficients. Studies in Applied Mathematics, 129(3):300–308, October 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00551.x.

VanderJeugt:1987:UPA

[Van87] Joris Van der Jeugt. Unimodal polynomials arising from Lie superalgebras. Studies in Applied Mathematics, 77(1):47–59, August 1987. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm198777147.

Varlamov:1997:TON

[Var97] Vladimir V. Varlamov. The third-order nonlinear evolution equation governing wave propagation in relaxing media. Studies in Applied Mathematics, 99(1):25–48, July 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00055.

Varkonyi:2009:FBP

[Vár09] Péter L. Várkonyi. Floating body problems in two dimensions. Studies in Applied Mathematics, 122(2):195–218, February 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2008.00429.x.

Varkonyi:2013:NFO

[Vár13] P. L. Várkonyi. Neutrally floating objects of density 1/2 in three dimensions. Studies in Applied Mathematics, 130(3):295– 315, April 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9590.2012.00569.

Vassiliou:2011:MSM

[Vas11] Peter J. Vassiliou. Method for solving the multidimensional n-wave resonant equations and geometry of generalized Darboux-Manakov-Zakharov systems. Studies in Applied Mathematics, 126(3):203–243, April 2011. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00503.x.

Vermersch:1976:NTL

[VB76] John A. Vermersch, Jr. and Robert C. Beardsley. A note on the theory of low-frequency waves in a rotating stratified channel. Studies in Applied Mathematics, 55(4):281–292, December 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1976554281.

Varley:1983:LES

[VB83] E. Varley and P. A. Blythe. Long eddies in sheared flows. Studies in Applied Mathematics, 68(2):103–187, April 1983. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1983682103.

Vanden-Broeck:1997:MEW

[VBA97] J.-M. Vanden-Broeck and Y. Agnon. A model equation for water waves. Studies in Applied Mathematics, 98(1):1–18, January 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00038.

Venturini:1992:INT

[Ven92] G. Venturini. Iterates of number theoretic functions with periodic rational coefficients (generalization of the 3x + 1 problem). Studies in Applied Mathematics, 86(3):185–218, April 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1992863185.

Vakulenko:2015:RMC

[VGW15] S. Vakulenko, D. Grigoriev, and A. Weber. Reduction methods and chaos for quadratic systems of differential equations. Studies in Applied Mathematics, 135(3):225–247, October 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12083.

Villarroel:1990:IPW

[Vil90] Javier Villarroel. The inverse problem for Ward’s system. Studies in Applied Mathematics, 83(3):211–222, October 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1990833211.

Villarroel:1991:DPT

[Vil91] Javier Villarroel. The DBAR problem and the Thirring model. Studies in Applied Mathematics, 84(3):207–220, April 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991843207.

Villarroel:2004:SBE

[Vil04] Javier Villarroel. The stochastic Burger’s equation in Ito’s sense. Studies in Applied Mathematics, 112(1):87–100, January 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2004.01505.x.

Vaganan:2011:GBE

[VJ11] B. Mayil Vaganan and T. Jeyalakshmi. Generalized Burgers equations transformable to the Burgers equation. Studies in Applied Mathematics, 127(3):211–220, October 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00515.x.

Vaganan:2006:ELI

[VK06] B. Mayil Vaganan and M. Senthil Kumaran. Exact linearization and invariant solutions of the generalized Burgers equation with linear damping and variable viscosity. Studies in Applied Mathematics, 117(2):95–108, August 2006. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2006.00348.x.

vonKerczek:1972:SOS

[vKD72] C. von Kerczek and S. H. Davis. The stability of oscillatory Stokes layers. Studies in Applied Mathematics, 51(3):239–252, September 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1972513239.

Vladimirov:2008:VFH

[Vla08] V. A. Vladimirov. Viscous flows in a half space caused by tangential vibrations on its boundary. Studies in Applied Mathematics, 121(4):337–367, November 2008. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00418.x.

Vladimirov:2017:TTH

[Vla17] V. A. Vladimirov. Two-timing hypothesis, distinguished limits, drifts, and pseudo–diffusion for oscillating flows. Studies in Applied Mathematics, 138(3):269–293, April 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12152.

Villarroel:2011:IPD

[VM11] Javier Villarroel and Miquel Montero. On the integrability of the Poisson driven stochastic nonlinear Schrödinger equations. Studies in Applied Mathematics, 127(4):372–393, November 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00526.x.

Villarroel:2018:BWS

[VM18] Javier Villarroel and Miquel Montero. Breaking waves and spectral analysis of the two-dimensional KdV–Bogoyavlenskii equation. Studies in Applied Mathematics, 140(1):78–130, January 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12192.

Vasan:2021:ODM

[VMA21] V. Vasan, Manisha, and D. Auroux. Ocean-depth measurement using shallow-water wave models. Studies in Applied Mathematics, 147(4):1481–1518, November 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12418.

Vaganan:2010:LTA

[VP10] B. Mayil Vaganan and S. Padmasekaran. Large time asymptotic behaviors for periodic solutions of generalized Burgers equations with spherical symmetry or linear damping. Studies in Applied Mathematics, 124(1):1–18, January 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00456.x.

Villarroel:2009:DLS

[VPE09] Javier Villarroel, J. Prada, and P. G. Estévez. Dynamics of lump solutions in a 2 + 1 NLS equation. Studies in Applied Mathematics, 122(4):395–410, May 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00440.x.

Varley:1985:ESL

[VS85] Eric Varley and Brian Seymour. Exact solutions for large amplitude waves in dispersive and dissipative systems. Studies in Applied Mathematics, 72(3):241–262, June 1985. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1985723241.

Varley:1988:MOE

[VS88a] Eric Varley and Brian Seymour. A method for obtaining exact solutions to partial differential equations with variable coefficients. Studies in Applied Mathematics, 78(3):183–225, June 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1988783183.

Verde-Star:1988:ICF

[VS88b] Luis Verde-Star. Interpolation and combinatorial functions. Studies in Applied Mathematics, 79(1):65–92, August 1988. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198879165.

Verde-Star:1991:DDC

[VS91] Luis Verde-Star. Divided differences and combinatorial identities. Studies in Applied Mathematics, 85(3):215–242, October 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991853215.

Verde-Star:1993:PSI

[VS93] Luis Verde-Star. Polynomial sequences of interpolatory type. Studies in Applied Mathematics, 88(3):153–172, April 1993. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1993883153.

Verde-Star:1994:OIS

[VS94] Luis Verde-Star. Operator identities and the solution of linear matrix difference and differential equations. Studies in Applied Mathematics, 91(2):153–177, February 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1994912153.

Verde-Star:1995:BPB

[VS95a] Luis Verde-Star. Biorthogonal polynomial bases and Vandermondelike matrices. Studies in Applied Mathematics, 95(3):269–295, October 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1995953269.

Verde-Star:1995:DDL

[VS95b] Luis Verde-Star. Divided differences and linearly recursive sequences. Studies in Applied Mathematics, 95(4):433–456, November 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1995954433.

Varley:1996:SNS

[VS96] E. Varley and B. R. Seymour. Solutions to the Navier–Stokes equations describing local flows of stratified and compressible fluids. Studies in Applied Mathematics, 97(4):321–347, November 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1996974321.

Voronovich:1998:TDM

[VSS98] V. V. Voronovich, V. I. Shrira, and Yu. A. Stepanyants. Two-dimensional models for nonlinear vorticity waves in shear flows. Studies in Applied Mathematics, 100(1):1–32, January 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00067.

VanAssche:2023:MOP

[VW23] Walter Van Assche and Thomas Wolfs. Multiple orthogonal polynomials associated with the exponential integral. Studies in Applied Mathematics, 151(2):411–449, August 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12608.

Wan:1983:OFH

[WA83]  Frederic Y. M. Wan and Keith Anderson.  Optimal

forest harvesting with ordered site access. Studies in Applied Mathematics, 68(3):189–226, June 1983. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1983683189.

Wadati:1978:ICL

[Wad78] Miki Wadati. Invariances and conservation laws of the Korteweg–de Vries equation. Studies in Applied Mathematics, 59(2):153–186, October 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1978592153.

Wadey:1991:NDG

[Wad91] P. D. Wadey. On the nonlinear development of Görtler vortices in a compressible boundary layer. Studies in Applied Mathematics, 85(4):317–341, November 1991. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1991854317.

Wagner:1994:SDB

[Wag94] Carl G. Wagner. Surjections, differences, and binomial lattices. Studies in Applied Mathematics, 93(1):15–27, September 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199493115.

Walton:1982:ORB

[Wal82] I. C. Walton. On the onset of Rayleigh–Bénard convection in a fluid layer of slowly increasing depth. Studies in Applied Mathematics, 67(3):199–216, December 1982. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1982673199.

Waleffe:1995:HST

[Wal95] Fabian Waleffe. Hydrodynamic stability and turbulence: Beyond transients to a self-sustaining process. Studies in Applied Mathematics, 95(3):319–343, October 1995. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995953319.

Walton:2001:ENR

[Wal01] Andrew G. Walton. Existence of neutral Rayleigh waves in the Hagen–Poiseuille flow through a pipe of circular cross section. Studies in Applied Mathematics, 106(3):315–335, April 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00168.

Wan:1969:SFN

[Wan69a] F. Y. M. Wan. Stress functions for nonlinear shallow Shell theory. Studies in Applied Mathematics, 48(2):177–179, June 1969. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1969482177.

Wan:1969:ERE

[Wan69b] Frederic Y. M. Wan. Exact reductions of the equations of linear theory of shells of revolution. Studies in Applied Mathematics, 48(4):361–375, December 1969. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1969484361. See errata [Wan70a].

Wan:1970:EER

[Wan70a] Frederic Y. M. Wan. Errata: Exact reductions of the equations of linear theory of shells of revolution. Studies in Applied Mathematics, 49(4):405, December 1970. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1970494405. See [Wan69b].

Wan:1970:ELT

[Wan70b] Frederic Y. M. Wan. On the equations of the linear theory of elastic conical shells. Studies in Applied Mathematics, 49(1):69–83, March 1970. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm197049169.

Wan:1970:RSS

[Wan70c]  Frederic Y. M. Wan.  Rotationally symmetric shear-

ing and bending of helicoidal shells. Studies in Applied Mathematics, 49(4):351–369, December 1970. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1970494351.

Wan:1972:LPD

[Wan72] Frederic Y. M. Wan. Linear partial differential equations with random forcing. Studies in Applied Mathematics, 51(2):163– 178, June 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1972512163.

Wan:1973:DML

[Wan73a] Frederic Y. M. Wan. A direct method for linear dynamical problems in continuum mechanics with random loads. Studies in Applied Mathematics, 52(3):259–276, September 1973. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1973523259.

Wan:1973:CFD

[Wan73b] Frederic Y. M. Wan. An in-core finite difference method for separable boundary value problems on a rectangle. Studies in Applied Mathematics, 52(2):103–113, June 1973. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1973522103.

Wan:1975:EPS

[Wan75] Frederic Y. M. Wan. An eigenvalue problem for a semiinfinite pretwisted strip. Studies in Applied Mathematics, 54(4):351–358, December 1975. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1975544351.

Wang:1976:CME

[Wan76a] D. L. Wang. Construction of a maximally edge-connected graph with prescribed degrees. Studies in Applied Mathematics, 55(1):87–92, March 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm197655187.

Wang:1976:NUI

[Wan76b] Da-Lun Wang. A note on uniquely intersectable graphs. Studies in Applied Mathematics, 55(4):361–363, December 1976. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1976554361.

Wan:1977:PAS

[Wan77] Frederic Y. M. Wan. Perturbation and asymptotic solutions for problems in the theory of urban land rent. Studies in Applied Mathematics, 56(3):219–239, June 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1977563219.

Wan:1980:CSC

[Wan80] Frederic Y. M. Wan. Constant sustainable consumption rate in optimal growth with exhaustible resources. Studies in Applied Mathematics, 63(1):47–66, August 1980. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198063147.

Wan:1985:OSA

[Wan85] Frederic Y. M. Wan. Ordered site access and optimal forest rotation. Studies in Applied Mathematics, 73(2):155–175, October 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1985732155.

Wan:1993:TPB

[Wan93] Frederic Y. M. Wan. On torsion of prismatic bodies. Studies in Applied Mathematics, 89(1):63–94, May 1993. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199389163.

Wang:2012:RON

[Wan12] Jing Ping Wang. Recursion operator of the narita–Itoh– Bogoyavlensky lattice. Studies in Applied Mathematics, 129(3):309–327, October 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2012.00556.x.

Wan:2014:CSB

[Wan14] Frederic Y. M. Wan. Cell-surface bound nonreceptors and signaling morphogen gradients. Studies in Applied Mathematics, 133(2):151–181, August 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12030.

Wan:2019:NBM

[Wan19] Frederic Y. M. Wan. The necessity and benefits of multiple feedback for robust biological development. Studies in Applied Mathematics, 143(1):3–41, July 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12258.

Wan:2021:SVM

[Wan21] Frederic Y. M. Wan. Spatially varying multifeedback for robust signaling. Studies in Applied Mathematics, 147(3):1058–1088, October 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12424.

Warren:1978:HFE

[War78] F. W. Warren. Hermitian forms and eigenvalue bounds. Studies in Applied Mathematics, 59(3):249–281, December 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1978593249.

Warn:1983:EFA

[War83] T. Warn. The evolution of finite amplitude solitary Rossby waves on a weak shear. Studies in Applied Mathematics, 69(2):127–133, October 1983. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1983692127.

Ward:1992:EIS

[War92] Michael J. Ward. Eliminating indeterminacy in singularly perturbed boundary value problems with translation invariant potentials. Studies in Applied Mathematics, 87(2):95–134, August 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm199287295.

Ward:1994:MPL

[War94] Michael J. Ward. Metastable patterns, layer collapses, and coarsening for a one-dimensional Ginzburg–Landau equation. Studies in Applied Mathematics, 91(1):51–93, January 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm199491151.

Ward:1996:AAL

[War96] Michael J. Ward. An asymptotic analysis of localized solutions for some reaction–diffusion models in multidimensional domains. Studies in Applied Mathematics, 97(2):103–126, August 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1996972103.

Waterman:1979:CRH

[Wat79] Michael S. Waterman. Combinatorics of RNA hairpins and cloverleaves. Studies in Applied Mathematics, 60(2):91–96, April 1979. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm197960291.

Waxman:1980:MJO

[Wax80] Allen M. Waxman. On Maclaurin and Jacobi objects embedded in halos: The shift in the point of bifurcation. Studies in Applied Mathematics, 62(3):263–271, June 1980. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1980623263.

Waxman:1984:DCS

[Wax84] Allen M. Waxman. Dynamics of a couple-stress fluid membrane. Studies in Applied Mathematics, 70(1):63–86, February 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198470163.

Witelski:1998:SSA

[WB98] Thomas P. Witelski and Andrew J. Bernoff. Self-similar asymptotics for linear and nonlinear diffusion equations. Studies in Applied Mathematics, 100(2):153–193, February 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00074.

Wu:2021:SPF

[WBZ21] Fan Wu, Xingzhi Bian, and Lixian Zhao. Solutions to a phase-field model for martensitic phase transformations driven by configurational forces. Studies in Applied Mathematics, 146(3):730–752, April 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12365.

Witelski:1995:FRS

[WC95] Thomas P. Witelski and Donald S. Cohen. Forbidden regions for shock formation in diffusive systems. Studies in Applied Mathematics, 95(3):297–317, October 1995.

CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995953297.

Wang:2022:SDS

[WCTZ22] Ning Wang, Wei Chen, Zhidong Teng, and Long Zhang. Spatial dynamics for an SIRE epidemic model with diffusion and prevention in contaminated environments. Studies in Applied Mathematics, 149(2):487–522, August 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12499.

Wu:2019:CIP

[WCYW19] Bin Wu, Qun Chen, Jun Yu, and Zewen Wang. A coefficient identification problem for a mathematical model related to ductal carcinoma in situ. Studies in Applied Mathematics, 143(4):356–372, November 2019. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12281.

Wan:2017:OPD

[WE17] Frederic Y. M. Wan and Germán A. Enciso. Optimal proliferation and differentiation of Chlamydia Trachomatis. Studies in Applied Mathematics, 139(1):129–178, July 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12175.

Wegner:2003:FBE

[Weg03] Franz Wegner. Floating bodies of equilibrium. Studies in Applied Mathematics, 111(2):167–183, August 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.t01-1-00231.

Weinitschke:1984:IEA

[Wei84] H. J. Weinitschke. An integrodifferential equation arising in thermo-viscoelasticity. Studies in Applied Mathematics, 71(1):79–94, August 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm198471179.

Weinitschke:1985:VIL

[Wei85] H. J. Weinitschke. Vortex induced lift on two dimensional low speed wings. Studies in Applied Mathematics, 73(2):177–182, October 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1985732177.

West:1978:PAH

[Wes78] Douglas B. West. Pairs of adjacent Hamiltonian circuits with small intersection. Studies in Applied Mathematics, 59(3):245–248, December 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1978593245.

Witte:2012:FDE

[WF12] N. S. Witte and P. J. Forrester. Fredholm determinant evaluations of the Ising model diagonal correlations and their λ generalization. Studies in Applied Mathematics, 128(2):183–223, February 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2011.00534.x.

Witte:2017:SVP

[WF17] N. S. Witte and P. J. Forrester. Singular values of products of Ginibre random matrices. Studies in Applied Mathematics, 138(2):135–184, February 2017. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12147.

Wang:2017:EEA

[WHS17] Liang Wang, Berthold K. P. Horn, and Gilbert Strang. Eigenvalue and eigenvector analysis of stability for a line of traffic. Studies in Applied Mathematics, 138(1):103–132, January 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12144.

Willke:1972:DTA

[Wil72] H. L. Willke, Jr. A distribution- theoretic approach to a stability problem of Orr. Studies in Applied Mathematics, 51(1):85– 98, March 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm197251185.

Witelski:1996:SIL

[Wit96] Thomas P. Witelski. The structure of internal layers for unstable nonlinear diffusion equations. Studies in Applied Mathematics, 97(3):277–300, October 1996. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1996973277.

Wycoff:1989:TES

[WK89] D. Wycoff and D. J. Kaup. Time evolution of the scattering data for the forced Toda lattice. Studies in Applied Mathematics, 81(1):7–19, August 1989. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19898117.

Ward:1991:NEP

[WK91] Michael J. Ward and Joseph B. Keller. Nonlinear eigenvalue problems under strong localized perturbations with applications to chemical reactors. Studies in Applied Mathematics, 85(1):1–28, July 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm19918511.

Wong:1992:AES

[WL92] R. Wong and H. Li. Asymptotic expansions for secondorder linear difference equations, II. Studies in Applied Mathematics, 87(4):289–324, November 1992. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1992874289.

Wang:2022:SPW

[WLR22] Xinjian Wang, Guo Lin, and Shigui Ruan. Spatial propagation in a within-host viral infection model. Studies in Applied Mathematics, 149(1):43–75, July 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12490.

Wang:2017:ASW

[WLS17] Xia Wang, Yijun Lou, and Xinyu Song. Age-structured within-host HIV dynamics with multiple target cells. Studies in Applied Mathematics, 138(1):43–76, January 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12135.

Woodruff:1994:PAI

[WM94] S. L. Woodruff and A. F. Messiter. A perturbation analysis of an interaction between long and short surface waves. Studies in Applied Mathematics, 92(2):159–189, June 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1994922159. See comment [TNM96].

Weidman:2012:LWN

[WNH12] P. D. Weidman, M. Nitsche, and L. Howard. Linear waves and nonlinear wave interactions in a bounded three-layer fluid system. Studies in Applied Mathematics, 128(4):385–406, May 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2011.00540.x.

Wood:1984:SFS

[Woo84] David Wood. The self-focusing singularity in the nonlinear Schrödinger equation. Studies in Applied Mathematics, 71(2):103–115, October 1984. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1984712103.

Woodruff:1995:UVA

[Woo95] S. L. Woodruff. A uniformly-valid asymptotic solution to a matrix system of ordinary differential equations and a proof of its validity. Studies in Applied Mathematics, 94(4):393–413, May 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1995944393.

Worster:1986:ALP

[Wor86] M. Grae Worster. The axisymmetric laminar plume: Asymptotic solution for large Prandtl number. Studies in Applied Mathematics, 75(2):139–152, October 1986. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1986752139.

Wang:2022:EUS

[WPO22] Hongyong Wang, Chaohong Pan, and Chunhua Ou. Existence, uniqueness and stability of forced waves to the Lotka–Volterra competition system in a shifting environment. Studies in Applied Mathematics, 148(1):186–218, January 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12438.

Wang:2013:GAM

[WQW13] Jun Wang, Weiyuan Qiu, and R. Wong. Global asymptotics for Meixner–Pollaczek polynomials with a varying parameter. Studies in Applied Mathematics, 130(4):345–392, May 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00570.x.

Wright:1985:NWR

[Wri85] T. W. Wright. Nonlinear waves in a rod: Results for incompressible elastic materials. Studies in Applied Mathematics, 72(2):149–160, April 1985. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1985722149.

Wright:2006:HCU

[Wri06] O. C. Wright III. Homoclinic connections of unstable plane waves of the long-wave–short-wave equations. Studies in Applied Mathematics, 117(1):71–93, July 2006. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2006.00345_117_1.x.

Wright:2020:MAH

[Wri20] Otis C. Wright III. Maximal amplitudes of hyperelliptic solutions of the derivative nonlinear Schrödinger equation. Studies in Applied Mathematics, 144(3):327–356, April 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12299.

Ward:1999:MDS

[WS99] Michael J. Ward and Douglas Stafford. Metastable dynamics and spatially inhomogeneous equilibria in dumbbell-shaped domains. Studies in Applied Mathematics, 103(1):51–73, July 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00120.

Wan:2010:GIC

[WSK10] Frederic Y. M. Wan, Alexander V. Sadovsky, and Natalia L. Komarova. Genetic instability in cancer: an optimal control problem. Studies in Applied Mathematics, 125(1):1–38, July 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00476.x.

Wang:2010:SSG

[WTL^{+}10] Pan Wang, Bo Tian, Wen-Jun Liu, Min Li, and Kun Sun. Soliton solutions for a generalized inhomogeneous variable-coefficient Hirota equation with symbolic computation. Studies in Applied Mathematics, 125(2):213–222, August 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00486.x.

Wang:2013:BPA

[WTW13] Yu-Feng Wang, Bo Tian, and Ming Wang. Bell-polynomial approach and integrability for the coupled Gross–Pitaevskii equations in Bose–Einstein condensates. Studies in Applied Mathematics, 131(2):119–134, August 2013. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12003.

Wu:2021:IAC

[Wu21] Xinglong Wu. Isentropic approximation of the compressible Euler equations in Besov spaces. Studies in Applied Mathematics, 146(4):1050–1062, May 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12375.

Weidman:1992:ISW

[WV92] P. D. Weidman and M. G. Velarde. Internal solitary waves. Studies in Applied Mathematics, 86(2):167–184, February 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1992862167.

Wang:2014:QPM

[WVBM14] Z. Wang, J.-M. Vanden-Broeck, and H. Meng. A quasiplanar model for gravity-capillary interfacial waves in deep water. Studies in Applied Mathematics, 133(2):232–256, August 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12043.

Warn:1978:ENC

[WW78] T. Warn and H. Warn. The evolution of a nonlinear critical level. Studies in Applied Mathematics, 59(1):37–71, July 1978. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm197859137.

Ward:2002:ESA

[WW02] Michael J. Ward and Juncheng Wei. The existence and stability of asymmetric spike patterns for the Schnakenberg model. Studies in Applied Mathematics, 109(3):229–264, October 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00223.

Wei:2003:ASP

[WW03] Juncheng Wei and Matthias Winter. Asymmetric spotty patterns for the Gray–Scott model in R^{2}. Studies in Applied Mathematics, 110(1):63–102, January 2003. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00231.

Wang:2005:UAO

[WW05] Z. Wang and R. Wong. Uniform asymptotics for orthogonal polynomials with exponential weights — the Riemann–Hilbert approach. Studies in Applied Mathematics, 115(1):139–155, July 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2005.01567. See erratum

[Ano05b].

Wong:2021:SPD

[WW21] Tony Wong and Michael J. Ward. Spot patterns in the 2-D Schnakenberg model with localized heterogeneities. Studies in Applied Mathematics, 146(4):779–833, May 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12361.

Wang:2010:GAP

[WWZ10] X. Wang, W. Wang, and G. Zhang. Global analysis of predator–prey system with hawk and dove tactics. Studies in Applied Mathematics, 124(2):151–178, February 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00466.x.

Wang:2022:RSS

[WX22] Zhi-An Wang and Xin Xu. Radial spiky steady states of a flux-limited Keller–Segel model: Existence, asymptotics, and stability. Studies in Applied Mathematics, 148(3):1251–1273, April 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12474.

Wu:2018:GUE

[WXZ18] Xiao-Bo Wu, Shuai-Xia Xu, and Yu-Qiu Zhao. Gaussian unitary ensemble with boundary spectrum singularity and σ-form of the Painlevé II equation. Studies in Applied Mathematics, 140(2):221–251, February 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12197.

Wong:2002:BLP

[WY02] R. Wong and Heping Yang. On a boundary layer problem. Studies in Applied Mathematics, 108(4):369–398, May 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.01430.

Wong:2003:AOR

[WY03] R. Wong and Heping Yang. On the Ackerberg–O’Malley resonance. Studies in Applied Mathematics, 110(2):157–179, February 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9590.00235.

Wong:2007:ASF

[WZ07] R. Wong and H. Y. Zhang. Asymptotic solutions of a fourth order differential equation. Studies in Applied Mathematics, 118(2):133–152, February 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2007.00369.x.

Wong:2008:NSC

[WZ08] R. Wong and Y. Zhao. On the number of solutions to Carrier’s problem. Studies in Applied Mathematics, 120(3):213–245, April 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2008.00395.x.

Wang:2022:DAC

[WZ22] Xueru Wang and Junyi Zhu. Dbar-approach to coupled nonlocal NLS equation and general nonlocal reduction. Studies in Applied Mathematics, 148(1):433–456, January 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12445.

Xu:2022:SIS

[XB22] Xiao-Chuan Xu and Natalia Pavlovna Bondarenko. Stability of the inverse scattering problem for the self-adjoint matrix Schrödinger operator on the half line. Studies in Applied Mathematics, 149(3):815–838, October 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12522.

Xu:2016:IBV

[XF16] Jian Xu and Engui Fan. Initial-boundary value problem for integrable nonlinear evolution equation with 3 × 3 Lax pairs on the interval. Studies in Applied Mathematics, 136(3):321–354, April 2016. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12108.

Xiao:2023:DHP

[XF23a] Wei Xiao and Tian Feng. The degenerate hyperbolic problem for the two-dimensional nonlinear wave system. Studies in Applied Mathematics, 150(4):1304–1328, May 2023. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12570.

Xu:2023:LTA

[XF23b] Taiyang Xu and Engui Fan. Large-time asymptotics to the focusing nonlocal modified kortweg-de Vries equation with step-like boundary conditions. Studies in Applied Mathematics, 150(4):1217–1273, May 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12568.

Xiang:2022:LBA

[XHW22] Chuang Xiang, Jicai Huang, and Hao Wang. Linking bifurcation analysis of Holling–Tanner model with generalist predator to a changing environment. Studies in Applied Mathematics, 149(1):124–163, July 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12492.

Xu:2021:ISM

[XJN21] Xiaoxue Xu, Mengmeng Jiang, and Frank W. Nijhoff. Integrable symplectic maps associated with discrete Korteweg– de Vries-type equations. Studies in Applied Mathematics, 146(1):233–278, January 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12346.

Xie:1993:TBM

[XK93] Shisheng Xie and Charles Knessl. On the transient behavior of the M / M / 1 and M / M / 1 / M queues. Studies in Applied Mathematics, 88(3):191–240, April 1993. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1993883191.

Xue:2015:SAP

[XL15] Lingling Xue and Q. P. Liu. A supersymmetric AKNS problem and its Darboux–Bäcklund transformations and discrete systems. Studies in Applied Mathematics, 135(1):35–62, July 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12080.

Xia:2015:STC

[XQZ15] Baoqiang Xia, Zhijun Qiao, and Ruguang Zhou. A synthetical two-component model with peakon solutions. Studies in Applied Mathematics, 135(3):248–276, October 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12085.

Xu:1990:ATS

[Xu90] Jian-Jun Xu. Asymptotic theory of steady axisymmetrical needlelike crystal growth. Studies in Applied Mathematics, 82(1):71–91, January 1990. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm199082177.

Xu:2021:MPC

[Xu21] Meng Xu. On mixed problem of Chaplygin’s hodograph equation in hyperbolic domain near the parabolic degenerate line. Studies in Applied Mathematics, 147(2):675–693, August 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12395.

Xia:2021:ISA

[XXZ21] Jun Xia, Shuai-Xia Xu, and Yu-Qiu Zhao. Isomonodromy sets of accessory parameters for Heun class equations. Studies in Applied Mathematics, 146(4):901–952, May 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12370.

Xu:2022:NOC

[XYZY22] Jing Xu, Zhensheng Yu, Tonghua Zhang, and Sanling Yuan. Near-optimal control of a stochastic model for mountain pine beetles with pesticide application. Studies in Applied Mathematics, 149(3):678–704, October 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12517.

Xu:2011:PXA

[XZ11] Shuai-Xia Xu and Yu-Qiu Zhao. Painlevé XXXIV asymptotics of orthogonal polynomials for the Gaussian weight with a jump at the edge. Studies in Applied Mathematics, 127(1):67–105, July 2011. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2010.00512.x.

Xu:2013:ADP

[XZ13] S.-X. Xu and Y.-Q. Zhao. Asymptotics of discrete Painlevé V transcendents via the Riemann–Hilbert approach. Studies in Applied Mathematics, 130(3):201–231, April 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00573.x.

Yang:2003:CFE

[YA03] J. Yang and T. R. Akylas. Continuous families of embedded solitons in the third-order nonlinear Schrödinger equation. Studies in Applied Mathematics, 111(3):359–375, October 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.t01-1-00238.

Yang:1995:SMC

[Yan95a] Jianke Yang. The sloshing motion of cross-waves. Studies in Applied Mathematics, 95(2):171–191, August 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995952171.

Yang:1995:SNE

[Yan95b] Jianke Yang. The stability and nonlinear evolution of edge waves. Studies in Applied Mathematics, 95(3):229–246, October 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1995953229.

Yang:1996:CSW

[Yan96] Jianke Yang. Coherent structures in weakly birefringent nonlinear optical fibers. Studies in Applied Mathematics, 97(2):127–148, August 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1996972127.

Yang:1997:VST

[Yan97] Jianke Yang. Vector solitons and their internal oscilliations in birefringent nonlinear optical fibers. Studies in Applied Mathematics, 98(1):61–97, January 1997. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00041.

Yang:1998:MPW

[Yan98] Jianke Yang. Multiple permanent-wave trains in nonlinear systems. Studies in Applied Mathematics, 100(2):127–152, February 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9590.00073.

Yang:2001:DES

[Yan01] Jianke Yang. Dynamics of embedded solitons in the extended Korteweg–de Vries equations. Studies in Applied Mathematics, 106(3):337–365, April 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00169.

Yang:2002:ELI

[Yan02] Jianke Yang. Eigenfunctions of linearized integrable equations expanded around an arbitrary solution. Studies in Applied Mathematics, 108(1):145–159, January 2002. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.01428.

Yang:2012:CSW

[Yan12] Jianke Yang. Classification of solitary wave bifurcations in generalized nonlinear Schrödinger equations. Studies in Applied Mathematics, 129(2):133–162, August 2012. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00549.x.

Yan:2014:LAS

[Yan14a] Zhenya Yan. Localized analytical solutions and parameters analysis in the nonlinear dispersive Gross–Pitaevskii meanfield GP (m, n) model with space-modulated nonlinearity and potential. Studies in Applied Mathematics, 132(3):266–284, April 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12028.

Yang:2014:CPT

[Yan14b] Jianke Yang. Can parity-time-symmetric potentials support families of non-parity-time-symmetric solitons? Studies in Applied Mathematics, 132(4):332–353, May 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12032.

Yang:2015:NMC

[Yan15] Jianke Yang. A numerical method for computing timeperiodic solutions in dissipative wave systems. Studies in Applied Mathematics, 134(4):420–455, May 2015. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12071.

Yang:2021:ACS

[Yan21] Jianke Yang. Analytical construction of soliton families in one- and two-dimensional nonlinear Schrödinger equations with nonparity-time-symmetric complex potentials. Studies in Applied Mathematics, 147(1):4–31, July 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12383.

Yang:1996:SPN

[YB96] J. Yang and D. J. Benney. Some properties of nonlinear wave systems. Studies in Applied Mathematics, 96(1):111–139, January 1996. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1002/sapm1996961111.

Yue:2008:MASa

[YBSM08a] Zengyuan Yue, Holger Broich, Florian Seifriz, and Joachim Mester. Mathematical analysis of a soccer game. Part I: Individual and collective behaviors. Studies in Applied Mathematics, 121(3):223–243, October 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2008.00413.x.

Yue:2008:MASb

[YBSM08b] Zengyuan Yue, Holger Broich, Florian Seifriz, and Joachim Mester. Mathematical analysis of a soccer game. Part II: Energy, spectral, and correlation analyses. Studies in Applied Mathematics, 121(3):245–261, October 2008. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00414.x.

Yu:2015:HIM

[YC15] Jie Yu and Siyu Chen. Hydrodynamic instability mechanism for rip currents. Studies in Applied Mathematics, 135(2):196– 223, August 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12074.

Yeh:1995:REI

[Yeh95] Yeong-Nan Yeh. A remarkable endofunction involving compositions. Studies in Applied Mathematics, 95(4):419–432, November 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1995954419.

Yanik:2005:DPS

[YF05] Mehmet F. Yanik and Shanhui Fan. Dynamic photonic structures: Stopping, storage, and time reversal of light. Studies in Applied Mathematics, 115(2):233–253, August 2005. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2005.00327.x.

Yin:2021:LIE

[YF21] Yue Yin and Wei Fu. Linear integral equations and twodimensional Toda systems. Studies in Applied Mathematics, 147(3):1146–1193, October 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12427.

Yu:2001:OLI

[YG01] J. Yu and L. K. Gross. The onset of linear instabilities in a solid combustion model. Studies in Applied Mathematics, 107(1):81–101, July 2001. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
1071179.

Yuan:2018:EIU

[YGJ18] Chunxin Yuan, Roger Grimshaw, and Edward Johnson. The evolution of internal undular bores over a slope in the presence of rotation. Studies in Applied Mathematics, 140(4):465–482, May 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12207.

Yang:1994:UST

[YK94a] Yongzhi Yang and Charles Knessl. The unconditional sojourn time distribution in the queue GI/M/1-K with processor sharing service. Studies in Applied Mathematics, 93(1):29–91, September 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm199493129.

Yu:1994:ISB

[YK94b] J. Yu and J. Kevorkian. The interaction of a strong bore with small disturbances in shallow water. Studies in Applied Mathematics, 91(3):247–273, March 1994. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1994913247.

Yang:1999:HSC

[YK99] Huijun Yang and Ruby Krishnamurti. Hele Shaw convection with imposed shear flows: Boundary-layer formulation. Studies in Applied Mathematics, 103(1):1–24, July 1999. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00117.

Yong:2002:IBV

[YK02a] Darryl H. Yong and J. Kevorkian. Initial boundary-value problems for a pair of conservation laws. Studies in Applied Mathematics, 108(4):351–367, May 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
01429.

Yong:2002:SBV

[YK02b] Darryl H. Yong and J. Kevorkian. Solving boundary-value problems for systems of hyperbolic conservation laws with rapidly varying coefficients. Studies in Applied Mathematics, 108(3):259–303, April 2002. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
01412.

Yao:2008:SQV

[YK08] Haishen Yao and Charles Knessl. On the shortest queue version of the Erlang loss model. Studies in Applied Mathematics, 120(2):129–212, February 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2007.00399.x.

Yue:2007:CEWb

[YKdM^{+}07] Z. Yue, H. Kleinöder, M. de Marées, U. Speicher, P. Wahl, and J. Mester. On the cardiovascular effects of wholebody vibration Part II. Lateral effects: Statistical analysis. Studies in Applied Mathematics, 119(2):111–125, August 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2007.00380.x.

Yu:2000:WNL

[YKH00] J. Yu, J. Kevorkian, and R. Haberman. Weak nonlinear long waves in channel flow with internal dissipation. Studies in Applied Mathematics, 105(2):143–163, August 2000. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.00146.

Yang:2007:UCS

[YL07] Jianke Yang and Taras I. Lakoba. Universally-convergent squared-operator iteration methods for solitary waves in general nonlinear wave equations. Studies in Applied Mathematics, 118(2):153–197, February 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2007.00371.x.

Yang:2008:AIT

[YL08] Jianke Yang and Taras I. Lakoba. Accelerated imaginarytime evolution methods for the computation of solitary waves. Studies in Applied Mathematics, 120(3):265–292, April 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00398.x.

Yu:1995:HAN

[YLP95] J. Yu, D. Lakin, and P. Penar. A hybrid asymptoticnumerical study of a model for intracranial pressure dynamics. Studies in Applied Mathematics, 95(3):247–267, October 1995. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1995953247.

Yu:2021:GFL

[YLW21] Xiao Yu, Kunquan Lan, and Jianhong Wu. Green’s functions, linear second-order differential equations, and one-dimensional diffusion advection models. Studies in Applied Mathematics, 147(1):319–362, July 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12384.

Yue:2004:MAR

[YM04] Z. Yue and J. Mester. A modal analysis of resonance during the whole-body vibration. Studies in Applied Mathematics, 112(3):293–314, April 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
0022-2526.2004.01513.x.

Yue:2007:CEWa

[YM07] Z. Yue and J. Mester. On the cardiovascular effects of wholebody vibration Part I. Longitudinal effects: Hydrodynamic analysis. Studies in Applied Mathematics, 119(2):95–109, August 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/j.1467-9590.2007.00379.x.

Yang:2004:DQS

[YMBC04] Jianke Yang, Igor Makasyuk, Anna Bezryadina, and Zhigang Chen. Dipole and quadrupole solitons in optically induced two-dimensional photonic lattices: Theory and experiment. Studies in Applied Mathematics, 113(4):389–412, November 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.0022-2526.2004.01539.x.

You:2013:USG

[You13] Y. You. Upper-semicontinuity of global attractors for reversible schnackenberg equations. Studies in Applied Mathematics, 130(3):232–263, April 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2012.00565.x.

Yang:2020:SBD

[YRZ20] Kai Yang, Svetlana Roudenko, and Yanxiang Zhao. Stable blow-up dynamics in the L^{2}-critical and L^{2}-supercritical generalized Hartree equation. Studies in Applied Mathematics, 145(4):647–695, November 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12328.

Yam:1998:EST

[YS98] Yeung Yam and Jingyan Song. Extending Shannon’s theorem to a general juggling pattern. Studies in Applied Mathematics, 100(1):53–66, January 1998. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9590.
00069.

Yang:2021:ESS

[YS21] Chen Yang and Huaqing Sun. Essential spectra of singular Hamiltonian differential operators of arbitrary order under a class of perturbations. Studies in Applied Mathematics, 147(1):209–229, July 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/sapm.12388.

Ye:2014:MHS

[YT14] Yiwei Ye and Chun-Lei Tang. Multiple homoclinic solutions for second-order perturbed Hamiltonian systems. Studies in Applied Mathematics, 132(2):112–137, February 2014. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12023.

Yu:2010:EFW

[Yu10] J. Yu. Effects of finite water depth on natural frequencies of suspended water tanks. Studies in Applied Mathematics, 125(4):373–391, November 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2010.00492.x.

Yu:2015:DAN

[Yu15] G.-F. Yu. Discrete analogues of a negative order AKNS equation. Studies in Applied Mathematics, 135(2):117–138, August 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12082.

Yueh:1981:NWS

[Yue81] T. Y. Yueh. Nonlinear waves in a self-gravitating medium. Studies in Applied Mathematics, 65(1):1–36, August 1981. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm19816511.

Yue:1992:TSK

[YW92] Z. Y. Yue and G. Winnewisser. On the temporal and the spatial Kelvin–Helmholtz instability of a hypersonic jet confined by a longitudinal magnetic field. Studies in Applied Mathematics, 87(4):325–349, November 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1992874325.

Yu:2003:ESP

[YY03] Jun Yu and Yi Yang. Evolution of small periodic disturbances into roll waves in channel flow with internal dissipation. Studies in Applied Mathematics, 111(1):1–27, July 2003. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9590.t01-2-00225.

Yang:2004:CMW

[YY04] Bo-Yin Yang and Yeong-Nan Yeh. A crowning moment for Wiener indices. Studies in Applied Mathematics, 112(4):333– 340, May 2004. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.0022-2526.2004.01516.
x.

Yan:2008:MDP

[YY08] Weigen Yan and Yeong-Nan Yeh. On the monomer–dimer problem of some graphs. Studies in Applied Mathematics, 120(4):351–360, May 2008. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2008.00404.x.

Yang:2018:TBN

[YY18] Bo Yang and Jianke Yang. Transformations between nonlocal and local integrable equations. Studies in Applied Mathematics, 140(2):178–201, February 2018. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12195.

Yang:2023:RWP

[YY23] Bo Yang and Jianke Yang. Rogue wave patterns associated with Okamoto polynomial hierarchies. Studies in Applied Mathematics, 151(1):60–115, July 2023. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12573.

Ye:2020:GSS

[YZ20] Rusuo Ye and Yi Zhang. General soliton solutions to a reverse-time nonlocal nonlinear Schrödinger equation. Studies in Applied Mathematics, 145(2):197–216, August 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12317.

Zakharov:2009:TIS

[Zak09] Vladimir E. Zakharov. Turbulence in integrable systems. Studies in Applied Mathematics, 122(3):219–234, April 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00430.x.

Zauderer:1972:BVP

[Zau72] Erich Zauderer. Boundary value problems for a second order differential equation with a turning point. Studies in Applied Mathematics, 51(4):411–413, December 1972. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1972514411.

Zhou:1992:ITD

[ZB92] Y. F. Zhou and D. J. Benney. Instability of three dimensional waves in two layered flows. Studies in Applied Mathematics, 86(4):269–280, May 1992. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1992864269.

Zhang:2010:SALa

[ZC10a] Da-Jun Zhang and Shou-Ting Chen. Symmetries for the Ablowitz–Ladik hierarchy: Part I. Four-potential case. Studies in Applied Mathematics, 125(4):393–418, November 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00493.x.

Zhang:2010:SALb

[ZC10b] Da-Jun Zhang and Shou-Ting Chen. Symmetries for the Ablowitz–Ladik hierarchy: Part II. Integrable discrete nonlinear Schrödinger equations and discrete AKNS hierarchy. Studies in Applied Mathematics, 125(4):419–443, November 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2010.00494.x.

Zaug:2021:DMS

[ZC21] Camille R. Zaug and John D. Carter. Dissipative models of swell propagation across the Pacific. Studies in Applied Mathematics, 147(4):1519–1537, November 2021. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12423.

Zeng:1991:EML

[Zen91] Jiang Zeng. The β-extension of the multivariable Lagrange inversion formula. Studies in Applied Mathematics, 84(2):167–182, February 1991. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1002/sapm1991842167.

Zhang:2012:ODB

[ZGQ12] Yunxin Zhang, Hao Ge, and Hong Qian. One-dimensional birth-death process and Delbrück-Gillespie theory of mesoscopic nonlinear chemical reactions. Studies in Applied Mathematics, 129(3):328–345, October 2012. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/j.
1467-9590.2012.00567.x.

Zhang:2019:DBS

[Zha19] Cheng Zhang. Dressing the boundary: On soliton solutions of the nonlinear Schrödinger equation on the half-line. Studies in Applied Mathematics, 142(2):190–212, February 2019. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12248.

Zheng:2020:VSC

[Zhe20] Junzhi Zheng. Viscosity solutions to a Cauchy problem of a phase-field model for solid-solid phase transitions. Studies in Applied Mathematics, 145(3):483–499, October 2020. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12325.

Zhou:2009:FDI

[Zho09] Ruguang Zhou. Finite-dimensional integrable Hamiltonian systems related to the nonlinear Schrödinger equation. Studies in Applied Mathematics, 123(4):311–335, November 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00452.x.

Zhou:2018:DTG

[Zho18] Zi-Xiang Zhou. Darboux transformations and global explicit solutions for nonlocal Davey–Stewartson I equation. Studies in Applied Mathematics, 141(2):186–204, August 2018. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12219.

Zhu:2009:SMA

[ZHY09] Yi Zhu, Richard Haberman, and Jianke Yang. Separatrix map analysis for fractal scatterings in weak interactions of solitary waves. Studies in Applied Mathematics, 122(4):449–483, May 2009. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2009.00442.x.

Zhen:2013:AAS

[ZK13] Qiang Zhen and Charles Knessl. Asymptotic analysis of spectral properties of finite capacity processor shared queues. Studies in Applied Mathematics, 131(2):179–210, August 2013. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12020.

Zhong:2017:ASK

[ZL17] X. X. Zhong and S. J. Liao. Analytic solutions of von Kármán plate under arbitrary uniform pressure — equations in differential form. Studies in Applied Mathematics, 138(4):371–400, May 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12158.

Zhang:2021:LSS

[ZL21] Zhi-Yong Zhang and Zhi-Xiang Lin. Local symmetry structure and potential symmetries of time-fractional partial differential equations. Studies in Applied Mathematics, 147(1):363–389, July 2021. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-

9590 (electronic). URL https://onlinelibrary.wiley.com/
doi/epdf/10.1111/sapm.12374.

Liu:2022:FLE

[zLWjZ22] Shu zhi Liu, Jing Wang, and Da jun Zhang. The Fokas– Lenells equations: Bilinear approach. Studies in Applied Mathematics, 148(2):651–688, February 2022. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12454.

Zhang:2017:DSE

[ZN17] F. W. Zhang and L. F. Nie. Dynamics of SIS epidemic model with varying total population and multivaccination control strategies. Studies in Applied Mathematics, 139(4):533–550, November 2017. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12173.

Zufiria:1986:ESE

[ZS86a] J. A. Zufiria and P. G. Saffman. An example of stability exchange in a Hamiltonian wave system. Studies in Applied Mathematics, 74(1):85–91, February 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm198674185.

Zufiria:1986:SIF

[ZS86b] J. A. Zufiria and P. G. Saffman. The superharmonic instability of finite-amplitude surface waves on water of finite depth. Studies in Applied Mathematics, 74(3):259–266, June 1986. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1002/sapm1986743259.

Zhang:2022:IBV

[ZS22] Hongwei Zhang and Xiao Su. Initial boundary value problem for a class of wave equations of Hartree type. Studies in Applied Mathematics, 149(3):798–814, October 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12521.

Zezyulin:2022:NCF

[ZSA22] D. A. Zezyulin, A. O. Slobodyanyuk, and G. L. Alfimov. On nonexistence of continuous families of stationary nonlinear modes for a class of complex potentials. Studies in Applied Mathematics, 148(1):99–124, January 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12432.

Zhang:2020:RMH

[ZTYH20] Yongshuai Zhang, Xiangxing Tao, Tengteng Yao, and Jingsong He. The regularity of the multiple higher-order poles solitons of the NLS equation. Studies in Applied Mathematics, 145(4):812–827, November 2020. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12338.

Zeng:2015:PIA

[ZXZ15] Zhao-Yun Zeng, Shuai-Xia Xu, and Yu-Qiu Zhao. Painlevé III asymptotics of Hankel determinants for a perturbed Jacobi weight. Studies in Applied Mathematics, 135(4):347–376, November 2015. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12090.

Zhang:2022:DAS

[ZYZ22] Shengqiang Zhang, Sanling Yuan, and Tonghua Zhang. Dynamic analysis of a stochastic eco-epidemiological model with disease in predators. Studies in Applied Mathematics, 149(1): 5–42, July 2022. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/sapm.12489.

Zhou:2007:IAE

[ZZ07] Jian-Rong Zhou and Yu-Qiu Zhao. An infinite asymptotic expansion for the extreme zeros of the Pollaczek polynomials. Studies in Applied Mathematics, 118(3):255–279, April 2007. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590

(electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2007.00373.x.

Zhao:2008:QAT

[ZZ08] Yan Zhao and Weiguo Zhang. Qualitative analysis and travelling wave solutions for the generalized Pochhammer– Chree equation with a dissipation term. Studies in Applied Mathematics, 121(4):369–394, November 2008. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2008.00420.x.

Zhang:2013:SAL

[ZZ13] D.-J. Zhang and S.-L. Zhao. Solutions to ABS lattice equations via generalized Cauchy matrix approach. Studies in Applied Mathematics, 131(1):72–103, July 2013. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/sapm.12007.

Zhang:2012:DLE

[ZZN12] D.-J. Zhang, S.-L. Zhao, and F. W. Nijhoff. Direct linearization of extended lattice BSQ systems. Studies in Applied Mathematics, 129(2):220–248, August 2012. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9590.2012.00552.x.

Zhu:2010:ESI

[ZZZ10] Z.-N. Zhu, H.-Q. Zhao, and F.-F. Zhang. Explicit solutions to an integrable lattice. Studies in Applied Mathematics, 125(1):55–67, July 2010. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-
9590.2010.00481.x.